Die Pakete

0.1 Package descriptions

This chapter contains all descriptions of packages located on our CD. Some packages are not described whereas other descriptions refer to ”older” versions than on our CD. This is due to time of going to press of the book which (of technical reasons) is approximately one week before that of the CD. We want to apologize for this inconvenience. Documentation on CD is more complete and more actual than the printed version anyway. In case of doubt you should look there first!

Moreover all descriptions of packages located in zq are missing. Series zq contains sources of those packages in series z. Names of source packages do correspond to the binary packages except that there is an s added to the end for making classification a little easier.

0.1.1 A: Linux Base System (You need it!)

aaa base S.u.S.E. Linux base package

This package installs several important configuration files. Central scripts like SuSEconfig are also in this package.

Contents: aaa base fillup skel

Copyright:  (c) 1996,97 S.u.S.E. GmbH Fuerth, Germany

Author:

Burchard Steinbild <bs@suse.de>, Florian La Roche <florian@suse.de>, Werner Fink <werner@suse.de>, Martin Scherbaum <maddin@suse.de>

Version:  98.3.30-0

aaa dir S.u.S.E. Linux directory structure

This package installs the directory structure. It also includes the home directories of the system users.

Contents: aaa dir

Copyright:  (c) 1997 S.u.S.E. GmbH Fuerth, Germany
Version:  98.3.26-0

base Some important GNU packages

This package normally also includes the program ”rmt”, which provides remote tape drive control. Since there is a compatible ’rmt’ in the ’dump’ package we didn’t put ’rmt’ into this package. If you are planning to use the remote tape features provided by tar you have to install the ’dump’ package as well.

Contents: freeramdisk grep-2.0 hdparm-3.3 ktools m4-1.4 sed-2.05 tar-1.11.8 time-1.7 which

Author:

Francois Pinard <pinard@iro.umontreal.ca>, Tom Lord <lord@cygnus.com>, David J. MacKenzie <djm@gnu.ai.mit.edu>, Akiko Matsushita <matusita@sra.co.jp>, Alan Magnuson <awm@osc.edu>, Alexander Lehmann <alex@hal.rhein-main.de>, Amos Shapira <amoss@cs.huji.ac.il>, Andreas Gustafsson <gson@niksula.hut.fi>, Andreas Schwab <schwab@issan.informatik.uni-dortmund.de>, Assar Westerlund <assar@nada.kth.se>, Ben A. Mesander <ben@piglet.cr.usgs.gov>, Bengt Mertensson <bengt@mathematik.uni-bremen.de>, Bernhard Daeubler <Bernhard.Daeubler@physik.uni-ulm.de>, Bjorn R. Bjornsson <brb@falcon.is>, Brendan Kehoe <brendan@cygnus.com>, Erez Zadok <ezk@cs.columbia.edu>, Eric Allman <sendmail@sendmail.ORG>, Eric Backus <ericb@lsid.hp.com>, Geoff Russell <grussell@guest.adelaide.edu.au>, Greg A. Woods <woods@web.apc.org>, Hal Peterson <hrp@pecan.cray.com>, Hoang Uong <hoang@ornews.intel.com>, Ian Lance Taylor <ian@cygnus.com>, Jason Merrill <jason@cygnus.com>, Jim Avera <jima@netcom.com>, Jim Kingdon <kingdon@cygnus.com>, Jim Meyering <meyering@na-net.ornl.gov>, Joel Sherrill <jsherril@uahcs2.cs.uah.edu>, John David Anglin <dave@hiauly1.hia.nrc.ca>, Joseph E. Sacco <jsacco@ssl.com>, Karl Berry <karl@cs.umb.edu>, Karl Vogel <vogelke@c-17igp.wpafb.af.mil>, Kaveh R. Ghazi <ghazi@noc.rutgers.edu>, Kristine Lund <lund@lpnaxp.in2p3.fr>, Krste Asanovic <krste@icsi.berkeley.edu>, Marion Hakanson <hakanson@cse.ogi.edu>, Mark Seiden <mis@seiden.com>, Matthias Rabe <rabe@mathematik.unibielefeld.de>, Michael Fetterman <mafetter@ichips.intel.com>, Michael L. Welcome <welcome@bigbird.llnl.gov>, Mike Lijewski <lijewski@theory.tc.cornell.edu>, Nick S. Kanakakor <skanan@otl.scu.edu>, Nicolas Pioch <pioch@inf.enst.fr>, Noah Friedman <friedman@gnu.ai.mit.edu>, Pete Chown <pete.chown@dale.dircon.co.uk>, Pierre Gaumond <gaumondp@ere.umontreal.ca>, Pierre Mathieu <mathieu@geod.emr.ca>, Robert Bernstein <rocky@panix.com>, Rene’ Seindal <seindal@kb.dk>, Richard Stallman <rms@gnu.ai.mit.edu>, Roland H. Pesch <roland@wrs.com>, Roland McGrath <roland@gnu.ai.mit.edu>, Scott Bartram , Simon Leinen <simon@lia.di.epfl.ch>, Skip Montanaro <skip@automatrix.com>, Stephen Perkins <perkins@cps.msu.edu>, Steve Williamson <willy@uinpla.npl.uiuc.edu>, Thorsten Ohl <ohl@physics.harvard.edu>, Tom Mc-Connell <tmcconne@sedona.intel.com>, Tom Quinn <trq@dionysos.thphys.ox.ac.uk>, Tom Tromey <tromey@cns.caltech.edu>, Ulrich Drepper <drepper@gnu.ai.mit.edu>, Vern Paxson <vern@ee.lbl.gov>, Vivek P. Singhal <singhal@cs.utexas.edu>, Walter Wong <wcw+@cmu.edu>

Version:  98.2.10-2

bash The GNU Bourne-Again Shell

Bash is a sh-compatible command interpreter that executes commands read from standard input or from a file. Bash also incorporates useful features from the Korn and C shells (ksh and csh). Bash is ultimately intended to be a conformant implementation of the IEEE Posix Shell and Tools specification (IEEE Working Group 1003.2). Contents: bash-2.01.1 readline-2.1

Copyright:  (c) 1989 Free Software Foundation, Inc.
Author:  Chet Ramey <chet@po.cwru.edu>
Version:  2.01.1-8

bash1 The GNU Bourne-Again Shell (older version)

Since the new bash, version 2.x, is somewhat stricter in its syntax checking than version 1.x, we are keeping the older version as bash1 in the distribution. If you have scripts that don’t satisfy the stricter rules of 2.x and you don’t want to look for the errors, simply replace ”/bin/sh” with ”/bin/bash1” in the first line. The script will then be run by the old bash shell. Contents: bash-1.14.7

Copyright:  (c) 1989 Free Software Foundation, Inc.
Author:  Chet Ramey <chet@po.cwru.edu>
Version:  1.14.7-8

bdflush Linux update daemon

Bdflush is used to start the kernel daemon to flush dirty buffers back to disk.

A: Linux Base System (You need it!)

Contents: bdflush-1.5

Author:

Eric Youngdale <eric@aib.com>, Phil Bostley <bostley@cs.colorado.edu>, Daniel Quinlan <quinlan@yggdrasil.com>

Version:  1.5-8

compress Standard Un*x compression program

Author:  Peter Jannesen <peter@ncs.nl>
Version:  4.2.4-8

cpio The GNU cpio backup and archiving utility

This is GNU cpio, a program to manage archives of files. This package also includes mt, a tape drive control program. Cpio copies files into or out of a cpio or tar archive. An archive is a file that contains other files plus information about them, such as their pathname, owner, timestamps, and access permissions. The archive can be another file on the disk, a magnetic tape, or a pipe. This package normally also includes the program ”rmt”, which provides remote tape drive control. Since there is a compatible ’rmt’ in the ’dump’ package we didn’t put ’rmt’ into this package. If you are planning to use the remote tape features provided by cpio you have to install the ’dump’ package as well. Contents: cpio-2.4.2

Copyright:  (c) 1989, 1990, 1991, 1992, 1993 Free Software Foundation, Inc.

Author:

David J. MacKenzie <djm@gnu.ai.mit.edu>, Jim Meyering <meyering@nanet.ornl.gov>

Version:  2.4.2-10

cron cron daemon

Automatically starts programs at specific times. Add new entries with ”crontab -e”. (See ”man 5 crontab” and ”man 1 crontab” for documentation.) Contents: cron-3.0.1

Author:  Paul Vixie <paul@vix.com>
Version:  3.0.1-10

db BSD database

Static and shared libraries for the BSD database routines. Contents: db.1.85

Version:  1.85-8

devs Device files

This package creates special files in the /dev directory, so that your system’s hardware can be accessed.

Author:  Hubert Mantel <mantel@suse.de>
Version:  98.2.5-2

diff GNU diffutils

The GNU diff utilities find differences between files (to make source code patches, for instance). Contents: diffutils-2.7

Copyright:  (c) 1985-94 Free Software Foundation, Inc.

Author:

Paul Eggert <eggert@twinsun.com>, David J. MacKenzie <djm@gnu.ai.mit.edu>, Richard Stallman <rms@gnu.ai.mit.edu>, Roland McGrath <roland@gnu.ai.mit.edu>

Version:  2.7-8

eazy Compatibility links

This package contains some links and enhancements of the /etc files which should make it easier to install software packages from the Internet. A typical example of where you can use this package is the link from /usr/sbin/sendmail to /usr/lib/sendmail.

Version:  98.2.9-2

elvis A vi-clone called elvis

A vi clone, the standard text editor on Unix systems. Contents: elvis-2.0

Copyright:  (c) Steve Kirkendall
Author:  Steve Kirkendall <kirkenda@cs.pdx.edu>
Version:  2.0-9

ext2fs Utilities for the second extended file system

Utilities needed to create and maintain ext2 filesystems under Linux. Included in this package are: chattr, lsattr, mke2fs, mklost+found, tune2fs, e2fsck, and badblocks.

Copyright:  (c) Remy Card, Theodore Ts’o
Author:  Remy Card <card@masi.ibp.fr>, Theodore Ts’o <tytso@mit.edu>
Version:  1.10-10

file determine file type

With ”file” you can obtain information on the file type of a specified file. File type recognition is controlled by the file /etc/magic, which contains the classification criteria. This command is used by apsfilter to permit automatic printing of different file types. Contents: file-3.22

Author:  Mark Moraes <moraes@deshaw.com>, Christos Zoulas <christos@deshaw.com>
Version:  3.22-9

fileutil Basic programs such as cp, mkdir,ln, ls

GNU fileutils: the basic file manipulation programs Contents: fileutils-3.16

Copyright:  (c) Patrick Volkerding
Author:  Jim Meyering <meyering@na-net.ornl.gov>
Version:  3.16-9

find GNU find

This package contains the GNU find, xargs, and locate programs. The find and xargs programs comply with POSIX 1003.2. They also support some additional options, some borrowed from Unix and some unique to GNU. Contents: findutils-4.1

Copyright:  (c) 1990,91,92,93,94 Free Software Foundation, Inc.

Author:

David J. MacKenzie <djm@gnu.ai.mit.edu>, Jan Brittenson <bson@gnu.ai.mit.edu>, Brian L. Matthews <blm@6sceng.UUCP>, Mike Rendell <michael@cs.mun.ca>, James A. Woods <jwoods@adobe.com>, Jay Plett <jay@silence.princeton.nj.us>, Tim Wood <tim@toad.com>

Version:  4.1-10

gawk GNU gawk

GNU gawk. It is upwardly compatible with the System V Release 4 awk. It is almost completely compliant with POSIX 1003.2. Contents: gawk-3.0.3

Copyright:  (c) 1989, 1991 - 1995 Free Software Foundation

Author:

David Trueman <david@cs.dal.ca>, Arnold Robbins <arnold@skeeve.atl.ga.us>, Michal Jaegermann <michal@gorte.phys.ualberta.ca>, Scott Deifik <scottd@amgen.com>, Darrel Hankerson <hankedr@mail.auburn.edu>, Kai Uwe Rommel <rommel@ars.de>, Pat Rankin <rankin@eql.caltech.edu>

Version:  3.0.3-13

gdbm GNU database routines

Static and dynamic library for the GNU database routines. Contents: gdbm-1.7.3

Version:  1.7.3-8

gppshare C++ shared libraries

Needed for dynamically linked C++ programs. The sources are available in the libgpp s package.

Author:  H.J. Lu <hjl@nynexst.com>, Per Bothner <bothner@cygnus.com>
Version:  2.7.2.1-1

groff GNU troff document formatting system.

The groff package provides compatible versions of troff, nroff, eqn, tbl, and other Unix text-formatting utilities. Groff is used to ”compile” man-pages stored in groff/nroff format for different output devices; e.g., for displaying to a screen or in PostScript(tm) format for printing on a PostScript(tm) printer. Most programs store their man-pages in either /usr/man/ or /usr/X11R6/man/.

Contents: groff-1.10

Copyright:  (c) 1989-94 Free Software Foundation, Inc.
Author:  James Clark <jjc@clark.com>
Version:  1.10-9

gzip GNU zip compression utilities

Gzip reduces the size of the named files using Lempel-Ziv coding LZ77. Whenever possible, each file is replaced by one with the extension .gz, while keeping the same ownership modes, access and modification times. Contents: gzip-1.2.4

Copyright:  (c) 1992-1993 Jean-loup Gailly
Author:  Jean-loup Gailly <gzip@prep.ai.mit.edu>
Version:  1.2.4-9

kbd Keyboard and font utilities

Load and save keyboard mappings. Needed if you are not using the US keyboard map. This package also contains utilities for changing your console fonts. If you install this package, YaST will include an extra menu to allow you to choose between the different fonts. This package additionally includes fonts from the kbd fonts.tar.gz package on Sunsite. Contents: kbd-0.93 kbd fonts

Copyright:  (c) 1994 Andries Brouwer
Author:  Andries Brouwer <aeb@cwi.nl>
Version:  0.93-7

kernmod Loadable kernel modules

Most Linux drivers are available as modules, they can be loaded at run-time. So, if you have installed a modular kernel the installation of this package is highly suggested. Please refer to the chapter about the kernel daemon in the accompanying handbook. Contents: VideoteXt-0.6 ibcs-2.0.970513 jazz-2.1d joystick-0.8.0 xosview-1.5.1

Copyright:  (c) Linus Torvalds et al
Author:  Linus Torvalds <torvalds@cs.helsinki.fi>
Version:  2.0.33-5

ldso ld.so, the dynamic linker/loader.

ld.so completes the final process of linking all necessary references to sharable objects and unreferenced symbols in an impure executable, (usually a dynamically linked executable), to produce a runnable file. Nearly all Linux binaries are ”incomplete” in the sense that they require further linking at run time. ld.so’s job is to thus complete the linking process started at compile time. Contents: ld.so-1.9.6

Copyright:  (c) 1995 David Engel, Rickard E. Faith,Yggdrasil Computing Inc.
Version:  1.9.6-7

less less - opposite of more

Less is a pager program similar to more (1), but allows backward as well as forward movement within a file. Also, less does not have to read the entire input file before starting. It is also possible to start an editor at any time from this program. Contents: less-332

Author:  Mark Nudelman <markn@3do.com>
Version:  332-10

libaout Libraries for the old a.out format.

Linux has changed its default binary format from a.out to elf. To be able to run older binary programs, you will need to install these old a.out shared libraries. Contents: libc-4.7.5.bin libc-4.7.5 libc-4.7.6

Author:  H.J. Lu <hjl@nynexst.com>
Version:  4.7.6-2

lilo LILO

The LInux-LOader: LILO boots Linux from your hard drive. It can also boot other operating systems such as MS-DOS and OS/2, and can even boot DOS from the second hard drive. The configuration file is /etc/lilo.conf.

Contents: lilo-20

Copyright:  (c) 1992-1997 Werner Almesberger
Author:  Werner Almesberger <almesber@di.epfl.ch>
Version:  20-11

loadlin LOADLIN

LOADLIN.EXE is a loader (running under DOS) for LINUX kernel images which fully supports the kernel’s command line feature. LOADLIN (and BOOTLIN) are the safest ways to boot Linux from your hard disk. If you have DOS (in addition to Linux) on your machine, LOADLIN is the right tool for booting Linux.

In contrast to LILO, LOADLIN does not modify the master boot record. Also, with LOADLIN, the physical location of the kernel image is not important. With LILO, an update must be run each time the kernel file is moved on the media (i.e. whenever you recompile the kernel or restore the kernel image from a backup medium). This means even if you have to restore your whole Linux filesystem because of a crash, you can still reboot the system with LOADLIN.

All you need is LOADLIN.EXE and a kernel image (zImage) copied to your DOS-partition.

Another good reason to install LOADLIN is that you can install a mini emergency-system on a UMSDOS partition (such as MiniLinux). Then, you can always boot this system with LOADLIN and you’ll be able to repair your main Linux system in those rare cases where you can’t boot (i.e. problems with your rc scripts or you simply forgot the root password).

If you haven’t yet installed LOADLIN by means of the DOS-Setup, you have to unpack the file lodlin16.zip into your DOS-partition and copy the kernel image to the same place.

Assuming your DOS C: partition is mounted as /dos, proceed as follows:

cd /dos

mkdir loadlin

cd loadlin

unzip -d /root/lodlin15.zip

cp -p /vmlinuz .

Copyright:  (c) 1994,95,96 Hans Lermen
Author:  Hans Lermen <lermen@suse.de>
Version:  1.6a-1

localedb database for ’locale’

Version:  1.99-2

man Program for displaying man pages.

Program for displaying man-pages on the screen or sending them to a printer (using groff).

Contents: man db-2.3.10

Author:  G. Wilford <G.Wilford@ee.surrey.ac.uk>
Version:  2.3.10-9

modules Load modules into the kernel

Utilities for the loading of kernel modules. Included are ’insmod’, ’lsmod’, ’rmmod’, ’depmod’, ’modprobe’. The configuration file /etc/conf.modules can be used to pass parameters to the modules; for this reason ’modprobe’ is recommended for loading modules.

Contents: modules-2.1.13 modutils-2.1.85

Author:

Bjorn Ekwall <bj0rn@blox.se> , Jon Tombs <jon@gtex02.us.es> , Bas Laarhoven <bas@vimec.nl>

Version:  2.1.85-1

ncurses New curses libraries

As soon as a text application needs to directly control its output to the screen (e.g. wants to place the cursor at location x y and then write text), ncurses is used. The panel and the forms libraries are included in this package. These new libraries support color, special characters and panels.

Contents: ncurses-4.2

Copyright:  (c) 1992-1995 Zeyd M. Ben-Halim, Eric S. Raymond

Author:

Zeyd M. Ben-Halim <zmbenhal@netcom.com> , Eric S. Raymond <esr@thyrsus.com>, Thomas Dickey <dickey@clark.net>, Juergen Pfeifer <Juergen.Pfeifer@t-online.de>

Version:  4.2-1

net tool Important programs for networking.

This package contains essential programs for network administration and maintenance: netstat, hostname, arp, ifconfig, rarp and route. Contents: net-tools-1.32-alpha

Copyright:

(c) 1988-1993 MicroWalt Corporation, 1993-1994 Swansea University Computer Society, 1995-1996 Bernd Eckenfels

Author:

Fred N. van Kempen <waltje@uwalt.nl.mugnet.org>, Bernd Eckenfels <nettools@lina.inka.de>, Fred Baumgarten <dc6iq@insu1.etec.uni-karlsruhe.de>, Peter Tobias <tobias@et-inf.fho-emden.de>

Version:  1.32alpha-7

netcfg Network configuration files in /etc

All of the basic configuration files for the network programs including /etc/inetd.conf, /etc/protocols and /etc/services. These are often used by network routines in the C-library and therefore must be installed for all network programs. Contents: netcfg

Version:  98.3.3-0

nkita A collection of basic network programs.

Attention! Former releases of this package included dip and sliplogin. These programs have moved to package slip. Contents: bootp-DD2.4.3 bootpc.v061 bwnfsd icmpinfo-1.10 linux pcnfsd2 netdate nfs-server-2.2beta29 pidentd-2.7.4 tcp wrappers 7.6 traceroute-1.4a5 ytalk-3.0.3-8bit

Author:

Florian La Roche <florian@suse.de>, Gordon W. Ross <gwr@mc.com>, Jon Peatfield <J.S.Peatfield@damtp.cam.ac.uk>, Fred N. van Kempen <waltje@uwalt.nl.mugnet.org>, Laurent Demailly <dl@hplyot.obspm.fr>, Alexander Bezprozvanny <bez@parsy.spb.su>, Donald Becker <becker@cesdis.gsfc.nasa.gov>, Peter Eriksson <pen@lysator.liu.se>, Sven Goldt <goldt@pegasus.in-berlin.de>, Jan Wedekind <jan@todonix.ping.de>, Wietse Venema <wietse@wzv.win.tue.nl>

Version:  98.3.25-0

nkitb A collection of net-programs like ftp, telnet,...

A collection of network Utilities. Contents: NetKit-0.09 NetKit-B-0.06 linux-ftpd-0.09-shadow portmap 5beta routed-8.10

Author:  Florian La Roche <florian@suse.de>
Version:  97.10.24-4

nvi ’vi’ clone

A freely redistributable replacement for the Berkeley ex and vi text editors. Contents: nvi-1.79

Author:  Keith Bostic <bostic@cs.berkeley.edu>
Version:  1.79-9

pcmcia PCMCIA support for Linux

This package contains the kernel modules for supporting the following PCMCIA cards: Ethernet: 3Com 3c589, 3c589B, 3c589C, Accton EN2212, EN2216 EtherCard, Allied Telesis Ethernet, AST 1082 Ethernet, CNet CN30BC, CN40BC Ethernet, Datatrek NetCard, Digital EtherWORKS Turbo Ethernet, D-Link DE-650, Edimax Technology Ethernet Combo, EFA InfoExpress SPT EFA 205 10baseT, Eiger Labs EPX-ET10T2 Combo, EP-210 Ethernet, Epson Ethernet, Explorer NE-10000 Ethernet, Farallon Etherwave, Genius ME3000II Ethernet, Grey Cell Ethernet, GVC NIC-2000P Ethernet Combo, Hypertec HyperEnet, IBM CreditCard Ethernet Adapter, IC-Card Ethernet, Infotel IN650ct Ethernet, Katron PE-520 Ethernet, Kingston KNE-PCM/M, LANEED Ethernet, Linksys EtherCard, Longshine Ethernet, Maxtech PCN2000 Ethernet, NDC Instant-Link, Network General “Sniffer”, New Media Ethernet, Novell/National NE4100 InfoMover, PreMax PE-200 Ethernet, Proteon Ethernet, RPTI EP400 Ethernet, SCM Ethernet, Socket Communications Socket EA LAN Adapter, Thomas-Conrad Ethernet, Trust Ethernet Combo, Volktek Ethernet Modem cards: All should work. Memory cards: New Media SRAM, Epson 2MB SRAM, Intel Series 2 and Series 2+ Flash SCSI adapter: Qlogic FastSCSI, New Media Bus Toaster SCSI, Adaptec APA-1460 SlimSCSI The following cards don’t work: Xircom ethernet and ethernet/modem cards, Canon/Compaq PCMCIA floppy drive Under construction: Adaptec/Trantor APA-460 SlimSCSI (David Hinds <dhinds@allegro.stanford.e du>), New Media .WAVjammer (Roger Pao <rpao@paonet.org>), Nat’l Inst DAQ-700 (Tim Raymund <traymund@arlut.utexas.edu>), PreMax SC100 SCSI (W.Kevin Pedigo <pedigo@cyberspace.com>), IBM Token Ring (Steve Kipisz <kipisz@vnet.ibm.com>), AT&T GIS / NCR WaveLAN (Anthony Joseph <adj@pdos.lcs.mit.edu>)

Contents: pcmcia-cs-3.0.0

Copyright:  (c) 1994, 1995, 1996 David A. Hinds
Author:  David Hinds <dhinds@hyper.stanford.edu>
Version:  3.0.0-0

perl Perl interpreter

What would Linux/Unix be without Perl ;-) To be able to use perl, you need to generate the system-dependent perl-include files. As root, on an installed system, simply enter: cd /usr/include ; h2ph *.h sys/*.h linux/*.h

Contents: perl5.004 04

Author:  Larry Wall, Louis J. LaBash, Jr. <llabash@siue.edu>
Version:  5.004 04-6

ps procps

Utilities for displaying process and memory information. Also includes Steve ”Mr. Bassman” Bryant’s enhanced ’w’.

Contents: procinfo-12 procps-1.2.6 psmisc

Author:

Helmut Geyer <Helmut.Geyer@iwr.uni-heidelberg.de>, Charles L. Blake <cblake@bbn.com>, Michael K. Johnson <johnsonm@merengue.oit.unc.edu>, Werner Almesberger <almesber@di.epfl.ch>, Branko Lankester <branko@hacktic.nl>, Robert Nation <nation@rocket.sanders.lockheed.com>, David Engel, Larry Greenfield <greenfie@gauss.rutgers.edu>, Roger Binns, Brian Edmonds, Steve ”Mr. Bassman” Bryant <bassman@muttley.soc.staffs.ac.uk>

Version:  98.3.13-0

rpm rpm - Red Hat Package Manager

rpm (Red Hat Package Manager) is the main tool for managing software packages of the S.u.S.E Linux distribution. rpm can be used to install and remove software packages; with rpm it’s easy to update packages. rpm keep track of all these manipulations in a central database. This way it is possible to get an overview of all installed packages; rpm also supports database queries.

Contents: RPM-HOWTO.html RPM-Tips.html rpm-2.4.12

Copyright:  (c) GPL
Author:  Erik Troan <ewt@redhat.com>, Marc Ewing <marc@redhat.com>
Version:  2.4.12-3

sh utils GNU shellutils

The programs in this package include: basename date dirname echo env expr false logname nice hohup pathchk printenv printf sleep stty tee test true who whoami yes tty uname Contents: sh-utils-1.12 sh-utils-1.16

Copyright:  (c) 1985-94 Free Software Foundation, Inc.

Author:

David J. MacKenzie <djm@gnu.ai.mit.edu>, Jim Meyering <meyering@nanet.ornl.gov>

Version:  1.16-8

shadow Shadow password suite

This package is configured to prevent root login from other then local consoles. If you want to enable root login via TCP/IP comment out ”CONSOLE” in /etc/login.defs. Please also have a look at our Support Database (sdb.rpm). Contents: shadow-961025

Version:  961025-10

shlibs The shared C libraries libc.so and libm.so

Shared C libraries - needed to run programs linked with the /lib/libc.so.5.*.* and /lib/libm.so.5.*.* libraries. You can find the sources in the libc s package.

Author:  H.J. Lu <hjl@gnu.ai.mit.edu>
Version:  5.4.44-1

syslogd Sysklogd

The sysklogd package implements two system log daemons. The syslogd daemon is the general system logging daemon which is responsible for handling requests for syslog services. This version of syslogd is similar to the standard Berkeley product but with a number of compatible extensions. The klogd daemon runs either standalone or as a client of syslogd. Klogd ’listens’ to kernel log messages, prioritizes them and routes them to either output files or to syslogd. This version of klogd will optionally translate kernel addresses to their symbolic equivalents if provided with a system map. Contents: sysklogd-1.3

Copyright:

(c) 1983, 1988 Regents of the University of California, 1994 Dr. Greg Wettstein, Enjellic Systems Development

Author:

Dr. Greg Wettstein <greg@wind.rmcc.com>, Stephen Tweedie <sct@dcs.ed.ac.uk>, Juha Virtanen <jiivee@hut.fi> , Shane Alderton <shane@scs.apana.org.au>

Version:  1.3-9

sysvinit SysV-Style init

System V style init programs by Miquel van Smoorenburg that controls the booting and shutdown of your system. These support a number of system runlevels, each one associated with a specific set of utilities. For example, normal system runlevel is 2, which starts a getty on virtual consoles tty1 - tty6. Runlevel 3 starts xdm. Runlevel 0 shuts down the system. Please see the individual man pages for: inittab, initscript, halt, init, kill5, powerd, reboot, runlevel, shutdown, and telinit. Contents: mingetty-0.9.4 sysvinit 2.65

Copyright:  (c) 1991-1995 Miquel van Smoorenburg

Author:

Miquel van Smoorenburg <miquels@drinkel.ow.nl>, Florian La Roche <florian@suse.de>

Version:  2.65f-9

termcap The libtermcap

Contents: curses-1.0.0 termcap-2.0.8

Copyright:  (c) 1985, 1987, 1993 Free Software Foundation, Inc.
Author:  H.J. Lu <hjl@gnu.ai.mit.edu>, David J. MacKenzie <djm@gnu.ai.mit.edu>
Version:  2.0.8-10

terminfo Complete /usr/lib/terminfo database

This is the terminfo reference database, maintained by Eric Raymond. This database is the official successor of the 4.4BSD termcap file and contains information about any known terminal. The ncurses library makes use of this database in order to use the terminals correctly. If you just use the Linux console, xterm and VT100, you probably will not need this database - a minimal /usr/lib/terminfo tree for these terminal is already included in the ncurses package.

Author:  Eric S. Raymond <esr@thyrsus.com>
Version:  4.2-1

texinfo GNU texinfo

Texinfo is a documentation system that uses a single source file to produce both on-line information and printed output. Using Texinfo, you can create a printed document with the normal features of a book, including chapters, sections, cross references, and indices. From the same Texinfo source file, you can create a menu-driven, on-line Info file with nodes, menus, cross references, and indices. This package is needed to read the documentation files in /usr/info. Contents: texi2html-1.51 texi2roff-2.0 texinfo-3.12

Copyright:  (c) 1987-95 Free Software Foundation, Inc.

Author:

Brian Fox <bfox@gnu.ai.mit.edu>, Jim Meyering <meyering@nanet.ornl.gov>, Noah Friedman <friedman@prep.ai.mit.edu>, Richard Stallman <rms@gnu.ai.mit.edu>, Robert J. Chassell <bob@gnu.ai.mit.edu>, Daniel Hagerty <hag@gnu.ai.mit.edu>, Roland McGrath <roland@gnu.ai.mit.edu>, David J. MacKenzie <djm@gnu.ai.mit.edu>, Charles Hannum <mycroft@gnu.ai.mit.edu>

Version:  3.12-0

textutil GNU textutil

These are the GNU text file (actually, file contents) processing utilities. Most of these programs have significant advantages over their Unix counterparts, such as greater speed, additional options, and fewer arbitrary limits. The programs in this package are: cat, cksum, comm, csplit, cut, expand, fold, head, join, nl, od, paste, pr, sort, split, sum, tac, tail, tr, unexpand, uniq, and wc. Contents: textutils-1.22

Version:  1.22-8

timezone Time zone descriptions

These are configuration files that describe possible time zones. With YaST you can select an appropriate time zone for your system. Contents: tz97e

Copyright:  (c) 1985-95 Regents of the University of California

Author:

Keith Bostic, Bob Devine, Paul Eggert, Robert Elz, Guy Harris, Mark Horton, John Mackin, Bradley White, <tz@elsie.nci.nih.gov>

Version:  97e-8

umsprogs File system utilities

Miscellaneous utilities which are needed for Jacques Gelinas’ UMSDOS file system. Contents: umsdos progs-0.9.src

Copyright:  (c) 1993 Jacques Gelinas
Author:  Jacques Gelinas <jacques@solucorp.qc.ca>
Version:  0.9-12

util Selected utilities

Selected utilities compiled from Rik Faith’s huge utility collection: agetty arch banner cal chroot chsh clear clock col colcrt colrm column ctrlaltdel dmesg domainname fdformat fdisk frag fsck hexdump ipcrm ipcs kbdrate kill last logger login look mesg mkfs.minix mkswap more mount newgrp passwd pwd rdev renice reset rev script setfdprm setserial setsid setterm simpleinit sln strings swapoff swapon sync syslogd tsort tunelp ul umount update wall whereis write zdump zic zones Contents: mount-2.6g util-linux-2.6

Author:  Rik Faith <faith@cs.unc.edu>
Version:  2.6-9

vim Vim: Vi IMproved

Vim (Vi Improved) is an almost compatible version of the UNIX editor vi whereby almost every possible command can be performed using only ASCII characters. Only the ’Q’ command is missing (you don’t need it). Many new features have been added: multi level undo, command line history, filename completion, block operations, editing of binary data, etc. Vi is available for the AMIGA, MS-DOS, Windows NT, and various versions of UNIX.

For the SuSE Distribution, Vim is used as /usr/bin/vi.

Contents: vim-4.6

Author:  Bram Moolenaar <mool@oce.nl>
Version:  4.6-8

yast Yet another Setup Tool

YaST is a tool for installing and configuring the S.u.S.E. Linux system. It provides easy administration capabilities. Installation and configuration of the system by means of YaST is explained detailed in the accompanying handbook of S.u.S.E Linux.

Contents: misc-yast-bins sel yast-0.89.1

Copyright:  (c) 1995,96 S.u.S.E. GmbH
Author:  Michael Andres <ma@suse.de>, Thomas Fehr <fehr@suse.de>
Version:  0.89.1-1

oldshlbs Old Version of libc

This package includes older versions of libc, along with the corresponding include-files.

Contents: libc-5.4.7

Copyright:  (c) 1981-95 Regents of the University of California. , Free Software Foundation, Inc.

Author:

H.J. Lu <hjl@gnu.ai.mit.edu>, Alan Cox <Alan.Cox@linux.org>, Adam J. Richter <adam@yggdrasil.com>, Alan Modra <alan@spri.levels.unisa.edu.au>, Andreas Schwab <schwab@issan.informatik.uni-dortmund.de>, Andrew Walker <andy@keo.kvaerner.no>, Andries Brouwer <aeb@cwi.nl>, Charles Lopes <tjarls@infm.ulst.ac.uk>, Chip Salzenberg <chip@fin.uucp>, Christoph Kern <kern@informatik.uni-wuerzburg.de>, Corey Minyard <minyard@wf-rch.cirr.com>, Craig Smith <craig@rezrov.natinst.com>, Daniel Quinlan <quinlan@yggdrasil.com>, Delman Lee <delman@mipg.upenn.edu>, Dominik Kubla <dominik.kubla@unimainz.de>, Dsouza <m.dsouza@mrc-applied-psychology.cambridge.ac.uk>, Eric Youngdale <eric@aib.com>, Peter Eriksson <pen@signum.se>, Florian La Roche <florian@suse.de>, Fred N. Van Kempen <waltje@uwalt.nl.mugnet.org>, Gerhard Koerting <g.koerting@koerting.techem.ruhr-uni-bochum.de>, H. Peter Anvin <hpa@yggdrasil.com>, Harald Koenig <koenig@tat.physik.unituebingen.de>, Igor Sharfmesser <igor@irbis.alma-ata.su>, J. Alan Eldridge <alane@wozzle.linet.org>, Jaakko Hyvatti <jaakko.hyvatti@www.fi>, James H. Cloos Jr. <cloos@rahul.net>, Jeschke <i6210109@ws.rz.tu-bs.de>, Jim Van Zandt <jrv@vanzandt.mv.com>, Jochen Hein <Hein@Informatik.TU-Clausthal.de>, Jochen Kiemes <jochen kiemes@s2.maus.de>, John F. Haugh Ii <jfh@rpp386>, Jonathan Kamens <jik@cam.ov.com>, Karl Hakimian <hakimian@eecs.wsu.edu>, Kaz Sasayama <kaz@lilia.iijnet.or.jp>, Marc Gingold <gingold@amoco.saclay.cea.fr>, Martin Kuck <martin@unifix.de>, Martin Pirker <pirker@eiunix.tuwien.ac.at>, Mathew G. Monroe <mmex@shadowland.pc.cc.cmu.edu>, Matthias Urlichs <urlichs@smurf.noris.de>, Matthias Urlichs <urlichs@test.smurf.noris.de>, Michael Riepe <riepe@ifwsn4.ifw.uni-hannover.de>, Mike Kennedy <kmb1@gate.net>, Mike Meissner <meissner@cygnus.com>, Miquel van Smoorenburg <miquels@drinkel.ow.nl>, Mitchum Dsouza <m.dsouza@mrc-applied-psychology.cambridge.ac.uk>, Nick Holloway <nick.holloway@alfie.demon.co.uk>, Nickolay Saukh <nms@ussr.eu.net>, Nicolai Langfeldt <janl@ifi.uio.no>, Niibe Yutaka <gniibe@mri.co.jp>, Olaf Flebbe <o.flebbe@science-computing.uni-tuebingen.de>, Patrick D’Cruze <pdcruze@orac.iinet.com.au>, Paul Eggert <eggert@twinsun.com>, Per Bothner <bothner@cygnus.com>, Peter Himmler <himmler@sparcy.mayn.sub.de>, Philip Gladstone <philip.gladstone@mail.citicorp.com>, Remy Card <remy.card@masi.ibp.fr>, Rick Sladkey <jrs@world.std.com>, Robert Andrew Ryan <rr2b+@andrew.cmu.edu>, Roland Mcgrath <roland@gnu.ai.mit.edu>, Rolf Rossius <rossius@hrz.tu-chemnitz.de>, Ross Biro <biro@yggdrasil.com>, Sean Mcneil <sean@mcneil.com>, Stephen L Moshier <moshier@world.std.com>, Stephen Tweedie <sct@dcs.ed.ac.uk>, Steve Losen <losen@delphi.com>, Steve M. Robbins <steve@nyongwa.montreal.qc.ca>, Swen Thuemmler <swen@uni-paderborn.de>, Theodore Ts’O <tytso@mit.edu>, Thomas Pundt <pundtt@goedel.uni-muenster.de>, Ulrich Drepper <drepper@gnu.ai.mit.edu>, Uwe Ohse <uwe@tirka.gun.de>, Vadim V. Vlasov <vvv@ms3.inr.ac.ru>, Wes Cowley <wcowley@cftnet.com>,

Version:  97.10.9-4

0.1.2 AP: Programs that don’t need X

a2ps Converts ASCII text into PostScript

Converts ASCII text into PostScript. This feature is used by e.g. apsfilter, in order to pretty-print ASCII text.

Contents: a2ps-V4.3

Author:  Miguel Santana <M.Santana@frgu.bull.fr>
Version:  4.3-9

alien Alien - Perl script to convert packages

Experimental Perl script to convert software packages. The following formats are supported: deb (Debian), rpm (Red Hat), and tgz (Slackware).

AP: Programs that don’t need X

Contents: alien-5.19

Copyright:  (c) GPL

Author:

Joey Hess <joey@kite.ml.org>, Christoph Lameter <clameter@debian.org>, Randoph Chung <rc42@cornell.edu>

Version:  5.19-4

apmd APM daemon

1) Daemon for monitoring battery status via APM

2) Application to query /proc/apm Includes: apm, apmd, xapm, and tailf

Contents: apmd-2.4

Copyright:  (c) 1996 Rickard E. Faith
Author:  Rik Faith <faith@cs.unc.edu>
Version:  2.4-3

aps The ultimate printer filter: apsfilter

This printer filter lets you print arbitrary files on your printer. The file type is recognized automatically. When needed ghostscript is run to print PostScript files on non-PostScript printers. In order to configure the filter, start YaST in your installed system. Choose ”System administration” in the main menu, then ”Integrate hardware into system”, and finally ”Configure printers”. Fill out the mask, and YaST will configure apsfilter for you. The conventional SETUP of apsfilter is sure avaible as /var/lib/apsfilter/SETUP This revised script uses the program dialog to make the usage more easier for you. The documentation of this version of apsfilter revised by S.u.S.E. is placed in /usr/doc/packages/apsfilter/. Contents: susefilter-1.1 Please note: Ghostscript is required if you will be using apsfilter with a non-PostScript printer.

Copyright:  (c) 1995 Andreas Klemm, Thomas Bueschgens, 1996,97 S.u.S.E.-GmbH

Author:

Andreas  Klemm  <andreas@knobel.gun.de>,  Thomas  Bueschgens

<sledge@hammer.oche.de>, werner@suse.de <werner@suse.de>

Version:  1.1.SuSE-2

arc a packer

ARC is a packer that was used on various operating systems. This package should allow you to unpack *.arc files under Linux. You shouldn’t use it for packing Linux-Software. Use gzip instead.

Contents: arc

Author:  Howard Chu <hyc@umix.cc.umich.edu>
Version:  5.21-8

ash ash shell

NetBSD’s ash (Almquist sh) for Linux is a small (62K - no job control) Bourne-compatible shell. Great for machines with low memory, but does not provide all the extras of shells like bash, tcsh, and zsh. Most shell scripts are run compatible to bash. Note that most scripts written for Linux use some bash-specific syntax. The Slackware setup scripts are a notable exception, since ash is the shell used during installation.

Contents: ash-linux-0.2

Copyright:  (c) 1989 by Kenneth Almquist, 1991 The Regents of the University of California

Author:

Kenneth Almquist, Florian La Roche <florian@suse.de>, Branko Lankester <branko@hacktic.nl>, Sunando Sen <sens@fasecon.econ.nyu.edu>, Arjan de Vet <devet@adv.win.tue.nl>

Version:  0.2-8

at Job Manager

With this program you can run jobs at a specified time.

Contents: at-3.1.8

Copyright:  (c) 1993 David Parsons, 1993, 1994 Thomas Koenig
Author:  Thomas Koenig <ig25@rz.uni-karlsruhe.de>, David Parsons <orc@pell.chi.il.us>
Version:  3.1.8-6

autolog Auto logout

This programs automatically logs out ”forgotten” logins.

Contents: autolog-0.31

Copyright:  (c) 1995 Kyle Bateman
Author:  Kyle Bateman <kyle@actarg.com>
Version:  0.31-8

bc GNU bc - command line calculator

bc is an interpreter that supports numbers of arbitrary precision and the interactive execution of statements. There are some similarities in the syntax to the C programming language. A standard math library is available through command line options. When used, the math library is read in before any other input files. bc then reads in all other files from the command line, evaluating their contents. Then, bc reads from standard input (usually the keyboard). The dc program is also included. Dc is a calculator which supports reverse-polish notation and allows unlimited precision arithmetic. Macros can also be defined. Normally, dc reads from standard in but can also read in files specified on the command line. A calculator with reverse-polish notation saves numbers to a stack. Arguments to mathematical operations (operands) are ”pushed” onto the stack until the next operator is read in which ”pops” its arguments off the stack and ”pushes” its results back onto the stack.

Contents: bc-1.04

Copyright:  (c) 1991, 1992, 1993, 1994 Free Software Foundation, Inc.
Author:  Philip A. Nelson <phil@cs.wwu.edu>
Version:  1.04-8

buffer Buffering stdin and stdout

This is a program designed to speed up writing tapes on remote tape drives. When this program is put ”in the pipe”, two processes are started. One process reads from standard-in and the other writes to standard-out. Both processes communicate via shared memory.

Copyright:  (c) 1990,1991 Lee McLoughlin
Author:  Lee McLoughlin <L.McLoughlin@doc.ic.ac.uk>
Version:  1.17-8

bzip Program for compression of files

Bzip is a very powerful program for compressing files.

Contents: bzip2-0.1pl2

Copyright:  (c) 1996,1997 Julian Seward
Author:  Julian Seward <jseward@acm.org>
Version:  0.1p2-5

cdcons Console CD Player

A small CD Player that runs on the text console. Documentation: /usr/doc/packages/cdcons

Contents: cd-console-1.1

Copyright:  (c) 1996, Osku Salerma
Author:  Osku Salerma <osku@iki.fi>
Version:  1.1-9

cdrecord Tool to write CD-Rs

Cdrecord has been implemented with a Philips CDD 521 and is now maintained with a Philips CDD-2000. It should run also on lots of Hardware, see /usr/doc/packages/cdrecord/README

Contents: cdrecord-1.6a9

Copyright:  (c) Copyright 1996 J. Schilling
Author:  Joerg Schilling <joerg@schily.isdn.cs.tu-berlin.de>
Version:  1.6-0

cdwrite Writing CDRs

Program to write data and audio CDs using a CD writer.

Contents: cdwrite-2.0 Be very careful using this program. If you enter the wrong SCSI device, you could irrevocably overwrite a hard disk (including partition table and foreign partitions).

Copyright:  (c) 1995, Adam J. Richter
Author:  Adam J. Richter <adam@yggdrasil.com>
Version:  2.0-8

cdwtools CDW tools

Frontend for various programs used to create CDs Contents: cdda2wav-0.9beta11 cdwtools-0.93

Copyright:  (c) 1996, Thomas Niederreiter
Author:  Thomas Niederreiter <p7003ad@sunmail.lrz-muenchen.de>
Version:  0.93-8

cfengine GNU cfengine

cfengine is an abstract programming language for system administrators of huge heterogeneous networks. With cfengine, system administrators have an easy and elegant way to maintain complicated networks. For introduction and documentation please look at the files doc/cfengine.tex in TeX mode or at /usr/info/cfengine.info.gz in TeXInfo mode. Contents: cfengine-1.4.10

Copyright:  (c) 1995 Mark Burgess, Oslo
Author:  Mark Burgess
Version:  1.4.10-1

chipmunk Hardware ’debugging’.

Documentation: /usr/doc/packages/chipmunk. Contents: chipmunk-1.4

Author:  R.E. Wolff <R.E.Wolff@et.tudelft.nl>, R.D.L. Stout <stout@dutecai.et.tudelft.nl>
Version:  1.4-10

compat some libraries in compatibility-versions

Contains libncurses.so.1.9.7a (from the old sources) libgdbm.so.2 and libdb.so.2 (originals with changed major-release) Contents: db.1.85 gdbm-1.7.3 ncurses-1.9.7a

Version:  97.11.27-2

ctags Generate C language tag files for use with vi

Copyright:  (c) 1996 Darren Hiebert
Author:  Darren Hiebert <darren@sirsi.com>
Version:  1.5-8

deco Demos Commander

DECO or Demos Commander is a visual front-end to the UNIX operating system, similar to the Norton Commander. Usually, DECOS displays two fields. In each of these, the contents of a directory is listed. The program is controlled using the keyboard. Use the cursor keys to choose a file, press return to select it; TAB changes between the two fields. Commands can be executed by using the F-keys or the menu. Additional commands can be added by the user. Unfortunately, the sources for this program are not freely available. Contents: deco-3.5

Copyright:  (c) 1989-1993 Serge Vakulenko
Author:  Serge Vakulenkom <vak@kiae.su>
Version:  3.5-4

dialog Menus and input boxes for shell scripts

This program lets you use menus and dialog boxes in shell scripts. Contents: dialog-0.62

Author:

Savio Lam <lam836@cs.cuhk.hk>, Alessandro Rubini <rubini@ipvvis.unipv.it>, Florian La Roche <florian@suse.de>, Marc Ewing <marc@redhat.com>

Version:  0.62-8

dosfsck Filesystem Check for DOS Filesystems

dosfsck is the Linux equivalent of PC/MS-DOS’ CHKDSK. It checks the consistency of PC/MS-DOS file systems and optionally tries to repair them. The tests dosfsck performs are described in the man page. Contents: dosfsck.1 Attention: This Package is still in ALPHA state, i.e. not absolutely error free and usage may result in serious damage of your DOS partition.

Author:  Werner Almesberger <almesber@bernina.ethz.ch>
Version:  beta1-6

duconv Unix <-> DOS Converter

A DOS <--> UNIX converter which even correctly converts German umlauts. Something everybody(!) is looking for..;) Contents: duconv-0.1

Copyright:  (c) 1994 Torsten F”ahrmann, TortiSoft
Author:  Torsten F”ahrmann <tfaehr@aol.com>
Version:  0.1-9

dump Backup program

a Backup program. Contents: dump-0.3 e2fsprogs-1.10

Copyright:  (c) Copyright (c) 1980, 1993 The Regents of the University of California
Author:  Remy Card <Remy.Card@freenix.fr>
Version:  0.3-7

ed standard old un*x line editor

Author:  Andrew Moore <alm@worm.talke.org>
Version:  0.2-8

edy An editor with menus

Simple ASCII editor. Several files can be edited simultaneously. Commands can be accessed via pull-down menus. Contents: edy-2.0

Copyright:  (c) 1993 Markus Bulik
Version:  2.0-8

enscript ASCII to PostScript(tm) converter

Enscript converts ASCII files to PostScript and writes the generated output to a file or sends it directly to the printer. Enscript is configured in /etc/enscript.cfg. Contents: enscript-1.5.5

Author:  Markku Rossi <mtr@iki.fi>
Version:  1.5.5-1

faxprint Printer-filter for use with SuSEFax

This printer-filter creates a printer called ’fax’. If you use this one with any application, you can send faxes with the spooling mechanism of SuSEFax. Contents: faxprint-filter

Copyright:  (c) 1998, S.u.S.E. GmbH, Fuerth, Germany
Author:  Carsten Hoeger <choeger@suse.de>
Version:  0.1-2

gcal A program for printing calendars

Gcal is a program for printing calendars. Gcal displays a calendar for a month or a year, eternal holiday lists and fixed date lists, in many ways. The program correctly omits the dates that were skipped when the current Gregorian calendar replaced the earlier Julian calendar. Contents: gcal-2.40

Copyright:  (c) GPL
Author:  Thomas Esken <esken@uni-muenster.de>
Version:  2.40-6

gdtmon Administrate a RAID array with a Vortex GDT controller

Version:  1-4

git GNU Interactive Tools

A complete system shell, for those who don’t prefer the command line :-) Contents: git-4.3.15

Copyright:  (c) 1994 Tudor Hulubei & Andrei Pitis
Author:  Tudor Hulubei <tudor@cs.unh.edu>, Andrei Pitis <pink@pub.ro>
Version:  4.3.15-8

glimpse Text search in huge databases

You can index lots of ASCII files very easily. Glimpse creates hash tables which enable you to search (and find!) any given text. In general glimpse does the same as grep but far more efficient and faster. Contents: glimpse-4.1

Copyright:

(c) Udi Manber and Burra Gopal of the University of Arizona, and Sun Wu of the National Chung-Cheng University, Taiwan.

Author:  Udi Manber <glimpse@cs.arizona.edu>
Version:  4.1-4

gpm General purpose mouse support

This program is based on the ”selection” package, and allows you to use the mouse to cut and paste text from the screen. It also acts as a mouse server for applications running on the Linux console, such as the Midnight Commander and dialog. A second program, gpm-root, provides pull down menus to be called up by pressing Ctrl+mouse-key. NOTE: This program may cause problems when you start X on a system that does not use a serial mouse. If you get an ’unable to open mouse device’ error from X, disable gpm. Contents: gpm-1.13

Copyright:  (c) 1994,1995 Alessandro Rubini
Author:  Alessandro Rubini <rubini@ipvvis.unipv.it>
Version:  1.13-5

gs both Ghostscript with X11 and with VGA features

This gs version is for X11 machines and vga terminal. Currently there is the newer version 5.10 of ghostscript available. Due to a change in the licence of this version we are not allowed to put this version onto this cdrom. Nevertheless this version you will find this version on our FTP server.

Copyright:  (c) 1989-96 Aladdin Enterprises.
Author:  Peter Deutsch <ghost@aladdin.com>
Version:  5.10-0

gs fonto Optional fonts for Ghostscript

Some useful fonts for the PostScript interpreter Ghostscript.

Version:  5.10-0

gs fonts Standard fonts for Ghostscript

The necessary fonts for the PostScript interpreter Ghostscript.

Version:  5.10-0

gs serv Ghostscript without X11 and without VGA features

This gs version should only be installed on print servers without X11. Currently there is the newer version 5.10 of ghostscript available. Due to a change in the licence of this version we are not allowed to put this version onto this cdrom. Nevertheless this version you will find this version on our FTP server.

Copyright:  (c) 1989-96 Aladdin Enterprises.
Author:  Peter Deutsch <ghost@aladdin.com>
Version:  5.10-0

gs vga Ghostscript without X11 but with VGA

This gs version is for X11-less machines and vga terminal or graphic card. The graphic card should be supported by the SVGA lib.

Currently there is the newer version 5.10 of ghostscript available. Due to a change in the licence of this version we are not allowed to put this version onto this cdrom. Nevertheless this version you will find this version on our FTP server.

Copyright:  (c) 1989-96 Aladdin Enterprises.
Author:  Peter Deutsch <ghost@aladdin.com>
Version:  5.10-0

gs x11 Ghostscript for X11

If you plan to preview PostScript documents under X, you need to install this. Currently there is the newer version 5.10 of ghostscript available. Due to a change in the licence of this version we are not allowed to put this version onto this cdrom. Nevertheless this version you will find this version on our FTP server.

Copyright:  (c) 1989-96 Aladdin Enterprises.
Author:  Peter Deutsch <ghost@aladdin.com>
Version:  5.10-0

hello A friendly greeting program

The GNU ‘hello’ program produces a familiar, friendly greeting. It allows nonprogrammers to use a classic computer science tool which would otherwise be unavailable to them. Because it is protected by the GNU General Public License, users are free to share and change it. Contents: hello-1.3 Hello, world!

Copyright:  (c) GPL

Author:

The  King  <elvis@graceland.gnu.ai.mit.edu>,  Charles  Han-
num  <mycroft@hill.gnu.ai.mit.edu>,  David  J.  MacKen-
zie  <djm@apple-gunkies.gnu.ai.mit.edu>,  Jan  Brittenson

<bson@kropotkin.gnu.ai.mit.edu>, Michael I Bushnell <mib@geech.gnu.ai.mit.edu>,

Mike  Haertel  <mike@churchy.gnu.ai.mit.edu>,  Noah  Fried-
man  <friedman@nutrimat.gnu.ai.mit.edu>,  Roland  McGrath

<roland@geech.gnu.ai.mit.edu>

Version:  1.3-6

id utils Language-independent identifier database tool

mkid is a simple, fast, high-capacity, language-independent identifier database tool. Actually, the term ‘identifier’ is too limiting -- mkid stores tokens, be the program identifiers of any form, literal numbers, or words of human-readable text. Database queries can be issued from the command-line, or from within emacs, serving as an augmented tags facility.

Contents: id-utils-3.2

Copyright:  (c) GPL
Author:  McGary <gkm@gnu.ai.mit.edu>, <Tom.Horsley@mail.hcsc.com>
Version:  3.2-7

idled Idle daemon

This program automatically logs out users who have been idle for some time; this time, of course, can be set by the system administrator.

Contents: idled-1.12

Author:  Michael P. Crider <crider@cs.hope.edu>
Version:  1.12-9

ingres A relational database system

This powerful relational database comes with the full sources and its own query language. This query language can be easily integrated into your own C-programs using the Ingres-internal preprocessor. It comes without saying that the system provides concepts of transaction and recovery to ensure data consistency.

Copyright:  (c) The Regents of the University of California 1977
Version:  8.9.4-6

ingrtool Tool package for Ingres

This add-on for Ingres gives you a Tcl/Tk Interface to the good old ingres database. For installation change to /usr/lib/ingrestools-1.0 and call ”Install”.

Documentation can be found under /usr/lib/ingrestools-1.0/doc. (German and English) Contents: ingrestools-1.0

Copyright:  (c) 1994 Gert Reinauer
Author:  Gert Reinauer
Version:  1.0-7

inlinem inlinemath

With this package you get inlined assembler functions of the mathematic library under /usr/include/i386/. To use this functions the system needs a math coprocessor (386 combined with 387, 486 with builtin coprocessor, 586). Documentation is stored under /usr/doc/packages/inline-math-1.6/ and the header file math.h and extensions.h under /usr/include/i386/. To include this files you have to patch the file /usr/include/math.h with /usr/doc/packages/inline-math-1.6/math.h-patch. You can control the usage of /usr/include/i386/ math.h with the gcc option ‘-D NO MATH INLINES’ and the usage of /usr/include/i386/extensions.h with the gcc option ‘-D MATH EXTENSIONS’. Note: If the package is removed the patched version of /usr/include/math.h has to be replaced by the original version. Contents: inlinem-1.9 For this package you really need a math coprocessor!!! After installation you have to patch the header file /usr/include/math.h with /usr/doc/packages/inlinemath-1.6/math.h-patch!

Copyright:  (c) (C) 1995 John C. Bowman
Author:  John C. Bowman <bowman@hagar.ph.utexas.edu>
Version:  1.9-8

isapnp ISA plug and play configuration utility

Two programs - one allows the dumping of resource data and generation of a skeleton configuration file, the other configures ISA PnP hardware using a configuration file. More Info: /usr/doc/packages/isapnp/README.SuSE Contents: isapnp

Copyright:  (c) GNU
Author:  Peter Fox <fox@roestock.demon.co.uk>
Version:  1.10-9

jed JED Editor

JED is an extremely powerful but small emacs programmer’s editor that is extensible in a C-like macro language and is able to perform color syntax highlighting. Among the many features: Emacs, wordstar, EDT emulation; C, Fortran, TeX, text editing modes; Full undo; GNU Emacs compatible info reader, and lots more. 8 bit clean, so you can even edit binary files! Contents: jed097-12 slang0.99-23

Copyright:  (c) 1992, 1993, 1995 John E. Davis
Author:  John E. Davis <davis@amy.tch.harvard.edu>
Version:  097.12-7

joe Joe Text editor

JOE (Joe’s Own Editor) is a freeware ASCII editor for UNIX. Joe is similar to most IBM PC text editors. The keyboard combinations are similar to WordStar[tm] and Turbo C. Some of Joe’s features include: -full support of termcap/terminfo (useful for device independent description to control screen output). -optimizing screen refresh even as GNU Emacs (Joe can even be used at 2400 baud). -easy installation. All features for UNIX integration of the VI: a marked block of text can be filtered through a UNIX command; and wherever JOE accepts a filename parameter the following can also be substituted: !command :to redirect from/to another command >>filename :to append onto an existing file filename,start,size :to edit a part of a file/device

- :to use standard input/output When JOE has several files opened at the same time, each file is displayed in its own window. Additionally, JOE allows for shell-windows whereby the output of the executed commands are saved in a buffer, automatic filename completion (via TAB), help windows, undo-redo, search and replace using regular expressions.

Contents: joe2.8

Copyright:  (c) 1992 Joseph H. Allen
Author:  Joseph H. Allen <jhallen@world.std.com>
Version:  2.8-9

jove Jonathan’s Own Version of Emacs

JOVE (Jonathan’s Own Version of Emacs) is a text editor based mainly on the original Emacs editor written by Richard Stallman at M.I.T. Although jove is mostly Emacs compatible there are several differences so that you shouldn’t rely on it. Jove offers you e.g. automatic indentation, multiple view of more than one file, and shorthands in form of keys, words, lines and paragraphs A mostly Emacs compatible text editor, with man pages and the ’teachjove’ tutorial program. If you’d rather have the real thing, the actual Emacs editor/environment is available as disk series ’E’. JOVE is configured by /var/lib/jove/jove.rc.

Contents: jove

Copyright:  (c) 1986, 1987, 1988 by Jonathan Payne
Author:  Jonathan Payne <jpayne@sun.com>
Version:  4.16-9

lha packer

Lha is a packer comparable to ZIP (PKZIP), ZOO etc. It has been included for compatibility reasons only. For general archiving purposes use GZIP, because this is standard for Linux.

Contents: lha

Copyright:  (c) 1992 Masaru Oki
Author:  Masaru Oki <oki@wbg.telcom.oki.co.jp>
Version:  1.0-9

lsof Lsof lists information about files opened by processes

An open file may be a regular file, a directory, a block special file, a character special file, an executing text reference, a library, a stream or a network file (Internet socket, NFS file or UNIX domain socket.) A specific file or all the files in a file system may be selected by path.

Contents: lsof 4.16

Copyright:  (c) 1994 Purdue Research Foundation, West Lafayette, Indiana
Author:  Victor A. Abell <abe@purdue.edu>
Version:  4.16-7

makewhat makewhatis - create a whatis database

The man system in S.u.S.E. Linux (package man) does not need a whatis database. Nevertheless some manual browser (e.g. tkman) still need this database. For this reason we included this package.

Contents: makewhat

Copyright:  (c) 1992 Rickard E. Faith
Author:  Werner Fink <werner@suse.de>
Version:  98.3.23-0

mbedit Mutil platform editor

mbedit is a full screen text editor with macro option, online calculator, command history buffer, hex editor and many other features. It is available for some common operating systems as: MS-DOS, os/9, Unix.

Contents: mbedit.7.71

Copyright:  (c) 1992-1997 / Michael Braun
Author:  Michael Braun <braun@hydro2.atlas.de>
Version:  7.71-8

mc Midnight Commander

The Midnight Commander is a Norton Commander clone; a program that manipulates and manages files and directories. It is useful, fast, and has color displays on the Linux console. It also has mouse support if you run the ”gpm” mouse server. This program requires the terminal description files in /usr/lib/terminfo, which are found in ncurses.tgz or terminfo.tgz. You can also use the Midnight Commander under X with your mouse. If you enter ”mc -c”, colors will be used. Under the Midnight Commander, the screen is divided into four sections, whereby the majority of the screen is covered by two directory panels. The second to last line on the screen is the shell command line; the last line displays the function key assignments. At the very top, the menu list is shown. One of the directories displayed is the current working directory. This is where most of the commands are found. For certain commands, like copy and move, the second directory is used as the target directory. Contents: mc-4.1.11

Copyright:  (c) 1994, 1995 Miguel de Icaza, Janne Kukonlehto

Author:

Miguel de Icaza <miguel@roxanne.nuclecu.unam.mx>, Janne Kukonlehto <jtklehto@paju.oulu.fi>, Dugan Orlando Porter <dugan@b011.eunet.es>, Radek Doulik <rodo@earn.cvut.cz>, Fred Leeflang <fredl@nebula.ow.nl>, Mauricio Plaza <mok@roxanne.nuclecu.unam.mx>, Jakub Jelinek <jj@sunsite.mff.cuni.cz>

Version:  4.1.11-4

mdutils Utilities for ’Multiple Devices’

You can combine different partitions (including SCSI and IDE) to one partition under Linux. More information can be found under /usr/doc/packages/mdutils an in the man pages for add, mdcreate and mdtab. Please be aware that the author himself considers this software as ALPHA. We at S.u.S.E. have not encountered any problems however. Contents: md035

Author:  Marc Zyngier <zyngier@ufr-info-p7.ibp.fr>
Version:  0.35-8

mkdosfs tool to format dos partitions

This is a tool to create the filesystem on dos partitions which works similar to mke2fs. Documentation is under /usr/doc/packages/mkdosfs This programm is not fully stable yet! Use at your own risk!!

Copyright:  (c) GPL
Author:  Dave Hudson <dave@humbug.demon.co.uk>
Version:  0.4-1

mkisofs Creating CDs under Linux

Actually, mkisofs is a pre-mastering program that generates an iso9660 filesystem . It takes a snapshot of a given directory tree, and generates a binary image which corresponds to an iso9660 filesystem which can be written to a block device. Contents: mkisofs-1.11.1 mkisofs-1.11

Copyright:  (c) 1993 Yggdrasil Computing, Incorporated

Author:

Eric Youngdale <eric@aib.com>, H. Peter Anvin <hpa@yggdrasil.com>, John Brezak <brezak@ch.hp.com>, Paul Eggert <eggert@twinsun.com>, Peter van der Veen <peterv@qnx.com>, R.-D. Marzusch <marzusch@odiehh.hanse.de>

Version:  1.11-4

mktemp Utility for tempfiles

mktemp is a small utility that interfaces to the mktemp() function call to allow shell scripts and other programs to use files in /tmp safely. Contents: mktemp-1.4

Copyright:  (c) BSD
Author:  Todd C. Miller <Todd.Miller@courtesan.com>
Version:  1.4-6

mmv Multiple Move

This mv command allows you something like ”mv *.c *.c.old”.

Documentation: man mmv Contents: mmv-1.01b

Copyright:  (c) 1989 by Vladimir Lanin
Author:  Vladimir Lanin <lanin@csd2.nyu.edu>
Version:  1.01b-8

mpegutil MPEG (de-)compression utility

Mpeg is an utility for the compression and decompression of MPEG files. MPEG animations can be composed out of single images, and decomposed into single images (in YUV format). Contents: mpeg-utils-1.2.1

Copyright:  (c) 1990-93 Andy C. Hung
Author:  Andy C. Hung <achung@cs.stanford.edu>
Version:  1.2.1-8

mt st mt-st controls magnetic tape drive operation

Although this mt program is tailored for SCSI tape, it can also be used with the QIC-02 driver and hopefully with other Linux tape drivers that use the same ioctls (some of the commands may not work with all drivers). Contents: mt-st-0.4

Copyright:  (c) 1980 The Regents of the University of California
Author:  Kai Makisara <Kai.Makisara@vtt.fi>
Version:  0.4-8

mtools Access files on an MS-DOS filesystem

Allows uncomplicated access to a MS-DOS filesystem on Disk, without mounting it. Included commands for working with MS-DOS Files mdir, mcd, mcopy, mformat. XDF support for OS/2 is also provided. Contents: mtools

Author:  Alain Knaff <Alain.Knaff@imag.fr>, David C. Niemi <niemidc@erols.com>
Version:  3.6-12

nc Northern Commander

It is not only the name which is similar to the well-known DOS program. Contents: nc-0.74

Copyright:  (c) 1993-94 Andreas Matthias
Author:  Andreas Matthias <matthias@hrz.uni-kassel.de>
Version:  0.74-11

obst The OBject system of STONE

OBST is the Objekt Management System of the STONE project. It was designed as a tool for saving persistent objects. OBST can be seen as an object-oriented database written in C. Contents: OBST3-4.3

Copyright:  (c) 1992-1994 by Forschungszentrum Informatik (FZI)
Author:  Forschungszentrum Informatik, OBST Projekt <obst@fzi.de>
Version:  3.4.3-6

octave Octave

Octave is a high level programming language. It’s designed for the solution of numeric problems. There is a command line interface supplied. Contents: octave-2.0.9

Copyright:  (c) 1992, 1993, 1994, 1995 John W. Eaton
Author:  John W. Eaton <jwe@bevo.che.wisc.edu>
Version:  2.0.9-7

onyx 4GL-Engine

The main goal in the development of Onyx was to create a tool for the fast development of database applications. The main idea is that all application are build in a similar way. Internal tables are used to store the data, and masks are used to view and change the data Transactions define the connection of the internal data with the rest of the world.

Copyright:  (c) 1993 M.Koehne
Author:  Michael Kraehe <kraehe@bakunin.north.de>
Version:  1.29-10

pdksh Public Domain Korn Shell

A public domain Korn Shell clone.

Contents: pdksh-5.2.13

Version:  5.2.13-4

pico Small, easy to use editor.

The editor normally used by pine. You will find the Sources in pine.spm.

Version:  3.96-10

postgres Postgres Database

DBMS (DataBase Management System) PostgreSQL is the successor of the relational database system INGRES. It has been developed at Berkeley, University of California. The main advantages of the new system are:

- better support of complex objects.

- improved user-side adaption of data types, operators and access methods.

- Integration of active databases (alerts and triggers), conclusions including forward- and backward chaining.

- simplification of dbms code for restoration after a crash.

- Provision of a design which concatenates the advantages of optical disks and workstations which consist of many compound processors and customer-specific VLSI chips.

- Avoidance of too many changes according to relational model. The query language for this system is SQL. This system offers you a wide range of additional features such as transparent history-system, embedding in your own C-routines, shell scripts and more.

Contents: appgen-0.2-alpha dbbrowser-0.9.src postgresql-6.2.1

Copyright:  (c) 1994-1997 Regents of the University of California
Author:  Andrew Yu <andrew@CS.Berkeley.edu>, Jolly Chen <jolly@CS.Berkeley.edu>
Version:  6.2.1-10

quota Disk Quota System

Allows to assign disk quotas to the individual users of the system. The kernel must be recompiled with disk quota support enabled. Contents: quota-1.55 This program only runs if an kernel with quota support has been installed!

Copyright:  (c) 1980, 1990 Regents of the University of California.
Author:  Marco van Wieringen <mvw@planets.ow.nl>
Version:  1.55-9

recode THE Character Set Converter

This filter converts various character sets.

Contents: recode-3.4.1

Copyright:  (c) 1987-94 Free Software Foundation, Inc.
Author:  Francois Pinard <pinard@iro.umontreal.ca>
Version:  3.4.1-8

rman rman

RosettaMan is a filter for UNIX manual pages needed by TkMan. It takes as input man pages formatted for a variety of UNIX flavors (not [tn]roff source) and produces as output a variety of file formats. Also included is http-rman which provides access to the man-pages via a normal web-browser.

Contents: http-rman rman-3.0.2a9

Copyright:  (c) T.A. Phelps
Author:  T.A. Phelps <phelps@cs.Berkeley.EDU>
Version:  3.0.2a9-9

rpm2html RPM database info to HTML converter

from the original spec file: rpm2html tries to solve 2 big problems one face when grabbing a RPM package from a mirror on the net and trying to install it:

- it gives more information than just the filename before installing the package. - it tries to solve the dependancy problem by analyzing all the Provides and Requires of the set of RPMs. It shows the cross references by the way of hypertext links.

Contents: rpm2html-0.70p1

Author:  Daniel Veillard <veillard@w3.org>
Version:  0.70p1-1

sc Spreadsheet

SC is a spreadsheet program. After invoking up sc a spreadsheet is shown which is arranged into rows and columns of cells. The spreadsheet is empty if the program is called without any parameters. Normally though, the filename parameter is read in by sc as a spreadsheet file. Each cell can contain a number, character string and/or expression (formula) which is based on numerical values, or character strings of other cells.

Contents: sc-6.21

Author:  Jeff Buhrt , Mark Weiser , James Gosling
Version:  6.21-9

screen Multi Screen on a VT100/ANSI Terminal

With this program you can take advantage of the multi-tasking abilities of your Linux system by opening several sessions over one terminal. The sessions can also be detached and resumed from another login terminal. Documentation: man screen

Contents: screen-3.7.3

Copyright:  (c) 1987-93 Oliver Laumann
Version:  3.7.3-8

scsiinfo Low level SCSI configuration

With this comfortable graphical user interface you may easily make ”disastrous” configurations to your SCSI subsystem. Please be very careful! You may destroy both your hardware and installed system by using this tool! Documentation: /usr/doc/packages/scsiinfo Contents: scsiinfo-1.7 Use with utmost care. Improper configuration can lead to severe hard and software damages!!!

Copyright:  (c) 1993, Yggdrasil Computing, Inc.
Author:  Michael Weller <eowmob@exp-math.uni-essen.de>, Eric Youngdale <eric@aib.com>
Version:  1.7-10

seejpeg An image viewer (libvga)

Seejpeg is Evan Harris’s JPEG viewer which utilizes svgalib and contains limited GIF, PPM and TARGA viewing support. The most useful feature of seejpeg is that it automatically decides on the ”best” video mode to use based on the image being displayed and the video card being used. If the image is larger than the screen the cursor keys can be used to scroll.

Contents: seejpeg-1.6.1

Copyright:  (c) 1993, 1994 Evan Harris
Author:  Evan Harris <evan@cs.mu.oz.au>
Version:  1.6.1-12

sharutil GNU shar utilities

shar, uuencode, uudecode and others...

Contents: sharutils-4.2c uudeview-0.5.13

Version:  4.2c-12

splitvt Split terminals

splitvt splits a VT100 compatible (Linux console, xterm) terminal‘s screen into two.

Contents: splitvt-1.6.3

Copyright:  (c) 1993 Sam Lantinga
Author:  Sam Lantinga <slouken@cs.ucdavis.edu>
Version:  1.6.3-8

sudo sudo

’sudo’ is a command that allows users to execute some commands as root. The /etc/sudoers file (edited with ’visudo’) specifies which users have access to sudo and which commands they can run. ’sudo’ logs all its activities to syslogd, so the system administrator can keep an eye on things. Sudo asks for the password for initializing an check period of a given time N (where N is defined at installation and is set to 5 minutes by default). Contents: sudo.v1.5.2

Copyright:  (c) 1991 The Root Group, Inc.

Author:

Jeff Nieusma <nieusma@rootgroup.com>, David Hieb <davehieb@rootgroup.com>, Ian McCloghrie <ian@ucsd.edu>

Version:  1.5.2-8

svgatext Extended SVGA text modes

No more place on your screen in text mode? Huge letters on the screen? You can change this by calling SVGATextMode which extends a SVGA text display up to 160x60 characters on good SVGA cards. Before use you have to configure the program by editing the file /etc/TextConfig to the capabilities of your graphics hardware. Furthermore you may create your own modelines for /etc/TextConfig with SVGATextCalc. Broad documentation can be found under /usr/doc/packages/svgatext and the respective man pages. Contents: SVGATextCalc-0.9 SVGATextMode-1.8 The author of SVGATextCalc would like to get a short mail message from you to tell him how you like the program.

Copyright:  (c) 1995,1996 Koen Gadeyne, 1996 Marco Paganini
Author:  Koen Gadeyne <kmg@barco.be>, Marco Paganini <paganini@ism.com.br>
Version:  1.8-0

taper Backup program taper

This is a tape backup and restore program that provides a friendly user interface to allow backing/restoring files to a tape drive. Alternatively, files can be backed up to archive files. Selecting files for backup and restore is very similar to the Mouseless Commander interface and allows easy traversal of directories. Recursively selected directories are supported. Incremental backup and automatic most recent restore are default settings. Both SCSI and floppy tape drives are supported. Information about archives is stored on the hard disk allowing information about archives to be obtained without having to have access to the tape. Contents: taper-6.7.4

Copyright:  (c) 1995 Yusuf Nagree
Author:  Yusuf Nagree <yusuf@mail.bogo.co.uk>
Version:  6.7.4-9

tcsh tcsh

tcsh is an enhanced but completely compatible version of the Berkeley UNIX C shell, csh(1). It is a command language interpreter usable both as an interactive login shell and a shell script command processor. It includes a command-line editor, programmable word completion, spelling correction, a history mechanism, job control, and a C-like syntax. Contents: tcsh-6.07.02

Copyright:  (c) 1980, 1991 The Regents of the University of California
Author:  Christos Zoulas <christos@deshaw.com>, Scott Krotz <krotz@mot.com>
Version:  6.07.02-8

teapot A three dimensional spreadsheet

Teapot is a spread sheet program for UNIX, which uses modern concepts besides being portable and extensible, for instance three dimensional tables and iterative expressions. It provides various output formats e.g. LaTeX and HTML.

There are executables in english, dutch and german available. They are started by typing ”teapot en”, ”teapot nl” or ”teapot de”. Just typing ”teapot” starts the english version. Contents: teapot-1.0

Copyright:  (c) Michael Haardt, 1995, 1996, 1997
Author:  Michael Haardt <michael@cantor.informatik.rwth-aachen.de>
Version:  1.0-0

unarj Pack program

Allows for unpacking of .arj files, which are widely spread in the DOS world. It is included for compatibility reasons to extract packages from DOS. It should not be used for compressing packages under Linux; use gzip which is the standard, instead. Contents: unarj

Copyright:  (c) 1991-93 Robert K Jung
Author:  Robert Jung <robjung@world.std.com>
Version:  2.41a-8

units units conversion utility

Copyright:  (c) 1989, 90, 91, 92, 93, 94 Free Software Foundation, Inc.
Author:  Adrian Mariano <adrian@cam.cornell.edu>
Version:  1.53-8

unzip Decompresser

UnZip is an extraction utility for archives compressed in .zip format (known as ”zip files”). Although highly compatible both with PKWARE’s PKZIP and PKUNZIP utilities for MS-DOS and with Info-ZIP’s own Zip program, our primary objectives have been portability and non-MSDOS functionality. Contents: unzip-5.32

Version:  5.32-3

xdelta Binary delta generator and RCS replacement library

Xdelta is a binary delta generator and RCS replacement library. The delta algorithm is faster than GNU diff for text files and works on binary files too. Contents: xdelta-0.14

Copyright:  (c) 1997 Josh MacDonald, Paul Eggert
Author:  Josh MacDonald <jmacd@cs.berkeley.edu>
Version:  0.14-5

yamm Yet Another Micro Monitor

similar to the program ”top”. Contents: yamm-2.5.6

Copyright:  (c) 1994, 1995 Riccardo Facchetti
Author:  Riccardo Facchetti <riccardo@cdc8g5.cdc.polimi.it>
Version:  2.5.6-8

ytree YTree - A filemanager like XTree

A (DOS) XTree-clone Contents: ytree-1.50

Copyright:  (c) Werner Bregulla
Author:  Werner Bregulla <werner@frolix.han.de>
Version:  1.50-8

z88 z88

Finite Elements programme with several modules: The FE processor reads the general structure data and the restrictions. Basicly, the Z88 entry files can be generated by a CAD convertor Z88X, by an editor or a text processor system or with a mixed procedure. Z88X outputs the processed structure data and restrictions, calculates the stiffness matrices of the elements, compiles the stiffness matrix of the entire system, scales the equation system, solves the system with an in situ Cholesky method and displays the displacements in ASCII format. This is why there are modules for calculating the stress, for the node strength and a net generator. You can graphically display and plot the net generator’s data.

D: Development (C, C++, Lisp, Perl, etc.)

Contents: z88

Copyright:  (c) Dr. Frank Rieg
Author:  <dr.frank.rieg@t-online.de>
Version:  98.1.24-2

zip compression program

Zip 2.0.1 is a compression and file packaging utility. It is compatible with PKZIP 2.04g (Phil Katz ZIP) for MSDOS systems. Contents: zip

Version:  2.0.1-8

ziptool Tool for the Iomega ZIP and JAZ drives

With this tool, configuration of the Iomega ZIP and JAZ drives is possible. Automatic eject and write protection (password too) may be configured. Contents: ziptool

Copyright:  (c) 1997 Mirko Kraft, Grant R. Guenther, Itai Nahson
Author:  Mirko Kraft <Mirko.Kraft@ixos.de>, Grant R. Guenther <grant@torque.net>
Version:  1.1b-9

zoo Pack program

Zoo is a packer based on the Lempel-Ziv algorithm. Lots of files on DOS/AmigaDOS and TOS systems used this packer for their archives. The compression rate of gzip are not reached at all. Thus zoo should only be used for decompressing old archives. Contents: zoo

Copyright:  (c) 1991 Rahul Dhesi

Author:

Ian Phillipps <igp@camcon.co.uk>, J. Brian Waters <jbwaters@bsu-cs.bsu.edu>, Paul Homchick <rutgers!cgh!paul>, Mark Alexander <amdahl!drivax!alexande>, Randal L. Barnes <rlb@skyler.mavd.honeywell.com>

Version:  2.10-8

zsh zsh version

Zsh is a UNIX command interpreter (shell) which of the standard shells most resembles the Korn shell (ksh), although it is not completely compatible. It includes enhancements of many types, notably in the command-line editor, options for customizing its behavior, filename globbing, features to make C-shell (csh) users feel more at home and extra features drawn from tcsh (another ‘custom’ shell). Zsh was written by Paul Falstad. Contents: zsh-3.0.5

Author:  Paul Falstad
Version:  3.0.5-6

gs lib Necessary files for running gs (ghostscript)

This package contains all startup files for calling gs (ghostscript). You will find some useful documentations about few printer drivers in the directory: /usr/share/ghostscript/4.03/doc/ Please read the files use.txt, devices.txt and for the experimental driver cdj850 hp850.txt. Currently there is the newer version 5.10 of ghostscript available. Due to a change in the licence of this version we are not allowed to put this version onto this cdrom. Nevertheless this version you will find this version on our FTP server.

Copyright:  (c) 1989-96 Aladdin Enterprises.
Author:  Peter Deutsch <ghost@aladdin.com>
Version:  5.10-0

0.1.3 D: Development (C, C++, Lisp, Perl, etc.)

autoconf An m4 macro package

This package is required for the configuration of many GNU source packages. Contents: autoconf-2.12

Copyright:  (c) 1994 Free Software Foundation, Inc.
Version:  2.12-8

automake Tool for automatically generating GNU style Makefile.in’s

Automake is a tool for automatically generating ”Makefile.in”s from files called ”Makefile.am”. ”Makefile.am” is basically a series of ”make” macro definitions (with rules being thrown in occasionally). The generated ”Makefile.in”s are compatible to the GNU Makefile standards.

Contents: automake-1.2

Copyright:  (c) 1994, 1995, 1996 Free Software Foundation, Inc.
Author:  Tom Tromey <tromey@cygnus.com>
Version:  1.2-10

beg Back End Generator

This package contains only the BEG executable to generate Source code for the GMD MOCKA Compiler. The Sources are not free.

For further information: /usr/doc/beg-1.75/

Contents: beg-1.75

Copyright:  (c) 1993 GMD Karlruhe, Germany
Author:  Helmut Emmelmann <emmel@ipd.info.uni-karlsruhe.de>
Version:  1.75-5

bin86 8086 assembler and linker

Needed to compile the kernel.

Contents: bin86-0.4

Author:

Bruce Evans <bde@runx.oz.au> , Linus Torvalds <Linus.Torvalds@cs.helsinki.fi> , H.J. Lu <hjl@nynexst.com>

Version:  0.4-8

binuti o GNU binutils (official binary release)

C compiler utilities: ar, gprof, ld, libconvert, nm, objdump, ranlib, size, strace, strip.

These utilities are needed whenever you want to compile a program or a kernel.

This is the official binutils version of H.J. Lu: ftp://tsx-11.mit.edu/pub/linux/packages/GCC/ Unlike the S.u.S.E. version, cross-development to other architectures is supported as long as you copy and install the appropriate include and library files.

Copyright:  (c) 1987-94 Free Software Foundation, Inc.
Version:  2.7.0.3-1

binutils GNU binutils

C compiler utilities: ar, gprof, ld, libconvert, nm, objdump, ranlib, size, strace, strip.

These utilities are needed whenever you want to compile a program or a kernel.

This version has been compiled by S.u.S.E. with the corresponding C library. They are only designed for i486-linux (ELF) and i486-linuxaout (a.out) but not for cross-compiling to other architectures such as Sparc, Alpha, and m68k.

Contents: binutils-2.8.1

Copyright:  (c) 1987-94 Free Software Foundation, Inc.

Author:

H.J. Lu <hjl@nynexst.com>, Ken Raeburn <raeburn@cygnus.com>, Ian Lance Taylor <ian@cygnus.com>, J.T. Conklin <jtc@cygnus.com>, Stan Shebs <shebs@cygnus.com>, Steve Chamberlain <sac@cygnus.com>, David J. Mackenzie <djm@cygnus.com>, Jeffrey A. Law <law@cygnus.com>, Sean Eric Fagan <sef@cygnus.com>, Jeffrey Osier <jeffrey@cygnus.com>, Jeffrey Wheat <cassidy@cygnus.com>, Mark Eichin <eichin@cygnus.com>, Roland H. Pesch <pesch@wrs.com>, Per Bothner <bothner@cygnus.com>, David D ‘Zoo’ Zuhn <zoo@cygnus.com>, Rob Savoye <rob@cygnus.com>, Jim Kingdon <kingdon@cygnus.com>, Mike Werner <mtw@cygnus.com>, Fred Fish <fnf@cygnus.com>, John Gilmore <gnu@cygnus.com>, Brendan Kehoe <brendan@cygnus.com>, D. V. Henkel-Wallace <gumby@cygnus.com>, K. Richard Pixley <rich@cygnus.com>, Bill Cox <bill@cygnus.com>, Peter Schauer <pes@regent.e-technik.tu-muenchen.de>, Eric Youngdale <eric@aib.com>, Richard Earnshaw <rwe@pegasus.esprit.ec.org>, Doug Evans <dje@cygnus.com>, Ian Dall <dall@hfrd.dsto.gov.au>, Jason Merrill <jason@cygnus.com>, Jim Wilson <wilson@cygnus.com>, Kung Hsu <kung@cygnus.com>, Michael Meissne <meissner@osf.org>, Michael Tiemann <tiemann@cygnus.com>, Richard Stallman <rms@gnu.ai.mit.edu>, Stu Grossman <grossman@cygnus.com>, Torbjorn Granlund <tege@cygnus.com>

Version:  2.8.1-12

bison GNU bison parser generator

Bison is a parser generator in the style of yacc(1). Contents: bison

Copyright:  (c) 1984-93 Free Software Foundation, Inc.

Author:

David J. MacKenzie <djm@gnu.ai.mit.edu>, Noah Friedman <friedman@ai.mit.edu>, Richard Stallman <rms@ai.mit.edu>

Version:  1.25-8

bissawt A java framework of java classes with its own look-and-feel

BISS-AWT is a java framework of about 150 Java classes for building graphical applications with stateof-the-art user interfaces. Instead of following the ”OS-native” look-and-feel (like Sun’s java.awt), it implements its own ”Java-native” look-and-feel. The framework can be extended without native library support (i.e. you can write your specialized versions of Text-Fields, Check-boxes, Lists etc. without being forced to write your own, non-portable, native libraries). !!! The actual version requires JDK1.0.x !!! Please read /usr/doc/packages/bissawt/README.SuSE! Documentation is available in LyX-format at /usr/doc/packages/bissawt/doc. Contents: biss-0.87 biss-awt-doc-0.87 biss

Copyright:  (c) BISS GmbH Germany
Author:  Peter Mehlitz <pcm@biss-net.com>, Joerg Mehlitz <jhm@biss-net.com>
Version:  0.87pl7-4

bwbasic There is Basic under Linux, too

A Basic interpreter. Example programs are in /usr/local/lib/bwbtest . Documentation: /usr/doc/misc/bwbasic-2.10.* Contents: bwbasic-2.10

Copyright:  (c) 1993, Ted A. Campbell
Author:  Ted A. Campbell <tcamp@delphi.com>
Version:  2.10-8

clisp A Common-Lisp interpreter

Common Lisp is a high-level, all-purpose programming language. CLISP is an implementation of Common Lisp that closely follows the book ”Common Lisp - The Language” by Guy L. Steele Jr. This package includes an interactive programming environment with interpreter, compiler and debugger. This environment will be started by the command clisp The documentation of CLISP is placed in the directories /usr/doc/packages/clisp/ /usr/doc/packages/clisp/doc/ As well as the conventional CLISP, the package also includes CLX, an extension of CLISP for the X Window System. Of course X11 has to be installed before the command clx is called. The description of this CLX version (new-clx) is placed in /usr/doc/packages/clisp/clx/ with the file README. The subdirectory /usr/doc/packages/clisp/clx/demos/ contains two nice applications. Contents: clisp-1997-12-06

Copyright:  (c) 1992,93 Bruno Haible und Michael Stoll, 1994-97 Bruno Haible und Marcus Daniels

Author:

Bruno Haible <haible@ma2s2.mathematik.uni-karlsruhe.de>, Michael Stoll <michael@rhein.iam.uni-bonn.de>, Marcus Daniels <marcus@sysc.pdx.edu>, Gilbert Baumann <gilbert@ma2s2.mathematik.uni-karlsruhe.de>

Version:  97.12.06-1

cvs Concurrent Versions System

CVS is a frontend to the Revision Control System, which is included in the standard Linux distribution. pcl-cvs, an Emacs frontend to CVS, is also included. Contents: cvs-1.9.24

Copyright:  (c) 1989-92 Brian Berliner, Jeff Polk
Author:  Brian Berliner <berliner@sun.com>, Jeff Polk <polk@bsdi.com>
Version:  1.9.24-1

efence ElectricFence

Tool for debugging your C-programs (malloc debugger). Contents: ElectricFence-2.0.5

Version:  2.0.5-8

f2c A Fortran-77 to C translator

Contains libf2c.so.0.11, ported by Rob Hooft. This package uses an ’f77’ script that hides the C translation process from the user. Contents: f2c libf2c-0.11

Copyright:  (c) 1990-95 AT&T Bell Laboratories and Bellcore.

Author:

David Gay <dmg@bell-labs.com>, Stu Feldman , Mark Maimone , Norm Schryer , Rob Hooft (maintainer for Linux) <Hooft@EMBL-Heidelberg.DE>

Version:  0.11-8

f2caout a.out version of libf2c

Copyright:  (c) 1990-95 AT&T Bell Laboratories and Bellcore.

Author:

David Gay <dmg@bell-labs.com>, Stu Feldman, Mark Maimone , Norm Schryer, Rob Hooft (maintainer for Linux) <Hooft@EMBL-Heidelberg.De>

Version:  0.10-1

flex flex - fast lexical analyzer generator

flex is a tool for generating scanners: programs which recognize lexical patterns in text. Contents: flex-2.5.4

Copyright:  (c) 1990 The Regents of the University of California.
Author:  Vern Paxson <vern@ee.lbl.gov>
Version:  2.5.4-8

fpk Free Pascal Compiler

FPK is a 32-bit compiler for the i386 processor. It is designed to be, as much as possible, source compatible with Turbo Pascal 7.0 and Delphi II, but it also enhances these languages with elements like function overloading.

Copyright:  (c) 1997 Michael Van Canneyt
Author:  Michael Van Canneyt <michael@tfdec1.fys.kuleuven.ac.be>
Version:  0.99.0-0

g77 GNU Fortran Compiler

This version of the GNU Fortran compiler needs GNU gcc 2.7.2.1. The binaries of the C compiler must be under /usr/lib/gcc-lib/i486-linux/2.7.2.1.

Copyright:  (c) 1995 Free Software Foundation, Inc.
Author:  James Craig Burley <burley@gnu.ai.mit.edu>
Version:  0.5.19.1-1

gcc The GNU C compiler and support files

Compiled with 486 optimization, to maximize performance on 486 processors, but works fine with 386.

NOTE: Be sure to install at least the following packages besides this one, or you won’t be able to compile: binutils, include, libc, and the third part of the kernel source. (the include files)

Original Version from FSF. Compiled with support for g77 and gnat.

Contents: g77-0.5.19.1 gcc-2.7.2.1 gnat-3.09-src gpc-2.0

Copyright:  (c) 1987-95 Free Software Foundation, Inc.

Author:

Adam Fedor <fedor@colorado.edu>, Alan Modra <alan@spri.levels.unisa.edu.au>, Allen Briggs <briggs@rrinc.com>, Andreas Schwab <schwab@issan.informatik.unidortmund.de>, Andrew Cagney <cagney@highland.com.au>, Andrew McCallum <mccallum@leopard.cs.rochester.edu>, Arne H. Juul <arnej@idt.unit.no>, Craig Burley <burley@gnu.ai.mit.edu>, DJ Delorie <dj@delorie.co)m, David Edelsohn <edelsohn@mhpcc.edu>, Doug Evans <dje@cygnus.com>, Douglas Rupp <drupp@cs.washington.edu>, Fergus Henderson <fjh@cs.mu.oz.au>, Gran Uddeborg <uddeborg@carmen.se>, Ian Lance Taylor <ian@cygnus.com>, J”orn Rennecke <amylaar@meolyon.hanse.de>, J.T. Conklin <jtc@cygnus.com>, Jason Merrill <jason@cygnus.com>, Jeffrey A. Law <law@cygnus.com>, Jim Meyering <meyering@na-net.ornl.gov>, Jim Wilson <wilson@cygnus.com>, John F. Carr <jfc@mit.edu>, Ken Raeburn <raeburn@wombat.gnu.ai.mit.edu>, Lee Iverson <leei@ai.sri.com>, Linus Torvalds <Linus.Torvalds@cs.Helsinki.FI>, Michael Gschwind <mike@lanai.vlsivie.tuwien.ac.at>, Michael Meissner <meissner@cygnus.com>, Mike Stump <mrs@cygnus.com>, Niklas Hallqvist <niklas@appli.se>, Oliver Kellogg <oliver.kellogg@space.otn.dasa.de>, Pat Rankin <rankin@eql.caltech.edu>, Paul Eggert <eggert@twinsun.com>, Per Bothner <bothner@cygnus.com>, Peter Flass <flass@lbdrscs.bitnet>, Philippe De Muyter <phdm@info.ucl.ac.be>, Richard Earnshaw <rearnsha@armltd.co.uk>, Richard Kenner <kenner@vlsi1.ultra.nyu.edu>, Richard Stallman <rms@mole.gnu.ai.mit.edu>, Roland McGrath <roland@churchy.gnu.ai.mit.edu>, Stan Cox <coxs@dgrtp.dg.com>, Stephen L Moshier <moshier@world.std.com>, Ted Lemon <mellon@toccata.fugue.com>, Tom Quiggle <quiggle@lovelace.engr.sgi.com>, Torbjorn Granlund <tege@cygnus.com>, Uwe Seimet <seimet@chemie.uni-kl.de>

Version:  2.7.2.1-6

gcc o The GNU C compiler (official binary release)

This is the official binary distribution of the GNU C Compiler for Linux. It includes gcc, g++ and gobjc.

Attention - This version does not work with g77 and gnat packages.

Copyright:  (c) 1987-95 Free Software Foundation, Inc.
Version:  2.7.2.1-1

gcl GNU Common Lisp

A Common Lisp interpreter with a Tcl/Tk interface.

Contents: gcl-2.2.2

Copyright:  (c) 1994 M. Hagiya, W. Schelter, T. Yuasa,
Author:  Bill Schelter <wfs@ma.utexas.edu>
Version:  2.2.2-7

gdb The GNU debugger

You can use GDB to debug programs written in C, C++, and Modula-2. GNU Fortran 77 is also partially included.

Contents: gdb-4.16

Copyright:  (c) 1986-95 Free Software Foundation, Inc.

Author:

Doug Evans <dje@cygnus.com>, Fred Fish <fnf@cygnus.com>, Ian Lance Taylor <ian@cygnus.com>, J.T. Conklin <jtc@cygnus.com>, Jason Molenda <crash@cygnus.com>, Jeffrey A. Law <law@cygnus.com>, Jim Kingdon <kingdon@cygnus.com>, Kung Hsu <kung@cygnus.com>, Michael Meissner <meissner@cygnus.com>, Per Bothner <bothner@cygnus.com>, Peter Schauer <pes@regent.e-technik.tu-muenchen.de>, Rob Savoye <rob@cygnus.com>, Stan Shebs <shebs@cygnus.com>, Stan Shebs <shebs@cygnus.com>, Steve Chamberlain <sac@cygnus.com>, Stu Grossman <grossman@cygnus.com>, Steve Chamberlain <sac@cygnus.com>

Version:  4.16-10

gdlib Gif-Manipulating Library

This is the gd gif-manipulating library, version 1.2. It was created to allow graphs, charts and the like to be generated on the fly for use on the World wide Web, but is useful for any application in which custom .GIFs are useful. It is not a paint program; it is a library. If you are looking for a paint program for the X Window System, check out David Koblas ”xpaint”, available by anonymous FTP from ftp.netcom.com in pub/ko/koblas. (I’m just a fan of Dave’s code.)

Author:  Tom Boutell <boutell@boutell.com>
Version:  1.2-2

gettext National Language Support, Message Translation

This Library stands under the GPL License, not under LGPL. So please remember to read /usr/doc/packages/gettext/COPYING and remember that you have to publish the sources of a program linked against this library.

Contents: gettext-0.10.32

Version:  0.10.32-6

gforth GNU Forth

Copyright:  (c) 1997 Free Software Foundation, Inc.

Author:

Anton  Ertl  <anton@mips.complang.tuwien.ac.at>,  Bernd  Paysan

<paysan@informatik.tu-muenchen.de>, Jens Wilke <wilke@jwdt.com>

Version:  0.3.0-4

gmp GNU MP library

Library for calculating huge numbers (integer and floating point).

Contents: gmp-2.0.2

Copyright:  (c) 1992-96 Free Software Foundation, Inc.
Author:  Torbjorn Granlund <tege@zevs.sics.se>
Version:  2.0.2-8

gnat GNU Ada Translator

The ”GNU Ada Translator”, an Ada83 and Ada9X front end to gcc-2.7.2. GCC 2.7.2 is required if you want to use this version of Ada.

Version:  3.09-2

gnats GNATS - Problem Report Management System

GNATS is a bug-tracking tool designed for use at a central support site. Software users who experience problems use electronic mail to communicate these problems to the the maintainers of that software; GNATS partially automates the tracking of these problems by: organizing problem reports into a database and notifying responsible parties of suspected bugs; allowing support personnel and their managers to edit, query and report on accumulated bugs; and providing a reliable archive of problems with a given program and a history of the life of the program by preserving its reported problems and their subsequent solutions. GNATS offers many of the same features offered by more generic databases. You can query and edit existing problem reports (PRs) as well as obtain reports on groups of PRs. The database itself is simply an ordered repository for problem reports; each PR receives a unique, incremental PRnumber, which identifies it throughout its lifetime. Contents: gnats-3.102-beta

Copyright:  (c) GPL

Author:

Brendan  Kehoe  <brendan@lisa.cygnus.com>,  Ian  Lance  Taylor

<ian@cygnus.com>, Jason Molenda <crash@godzilla.cygnus.co.jp>, Doug Evans <dje@canuck.cygnus.com>

Version:  3.102beta-9

gobjc GNU objective C Compiler

The source code is available in the gcc s package.

Version:  2.7.2.1-2

gpc GNU Pascal Compiler

A Pascal compiler based on the GCC 2.7.2.1 compiler.

Author:  Jukka Virtanen <jtv@hut.fi>, Peter Gerwinski <peter.gerwinski@uni-essen.de>
Version:  2.0-2

gpp GNU C++ Compiler

You can find the source code for gpp in the gcc s package.

Version:  2.7.2.1-3

guavac A free compiler for the Java language

Guavac is a free compiler for the Java language. Contents: guavac-0.3.1

Copyright:  (c) 1989, 1991 Free Software Foundation, Inc.
Author:  Effective Edge <info@eecs.com>
Version:  0.3.1-8

guile GNU’s Ubiquitous Intelligent Language for Extension

This is Guile, a portable, embeddable Scheme implementation written in C. Guile provides a machine independent execution platform that can be linked in as a library when building extensible programs. Contents: guile-1.2

Copyright:  (c) GLP

Author:

Jim Blandy <jimb@floss.red-bean.com>, George Carrette, Aubrey Jaffer, Gary Houston <ghouston@actrix.gen.nz>, Tom Lord, Anthony Green, Mikael Djurfeldt <mdj@mdj.nada.kth.se>, Mark Galassi <rosalia@papageno.lanl.gov>, Tim Pierce <twp@twp.tezcat.com>, Marius Vollmer <mvo@zagadka.ping.de>, Tom Tromey <tromey@cygnus.com>

Version:  1.2-9

icmake A Maker for C-Fans

A Maker using a C-like syntax. I like that :-) Documentation can be found in /usr/doc/misc/icmake-6.17.ps; examples are in /usr/local/lib/icmake-6.17-examples/ Contents: icmake-6.17

Copyright:  (c) ICCE 1992-1994
Author:  Frank B. Brokken <F.B.Brokken@icce.rug.nl>, Karel Kubat <K.Kubat@icce.rug.nl>
Version:  6.17-9

indent formats C source code

Version:  1.9.1-8

java Java Developers Kit v1.1.3

With this package you are able to create your own platform-independent JAVA applications. You can compile your it with the Java compiler and start it with the Java interpreter or the Appletviewer. The authors of this package stipulated that the sources should not cross hands more than two times. For this reason we are unable to provide a source package for this program. To get the sources, look at: http://www.blackdown.org/java-linux/Mirrors.cgi For more information please refer to /usr/doc/packages/java.

Contents: linux-jdk.1.1.3-v2

Copyright:  (c) 1994, 1995, 1996, 1997 Sun Microsystems, Inc.
Author:  Steve Byrne <sbb@gnu.ai.mit.edu>
Version:  1.1.3v2-8

javadoc Java JDK1.1.3 documentation and demos

The documentation and demo apps and applets that belongs to the jdk1.1.3. ( includes the original sources of jdk1.1.3 at /usr/doc/packages/javadoc/src.zip ) All will be installed at /usr/doc/packages/javadoc. Contents: jdk113demo jdk113doc

Copyright:  (c) (c) 1994, 1995, 1996, 1997 Sun Microsystems, Inc.
Version:  1.1.3-5

javarunt Java Runtime Environment v1.1.3-v2

Copyright:  (c) 1997 Sun Microsystems, Inc.
Author:  Steve Byrne <sbb@gnu.ai.mit.edu>
Version:  1.1.3v2-7

jdk115 Java Developers Kit 1.1.5

The Java Developers Kit version 1.1.5 Contents: jdk1.1.5

Copyright:  (c) 1994, 1995, 1996, 1997 Sun Microsystems, Inc.
Author:  Steve Byrne <sbb@gnu.ai.mit.edu>
Version:  1.1.5-0

jresep Java Runtime Environment 1.1.3v2

The separated version of JRE. There are two components which make up JRE: the basic VM runtime, and the internationalization support ( /usr/lib/rt and /usr/lib/i18n ). Please read /usr/doc/packages/jresep/* Contents: linux-rt.1.1.3-v2 rt

Copyright:  (c) 1997 Sun Microsystems, Inc.
Author:  Steve Byrne <sbb@gnu.ai.mit.edu>
Version:  1.1.3v2-7

kaffe A virtual machine to run Java(tm) code

Kaffe is a virtual machine design to execute Java byte code. Unlike other virtual machines available, this machine performs ”just-in-time” code conversion from the abstract code to the host machine’s native code. *** Kaffe now supports JDK 1.1.1 Please take a look at /usr/doc/packages/kaffe/* Contents: kaffe-0.9.2-package-javasoft.com kaffe-0.9.2

Copyright:  (c) 1996 T. J. Wilkinson & Associates, London, UK
Author:  Tim Wilkinson <tim@tjwassoc.co.uk>
Version:  0.9.2-5

lclint LCLint - a tool for statically checking C programs

With minimal effort, LCLint can be used as a better lint. If additional effort is invested adding annotations to programs, LCLint can perform stronger checks than can be done by any standard lint. Contents: lclint-2.2a

Copyright:  (c) Massachusetts Institute of Technology 1994, 1995, 1996.
Author:  David Evans <evs@larch.lcs.mit.edu>
Version:  2.2a-9

leda LEDA 3.5.2 a library of efficient data types and algorithms

LEDA is a library of the data types and algorithms of combinatorial computing. The main features are: LEDA provides a sizable collection of data types and algorithms in a form which allows them to be used by non-experts. In the current version, this collection includes most of the data types and algorithms described in the text books of the area.

Any user of LEDA needs a license. This license is available free of charge for academic research and teaching. Any commercial use of this software requires a license which is distributed by LEDA Software GmbH.

Any use by a firm, a commercial company or a government institution is considered to be commercial use.

LEDA Software GmbH Postfach 151101 66041 Saarbruecken Germany email:leda@mpi-sb.mpg.de fax: +49 681 842502

For a research license please contact ftp://ftp.mpi-sb.mpg.de/pub/LEDA/www/ledaavailable.html

Contents: leda-3.5.2 Any user of LEDA needs a license. This license is available free of charge for academic research and teaching. Any commercial use of this software requires a license which is distributed by LEDA Software GmbH.

Any use by a firm, a commercial company or a government institution is considered to be commercial use.

LEDA Software GmbH Postfach 151101 66041 Saarbruecken Germany email:leda@mpi-sb.mpg.de fax: +49 681 842502

For a research license please contact ftp://ftp.mpi-sb.mpg.de/pub/LEDA/www/ledaavailable.html

Copyright:  (c) LEDA Software GmbH
Author:  LEDA Software GmbH <leda@mpi-sb.mpg.de>
Version:  3.5.2-8

libc Libraries for the C compiler

These libraries are needed to compile C code.

Contents: libc-5.4.44 libpthreads

Copyright:  (c) 1981-95 Regents of the University of California. , Free Software Foundation, Inc.

Author:

H.J. Lu <hjl@gnu.ai.mit.edu>, Alan Cox <Alan.Cox@linux.org>, Adam J. Richter <adam@yggdrasil.com>, Alan Modra <alan@spri.levels.unisa.edu.au>, Andreas Schwab <schwab@issan.informatik.uni-dortmund.de>, Andrew Walker <andy@keo.kvaerner.no>, Andries Brouwer <aeb@cwi.nl>, Charles Lopes <tjarls@infm.ulst.ac.uk>, Chip Salzenberg <chip@fin.uucp>, Christoph Kern <kern@informatik.uni-wuerzburg.de>, Corey Minyard <minyard@wf-rch.cirr.com>, Craig Smith <craig@rezrov.natinst.com>, Daniel Quinlan <quinlan@yggdrasil.com>, Delman Lee <delman@mipg.upenn.edu>, Dominik Kubla <dominik.kubla@unimainz.de>, Dsouza <m.dsouza@mrc-applied-psychology.cambridge.ac.uk>, Eric Youngdale <eric@aib.com>, Peter Eriksson <pen@signum.se>, Florian La Roche <florian@suse.de>, Fred N. Van Kempen <waltje@uwalt.nl.mugnet.org>, Gerhard Koerting <g.koerting@koerting.techem.ruhr-uni-bochum.de>, H. Peter Anvin <hpa@yggdrasil.com>, Harald Koenig <koenig@tat.physik.unituebingen.de>, Igor Sharfmesser <igor@irbis.alma-ata.su>, J. Alan Eldridge <alane@wozzle.linet.org>, Jaakko Hyvatti <jaakko.hyvatti@www.fi>, James H. Cloos Jr. <cloos@rahul.net>, Jeschke <i6210109@ws.rz.tu-bs.de>, Jim Van Zandt <jrv@vanzandt.mv.com>, Jochen Hein <Hein@Informatik.TU-Clausthal.de>, Jochen Kiemes <jochen kiemes@s2.maus.de> , John F. Haugh Ii <jfh@rpp386>, Jonathan Kamens <jik@cam.ov.com>, Karl Hakimian <hakimian@eecs.wsu.edu>, Kaz Sasayama <kaz@lilia.iijnet.or.jp>, Marc Gingold <gingold@amoco.saclay.cea.fr>, Martin Kuck <martin@unifix.de>, Martin Pirker <pirker@eiunix.tuwien.ac.at>, Mathew G. Monroe <mmex@shadowland.pc.cc.cmu.edu>, Matthias Urlichs <urlichs@smurf.noris.de>, Matthias Urlichs <urlichs@test.smurf.noris.de>, Michael Meskes <meskes@debian.org>, Michael Riepe <Michael.Riepe@stud.unihannover.de>, Mike Kennedy <kmb1@gate.net>, Mike Meissner <meissner@cygnus.com>, Miquel van Smoorenburg <miquels@drinkel.ow.nl>, Mitchum Dsouza <m.dsouza@mrc-applied-psychology.cambridge.ac.uk>, Nick Holloway <nick.holloway@alfie.demon.co.uk>, Nickolay Saukh <nms@ussr.eu.net>, Nicolai Langfeldt <janl@ifi.uio.no>, Niibe Yutaka <gniibe@mri.co.jp>, Olaf Flebbe <o.flebbe@science-computing.uni-tuebingen.de>, Patrick D’Cruze <pdcruze@orac.iinet.com.au>, Paul Eggert <eggert@twinsun.com>, Per Bothner <bothner@cygnus.com>, Peter Himmler <himmler@sparcy.mayn.sub.de>, Philip Gladstone <philip.gladstone@mail.citicorp.com>, Remy Card <remy.card@masi.ibp.fr>, Rick Sladkey <jrs@world.std.com>, Robert Andrew Ryan <rr2b+@andrew.cmu.edu>, Roland Mcgrath <roland@gnu.ai.mit.edu>, Rolf Rossius <rossius@hrz.tu-chemnitz.de>, Ross Biro <biro@yggdrasil.com>, Sean Mcneil <sean@mcneil.com>, Stephen L Moshier <moshier@world.std.com>, Stephen Tweedie <sct@dcs.ed.ac.uk>, Steve Losen <losen@delphi.com>, Steve M. Robbins <steve@nyongwa.montreal.qc.ca>, Swen Thuemmler <swen@uni-paderborn.de>, Theodore Ts’O <tytso@mit.edu>, Thomas Pundt <pundtt@goedel.uni-muenster.de>, Ulrich Drepper <drepper@gnu.ai.mit.edu>, Uwe Ohse <uwe@tirka.gun.de>, Vadim V. Vlasov <vvv@ms3.inr.ac.ru>, Wes Cowley <wcowley@cftnet.com>

Version:  5.4.44-1

libd Profiling and Debugging versions of libc.

Version:  5.4.44-1

libgpp C++-library

Version:  2.7.2.1-1

libobjc libobjects

The library for the GNU Objective C compiler. Contents: libobjects-0.1.19

Version:  0.1.19-9

linclude Include files for the kernel

The kernel include files aren’t only needed for rebuilding the kernel but are also required to compile system programs. So that you don’t have to re-install the complete kernel tree each time you build a program, we’ve placed the include files in an extra package. If you want to rebuild the kernel please install the ’linux’ package as well.

Author:  Linus Torvalds <torvalds@transmeta.com>
Version:  2.0.33-0

linux Rest of the kernel source code.

This package contains the rest of the Linux Kernel source files.

Contents: linux-2.0.33

Author:

Linus Torvalds <torvalds@transmeta.com>,  ,  Alain L. Knaff

<Alain.Knaff@poboxes.com>, Alan Cox <alan@lxorguk.ukuu.org.uk>, Alessandro Rubini <rubini@ipvvis.unipv.it>, Andrew Haylett <ajh@primag.co.uk>, Andrew Tridgell <Andrew.Tridgell@anu.edu.au>, Andries Brouwer <aeb@cwi.nl>, Angelo Haritsis <ah@doc.ic.ac.uk>, Arindam Banerji <axb@cse.nd.edu>, Avery Pennarun <apenwarr@bond.net>, Axel Boldt <boldt@math.ucsb.edu>, Bas Laarhoven <bas@vimec.nl>, Bernhard Kaindl <bkaindl@netway.at>, Bill Bogstad <bogstad@pobox.com>, Bjorn Ekwall <bj0rn@blox.se>, Bob Frey <bobf@advansys.com>, Bradley McLean <brad@bradpc.gaylord.com>, Bruno Haible <haible@ma2s2.mathematik.uni-karlsruhe.de>, Chih-Jen Chang <chihjenc@scf.usc.edu>, Chris Smith <csmith@convex.com>, Corey Minyard <minyard@wf-rch.cirr.com>, Craig Metz <cmetz@inner.net>, Craig Small <csmall@triode.apana.org.au>, Daniel Quinlan <quinlan@pathname.com>, Danny ter Haar <dth@cistron.nl>, Dario Ballabio <dario@milano.europe.dg.com>, Darren Senn <sinster@darkwater.com>, David C. Niemi <niemi@erols.com>, David Davies <davies@wanton.lkg.dec.com>, David Gentzel <gentzel@telerama.lm.com>, David Mosberger-Tang <davidm@azstarnet.com>, David S. Miller <davem@caip.rutgers.edu>, Derek Atkins <warlord@MIT.EDU>, Dirk Hohndel <hohndel@aib.com>, Dirk Melchers <dirk@merlin.nbg.sub.org>, Dirk Verworner, Dmitry S. Gorodchanin <begemot@bgm.rosprint.net>, Donald Becker <becker@cesdis.gsfc.nasa.gov>, Doug Evans <dje@cygnus.com>, Drew Eckhardt <drew@PoohSticks.ORG>, Drew Sullivan <drew@ss.org>, Eberhard Moenkeberg <emoenke@gwdg.de>, Ed Carp <ecarp@netcom.com>, Eddie C. Dost <ecd@skynet.be>, Eric S. Raymond <esr@thyrsus.com>, Eric Schenk <Eric.Schenk@dna.lth.se>, Eric Youngdale <eric@aib.com>, Florian La Roche <rzsfl@rz.uni-sb.de>, Frank Xia <qx@math.columbia.edu>, Fred Baumgarten <dc6iq@insu1.etec.uni-karlsruhe.de>, Fred N. van Kempen <waltje@uwalt.nl.mugnet.org>, Frederic Potter <Frederic.Potter@masi.ibp.fr >, Fritz Elfert <fritz@wuemaus.franken.de>, G”unter Windau <gunter@mbfys.kun.nl>, Gerard Roudier <groudier@iplus.fr>, Gero Kuhlmann <gero@gkminix.han.de>, Gordon Chaffee <chaffee@plateau.cs.berkeley.edu>, Grant Guenther <grant@torque.net>, Greg Hankins <gregh@cc.gatech.edu>, Greg Page <greg@caldera.com>, Greg Wettstein <greg@wind.rmcc.com>, H. Peter Anvin <hpa@zytor.com>, H.J. Lu <hjl@gnu.ai.mit.edu>, Hamish Coleman <hamish@zot.apana.org.au>, Hamish Macdonald <hamish@border.ocunix.on.ca>, Hannu Savolainen <hannu@voxware.pp.fi>, Hans Lermen <lermen@elserv.ffm.fgan.de>, Hans-Joachim Widmaier <hjw@zvw.de>, Harald Koenig <koenig@tat.physik.unituebingen.de>, Harald Milz <hm@seneca.linux.de>, Heiko Eissfeldt <heiko@colossus.escape.de heiko@unifix.de>, Hennus Bergman <hennus@cybercomm.nl>, Henrik Storner <storner@osiris.ping.dk>, Ian A. Murdock <imurdock@gnu.ai.mit.edu>, Ian Jackson <iwj10@cus.cam.ac.uk>, Ian Kluft <ikluft@thunder.sbay.org>, Jacques Gelinas <jacques@solucorp.qc.ca>, James B. MacLean <macleajb@ednet.ns.ca>, Jeff Tranter <Jeff Tranter@Mitel.COM>, Jeffrey A. Uphoff <juphoff@nrao.edu>, Jeremy Fitzhardinge <jeremy@sw.oz.au>, Jochen Hein <jochen.hein@delphi.central.de>, Joerg Reuter <jreuter@lykos.oche.de>, Johan Myreen <jem@vipunen.hut.fi>, John A. Martin <jam@acm.org>, John Aycock <aycock@cpsc.ucalgary.ca>, John Boyd <boyd@cis.ohio-state.edu>, John E. Gotts <jgotts@engin.umich.edu>, John Shifflett <john@geolog.com>, Jon Tombs <jon@gtex02.us.es>, Jonathan Naylor <g4klx@g4klx.demon.co.uk>, Jos Vos <jos@xos.nl>, Juan Jose Ciarlante <jjciarla@raiz.uncu.edu.ar>, Juergen Weigert <jnweiger@immd4.informatik.uni-erlangen.de>, J”urgen Fischer <fischer@etinf.fho-emden.de>, Kai Harrekilde-Petersen <khp@dolphinics.no>, Kai M”akisara <Kai.Makisara@metla.fi>, Kai Petzke <wpp@marie.physik.tu-berlin.de>, Karl Fogel <kfogel@cs.oberlin.edu>, Karl Keyte <kkeyte@koft.rhein-main.de>, Ken Pizzini <ken@halcyon.com>, Kenji Tsutomu Hollis <khollis@bitgate.com>, Kevin E. Martin <martin@cs.unc.edu>, Kevin Lentin <kevinl@cs.monash.edu.au>, Krishna Balasubramanian <balasub@cis.ohio-state.edu>, Lars Wirzenius <liw@iki.fi>, Laurence Culhane <loz@holmes.demon.co.uk>, Lawrence Foard <entropy@world.std.com>, Leo Spiekman <leo@netlabs.net>, Leonard N. Zubkoff <lnz@dandelion.com>, Marc Zyngier <maz@gloups.fdn.fr>, Marco van Wieringen <mvw@mercury.mcs.nl.mugnet.org>, Mark Lord <mlord@pobox.com>, Markus Kuhn <mskuhn@cip.informatik.uni-erlangen.de>, Martin Mares <mj@k332.feld.cvut.cz>, Martin Schulze <joey@linux.de>, Martin von Loewis <loewis@informatik.hu-berlin.de>, Matt Welsh <mdw@sunsite.unc.edu>, Matthias Urlichs <urlichs@noris.de>, Matti Aarnio <mea@utu.fi>, Michael A. Griffith <grif@cs.ucr.edu>, Michael Callahan <callahan@maths.ox.ac.uk>, Michael Hipp <mhipp@student.uni-tuebingen.de>, Michael K. Johnson <johnsonm@redhat.com>, Michael Meskes <meskes@topsystem.de>, Michael Neuffer <mike@i-Connect.Net>, Michael O’Reilly <michael@iinet.com.au>, Miguel de Icaza Amozurrutia <miguel@nuclecu.unam.mx>, Mike Jagdis <jaggy@purplet.demon.co.uk>, Mike McLagan <mike.mclagan@linux.org>, Mike Shaver <shaver@ingenia.com>, Miquel van Smoorenburg <miquels@cistron.nl>, Nick Holloway <Nick.Holloway@alfie.demon.co.uk>, Nigel Gamble <nigel@nrg.org>, Nigel Metheringham <Nigel.Metheringham@ThePLAnet.net>, Niibe Yutaka <gniibe@mri.co.jp>, Olaf Kirch <okir@monad.swb.de>, Orest Zborowski <orestz@eskimo.com>, Paal-Kristian Engstad <engstad@funcom.com>, Pat Mackinlay <pat@it.com.au>, Patrick Volkerding <volkerdi@ftp.cdrom.com>, Paul Gortmaker <gpg109@rsphy1.anu.edu.au>, Pauline Middelink <middelin@polyware.iaf.nl>, Peter Bauer <100136.3530@compuserve.com>, Peter De Schrijver <stud11@cc4.kuleuven.ac.be>, Peter MacDonald, Phil Lewis <beans@bucket.ualr.edu>, Philip Blundell <Philip.Blundell@pobox.com>, Philip Gladstone <philip@raptor.com>, Ralf Baechle <ralf@gnu.ai.mit.edu>, Ralf Flaxa <rfflaxa@immd4.informatik.uni-erlangen.de>, Randolph Bentson <bentson@grieg.seaslug.org>, Ray Burr <ryb@nightmare.com>, Ray Dassen <jdassen@wi.LeidenUniv.nl>, Raymond Chen <raymondc@microsoft.com>, Remy Card <Remy.Card@masi.ibp.fr>, Richard Henderson <richard@gnu.ai.mit.edu>, Rick Miller <rdmiller@execpc.com>, Rick Sladkey <jrs@world.std.com>, Rik Faith <faith@cs.unc.edu>, Rob W. W. Hooft <hooft@EMBL-Heidelberg.DE>, Robert Sanders <gt8134b@prism.gatech.edu>, Roger E. Wolff <wolff@dutecai.et.tudelft.nl>, Ron Holt <ron@novell.com>, Ross Biro <bir7@leland.Stanford.Edu>, Rudolf Koenig <rfkoenig@immd4.informatik.uni-erlangen.de>, Savio Lam <lam836@cs.cuhk.hk>, Scott Snyder <snyder@fnald0.fnal.gov>, Sebastian Hetze <she@lunetix.de>, Simmule Turner <sturner@tele-tv.com>, Simon Shapiro <shimon@i-Connect.Net>, Stefan Probst <snprobst@immd4.informatik.unierlangen.de>, Stefan Reinauer <stepan@home.culture.mipt.ru>, Stephen R. van den Berg (AKA BuGless) <berg@pool.informatik.rwth-aachen.de>, Stephen Rothwell <Stephen.Rothwell@canb.auug.org.au>, Stephen Tweedie <sct@dcs.ed.ac.uk>, Stuart Cheshire <cheshire@cs.stanford.edu>, Terry Dawson <terry@perf.no.itg.telecom.com.au>, Theodore Ts’o <tytso@mit.edu>, Thomas Bogendoerfer <tsbogend@alpha.franken.de>, Thomas Dunbar <tdunbar@vtaix.cc.vt.edu>, Thomas Sailer <sailer@ife.ee.ethz.ch>, Thomas Uhl <uhl@sun1.rz.fh-heilbronn.de>, Todd J. Derr <tjd@fore.com>, Tommy Thorn <Tommy.Thorn@irisa.fr>, Torsten Duwe <Torsten.Duwe@informatik.unierlangen.de>, Tsu-Sheng Tsao <tsusheng@scf.usc.edu>, Tuomas J. Lukka <Tuomas.Lukka@Helsinki.FI>, Volker Lendecke <lendecke@namu01.Num.Math.Uni-Goettingen.de>, Warner Losh <imp@village.org>, Wayne Davison <davison@borland.com>, Werner Almesberger <werner.almesberger@lrc.di.epfl.ch>, Werner Zimmermann <zimmerma@rz.fht-esslingen.de>, William (Bill) Metzenthen <billm@suburbia.net>, William E. Roadcap <roadcapw@cfw.com>, Willy Konynenberg <willy@xos.nl>, Winfried Tr”umper <winni@xpilot.org>, Yuri Per <yuri@pts.mipt.ru>

Version:  2.0.33-0

38 lthreads Linuxthreads - POSIX 1003.1c kernel threads for Linux

This is a release (BETA) of LinuxThreads, a BiCapitalized implementation of the Posix 1003.1c ”pthread” interface for Linux. Contents: linuxthreads-0.5

Copyright:  (c) 1996 Xavier Leroy
Author:  Xavier Leroy <Xavier.Leroy@inria.fr>
Version:  0.5-9

lx hack Source of the newest hacker’s kernel

The source tree of the newest development kernel. As these kernels contain the latest features and enhancements, you may run into problems when using this kernel. So we explicitly state that we do not give support for these kernels. Contents: linux-2.1.91 patches-until-2.1.91

Author:  Linus Torvalds <Linus.Torvalds@Helsinki.FI> et al.
Version:  2.1.91-0

lx suse Kernel sources with additional drivers

Kernel sources with additional drivers:

- New Tekram SCSI controller

- New BSD-NCR driver

- New driver for Tulip network cards

- Several drivers for parallel port CD-ROM drives Contents: linux-2.0.33.SuSE

Copyright:  (c) Linus Torvalds
Author:  Linus Torvalds <torvalds@transmeta.com>
Version:  2.0.33.SuSE-4

make GNU make command

Version:  3.76.1-6

minigcc The GCC Preprocessor without the C-Compiler

This Package contains just the Preprocessor which is used by the XFree86 packages.

Copyright:  (c) 1987-95 Free Software Foundation, Inc.
Version:  2.7.2.1-1

mocka Modula-2 compiler

mocka is a real modula2-compiler, now available as ELF version. The developers ask you to register to know how many of you use it since the money they get for developing depends on the amount of people using it. So please register! Documentation is to be found under: /usr/doc/packages/mocka Start with ’mocka’ (instead of ’mc’) Contents: mocka mocka9502main-linux mocka9502src

Copyright:  (c) 1993 by GMD Karlruhe, Germany
Version:  9502-9

mtc Modula-2 / C Converter

Unfortunately, there is no documentation available for this package. Contents: mtc-25sep92

Copyright:  (c) 1990 GMD - Universit”at Karlsruhe
Author:  Josef Grosch <grosch@karslruhe.gmd.de>
Version:  92.9.25-5

p2c A Pascal to C translator.

Translates PASCAL programs to C. These can then be compiled using gcc. Contents: p2c-1.20

Copyright:  (c) 1989-91 Free Software Foundation.
Author:  Dave Gillespie <daveg@csvax.cs.caltech.edu>
Version:  1.20-9

patch GNU patch program

Version:  2.1-8

pmake BSD make

This may be required if you’re going to port software with Makefiles designed for BSD. The binary is installed as usr/bin/pmake. Man pages are included.

Contents: pmake-1.0

Copyright:  (c) 1988-90 Regents of the University of California
Author:  Karl London <karl@borg.demon.co.uk>
Version:  1.0-8

prcs PRCS -- Project Revision Control System

The Project Revision Control System, PRCS, is the front end to a set of tools that (like CVS) provide a way to deal with sets of files and directories as an entity, preserving coherent versions of the entire set.

Contents: prcs-1.2.0

Copyright:  (c) 1997 Josh MacDonald
Author:  Josh MacDonald
Version:  1.2.0-7

python Python

Python is an interpreted object-oriented programming language, and is often compared to Tcl, Perl, Scheme or Java. A survey of Python you could find in the included documentation and tutorial. If you have web access, point your browser to http://www.python.org.

Contents: python-1.5

Copyright:  (c) 1991-1995 by Stichting Mathematisch Centrum, Amsterdam, The Netherlands.
Author:  Guido van Rossum <guido@python.org>
Version:  1.5-3

rcs Revision Control System

RCS, the Revision Control System, manages multiple revisions of files. RCS can store, retrieve, log, identify, and merge revisions. It is useful for files that are frequently revised e.g. programs, documentation, graphics, and papers.

Contents: rcs-5.7

Copyright:  (c) 1982-89 Walter F. Tichy, 1990-95 Paul Eggert

Author:

Walter F. Tichy <tichy@ira.uka.de>, Paul Eggert <eggert@twinsun.com>, Adam Hammer <hammer@cs.purdue.edu>, Guy Harris <guy@sun.com>, Jay Lepreau <lepreau@cs.purdue.edu>

Version:  5.7-8

scheme UMB Scheme

UMB Scheme is an implementation of the language described in the IEEE Standard for the Scheme Programming Language (December, 1990). All syntax, variables and procedures are implemented. Integers are implemented as fixnums and bignums, rationals as pairs of integers, (inexact) reals as double-precision floats, and (inexact) complex numbers as pairs of double-precision floats.

Contents: scheme-2.5

Copyright:  (c) Copyright: 1988
Author:  William R. Campbell <bill@cs.umb.edu>
Version:  2.5-10

slang Display control

Library with routines for display access. Similar to ncurses, but it has its own programming interface. Slang uses the entries in /usr/lib/terminfo.

Contents: slang

Copyright:  (c) 1993 John E. Davis
Author:  John E. Davis <davis@amy.tch.harvard.edu>
Version:  0.99.38-9

smaltalk GNU Smalltalk

GNU Smalltalk is an implementation which is almost Smalltalk-80 conform (^{TM}ParkPlace Systems) as described in Adele Goldberg and David Robson’s book ’Smalltalk-80: the Language and its Implementations’, which will be referred to as the ’Blue Book’.

Smalltalk is an object oriented programming language, meaning that the programmer not only has to keep the data in his head but the operations which can be applied to this object. The ability of an object of keeping not only track of the data but also to those operations that are connected to itself are inseparable glued together. The amount of actions that can be performed by the object is precisely defined by the amount of operations of this special object (called methods in Smalltalk). You can’t ’see’ the contents of an object from outside the object. Seen from the outside an object is a black box which has some particular operations and states, but that’s about all you know. In Smalltalk everything is an object. This includes variables, executable procedures (methods) and stack frames (context of invoked methods or context of a block). Every object is an instance of a class. You can see a class as a datatype and its incorporated functions. An instance is a particular variable of this datatype. If you want to perform an operation on one of these objects you have to send it a message and the object reacts by performing the correspondent action. To print something, e.g., you could send the message like this: someObject print The message you send is indeed the name of a method (procedure) which should be carried out. If you send a message to an object Smalltalk tries to find out the method that has been defined for this kind of object. First it starts looking at the objects class. If there is no such class it bothers the parent class, in the grandfather class etc.. At the top of this class-construction there is a class called Object which has no ancestors. If Smalltalk couldn’t find a method until then it produces an error code. For a complete explanation of Smalltalk there is a tutorial which is provided some extra example programs. Additionally the X Smalltalk interfaces ’stix (gst stix) as well as the TCP interface (gst tcp) are included.

Contents: smalltalk-1.2.alpha5

Copyright:  (c) 1990-1996 Free Software Foundation, Inc.
Author:  Steve Byrne <sbb@gnu.ai.mit.edu>
Version:  1.2.alpha5-8

smeiffel SmallEiffel

Implementation of the programming language Eiffel. If you want to use the programming environment, you should include the directory /usr/lib/SmallEiffel/bin into your PATH environment variable.

Contents: SmallEiffel

Copyright:  (c) Dominique COLNET, Suzanne COLLIN
Author:  Dominique Colnet <colnet@loria.fr>
Version:  0.03-9

strace Trace system calls of a program.

With strace you can trace the activity of a program. Information as to any system calls the program made and the signals it has received/processed can be seen. Even child processes can be tracked.

Contents: strace-3.1

Copyright:  (c) 1991-94 Paul Kranenburg, Branko Lankester, Rick Sladkey

Author:

Paul Kranenburg <pk@cs.few.eur.nl>, Branko Lankester <branko@hacktic.nl>, Rick Sladkey <rs@world.std.com>

Version:  3.1-8

svgaaout Svgalib Super-VGA Graphics Library (a.out)

This library is included for compatibility reasons. The ’real’ svgalib can be found in the package svgalib.

Copyright:  (c) 1993 Tommy Frandsen, 1993 Harm Hanemaayer

Author:

H. Hanemaayer <hhanemaa@cs.ruu.nl>, Michael Weller <eowmob@expmath.uni-essen.de>, Orest Zborowski <obz@Kodak.com>, David Wexelblat <dwex@XFree86.org>, Thomas Roell <roell@informatik.tu-muenchen.de>, Kevin E. Martin <martin@cs.unc.edu>, Rickard E. Faith <faith@cs.unc.edu>, Christopher Wiles <wileyc@fsr.net>

Version:  1.26-1

svgalib Svgalib Super-VGA Graphics Library

A video hardware graphics library, based on VGAlib. Supports most standard and non-standard VGA modes, and Super-VGA modes for Tseng ET4000, Cirrus 542x (including 15/16/24 bit modes), Mach32, Trident 8900c/9000, and other cards. Documentation is located in /usr/doc/packages/svgalib . Contents: svgalib-1.2.12

Copyright:  (c) 1993 Tommy Frandsen, 1993 Harm Hanemaayer

Author:

H. Hanemaayer <hhanemaa@cs.ruu.nl>, Michael Weller <eowmob@expmath.uni-essen.de>, Orest Zborowski <obz@Kodak.com>, David Wexelblat <dwex@XFree86.org>, Thomas Roell <roell@informatik.tu-muenchen.de>, Kevin E. Martin <martin@cs.unc.edu>, Rickard E. Faith <faith@cs.unc.edu>, Christopher Wiles <wileyc@fsr.net>

Version:  1.2.12-6

swi pl SWI-Prolog distribution

An extensive Prolog distribution. Please read the READMEs in /usr/doc/packages/swi-pl and the pl manpage. Contents: swi-pl

Copyright:  (c) 1990-1996 Jan Wielemaker
Author:  Jan Wielemaker <jan@swi.psy.uva.nl>
Version:  2.8.6-6

toolpack extensive collection of tools for Fortran programmers

An extensive collection of tools for Fortran programmers. Contents: toolpack-1.2

Copyright:  (c) Numerical Algorithms Group Ltd
Author:  Thomas Koenig (hacked for Linux) <ig25@rz.uni-karlsruhe.de>, Roman Augustyn
Version:  1.2-10

tvision tvision

Porting from Borland TurboVision for UNIX Contents: tvision-0.5

Copyright:  (c) Borland copyright and BSD-style copyright
Author:  Sergio Sigala <ssigala@globalnet.it>
Version:  0.5-0

tya TYA is a 100% inofficial JIT-compiler

TYA is a “100% inofficial” JIT-compiler. Documentation, test- and example-programs can be found at /usr/doc/packages/tya. Contents: tya04

Copyright:  (c) 1997 Albrecht Kleine
Author:  Albrecht Kleine <kleine@ak.sax.de>
Version:  04-2

xclass Linux Explorer + library

This package contains a library for uniform presentation of fvwm95 programs and (as a demo) the Linux Explorer, a file viewer, a Run utility and a Win95 clock. Best with fvwm95. Documentation in: /usr/doc/packages/xclass /usr/X11R6/lib/xclass/tutorial

Copyright:  (c) 1996, Hector Peraza
Author:  Hector Peraza <peraza@mitac11.uia.ac.be>
Version:  970707-7

xscheme An object-oriented scheme

DOC: Documentation

XScheme is an implementation of the Scheme programming language with extensions to support object-oriented programming. There are currently implementations of XScheme running on the IBM-PC and clones under MS-DOS, on the Macintosh, the Atari-ST, and the Amiga. It is completely written in the programming language ’C’ and is easily extended with user written built-in functions and classes. It is available in source form to non-commercial users. XScheme follows the ”Revised 3 Report on the Algorithmic Language Scheme”. Contents: xscheme-0.16

Copyright:  (c) 1989 David Michael Betz
Author:  David Michael Betz
Version:  0.16-7

yacc BSD yacc

”bison -y” should react exactly the same. Contents: yacc

Version:  91.7.30-5

joystick joystick support (kernel module and binaries)

Kernel drivers for analog joysticks as a loadable module. Supports up to two joysticks. The kernel module is contained in package kernmod as /lib/modules/KERNEL-VERSION/misc/joystick.o Unless you are using the SuSE kernel sources, you might have to copy /usr/doc/packages/joystick/joystick.h to /usr/include/linux Contents: joystick-0.8.0

Copyright:  (c) Freely distributable
Author:  asmith@cbnewsd.att.com
Version:  0.8.0-8

0.1.4 DOC: Documentation

allman All man pages

This package contains duplicates of all man pages of every package on the CD (for those who don’t want to install every single package just to get the man pages online).

Version:  98.4.1-0

books Some books

This package contains some books, which are installed under /usr/doc/Books. The books are: The German ”Linux Anwender Handbuch” - Sebastian Hetze, Dirk Hohndel, Martin Mueller, Olaf Kirch a.o. ”Linux Installation and getting started” - Matt Welsh ”Linux Kernel Hackers’ Guide” - Michael K. Johnson ”Linux Programmers Guide” - Sven Goldt and Sven van der Meer ”Linux Network Administrators Guide” - Olaf Kirch ”Linux System Administrators Guide” - Lars Wirzenius ”Linux Users Guide” - Larry Greenfield ”DVIPS: A TeX Driver” ”The C Preprocessor” - Richard M. Stallman ”Using and porting GNU CC” - Richard M. Stallman ”The GNU C Library Reference Manual” - Sandra Loosemore, Roland McGrath, Andrew Oram and Richard M. Stallman The German ”Das kleine TeX Buch” - Fritz Cremer The English and German versions of the S.u.S.E. Linux handbook. As well, there is an extensive list of Unix books (”A Concise Guide to UNIX Books”) by Samuel Ko, and an introduction to the theme ”World Wide Web server Software”. Contents: install-guide-2.3 khg-0.6 lpg-0.4 nag-1.0.dvi nag-1.0.src sag-0.6 user-beta-1-src

Version:  98.3.24-0

booksfnt Pixmap fonts for the dvi files in the distribution

Installing this package will save you a lot of time waiting for the fonts to be computed for the dvi files in the distribution. Since not all fonts used in the S.u.S.E. handbooks are included in the teTeX package, it is recommended that you install this package for them. Contents: dvi-i386

Version:  98.4.1-0

cdb CDB - Components Database

With CDB (Component DataBase), you may get easily information about status of support for many hardware components. Start the Server with /sbin/init.d/cdb start. The frontend is done via WWW. Just start a browser on your computer with the URL: <your hostname>/cdb/english

Copyright:  (c) 1997 S.u.S.E. GmbH
Author:  Richard Jelinek <rj@suse.de>, Simon Pogarcic <sim@suse.de>
Version:  98.3.6-0

changes Change-Logs for the packages

Copyright:  (c) 1996 S.u.S.E. GmbH, F”urth, Germany
Version:  98.4.1-0

dns faq DNS-FAQ

The Domain-Name-Service (DNS) FAQ.

Copyright:  (c) 1997, Chris Peckham, cdp@pfmc.net
Author:  Chris Peckham <cdp@pfmc.net>
Version:  98.3.29-0

elispint Programming in Emacs Lisp: An Introduction

Programming in Emacs Lisp: An Introduction Contents: emacs-lisp-intro-1.05

Copyright:  (c) FSF, GPL
Author:  Robert J. Chassell <bob@gnu.ai.mit.edu>
Version:  1.05-6

elispman GNU Emacs Lisp Reference Manual 19 (Version 2.4.2)

GNU Emacs Lisp Reference Manual edition 2.4.2, corresponding to GNU Emacs version 19.34. Contents: elisp-manual-19-2.4.2

Copyright:  (c) Free Software Foundation, Inc., GPL
Author:  Bil Lewis, Dan LaLiberte, Richard Stallman <rms@gnu.ai.mit.edu>
Version:  2.4.2-9

fhs Filesystem Hierarchy Standard

This standard consists of a set of requirements and guidelines for file and directory placement under UNIX-like operating systems. The guidelines are intended to support interoperability of applications, system administration tools, development tools, and scripts as well as greater uniformity of documentation for these systems. Contents: fhs-2.0

Copyright:  (c) 1994, 1995, 1996, 1997 Daniel Quinlan

Author:

Daniel  Quinlan  <quinlan@pathname.com>,  Brandon  S.  Allbery

<bsa@kf8nh.wariat.org>, Keith Bostic <bostic@cs.berkeley.edu>, Drew Eckhardt <drew@colorado.edu>, Rik Faith <faith@cs.unc.edu>, Stephen Harris <sweh@spuddy.mew.co.uk>, Ian Jackson <ijackson@cus.cam.ac.uk>, John A. Martin <jmartin@acm.org>, Ian McCloghrie <ian@ucsd.edu>, Chris Metcalf <metcalf@lcs.mit.edu>, Ian Murdock <imurdock@debian.org>, David C. Niemi <niemidc@clark.net>, Eric S. Raymond <esr@thyrsus.com>, Mike Sangrey <mike@sojurn.lns.pa.us>, David H. Silber <dhs@glowworm.firefly.com>, Theodore Ts’o <tytso@athena.mit.edu>, Stephen Tweedie <sct@dcs.ed.ac.uk>, Fred N. van Kempen <waltje@infomagic.com>

Version:  2.0-0

gawkdoc Documentation for gawk

Version:  3.0.3-13

gccinfo GNU info-pages for gcc

Version:  2.7.2.1-1

ge info Info files for GNU emacs

This package contains all info files for GNU emacs. All these files can be read online with GNU emacs and describe emacs and some of its modes.

Copyright:  (c) 1985-96 Free Software Foundation, Inc.

Author:

Richard Stallman <rms@gnu.ai.mit.edu>, Karl Heuer <kwzh@gnu.ai.mit.edu>, Geoff Voelker <voelker@cs.washington.edu>, David J. MacKenzie <djm@gnu.ai.mit.edu>, Morten Welinder <terra@gnu.ai.mit.edu>, Paul Reilly <pmr@gnu.ai.mit.edu>, Roland McGrath <roland@gnu.ai.mit.edu>, Brian Fox <bfox@gnu.ai.mit.edu>, Jim Blandy <jimb@gnu.ai.mit.edu>, Eric S. Raymond <esr@thyrsus.com>

Version:  19.34b-9

ger man German man pages

Most part of these German man pages has been written by Sebastian Hetze - coauthor of the popular ”Linux Anwender Handbuch”. Meanwhile there are about two dozens of people translating man pages. Coordinator of this project has been Andries E. Brouwer <aeb@cwi.nl> and now is Martin Schulze <joey@infodrom.north.de>. Since this collection is considered kind of incomplete much work still has to be done. If you are interested to join in this group please contact Martin Schulze- he and of course all of the other ’translators’ will be happy hearing from you...;) Please note that these man pages are distributed under a variety of copyright licenses. Although these licenses permit free distribution of the nroff sources contained in this package, commercial distribution may impose other requirements (e.g., acknowledgment of copyright or inclusion of the raw nroff sources with the commercial distribution). If you distribute these man pages commercially, it is your responsibility to figure out your obligations. (For many man pages, these obligations require you to distribute nroff sources with any pre-formatted man pages that you provide.) Each file that contains nroff source for a man page also contains the author(s) name, email address, and copyright notice. Contents: man-pages-de-0.1

Copyright:  (c) 1993 Sebastian Hetze, diverse

Author:

Sebastian Hetze <she@soho.lunetix.de>, Andries E. Brouwer <aeb@cwi.nl>, Martin Schulze <joey@infodrom.north.de>, Rene Tschirley <gremlin@cs.tuberlin.de>, Norbert K”umin <norbert.kuemin@lugs.ch>, Elmar Jansen <ej@pumuckel.gun.de>, Aldo Valente <aldo@dagobar.rhein.de>, Lars J. Brandt <ljbrandt@jorma.ping.de>, Lutz Donnerhacke <lutz@as-node.jena.thur.de>, Michael Haardt <michael@cantor.informatik.rwth-aachen.de>, Christian Schmidt <c.schmidt@ius.gun.de>, Norbert Weuster <weuster@etecs0.uni-duisburg.de>, Ralf Baechle <ralf@waldorf-gmbh.de>, Patrick Rother <krd@burn.rhein-ruhr.de>, Jonas Rovan <jonas@blitz.de>, Wolfgang Jung <woju@keep.in-berlin.de>, Florian Brons <florian@merlin.pn.org>, Michael Burghardt <mb@suse.de>, Ralf Baumert <bau@heineken.chemie.uni-dortmund.de>, Heiko Schlittermann <heiko@lotte.sax.de>, Carsten Borchardt <carsten@bor.north.de>, Kai Dupke <dupke@koma.han.de>, Harald Wieland <dg9bcy@epc99.fbe.fh-wilhelmshaven.de>, Ralf W. Stephan <ralf@franken.de>, Jochen Hein <Hein@Informatik.TU-Clausthal.de>, Hanno Wagner <wagner@bidnix.bid.fh-hannover.de>

Version:  0.1-3

gnuhtml HTML Documentation for GNU-Software

The GNU-info-files converted to html.

Version:  97.11.18-4

gppinfo Documentation for libg++

Version:  2.7-1

howto Collection of HOWTOs from the ’Linux Documentation Project’.

Documentation on many tasks relating to the setup and operation of Linux software. These documents will be installed in /usr/doc/howto. You can read it with ‘less <filename>‘. Contents: howtos

Version:  98.3.31-0

howtodvi HOWTOs in DVI format

Version:  98.3.31-0

howtohtm HOWTOs in HTML format

Version:  98.3.31-0

howtoinf HOWTOs in info format

Version:  98.3.31-0

howtops HOWTOs as PostScript files

Version:  98.3.31-0

howtosgm SGML sources of the HOWTOs

Version:  98.3.31-0

i4ldoc Documentation for ISDN

More documentation about ISDN and Linux Includes:

- i4l-FAQ 1.1.1 (English version: pre 1.0.5)

- leafsite from Berhard Hailer (http://www.lrz-muenchen.de/˜ ui161ab/www/isdn) Version 2. Feb. 1997

Copyright:  (c) Berhard Hailer
Author:  Bernhard.Hailer@lrz.uni-muenchen.de
Version:  97.6.29-2

inf2htm info2html, converts INFO pages to HTML pages

This packages contains the CGI script info2html which creates html pages out of info documents. info2html generates by calling within a WWW browser a html page of the file /usr/info/dir You can easy click on the info references and nodes to get the on the fly generated html pages. Contents: arrows info2html1.1

Copyright:  (c) 1993-95 Karl Guggisberg
Author:  Karl Guggisberg <guggis@iam.unibe.ch>
Version:  1.1-8

kcs Kernel Change Summaries

Kernel Change Summaries for all the kernel versions since 1.3.42. Contents: kcs

Copyright:  (c) 1996 Michael Elizabeth Chastain
Author:  Michael Elizabeth Chastain <mec@duracef.shout.net>
Version:  98.2.3-2

ldp Linux Documentation Project HTML archive

A lot of Linux docu in HTML format. Contents: ldp

Copyright:  (c) GPL
Author:  A lot!, Greg Hankins <gregh@sunsite.unc.edu>
Version:  98.3.31-0

ldpman LDP Man-Pages

The man pages of the Linux Documentation Project. They are maintained by Andries Brouwer. Contents: man-pages-1.18

Author:

Alain L. Knaff <Alain.Knaff@imag.fr>, Alan Cox <Alan.Cox@linux.org>, Andries Brouwer <aeb@cwi.nl>, Arnt Gulbrandsen <agulbra@nvg.unit.no>, Bjorn Ekwall <bj0rn@blox.se>, Christopher Oliver <oliver@fritz.co.traverse.com>, Dan Miner <dminer@nyx.cs.du.edu>, Daniel Quinlan <quinlan@yggdrasil.com>, Darren Senn <sinster@darkwater.com>, David Metcalfe <david@prism.demon.co.uk>, Drew Eckhardt <drew@PoohSticks.ORG>, Giorgio Ciucci <giorgio@crcc.it>, Graeme W. Wilford <G.Wilford@ee.surrey.ac.uk>, Ian Jackson <iwj10@cus.cam.ac.uk>, Jens Schweikhardt <Jens.Schweikhardt@rus.uni-stuttgart.de>, Jim Van Zandt <jrv@vanzandt.mv.com>, Jochen Hein <Hein@Informatik.TU-Clausthal.de>, Jochen Karrer <cip307.cip.physik.uni-wuerzburg.de>, Luigi P. Bai <lpb@neosoft.com>, Mark D. Roth <roth@uiuc.edu>, Michael Elizabeth Chastain <mec@shell.portal.com>, Michael Haardt <michael@cantor.informatik.rwth-aachen.de>, Michael Shields <shields@tembel.org>, Mike Battersby <mike@starbug.apana.org.au>, Mitchum DSouza <Mitch.DSouza@NetComm.IE>, Nicolai Langfeldt <janl@ifi.uio.no>, Neil P. Soveran-Charley <athan@mordor.mersinet.co.uk>, Ralf Baechle <ralf@waldorf-gmbh.de>, Rik Faith <faith@cs.unc.edu>, Robert K. Nichols <Robert.K.Nichols@att.com>, Stephen Lee <sl14@cornell.edu>, Steven Greenland <stevegr@sugar.neosoft.com>, Thomas K. Dyas <tdyas@eden.rutgers.edu>, Thomas Koenig <ig25@rz.uni-karlsruhe.de>, Ulrich Drepper <drepper@gnu.ai.mit.edu>, Urs Thuermann <urs@isnogud.escape.de>

Version:  1.18-3

libcinfo Infofiles for the GNU C library and GDBM library

Version:  5.0.9-1

man9 Some man pages from section 9

This man pages are containing infos about kernel structures.

Contents: man9

Author:

Cyrus Durgin <cider@speakeasy.org>, Keith Owens <kaos@ocs.com.au>, Kirk Petersen <kirk@speakeasy.org>, Stephen Williams <steve@icarus.com>

Version:  98.2.14-1

manyfaqs A collection of frequently asked questions/answers

Contents of the package: The full set of Linux HOWTOs. The comp.lang FAQs for Fortran, GCC, G++, TCL, Objective-C, Lisp, Scheme, and Perl. The PostScript FAQs. The Linux FAQ. The comp.unix.question FAQs. The comp.windows.x X FAQ and many other FAQs relating to X. manyfaqs: These FAQs will be installed under /usr/doc/faq.

Contents: faq

Version:  98.2.3-2

odshtml HTML-Pages from Offenes Deutsches Schulnetz

Version:  97.11.18-4

perlref Perl 5 Reference Guide

Author:  Johan Vromans <jvromans@squirrel.nl>
Version:  5.001.2-1

ppp nt Linux connection to Windows NT

Documentation, but only in German: Dieses Dokument beschreibt, wie man eine PPP-Verbindung zwischen einem LINUX-Host und einem Windows NT Server unter Verwendung der PAP/CHAP Authentifizierung mit automatischem Rueckruf durch den Windows NT Server herstellt

Contents: ppp nt-1.1

Copyright:  (c) Siegfried Schoen <sigi@f30x.erl.scn.de>
Author:  Siegfried Schoen <sigi@f30x.erl.scn.de>
Version:  1.1-2

rfc The RFC’s (request for comments)

This is the reference concerning computer related questions. The discussed matter might be a little to theoretical for laymen but professionals do find all what they are looking for.

Copyright:  (c) many
Version:  98.3.27-0

sdb S.u.S.E. Support DataBase

A collection of frequently asked support questions. These are written in HTML and can be found under /usr/doc/support-db/sdb e. The English version is under construction.

Copyright:  (c) (c) 1996 S.u.S.E. GmbH, Fuerth, Germany
Author:  kfr@suse.de
Version:  98.4.1-0

sdb cgi Support Database search function

Search function in Support Database There will be installed two CGI-scripts for searching in Support Database via a local installed webserver (Apache). More info: /usr/doc/support-db/sdb e/sdb.html

Copyright:  (c) S.u.S.E. GmbH, Fuerth, Germany
Author:  Klaus Franken <kfr@suse.de>
Version:  0.70-9

selfhtml selfhtml

HTML3.2 Referenz and Tutorial Contents: selfhtml

Copyright:  (c) Stefan M”unz 1998
Author:  Stefan M”unz <muenz@compuserve.com>
Version:  6.1-0

suse lx Handbook S.u.S.E. Linux (german)

Copyright:  (c) S.u.S.E. GmbH, Fuerth
Version:  5.2.0.1-7

suse lxe Handbook S.u.S.E. Linux (english)

Copyright:  (c) S.u.S.E. GmbH, Fuerth
Version:  5.2.0.1-4

susebuch LaTeX macros for the S.u.S.E. Linux Handbook

Copyright:  (c) (c) S.u.S.E., GPL
Author:  Martin Scherbaum <maddin@suse.de>, Werner Fink <werner@suse.de>
Version:  1.0.2-1

susehilf S.u.S.E. Help-System (base)

This package contains the basics of the S.u.S.E. Help-System. To prevent unnecessary disk usage the man pages will be formatted ”on the fly”. Thus you can only view man pages that are already installed. Start ’susehelp’ by invoking help. susehelp --help shows a list of configuration possibilities. Contents: http-rman rman

Copyright:  (c) 1995,96 S.u.S.E. GmbH, F”urth, Germany
Author:  <suse@suse.de>
Version:  98.4.1-0

susepak S.u.S.E. Help-System (package descr. German)

This package contains German descriptions for all the packages in html format.

Copyright:  (c) 1995 S.u.S.E. GmbH, F”urth, Germany
Author:  <suse@suse.de>
Version:  98.4.1-0

susepake S.u.S.E. Help-System (package descr. English)

Copyright:  (c) 1995 S.u.S.E. GmbH, F”urth, Germany
Author:  <suse@suse.de>
Version:  98.4.1-0

toppp T-Online / PPP scripts

T-Online now offers PPP access. This package contains some PPP example scripts for T-Online. These scripts are copied to /usr/doc/packages/toppp. A description can be found under /usr/doc/packages/toppp/README.SuSE.

\underline{E: Emacs}

Contents: toppp

Copyright:  (c) 1997 Klaus Franken, S.u.S.E. GmbH, Fuerth, Germany
Author:  Klaus Franken <kfr@suse.de>
Version:  98.2.13-1

uucpdoc Taylor UUCP Documentation

Copyright:  (c) 1991-93 Ian Lance Taylor
Author:  Ian Lance Taylor <ian@cygnus.com>
Version:  1.06.1-8

wie dvi German HOWTOs in DVI format

Version:  98.3.27-0

wie geht Collection of German HOWTOs

Author:

Brian Ward <bri@blah.math.tu-graz.ac.at>, Grant Taylor <gtaylor@picante.com>, Jens Vonderheide <Jens Vonderheide@wharfrat.fido.de>, Marco Budde <Budde@tuharburg.d400.de>, Matt Welsh <mdw@sunsite.unc.edu>, Peter S”utterlin <pit@unisw.gwdg.de>

Version:  98.3.27-0

wie html German HOWTOs in HTML format

Version:  98.3.27-0

wie info German HOWTOs in info format

Version:  98.3.27-0

wie pdf German HOWTOs in Acrobat (PDF) format

Version:  98.3.27-0

wie ps German HOWTOs as PostScript files

Version:  98.3.27-0

wie sgml SGML sources of the German HOWTOs

Version:  98.3.27-0

xe info Info files for X emacs

This package contains all info files for X emacs. All these files can be read online with X emacs and describe xemacs and some of its modes.

Copyright:

(c) 1994, 1995 Board of Trustees, University of Illinois, 1994 Free Software Foundation, Inc.

Author:

Chuck Thompson <cthomp@cs.uiuc.edu>, Ben Wing <wing@666.com>, and many other contributors

Version:  20.4-1

xf86html HTML Pages from XFree86

Version:  98.3.25-0

0.1.5 E: Emacs

elib Elib, a Lisp library for the GNU-Emacs

This package contains the completed GNU-Emacs Lisp Library Elib in Version 1.0. Elib was developed as a library for Emacs-Lisp programs for the same reasons that libg++ was for C++ programs. Elib contains macros for:

- ‘container’ data structures (queues, stacks, AVL trees, etc.)

- string handling routines missing in standard emacs

- routines to handle so-called ‘cookies’ in a buffer

Contents: elib-1.0

Copyright:  (c) Inge Wallin, Per Cederqvist
Author:  Inge Wallin <inge@lysator.liu.se>, Per Cederqvist <ceder@lysator.liu.se>
Version:  1.0-10

ge auc TeX add-on for Emacs

These macros make it easier for you to write TeX documents with GNU emacs. Documentation is to be found under: /usr/doc/packages/auctex/ and in the dvi-files math-ref.dvi tex-ref.dvi (reference cards) and in info files auctex in emacs infomode. AuC-TeX is integrated in X-Emacs 19.15, therefore you may not install this packages. Note that the binary formats of the byte-compiled lisp files of the both emacs editors are incompatible, therefore you can not use this package for X-Emacs. Contents: auctex-9.7p

Copyright:  (c) 1994 AUC TeX development

Author:

Kresten Krab Thorup <auc-tex mgr@iesd.auc.dk>, Per Abrahamsen <abraham@dina.kvl.dk>

Version:  9.7p-10

ge exec GNU-Emacs

Eight Megabytes Always Continuously Swapping ;-) Love it or leave it. Contents: emacs-19.34b

Copyright:  (c) 1985-96 Free Software Foundation, Inc.

Author:

Richard Stallman <rms@gnu.ai.mit.edu>, Karl Heuer <kwzh@gnu.ai.mit.edu>, Geoff Voelker <voelker@cs.washington.edu>, David J. MacKenzie <djm@gnu.ai.mit.edu>, Morten Welinder <terra@gnu.ai.mit.edu>, Paul Reilly <pmr@gnu.ai.mit.edu>, Roland McGrath <roland@gnu.ai.mit.edu>, Brian Fox, Jim Blandy <jimb@gnu.ai.mit.edu>, Eric S. Raymond <esr@thyrsus.com>

Version:  19.34b-9

ge lisp Several lisp files for GNU emacs

Several lisp files not needed for running GNU emacs. Most of these files are pre byte compiled and therefore not necessary.

Copyright:  (c) 1985-96 Free Software Foundation, Inc.

Author:

Richard Stallman <rms@gnu.ai.mit.edu>, Karl Heuer <kwzh@gnu.ai.mit.edu>, Geoff Voelker <voelker@cs.washington.edu>, David J. MacKenzie <djm@gnu.ai.mit.edu>, Morten Welinder <terra@gnu.ai.mit.edu>, Paul Reilly <pmr@gnu.ai.mit.edu>, Roland McGrath <roland@gnu.ai.mit.edu>, Brian Fox <bfox@gnu.ai.mit.edu>, Jim Blandy <jimb@gnu.ai.mit.edu>, Eric S. Raymond <esr@thyrsus.com>

Version:  19.34b-9

ge site Site-lisp extensions for GNU emacs

This package contains extensions for GNU emacs of the site-lisp path of emacs. It puts together the AUC-TeX package (version 9.7p) and a system wide startup file (site-start.el) which uses a few lisp macro files combined by us. For copyright have a look on these files. Note, that GNU-Emacs and XEmacs are incompatible. They use different binary code for compiled lisp files and they have different builtin lisp functions, not only names of such functions are different. Therfore you will find in /etc/skel/ an .emacs for your personal ˜ /.emacs which will switch between different init files. Fore GNU emacs there is a /etc/skel/.gnu-emacs for a personal ˜ /.gnu-emacs and therefore your ˜ /.emacs or ˜ /.gnu-emacs files should be checked against the /etc/skel versions. You can always make your preferred ˜ /.gnu-emacs by copying the /etc/skel/.emacs and /etc/skel/.gnuemacs file in your HOME directory and incorporating your extensions or changes. Contents: site-lisp

Author:

Jonathan Stigelman <stig@hackvan.com>, Per Abrahamsen <abraham@iesd.auc.dk>, Werner Fink <werner@suse.de>, Alon Albert <alon@milcse.rtsg.ai.edu>, Wayne Mesard <wmesard@cs.stanford.edu>, Bruno Salvy <salvy@inria.inria.fr>, Marc Andreessen <marca@ncsa.uiuc.edu>, Inge Frick <inge@nada.kth.se>, Lawrence R. Dodd <dodd@roebling.poly.edu>, Mark B. Motl <motl@cssun.tamu.edu>

Version:  1.1c-1

EMU: Emulators

lacheck Syntax checker for LaTeX documents

Lacheck is a syntax checker for LaTeX documents. It is not any more a part of newer AuC-TeX packages. Therefore this packages is given as its own. You can use lacheck as standalone program too.

Contents: lacheck-1.23

Copyright:  (c) 1991,92 Kresten Krab Thorup, 199394 Per Abrahamsen

Author:

Kresten Krab Thorup <auc-tex mgr@iesd.auc.dk>, Per Abrahamsen <abraham@dina.kvl.dk>

Version:  1.23-9

psgml Emacs Add-On to edit SGML documents

‘psgml’ supports you while editing SGML documents. It respects the context of the used DTD (Document Type Definition) and offers the valid elements and attributes. Contents: psgml-1.0.1 Compiled for GNU Emacs (XEmacs has its own version!). Insert the content of the file /usr/doc/packages/psgml/dot.emacs in your own .emacs resp. .gnu-emacs!

Copyright:  (c) GPL

Author:

James Clark <jjc@clark.com>, Lennart Staflin <lenst@lysator.liu.se>, David Megginson <dmeggins@sprynet.com>, Matthias Clasen <mclasen@sun2.mathematik.unifreiburg.de>

Version:  1.0.1-11

xe exec X-Emacs

This is the current version of X-Emacs, former known as Lucid-Emacs. It’s got a nice GUI but is quite big and derived from GNU-emacs 19.13 (current is 19.31 or 19.34). Lisp macros are not necessarily interchangeable between GNU-emacs and Xemacs. This is mainly important for translated .elc files and the key macros.

Contents: xemacs-20.4-elc xemacs-20.4-info xemacs-20.4-mule xemacs-20.4

Copyright:

(c) 1994, 1995 Board of Trustees, University of Illinois, 1994 Free Software Foundation, Inc.

Author:

Chuck Thompson <cthomp@cs.uiuc.edu>, Ben Wing <wing@666.com>, and many other contributors

Version:  20.4-1

xe lisp Several lisp files for X emacs

Several lisp files not needed for running X emacs. Most of these files are pre byte compiled and therefore not necessary. The true X-emacs addict will install them nevertheless.

Copyright:

(c) 1994, 1995 Board of Trustees, University of Illinois, 1994 Free Software Foundation, Inc.

Author:

Chuck Thompson <cthomp@cs.uiuc.edu>, Ben Wing <wing@666.com>, and many other contributors

Version:  20.4-1

0.1.6 EMU: Emulators

adamem A Coleco Vision and Coleco ADAM emulator

Adamem is an emulator for the Z80 based systems ColecoVision and Coleco ADAM. To get the emulator working you have to install the necessary ROMs. The package ei helps with this task as well.

Contents: adamem

Copyright:  (c) 1996 Marcel de Kogel
Author:  Marcel de Kogel <m.dekogel@student.utwente.nl>
Version:  1.2-1

dosemu The DOS emulator

This package allows MS-DOS programs to be started in Linux. A virtual machine (the DOS box) provides the necessary BIOS functions. You will find the sources in the ZQ series. Documentation can be found in ’/usr/doc/packages/dosemu’ and in the man page, as well as in the sources.

If you wish to allow non-root users to run the emulator, you must enter their names in the file ’/etc/dosemu.users’. Even so, some DOS programs will run only if DOSEMU was started by root. Please note as well that you should only allow ’trusted users’ to use the DOS emulator, as it inherently permits hardware access that would otherwise be forbidden to ’normal’ users. The parameter HogThreshold in ’/etc/dosemu.conf’ states when an ’idling’ DOSEMU should return the CPU to Linux. It is set to use the CPU as much as it wants (0). The lowest useful value is 10. If that isn’t fast enough, you (as UID root) can get maximum performance with nice -19 dos -D-a 2>/dev/null Don’t be surprised though if the other Linux processes then run very sluggishly. Some graphic cards with S3 chips can only be run in 16 color VESA modes, even if you have inserted video vga console graphics chipset s3 memsize 1024 in ’/etc/dosemu.conf’. On sensitive systems you should therefore use the ’secure on’ option in ’/etc/dosemu.conf’, which prevents the use of ’system.com’. This option also closes another security hole occurring with (mainly Dos4gw based) DPMI clients, which allows them access to the complete user space (including the DOSEMU code). The Aout DOSEMU has a second security hole (the possibility of replacing libdosemu) which has disappeared in ELF, since the ELF version no longer generates libdosemu. Contents: dosemu-0.66.7 In order to start dosemu, you will need a bootable hdimage. Please read the handbook, cd to /var/lib/dosemu and start ./setup-hdimage. ATTENTION: if you have already used dosemu: the configuration has changed slightly, so please read the dosemu-part of the handbook even if you think you know what you are doing :-)

Copyright:  (c) Matthias Lautner, Robert Sanders, Carnegie Mellon University

Author:

Michael E. Deisher <deisher@enws99.EAS.ASU.EDU>, James B. Maclean <macleajb@ednet.ns.ca>, Mike Batchelor <mikebat@netcom.com>, Lam Lai Yin <lam836@cs.cuhk.hk>, Uwe Bonnes <bon@elektron.ikp.physik.thdarmstadt.de>, Scott Buchholz <scottb@eecs.nwu.edu>, J. Lawrence Stephan <jlarry@ssnet.com>, John Davis <davis@space.mit.edu>, David Etherton <etherton@netcom.com>, Pasi Eronen <peronen@vipunen.hut.fi>, Jason E Gorden <gorden@jegnixa.hsc.missouri.edu>, Matthew Grant <grantma@ritz.equinox.gen.nz>, David Hansen <dhansen@metapath.com>, Jochen Hein <Hein@Informatik.TU-Clausthal.de>, Alan Hourihane <alanh@fairlite.demon.co.uk>, John Kohl <jtk@atria.com >, Rob Janssen <pe1chl@rabo.nl>, Harald Koenig <koenig@nova.tat.physik.uni-tuebingen.de>, Vinod G Kulkarni <vinod@cse.iitb.ernet.in>, Kang-Jin Lee <lee@tengu.inberlin.de>, Kevin P Lawton <bochs@world.std.com>, Marty Leisner <leisner@sdsp.mc.xerox.com>, Hans Lermen <lermen@suse.de>, Dong Liu <dliu@faraday-gw.njit.edu>, Tim Van der Linden <timvdl@innet.be>, Martin Ludwig <Martin.Ludwig@ruba.rz.ruhr-uni-bochum.de>, Alistair MacDonald <alistair@slitesys.demon.co.uk>, Amit Margalit <amitm@netvision.net.il>, Lawrence K Mao <lkmao@unf.edu>, Rod May <stoke@melbourne.DIALix.oz.au>, WAYNE P MEISSNER <n1046128@student.fit.qut.edu.au>, Lutz Molgedey <lutz@summa.physik.hu-berlin.de>, Erik Mouw <jakmouw@et.tudelft.nl>, Christoph Niemann <Christoph.Niemann@linux.org>, Max Parke <mhp@light.lightlink.com>, Mark Rejhon <mdrejhon@magi.com>, Ronnie <ronnie@epact.se>, Ed Sirett <ed@cityscape.co.uk>, Larry Stephan <jlarry@ssnet.com>, Corey Sweeney <corey@bbs.xnet.com>, Andrew Tridgell <Andrew.Tridgell@anu.edu.au>, Theodore Ts’o <tytso@mit.edu>, Robert Sanders <gt8134b@prism.gatech.edu>, Matthias Lautner, Tim Bird, Joel N. Weber II

Version:  0.66.7-10

ei Installation utility for various emulators.

Ei provides a comfortable and unified way to install ROM images of various emulators. This can be done from a MS-DOS formatted disk or via a given path in the Filesystem. You may also install software for these emulators via Internet.

Contents: ei

Copyright:  (c) (c) 1996,97 S.u.S.E GmbH, Fuerth Germany
Author:  Richard Jelinek <rj@suse.de>
Version:  0.4-6

ibcs2 The iBCS2 emulator

The iBCS2 emulator allows you to run binaries of other UNIX system such as SCO-Unix[tm] under Linux. The emulator is a loadable module that can be loaded into the kernel using the command ’insmod iBCS’. If you have activated the kernel daemon, the ibcs2 emulator module will be loaded automatically on demand. CAUTION: This package contains the libc.so.1 for SYSV 4.x systems! If you install this package you may encounter strange things! If you compile a program and specify the linker option ”-L/usr/lib”, the linker tries to link THIS library which of course doesn’t work. In such a case, either specify the option ”-L/lib -L/usr/lib” or do not specify any path. Then the program will be successfully linked. Contents: ibcs-2.0.970513 libnsl-960201 protlib-960201 sco-libs-940731 svr4-shlib-961222 CAUTION: This package contains the libc.so.1 for SYSV 4.x systems! If you install this package you may encounter strange things! If you compile a program and specify the linker option ”-L/usr/lib”, the linker tries to link THIS library which of course doesn’t work. In such a case, either specify the option ”-L/lib -L/usr/lib” or do not specify any path. Then the program will be successfully linked.

Copyright:  (c) 1993-95 Eric Youngdale, Joe Portman, Mike Jagdis, Al Longyear
Version:  970513-13

ines Nintendo Entertainment System Emulator

This is a Emulator for the Nintendo Entertainment System. It is fully operational. You just need the Games (*.NES-Images).

Contents: ines

Copyright:  (c) 1996,97 Marat Fayzullin
Author:  Marat Fayzullin <fms@freeflight.com>
Version:  0.6-3

stella Stella - an Atari 2600VCS Emulator

The Atari 2600 Video Computer System (VCS), introduced in 1977, was the most popular home video game system of the early 1980’s. Now you can enjoy all of your favorite Atari 2600 games on your PC thanks to Stella!

Contents: stella

Copyright:  (c) 1996,1997 Bradford W. Mott
Author:  Bradford W. Mott <bwmott@eos.ncsu.edu>
Version:  0.6-9

stonx ST on X - an Atari emulator

This program allows you to emulate one (or more..;) ) ATARIs under Linux. You can run almost any GEM program as well as some ”hardware-close” games.

Contents: STonX-0.6.7-epsilon

Copyright:  (c) 1995-1997 Marinos Yannikos and Matrin D. Griffiths
Version:  0.6.7.epsilon-6

stonxsft Feed STonX with this ATARI software

A collection of selected software for the ATARI emulator. An overview of the enclosed programs you can find in suse.txt under /var/lib/stonx/software or on drive C: when you start stonx.

Contents: stonxsft

Version:  98.2.13-1

uae An Commodore Amiga emulator

UAE is an amiga emulator, that nearly completely simulates an A500. You’ll need a Kickstart ROM as well as a Workbench, to work with UAE. Some programs (demos/games) will work without these too.

Contents: uae-0.6.8

Copyright:  (c) Copyright 1995-1997 Bernd Schmidt & Co-Autoren
Author:  Berndt Schmidt <crux@Pool.Informatik.RWTH-Aachen.DE>
Version:  0.6.8-8

vgb A virtual Gameboy. This is real.

This emulator adds the whole world of Nintendo’s Gameboy to your PC. Call this depending on the color depth of your X-Server with vgb (8bit), vgb16 (16bit) and vgb32 (32bit).

Contents: vgb

Copyright:  (c) Marat Fayzullin 1994,1995,1996
Author:  Marat Fayzullin <fms@freeflight.com>
Version:  0.7-8

vice Commodore 8-bit Emulator (C64/C128/VIC20/PET)

This is a collection of Emulators for 8bit Commodore computers. The packet includes a C64, C128, VIC20 and a PET emulator. You’ll need the ROMS of these machines to get the emulators work. See the manual for this.

Contents: vice

Copyright:

(c) 1993-1996 Jarkko Sonninen, Teemu Rantanen, Jouko Valta, Ettore Perazzoli, Andre Fachat

Author:

Jarkko Sonninen <sonninen@lut.fi>, Teemu Rantanen <tvr@cs.hut.fi>, Jouko Valta <jopi@stekt.oulu.fi>, Ettore Perazzoli <ettore@comm2000.it>, Andre Fachat <a.fachat@physik.tu-chemnitz.de>

Version:  0.11.0-8

wine Wine - The MS Windows emulator (ALPHA!)

The Windows emulator. Please adapt /etc/wine.conf to your needs. Additionally you will need an installed MS-Windows on your system.

Please have a look at /usr/doc/packages/wine/README.SuSE. There is more documentation available under this directory. Also consider using ’man wine’.

Please be aware that this software still is ALPHA and might damage system files!

After having adapted all necessary files (/etc/wine.conf) you can invoke wine by entering: ’wine program’ (e.g.)

Contents: wine-980104

Copyright:

(c) 1993-95 John Brezak, Erik Bos, Eric Youndale, Alexandre Julliard, David W. Metcalfe, Martin Ayotte, Martin von Loewis, Robert J. Amstadt, David Metcalfe, Johannes Ruscheinski, Miguel de Icaza, John Burton

Author:

Albrecht  Kleine  <kleine@ak.sax.de>,  Alex  Korobka

<alex@phm30.pharm.sunysb.edu>, Alexandre Julliard <julliard@di.epfl.ch>, Anand Kumria <akumria@ozemail.com.au>, Andreas Kirschbaum <ank@rbg.informatik.th-darmstadt.de>, Andrew Bulhak , Andrew Taylor <andrew@riscan.com>, Bernd Schmidt <crux@pool.informatik.rwth-aachen.de>, Bob Amstadt <bob@amscons.com>, Cameron Heide <heide@ee.ualberta.ca>, Carl Williams <crw@harris.mlb.semi.harris.com>, Charles M. Hannum <mycroft@mit.edu>, Constantine Sapuntzakis <csapuntz@mit.edu>, Dag Asheim <dash@ifi.uio.no>, Daniel Schepler <daniel@shep13.wustl.edu>, David Metcalfe <david@prism.demon.co.uk>, Duncan C Thomson <duncan@spd.eee.strath.ac.uk>, Eric Warnke <ew2193@csc.albany.edu>, Eric Youngdale <eric@aib.com>, Erik Bos <erik@hacktic.nl>, Erik Svendsen <z3esv@kmd-ac.dk>, Fons Botman <botman@rabo.nl>, Frans van Dorsselaer <dorssel@rulhm1.leidenuniv.nl>, Goran Thyni , Graham Menhennitt <gfm@werple.mira.net.au>, Greg Kreider <kreider@natlab.research.philips.com>, Gregory Trubetskoy <grisha@mira.com>, Hans de Graaff <J.J.deGraaff@twi.tudelft.nl>, Huw D. M. Davies <h.davies1@physics.oxford.ac.uk>, Itai Nahshon <nahshon@vnet.ibm.com>, James Youngman <mbcstjy@afs.man.ac.uk>, Jan Willamowius <jan@janhh.sh.sub.de>, Jeffrey Hsu <hsu@freefall.cdrom.com>, Jim Peterson <jspeter@birch.ee.vt.edu>, Jimmy Tirtawangsa <j0t2527@tam2000.tamu.edu>, Jochen Hoenicke <Jochen.Hoenicke@arbi.Informatik.Uni-Oldenburg.de>, Jochen Karrer <cip307@wpax01.Physik.Uni-Wuerzburg.DE>, Johannes Ruscheinski , John Brezak <brezak@apollo.hp.com>, John Harvey <john@division.co.uk>, John Richardson <jrichard@cs.uml.edu>, Jon Konrath , Jon Tombs <jon@gte.esi.us.es>, Juergen Marquardt <marqu@lunar.advantest.de>, Jukka Iivonen <iivonen@cc.helsinki.fi>, Karl Guenter Wuensch <hz225wu@unidui.uni-duisburg.de>, Keith Reynolds <keithr@sco.COM>, Kenneth MacDonald <K.MacDonald@ed.ac.uk>, Linus Torvalds <linus.torvalds@helsinki.fi>, Manfred Weichel <Manfred.Weichel@mch.sni.de>, Marcus Meissner <msmeissn@faui01.informatik.uni-erlangen.de>, Martin Ayotte <martin2@trgcorp.solucorp.qc.ca>, Martin von Loewis <loewis@informatik.hu-berlin.de>, Michael Patra <micky@marie.physik.tuberlin.de>, Michael Veksler <s1678223@t2.technion.ac.il>, Miguel de Icaza <miguel@roxanne.nuclecu.unam.mx>, Morten Welinder <terra@diku.dk>, Niels de Carpentier <niels@cindy.et.tudelft.nl>, Olaf Flebbe <o.flebbe@science-com-puting.uni-tuebingen.de>, Onno Hovers <onno@stack.urc.tue.nl>, Paul Falstad <pf@zoof.cts.com>, Peter Galbavy , Peter MacDonald , Philippe De Muyter <phdm@info.ucl.ac.be>, Ramon Garcia <ramon@ie3.clubs.etsit.upm.es>, Rick Sladkey <jrs@world.std.com>, Robert Pouliot <krynos@qbc.clic.net>, Roman Dolejsi <roman@sorry.vse.cz>, Ross Biro <biro@yggdrasil.com>, Scott A. Laird , Steffen Moeller <smoe0024@rz.uni-hildesheim.de>, Sven Verdoolaege <skimo@dns.ufsia.ac.be>, Thomas Sandford <t.d.g.sandford@prdsgrn.demon.co.uk>, Tristan Tarrant <tst@sthinc.demon.co.uk>, Ulrich Schmid <uschmid@mail.hh.provi.de>, Uwe Bonnes <bon@elektron.ikp.physik.thdarmstadt.de>, William Magro <wmagro@tc.cornell.edu>, William Smith <wos@dcs.warwick.ac.uk>, Yngvi Sigurjonsson <yngvi@hafro.is>

Version:  980104-3

xzx Sinclair ZX-Spectrum emulator (with ROM)

This program emulates a ZX-Spectrum with 48K, 128K and Interface1. Files for this computer can be loaded in sna or z80 format. Alternatively you may type your own programs in BASIC - of course. :-) Amstrad has kindly given permission to include the original ROM images. Amstrad further has the copyright on the ROM images. Contents: xzx-2.1.3

Copyright:  (c) 1993-97 Des Herriott, Erik Kunze
Author:  Des Herriott <dnh@mfltd.co.uk>, Erik Kunze <Erik.Kunze@fantasy.muc.de>
Version:  2.1.3-7

z81 Sinclair ZX81 (Timex TS1000) emulator (with ROM)

Z81 is a ZX81-Emulator. The ZX81 is also known as Timex TS1000 and a small predecessor of the ZX Spectrum. It used to have 1KB RAM of memory. Amstrad kindly gave permission to include the ROM-image containing the original OS. Amstrad retain it’s copyright. Contents: z81

Copyright:  (c) 1995-1997 Russell Marks, Amstrad (ROM)
Author:  Russel Marks <>
Version:  0.2-8

0.1.7 FUN: Games and more

3d chess 3 dimensional chess

Rules are to be found under /usr/doc/packages/3d-chess/3Dc-rules.html. Contents: 3Dc-0.8.1

Copyright:  (c) 1995 Paul Hicks
Author:  Paul Hicks <paulh@euristix.ie>
Version:  0.8.1-7

antipoli X11 Game

This game can be played with 2 - 10 players in a network. You will find the game and all available documentation under /usr/games/AntipoliX-2.1. Contents: AntipoliX-2.1

Copyright:  (c) 1994 Bernard Bross
Author:  Bernard Bross <bross@essi.fr>
Version:  2.1-8

asclock AfterStep’s clock

asclock was originally the AfterStep window manager’s clock. Documentation in: /usr/doc/packages/asclock man asclock Contents: asclock-1.0

Copyright:  (c) 1996, Beat Christen
Author:  Beat Christen <bchriste@iiic.ethz.ch>
Version:  1.0-8

astrolog An astrology program

This is an astrology program that, among many other things, will produce PostScript graphics and display charts in X. You can display the program parameters by typing astrolog -H. Documentation: /usr/doc/astrolog-4.1.README Contents: astrolog-4.10

Copyright:  (c) 1994 Walter D. Pullen
Author:  Walter D. Pullen <Astara@msn.com>
Version:  4.10-7

batalion A mesa-based 3D game

Battalion is a good example of how 3D Games can be implemented using Mesa and how existing OpenGL applications may be ported to Linux. Thanks to Mesa the work involved is quite low. Contents: batalion

Copyright:  (c) 1994-1996 by Andrew Johnson
Author:  Andrew Johnson <ajohnson@eecs.uic.edu>
Version:  97.6.29-9

battball Not so realistic military soccer

Play against the computer or other players (via network) soccer with helicopters or tanks. Contents: battball-2.0

Copyright:  (c) 1996, 1997 Philip A. Hardin
Author:  Philip A. Hardin <pahardin@cs.utexas.edu>
Version:  2.0-1

bsdgames Several text mode games

This package copies these games into /usr/games: arithmetic, atc, backgammon, battlestar, bcd, bog, caesar, canfield, cfscores, cribbage, factor, fish, fortune, hangman, hunt, mille, monop, morse, number, paranoia, pom, ppt, primes, rain, robots, sail, snake, snscore, teachgammon, trek, wargames, worm, worms and wump.

FUN: Games and more

Contents: bsd-games-1.5

Copyright:  (c) 1989/1990 The Regents of the University of California.
Author:  Curt Olson <curt@sledge.mn.org>, Andy Tefft <teffta@erie.ge.com>
Version:  1.5-5

craft Craft - the vicious vikings

Craft is a strategy/war game similar to Warcraft and Dune II. Each player starts off with two workers and a knight, and tries to build a kingdom with which he can kill off all the other players. The workers can build buildings such as mills (which make archers), docks (for ships), and markets (to trade gold and lumber). The game takes a little while to figure out, but it’s worth it.

Contents: craft-3.05

Copyright:  (c) Uwe Beyer
Author:  Uwe Beyer <uwe.beyer@gmd.de>
Version:  3.05-8

crosfire Graphical multi-player arcade game for X11

Adventure game, like nethack plus graphical interface. Can be played over a network, includes an editor. Documentation: man crossfire

Contents: crossfire-0.93.5.client crossfire-0.93.5.maps crossfire-0.93.5

Copyright:  (c) 1992 Frank Tore Johansen, 1994 Mark Wedel
Author:  Frank Tore Johansen <frankj@ifi.uio.no>
Version:  0.93.5-7

daliclck Dali-Clock

An interesting clock, where the digits change smoothly into the next.

Contents: xdaliclock-2.05

Copyright:  (c) (C) 1991, 1992, 1993, 1994, 1995 by Jamie Zawinski.
Author:  Jamie Zawinski <jwz@netscape.com>
Version:  2.05-8

emiclock One more Clock

Once again a clock..;). But this one is special. Nicely designed clock with little girls on the clock-face. You can choose one out of 20 sweetly designed dresses for the girl. Furthermore it has sound support (requires a new compilation) and an alarm clock included (who wants to miss his/her personal leisure time). This is the nicest ;-) Take a look at it. Docu: man emiclock

Contents: EmiClock-1.0.2

Copyright:  (c) 1994 Masayuki Koba
Author:  Masayuki Koba <masa-k@sag.hitachi-sk.co.jp>
Version:  1.0.2-8

empire Well-known strategy game (text mode)

For more information, please read the man page or print /usr/doc/games/empire-5.0.dvi .

Contents: empire-5.0

Copyright:  (c) 1987, 1988 Chuck Simmons
Author:  Chuck Simmons
Version:  5.0-8

fishtank An aquarium in the root window

A nice little aquarium with funny fish. This program only runs when the X server is running in 256 color mode.

Contents: xfishtank-2.1

Copyright:  (c) 1988-89 Patrick J. Naughton, Eric Bina
Author:  Patrick J. Naughton <naughton@wind.sun.com>, Eric Bina <ebina@ncsa.uiuc.edu>
Version:  2.1-9

flying Billiard Game

You can play a lot of billiard-like games: Snooker, Pool-Billiard, Karambolage, Carrom, Air-Hockey and Curling.

Documentation: man flying

Contents: flying-6.20

Copyright:  (c) 1994 Helmut Hoenig
Author:  Helmut Hoenig <Helmut.Hoenig@hub.de>
Version:  6.20-9

fortune talk big...

Fortune displays a random text string from a set of files in a certain format. This occurs each time you start a login shell.

Contents: fortune

Copyright:  (c) 1986, 1993 The Regents of the University of California.
Author:  Ken Arnold
Version:  1.0-8

freeciv Free Civilization Clone

A clone of the well-known DOS-game Civilization by Microprose. Every player is the leader of an imaginary nation. The aim of the game can be to create a prospering civilization with commerce and knowledge exchange or (more often) the extinction of all other civilizations. To start a new game : First start the server civserver, then start the client civ. Have fun !

Copyright:  (c) 1997, Allan Ove Kjeldbjerg, Claus Leth Gregersen, Peter Joachim Unold

Author:

Allan Ove Kjeldbjerg <allan@daimi.aau.dk>, Claus Leth Gregersen <leth@daimi.aau.dk>, Peter Joachim Unold <pjunold@daimi.aau.dk>

Version:  1.5.1-0

frisk X Window System Risk

A version of the Parker Brother’s classic as an X11 application. Can also be played over a network. The package includes: risk (stand-alone version), friskserver (server) & xfrisk (client). Documentation: /usr/doc/packages/frisk, as well as extensive online help.

Contents: Frisk-0.99b4

Copyright:  (c) 1993, 1994 Elan Feingold
Author:  Elan Feingold <feingold@zko.dec.com>
Version:  0.99b4-9

glchess 3D frontend for gnuchess

glchess is a fine 3D mesa frontend for gnuchess

Contents: glchess-linux-bin

Copyright:  (c) GPL
Author:  Tomi.Sarvela@iki.fi
Version:  0.10b-1

gnuchess The GNU chess program

A worthy chess opponent. Runs in text mode. You can find an X11 interface in the xboard package.

Contents: gnuchess-4.0.pl75

Copyright:  (c) 1988-90 John Stanback, 1992 Free Software Foundation

Author:

Hans Eric Sandstrom <hes@log-sv.se>, Stuart Cracraft <cracraft@ricechex.ai.mit.edu>, Mike McGann <mwm@hslrswi.hasler.ascom.ch>, Philippe Schnoebelen <phs@lifia.imag.fr>, Tim Mann <mann@src.dec.com>

Version:  4.0.pl75-8

gnugo Chinese game ”Go”

/usr/doc/gnugo/Documentation

Contents: gnugo-1.2

Author:  Man L. Li <manli@cs.uh.edu>
Version:  1.2-8

gshogi A strategic game

Copyright:  (c) 1988-90 John Stanback, 1992 Free Software Foundation, Inc., 1993-95 Matthias Mutz
Author:  John Stanback, Matthias Mutz <mutz@fmi.uni-passau.de>
Version:  1.2p03-8

hextris Xhextris and CXhextris, hexagonal versions of Xtetris

This package provides two games under X Window System: Xhextris and the colored version CXhextris. Both are fast games involving dropping hexagonal blocks similar to xtetris. Both version are modified. They use the same font (but different letters) and they use the cursor keys. Contents: hextris

Copyright:  (c) Copyright 1990 David Markley, 1993 Khun Yee Fung (color version)
Version:  1.0-9

imaze X11/Xview - Multi-Player - 3D Game

a 3D TCP/IP-network-game under X11/XView. There are clients (imaze, ninja) working together with one server (imazesrv). To start the game, you’ve to fire up the server first: imazesrv /usr/games/imaze/labs/demolab.lab & Then you can start the clients with: ninja <Server> & imaze & Contents: imaze1.3

Copyright:  (c) 1993-1996 by Hans-Ulrich Kiel and Joerg Czeranski

Author:

Hans-Ulrich  Kiel  <Kiel@rz.tu-clausthal.de>,  Joerg  Czeranski

<Czeranski@informatik.tu-clausthal.de>

Version:  1.3-9

lincity lincity

Lincity is a well designed and easy to play SimCity clone. Besides the game itself it provides an editor for easily creating new levels and obstacles. Contents: lincity103

Copyright:  (c) I J Peters 1995,1997.
Author:  I J Peters <ijp@floot.demon.co.uk>
Version:  1.03-9

maelstr A high-quality Asteroids-Clone

You are on space patrol. Blast asteroids, fight UFOs, get extra-weapons. Beware of supernovae and black holes. Contents: maelstr

Copyright:  (c) 1995 by Ambrosia Software, Inc.
Author:  Sam Lantinga <slouken@cs.ucdavis.edu>
Version:  1.4.3-9

manix Pacman Clone

X11 pacman clone. Contents: manix-19feb94

Copyright:  (c) 1994 Fernando Pereira
Author:  Fernando Pereira <fjp@minerva.inesc.pt>
Version:  94.2.19-4

meltflip Two nice Demos

Try it! Your display will turn upside down or it just melts away. Contents: xmelt-xflip-07feb91

Copyright:  (c) 1990 David Lemke
Author:  David Lemke <lemke@ncd.com>
Version:  07feb91-8

mirrma Arcade style game for color X11 with stereo sounds.

A game like ”Mindbender” (Amiga) or ”Deflektor” (Atari ST). This is a port of a program I wrote for the Amiga in 1989. Comes with full source code. Contents: mirrormagic-1.3

Copyright:  (c) 1995 by Holger Schemel.
Author:  Holger Schemel <aeglos@valinor.owl.de>
Version:  1.3-8

moontool The Moon in focus

One of the most important programs existing :-) Using this program, you can display all important information about the moon constantly. At last... Contents: xmoontool-22sep94

Copyright:  (c) 1993 John Walker
Author:  John Walker <kelvin@Autodesk.com>
Version:  94.9.22-7

net3d A networked, multi-player, 3d combat game

Copyright:  (c) Copyright (C) 1994 by Jamie Cameron
Author:  Jamie Cameron <jcam2@ccds.cc.monash.edu.au>
Version:  0.08-8

nethack A character based RPG

This RPG is somewhat cryptic with its character based output. But a true fan knows and appreciates its complexity and possibilities. Contents: nethack

Author:  Stephen L. Ericksen <stevee@cc.usu.edu>
Version:  3.2.2-9

oneko Oneko, a cat catches your mouse

A nice program which changes your cursor into a cat playing with your mouse cursor. The manual page will show more possibilities to change your cursor. Contents: oneko-1.1b

Copyright:  (c) 1990 Masayuki Koba

Author:

Tatsuya Kato <kato@ntts.co.jp>, John Lerchey <jl4l+@andrew.cmu.edu>, Eric Anderson <eanders+@cmu.edu>, Rick Niles <niles@axp745.gsfc.nasa.gov>

Version:  1.1b-1

oonsoo A Solitaire Card Game

Oonsoo is a solitaire card game. The goal of the game is to arrange twelve suits of cards, in order, onto twelve playing decks. When the game starts, one card face down and one card face up is dealt to each of the playing decks. Contents: oonsoo-1.0 oonsoo-1.1

Copyright:  (c) 1994 Bradford W. Mott
Author:  Bradford W. Mott <bwmott@eos.ncsu.edu>
Version:  1.1-5

pacman Pacman

Another pacman-game. Contents: pacman

Copyright:  (c) Roar Thronaes
Author:  Roar Thronaes
Version:  1.195.07.18-9

pipeman A game of skill

A well known game of skill. Docu: man xpipeman. Contents: xpipeman-18jul91

Copyright:  (c) 1991 Nigel Paver
Author:  Nigel Paver <pavern@uk.ac.man.cs>
Version:  91.7.18-5

playone Game - similar to monopoly

A game -similar to monopoly. Contents: play1-0.2

Copyright:  (c) 1994 TortiSoft, M”unchen
Author:  Torsten F”ahrmann <tfaehr@aol.com>
Version:  0.2-9

puzzle A puzzle

a puzzle..;)

Contents: puzzle-21aug91

Copyright:  (c) 1988 Don Bennett
Author:  Don Bennett
Version:  91.8.21-4

rockdiam Arcade style game for color X11 with stereo sounds.

A game like ”Boulderdash” (C=64) or ”Emerald Mine” (AMIGA). Sound server takes advantage of some features of VoxWare(tm), but works on plain /dev/audio, too. Comes with full source code. Contents: rocks n diamonds-0.9b

Copyright:  (c) 1995 by Holger Schemel.
Author:  Holger Schemel <aeglos@valinor.owl.de>
Version:  0.9-8

sastroid Sasteroids

Sasteroids is a game similar to asteroids. The aim is (what else..;) to shoot every thing that appears on your monitor (asteroids, spaceships). You should avoid crashing into those asteroids. You collect points for every destroyed asteroid or spaceship. This started out as a direct port of xasteroids to run under svgalib (a wonderful super-vga graphics library by Harm Hanemaayer) Sasteroids was derived from Phil Goetz’s Xasteroids by Brad Pitzel. Contents: sast1.5-snd+gfx sast1.5-src

Copyright:  (c) 1994 Brad Pitzel
Author:  Brad Pitzel <pitzel@cs.sfu.ca>
Version:  1.5-8

space space

shooting game for console Contents: spaceboom-0.95.src

Copyright:  (c) FREEWARE
Author:  Joe Lee <hosxe@dds.nl>
Version:  0.95-0

thrust A gravity skill/action game

Nice action game. Thrust has some elements borrowed from xpilot. Contents: thrust-0.83c

Copyright:  (c) 1996,96 Peter Ekberg
Author:  Peter Ekberg <peda@lysator.liu.se>
Version:  0.83c-8

vgacard A small collection of card games (libvga)

The card games vga klondike, vga solitaire, vga ohhell & vga spider. X11 is not required to run these games, as they use libvga. However, especially S3 card tend to have problems using this library, so do not be disappointed if the programs do not run :-) Contents: vga cardgames-1.3.1

Copyright:  (c) 1991-94 Evan Harris
Author:  Evan Harris <evan@cs.mu.oz.au>
Version:  1.3.1-5

x3d 3D animation

3D animation. The displayed objects can be turned using the mouse. Some example object files can be found in /usr/X11R6/lib/X11/x3d/ . The package includes the program hull, which can be used to create object files. See the file cube.poly for an example of the input format of the compiler. Contents: x3d-2.2

Copyright:  (c) 1992 Mark Spychalla
Author:  Mark Spychalla <spy@castlab.engr.wisc.edu>
Version:  2.2-6

xabacus X-based abacus

There are books on how to use an abacus, but basically all it does is add and subtract, the rest you have to do in your head. Essentially, this is a proof by induction that a computer is more powerful than an abacus, since a computer program can contain an abacus. (But then again, you can simulate a computer within a computer, so what does that show). Actually, with a real abacus, one can move more than one row at a time with 10 fingers. But on the other hand, a real abacus does not have the current sum displayed as an integer.

Contents: xabacus-5.0

Copyright:  (c) 1994 David Albert Bagley
Author:  David Albert Bagley <bagleyd@bigfoot.com>
Version:  5.0-1

xancur animate cursor for x-window

animate cursor for x-window

Contents: xancur-1.4

Copyright:  (c) Christopher Alex.North-Keys Unix Systems Analyst Group Talisman
Author:  Christopher Alex <erlkonig@gnu.ai.mit.edu>
Version:  1.4-2

xaos Powerful fractal generator

XaoS is a fast portable real-time interactive fractal zoomer. It displays the Mandelbrot set (among other escape time fractals) and allows you zoom smoothly into the fractal. Various coloring modes are provided for both the points inside and outside the selected set. In addition, switching between Julia and Mandelbrot fractal types and displaying planes is provided.

Contents: XaoS-2.2

Copyright:  (c) 1997, Jan Hubicka
Author:  Jan Hubicka <hubicka@paru.cas.cz>, Thomas Marsh <tmarsh@austin.ibm.com>
Version:  2.2-8

xball Bouncing balls

This program simulates bouncing balls in a window. You create the balls in a variety of ways, and can set the gravity, elasticity, whether balls collide or not, etc.

Contents: xball-3.0

Copyright:  (c) 1993 David Nedde
Author:  David Nedde <daven@ivy.wpi.edu>
Version:  3.0-11

xbill Who is afraid of little Bill?

A small but mean hacker named Bill wants to infect some Workstations with his program. It is on you to save the world and to slap him for trying this.

Contents: xbill

Copyright:  (c) 1994-1996 Brian Wellington and Matias Duarte
Author:  Brian Wellington <bwelling@eng.umd.edu>, Matias Duarte <matias@eng.umd.edu>
Version:  2.0-9

xbl 3D Tetris

3D Tetris for X-Window

Copyright:  (c) GPL
Author:  Thierry EXCOFFIE <Thierry.EXCOFFIER@ligia.univ-lyon1.fr>
Version:  1.0h-0

xblast A multi-player arcade game for X11

XBlast is a multi-player arcade game for X11R5/R6. The game can be played with at least two players and up to four players. It was inspired by the video/computer game Bomberman(Dynablaster), which was to my knowledge first programmed for NEC’s PC Engine/Turbo Grafx. Other (commercial) versions of the original game exist for IBM-PC, Amiga, NES, GameBoy and Super NES.

Contents: XBlast-2.0

Copyright:  (c) 1993/94 Oliver Vogel
Author:  Oliver Vogel <vogel@ikp.uni-koeln.de>
Version:  2.0-8

xboard X11 Interface for GNU Chess

Copyright:

(c) 1991 by Digital Equipment Corporation, Maynard, Massachusetts. 1991-1995 Free Software Foundation, Inc.

Author:

Tim Mann <mann@src.dec.com>, Chris Sears, Dan Sears, Elmar Bartel <bartel@@informatik.tu-muenchen.de>

Version:  3.4.pl0-9

xboing X11 breakout game

This version was compiled with rplay support. If you want to get the sounds, you should install the rplay daemon.

Contents: xboing2.3

Copyright:  (c) 1993, 1994, Justin C. Kibell
Author:  Justin C. Kibell, <jck@citri.edu.au>
Version:  2.3-9

xbomb Another minesweeper clone

Copyright:  (c) 1994 1995 AMB Software
Author:  Andrew M. Bishop <amb@gedanken.demon.co.uk>
Version:  2.0-0

xbombs Minesweeper

Game of minesweeper. 30 on 16 fields :-)

Contents: xmines

Copyright:  (c) 1993 Andrew Zimmerman
Author:  Andrew Zimmerman <zimmer@calvin.stanford.edu>
Version:  03dec93-8

xdemine Not just another version of the minesweeper game...

The object of the game is to find the location of the mines and to mark their location by placing a small flag. You have exactly as many flags as there are mines, no more, no less. The number of flags that remain to be placed is displayed in the top left corner and a timer is displayed in the top right corner.

Contents: xdemineur-2.1

Copyright:  (c) 1993 Marc Baudoin
Author:  Marc Baudoin <babafou@agm-ita.ensta.fr>
Version:  2.1-1

xearth Our planet earth as animated background

Our planet earth as a background picture. Light and shades are time dependent. One can adjust lots of parameters..;)

Contents: xearth-1.0

Copyright:  (c) 1994 Kirk Lauritz Johnson
Author:  Kirk Lauritz Johnson <tuna@cag.lcs.mit.edu>
Version:  1.0-8

xengine reciprocationg engine for X

Xengine displays a reciprocationg engine in a window and can be abused as a benchmark for X.

Contents: xengine.beta1

Copyright:  (c) Kazuhiko Shutoh, YAMAHA Corp.
Author:  Kazuhiko Shutoh <shutoh@isl.yamama.co.jp>
Version:  beta1-5

xfract Fractals under X11

Xfractint draws a wide variety of fractals (fast!). Xfractint is a port of the DOS program Fractint.

Contents: xfractint-3.02

Author:

Bert Tyler, Timothy Wegner, Mark Peterson , Pieter Branderhorst, Ken Shirriff <shirriff@eng.sun.com>

Version:  3.0-8

xgammon A backgammon program for Linux

xgammon contains an X11 interface, a simple file selector and saving dialog, a edit position menu option. a rollout option. a mail dump option for newsgroup posting of positions. and it is quite good; even for the experienced backgammon player.

Contents: xgammon-0.96 xgammon 0.98.orig

Copyright:  (c) 1994 Lambert Klasen & Detlef Steuer

Author:

Lambert Klasen <klasen@asterix.uni-muenster.de> , Detlef Steuer <steuer@amadeus.statistik.uni-dortmund.de>

Version:  0.98-5

xgas animated simulation of an ideal gas

xgas is a physical simulation of an ideal gas in a heated box. The temperature of each box can be independently controlled by a scrollbar. Contents: xgas-29aug91

Copyright:  (c) 1989 Massachusetts Institute of Technology
Author:  Larry Medwin, Dave Sternlicht
Version:  91.8.29-5

xjewel X11 game - similar to tetris

X11 game - similar to tetris. Has a flavour of its own :-) Doku: /usr/doc/X11-misc/xjewel-1.6.README Contents: xjewel-1.6

Copyright:  (c) 1990 Yoshihiro Satoh
Author:  Yoshihiro Satoh
Version:  1.6-8

xlife Conway’s game of life

Nice version of ”game of life”. Lots of interesting startpositions you can find under /usr/X11R6/lib/xlife/patterns. Documentation: /usr/doc/packages/xlife Contents: xlife

Copyright:  (c) 1989 Jon Bennett

Author:

Jon Bennett <jcrb@cs.cmu.edu>, Daniel E Lovinger <dl2n@andrew.cmu.edu>, Charles Humphrey Silvers <Charles.Silvers@cs.cmu.edu>, Eric <esr@snark.thyrsus.com>

Version:  5.0-8

xlyap displays Lyapunov exponents

Copyright:  (c) 1991 Ronald Joe Record
Author:  Ronald Joe Record <rr@sco.com>
Version:  2.3-8

xmahjong mahjong

Mahjongg is a challenging chinese game similar to domino. It is usually played by four players. Xmahjongg is the solitair version designed for the X-Window System. More can be found in the appropriate manual page. Contents: xmahjong

Copyright:  (c) Jeff S. Young
Author:  Jeff S. Young <jsy@cray.com>
Version:  97.9.21-4

xmemory xmemory - a memory game

This memory game is playable by many users on different X11 displays simultanously. Contents: xmemory-2.0

Copyright:  (c) 1993+1995, Helmut Hoenig, Heiligenhaus
Author:  Helmut Hoenig <Helmut.Hoenig@hub.de>
Version:  2.0-9

xmine Yet another minesweeper...;)

Version:  1.0.3-8

xmines Minesweeper clone

Copyright:  (c) 1993 Ashley Roll
Author:  Ashley Roll <ash@dstc.edu.au>
Version:  3.0-8

xmountns fractal landscape in the background

Xmountains is a fractal terrain generator written by Stephen Booth. It can be configured to calculate new ’mountains’ in a certain period of time. Contents: xmountains-2.4

Copyright:  (c) 1994 by Stephen Booth, the University of Edinburgh
Author:  Stephen Booth <S.Booth@ed.ac.uk>
Version:  2.4-8

xmris A nice X11 game

In this game, you are walking around, collecting fruits. The package includes an editor for new screens. This means that in case you get hooked you can create more difficult screens yourself. Contents: xmris

Copyright:  (c) 1993, 1992 Nathan Sidwell
Author:  Nathan Sidwell <nathan@pact.srf.ac.uk>
Version:  4.04-7

xpat2 Patience card game

Patience (card game); nice graphics. Contents: xpat2-1.04

Copyright:  (c) 1994 Heiko Eissfeldt, Michael Bischoff

Author:

Heiko  Eissfeldt  <heiko@colossus.escape.de>,  Michael  Bischoff

<mbi@mo.math.nat.tu-bs.de>

Version:  1.04-8

xphoon Shows moon phases..;)

Program which lets you see the actual moon phase. You can accelerate time to see how the moon will look like in a couple of days..;) Contents: xphoon

Copyright:  (c) 1988, 1991 Jef Poskanzer, Craig Leres
Author:  Craig Leres <>, Jef Poskanzer <>
Version:  91.9.18-5

xpilot A multiplayer arcade game

A shoot’em up which can be played by several players in a network. Please note that you have to start the server (xpilots) before starting the client (xpilot, plus the name of the computer which runs the server). Several maps can be found in /usr/local/games/lib/xpilot/maps/ Doku: /usr/doc/packages/xpilot Contents: xpilot-3.6.2

Copyright:  (c) 1991-95 Bjoern Stabell, Ken Ronny Schouten, Bert Gijsbers

Author:

Bjoern Stabell <bjoerns@staff.cs.uit.no> , Ken Ronny Schouten <kenrsc@stud.cs.uit.no>, Bert Gijsbers <bert@mc.bio.uva.nl>

Version:  3.6.2-8

xpinguin The ’Logo’ of Linux 2.0

Based on xteddy - really cute ;-) Contents: xpinguin

Copyright:  (c) 1994 Stefan Gustavson

Author:

Stefan Gustavson <stefang@isy.liu.se>, Jens P”onisch <J.Poenisch@wirtschaft.tuchemnitz.de>

Version:  1.0.2-8

xpuzzles Some game of skill under X11

Copyright:  (c) 1994 David Albert Bagley
Author:  David Albert Bagley <bagleyd@bigfoot.com>
Version:  5.0.1-8

xroach Some cockroaches on your root window

Xroach displays disgusting cockroaches on your root window. These creapy crawlies scamper around until they find a window to hide under. Whenever you move or iconify a window, the exposed orthopteras again scamper for cover. See manpage (man xroach). Contents: xroach

Copyright:  (c) 1991 by J.T. Anderson
Author:  J.T. Anderson <jta@locus.com>
Version:  12.6.97-8

xrobots A game of skill

Try to escape from a bunch of robots. Use your skill! Contents: xrobots-17nov89

Copyright:  (c) 1989 Brian Warkentine
Author:  Brian Warkentine <brianw@Sun.COM>
Version:  89.11.17-4

xshogi X11-interface for GNU Shogi

Copyright:  (c) 1992 Free Software Foundation, Inc., 1993-95 Matthias Mutz
Author:  Matthias Mutz <mutz@fmi.uni-passau.de>
Version:  1.2p03-9

xskat Skat

Skat is a popular game in Germany. Now with ”Ramsch” and more... WWW page: http://www.draeger.com/˜ gerhardt/xskat-latest.html Contents: xskat-2.2

Copyright:  (c) Frei
Author:  <Gunter Gerhardt> gerhardt@draeger.com
Version:  2.2-0

xsnow Happy X-Mas!

A really nice animation. Santa clause drives across the screen, pulled by his reindeer. It snows, and the snow accumulates on top of the windows :-) Documentation: man xsnow Contents: xsnow-1.40

Copyright:  (c) 1984 - 1994 Rick Jansen
Author:  Rick Jansen <rick@sara.nl>
Version:  1.40-8

xsok strategy game

Sokoban Contents: xsok-1.02

Author:  Michael <mbi@mo.math.nat.tu-bs.de>
Version:  1.02-8

xsol Solitair

The game Solitair. Contents: xsol-30nov92

Copyright:  (c) 1988 Sun Microsystems
Author:  Dave Lemke <lemke@sun.com>
Version:  92.11.30-5

xspringi a mass and spring simulation system

Xspringies is a simulator which allows you to interactively create and edit a system of masses and springs. The parameters of the masses and springs (such as mass elasticity and spring K) as well as those of the surrounding system (such as air viscosity and gravity) can be changed. Several configuration files can be found in /usr/X11R6/lib/X11/xspringies/ . Contents: xspringies-1.1

Copyright:  (c) 1991, 1992 Douglas M. DeCarlo
Author:  Douglas M. DeCarlo <dmd@gradient.cis.upenn.edu>
Version:  1.1-8

xtacy psychedelic animation program

This is a program similar to fractint. It may not contain as many mathematical functions, but it is command-line oriented and may be run as a root animation, in case your CPU load is too low :-) Contents: Xtacy

Copyright:  (c) 1994 Jer Johnson
Author:  Jer Johnson <mpython@gnu.ai.mit.edu>
Version:  1.13-7

xtartan xtartan

Xtartan offers you a wide variety of scottish tartans. They can be viewed or alternatively used as a background picture. One of the most exciting programs at all..;)) Contents: xtartan-2.1

Copyright:  (c) Jim McBeath Globetrotter Software, Saratoga, California
Author:  Jim McBeath <jimmc@hisoft.uucp>
Version:  2.1-8

GRA: All about graphics

xteddy A cuddly teddy bear for your X Windows desktop

Normally, xteddy just sits around doing nothing. After all, that’s what teddy bears are for. Look at him, talk to him, place heavy windows on top of him, zap him around until he becomes dizzy, do what you like; he will always be your true (albeit virtual) friend. You can move xteddy with the mouse by pointing at him and dragging him around. When clicked upon, he will pop up on top of all other windows. If you type ”q” on him, he will die (or, as I like to think of it, be tucked away in the file system until you need him next time). That’s it. But he’s cute. Contents: xteddy-1.0.3

Copyright:  (c) GPL
Author:  Stefan Gustavson <stefang@isy.liu.se>
Version:  1.0.3-7

xtetris X Tetris

Version:  2.6-7

xthing colorful images

This program creates psychadelic images. Sample files can be found in /usr/local/lib/xthing-sample . For more information see /usr/doc/X11-misc/xthing-15nov91.README . Contents: xthing-15nov91

Copyright:  (c) 1991 John L. Cwikla
Author:  John L. Cwikla <cwikla@uimrl7.mrl.uiuc.edu>
Version:  91.11.15-6

xtron A game of skill

Copyright:  (c) 1995, Rhett D. Jacobs
Author:  Rhett D. Jacobs <rhett@hotel.canberra.edu.au>
Version:  1.1a-8

xvier connect IV

xvier is a simple game where you and your opponent alternately throw stones into free columns. The stones pile up in the columns, and the goal is to get four stones in a row, in a column or diagonally. Contents: xvier-1.0

Author:  Norbert Jung <jung@dia.informatik.uni-stuttgart.de>
Version:  1.0-8

yahtzee Yahtzee in text mode

The well-known game of dice in text mode; up to 20 players (human or computer). Contents: colour-yahtzee yahtzee-1.0

Copyright:  (c) 1992 Orest Zborowski
Author:  Orest Zborowski
Version:  1.0-9

0.1.8 GRA: All about graphics

bttv BTTV - BT848 frame grabber driver

Driver for TV cards with a X11/Motifprogram for cards with the Bt848 chipset: Miro PCTV, Hauppauge Win/TV pci, STB PCI TV, Intel Smart Video III and the Diamond DTV2000. The driver may be allready stable, but it is a early version, so use it carefull, please! Have a lot fun with watching TV! ;-) You need a TV card for it. Read the README.SuSE and the original documentation in the directory /usr/packages/bttv-<VERSION>, please!

Copyright:  (c) (C) 1996,97 Ralph Metzler & Marcus Metzler
Author:  Ralph Metzler <rjkm@thp.uni-koeln.de>, Marcus Metzler <mocm@thp.uni-koeln.de>
Version:  0.4.3-2

cgmdraw CGM Draw - library for generating CGM files

CGM Draw is a freely available library for generating CGM files from a C program. CGM (Computer Graphics Metafile) is a vector graphics format that can be read by many popular packages. With CGM Draw your code can quickly draw images complete with lines, arcs, rectangles, polygons, and text. CGM Draw is ideal for creating CGM files on the fly when you have a rapidly changing data set (such as in response to database queries.)

Copyright:  (c) Freely Redistributable
Author:  G. Edward Johnson <lorax@nist.gov>
Version:  1.3-3

faxedit A simple editor for pbm-files

Faxedit is a tool for editing pbm files like they are usually output or expected by fax programs. Its purpose is the easy answering of faxes. Parts of them can be cut out, rotated or resized. Input of text and the printing of the picture is also possible.

Contents: faxedit-1.14

Copyright:  (c) 1998 S.u.S.E. GmbH
Author:  Kalle Dalheimer <kalle@kde.org>
Version:  1.14-0

films A few computer animation files

A small collection of FLI, GL, AVI, and MPEG files. Unfortunately, nobody knows where these films came from. In case you are one of the artists - please contact us; you will get the money back for the CD :-)

Contents: films

Version:  97.11.27-3

freetype TrueType Font Engine

Library for working with TrueType Fonts. Documentation at /usr/doc/packages/freetype

Contents: freetype-1.0

Copyright:  (c) 1996 David Turner / 1998 Robert Wilhelm / Werner Lemberg

Author:

David Turner <turner@enst.fr>, Robert Wilhelm <robert@physiol.med.tumuenchen.de>, Werner Lemberg <a7971428@unet.univie.ac.at>

Version:  1.0-0

gd an autoloadable interface module for libgd

This is a autoloadable interface module for libgd, a popular library for creating and manipulating GIF files. With this library you can create GIF images on the fly or modify existing files.

Contents: GD.pm

Author:  Lincoln Stein <lstein@w3.org>
Version:  1.15-3

gimp Extendable graphic manipulation program

GIMP is the abbreviation of General Image Manipulation Program. It is a tool for creating, manipulating and saving graphical images. TIFF, JPEG, PNG, GIF and XPM are supported.

Contents: gimp-0.99.18 gimp-data-extras-0.99a gimp-data-min-0.99b

Copyright:  (c) 1995-97 Spencer Kimball and Peter Mattis
Author:  Spencer Kimball <spencer@xcf.berkeley.edu>, Peter Mattis <petm@xcf.berkeley.edu>
Version:  0.99.18-4

gimpman manual for gimp

manual for the gimp program as compressed postscript file.

Copyright:  (c) 1998 Olof & Karin Kylander
Author:  Olof Kylander, Karin Kylander
Version:  0.9.01-4

gsview Ghostview

Ghostview offers you an X11 GUI for viewing PostScript files. This is an X11 interface to ghostscript.

Contents: ghostview-1.5

Copyright:  (c) 1992 Timothy O. Theisen
Author:  Tim Theisen <tim@cs.wisc.edu>
Version:  1.5-8

gtk library for creation of graphical user interfaces

This a library that is needed for running the GIMP or SANE.

Contents: gtk+-0.99.3

Copyright:  (c) 1995-1997 Peter Mattis, Spencer Kimball and Josh MacDonald

Author:

Peter Mattis (petm@xcf.berkeley.edu), Spencer Kimball (spencer@xcf.berkeley.edu), Josh MacDonald (jmacd@xcf.berkeley.edu)

Version:  0.99.3-2

gv GV -- a ghostview derivate

GV offers you an X11 GUI for viewing PostScript files. This is an X11 interface to ghostscript. Contents: gv-3.5.8

Copyright:  (c) 1992 Timothy O. Theisen; 1995-97 Johannes Plass

Author:

Tim Theisen <tim@cs.wisc.edu>, Johannes Plass <plass@dipmza.physik.unimainz.de>

Version:  3.5.8-6

html2ps HTML to PostScript converter

HTML to PostScript converter written in Perl. html2ps know about many HTML 4.0 features. From the Homepage (http://www.tdb.uu.se/˜ jan/html2ps.html):

* Many possibilities to control the appearance; this is mostly done using configuration files. * Support for processing multiple documents, also automatically by recursively following links. * A table of contents can be generated, either from the links in a document, or automatically from document headings. * Configurable page headers/footers, that for example can contain document title, URL, page number, current heading, and date. * Automatic hyphenation and text justification can be selected. Contents: html2ps-1.0b1

Copyright:  (c) Jan Karrman and Jean-Philippe Argaud, GNU
Author:  Jan Karrman <jan@tdb.uu.se>, Jean-Philippe Argaud <no@e.mail>
Version:  1.0b1-3

imagemag Image / film viewer

Image Magick allows you to view and edit images. The tools read and write images in MIFF format. Other formats can be converted into MIFF. Contents: ImageMagick-4.0.1

Copyright:  (c) 1996 E. I. du Pont de Nemours and Company
Author:  John Cristy <cristy@dupont.com>
Version:  4.0.1-2

imfilm Example film for ImageMagick

An example film with a rotating DNA for demonstration of ImageMagick. The pictures are in the directory /usr/ImageMagick/scenes.

Version:  4.0.1-2

imlib Shared Library for loading and rendering 3D images in X.

This is a general, all-purpose Image loading and rendering (Image data to screen data) library. It will only load PPM format images itself. If you wish to load other formats, you will need commands like ”giftopnm”, ”tgatoppm”, ”bmptoppm”, ”djpeg”, ”xpmtoppm”, ”ilbmtoppm”, etc. (contained in the S.u.S.E. Linux packages netpbm and jpeg). Via using these commands internally, imlib is capable of currently loading 25 different image formats. See the file /usr/include/X11/imlib.h for more details. This shared library is needed for the Enlightenment windowmanager. Contents: imlib DR-0.4

Copyright:  (c) 1997 Carsten Haitzler (Rasterman)
Version:  DR 0.4-6

impict Example pictures for ImageMagick

Some example pictures for demonstration of ImageMagick. The pictures are in the directory /usr/ImageMagick/images.

Version:  4.0.1-2

jpeg Independent JPEG Group’s JPEG Software.

Software to implement JPEG image compression and decompression. JPEG pronounced ”jay-peg”) is a standardized compression method for full-color and gray-scale images. JPEG is intended for compressing ”real-world” scenes (most of the time there are pictures that have been scanned-in with a scanner). Cartoons and other non-realistic images are not its strong suit. You should note that the output of JPEG is not necessarily the same as its input. If this is a factor for you, then don’t use it. With typical real-world scenes, JPEG can achieve high compression rates, without noticeable differences. If you can accept pictures of lower quality, JPEG can achieve amazingly high compression rates. There are some library functions available for reading and writing JPEG files. The ’cjpeg’ and ’djpeg’ applications use the library to make conversions between JPEG and other popular graphics file formats possible. The library is meant to be used within other applications. Cjpeg compresses the input file, or standard input if no filename was given, and produces a JPEG/JFIF to standard output. Currently supported input file formats include: PPM (PBMPLUS color format), PGM (PBMPLUS Grayscale format), BMP, GIF, Targa and RLE (Utah Raster Toolkit Format). (RLE is only supported if the RLE library is available.) Djpeg decompresses a JPEG file in one of the above mentioned formats.

Contents: jpeg-6

Copyright:  (c) 1991, 1992, 1993, 1994, Thomas G. Lane

Author:

Tom Lane <jpeg-info@uunet.uu.net>, Philip Gladstone <pjsg@ix.netcom.com>, Luis Ortiz <jpeg-info@uunet.uu.net>, Jim Boucher <jpeg-info@uunet.uu.net>, Lee Crocker <jpeg-info@uunet.uu.net>, George Phillips <phillips@cs.ubc.ca>, Davide Rossi <jpeg-info@uunet.uu.net>, Ge’ Weijers <jpeg-info@uunet.uu.net>

Version:  6a-10

libfbm Fuzzy Pixmap Manipulation Library

This is relese 1.0 of the Fuzzy Pixmap Manipulation (FBM) library. The sources are contained in the libgr source package

Copyright:  (c) Rob Hooft, Michael Mauldin, (see netpbm)
Author:  Rob Hooft <hooft@EMBL-Heidelberg.DE>, Michael Mauldin <mlm@cs.cmu.edu>
Version:  1.0.0-10

libgr libgr - documentation from graphics libraries

The shared graphics libraries with GIF, and JPEG support are contained in libfbm, libnetpb, libjpeg, librle and libz. These libraries are needed if you intend to run the XV image viewer. If you additionally need the tiff library, please install the package libtiff. The tools for converting graphics are contained in jpeg, netpbm and tiff. All sources may be found in the source package of libgr. See also /usr/doc/packages/libgr-*/README.SuSE

Contents: libgr-2.0.13 libjpeg-6a libpng-0.99d pnmtopng seejpeg-1.6.1 tiff-v3.4beta037

Copyright:  (c) Rob Hooft, Michael Mauldin, (see netpbm)
Author:  Rob Hooft <hooft@EMBL-Heidelberg.DE>, Michael Mauldin <mlm@cs.cmu.edu>
Version:  2.0.13-10

libgrout libgr - graphics libraries (a.out)

Shared graphics libraries with GIF, TIFF, JPEG support. These libraries are needed to run the XV image viewer when using the old a.out version.

Copyright:  (c) Rob Hooft, Michael Mauldin, (see netpbm)
Author:  Rob Hooft <hooft@EMBL-Heidelberg.DE>, Michael Mauldin <mlm@cs.cmu.edu>
Version:  1.3-1

libjpeg JPEG-Libraries

The libraries (static and dynamic) for the jpeg-graphics format. The sources are contained in the libgr source package

Copyright:  (c) Rob Hooft, Michael Mauldin, (see netpbm)
Author:  Rob Hooft <hooft@EMBL-Heidelberg.DE>, Michael Mauldin <mlm@cs.cmu.edu>
Version:  6.0.1-10

libmpeg Mpeg library

Library with functions for access to mpeg files.

Contents: mpeg lib-1.2.2

Copyright:  (c) 1992 The Regents of the University of California, 1994-95 Gregory P. Ward
Author:  Gregory P. Ward <greg@bic.mni.mcgill.ca>
Version:  1.2.2-5

libnetpb Libraries for the netpbm (NetPortableBitmap) gaphic formats

These are the libs for the netpbm graphic formats. Tools can be found in package netpbm. The sources are contained in the libgr source package

Copyright:  (c) Rob Hooft, Michael Mauldin, (see netpbm)
Author:  Rob Hooft <hooft@EMBL-Heidelberg.DE>, Michael Mauldin <mlm@cs.cmu.edu>
Version:  1.0.0-10

libpng Library for the PortableNetGraphics Format

This is the sixth (and hopefully last) beta release of libpng 1.0.

Copyright:  (c) Copyright (c) 1996, 1997 Andreas Dilger
Author:  Andreas Dilger <adilger@enel.ucalgary.ca>
Version:  1.0.99d-10

librle rle-graphics library

The dynamic and static versions of librle. The sources are contained in the libgr source package

Copyright:  (c) Rob Hooft, Michael Mauldin, (see netpbm)
Author:  Rob Hooft <hooft@EMBL-Heidelberg.DE>, Michael Mauldin <mlm@cs.cmu.edu>
Version:  1.0.0-10

libtiff Tiff Library (with jpeg and compression support)

To link a program with this libtiff, you will have to add -ljpeg and -lz to include the necessary libjpeg and libz in the linking process.

Copyright:  (c) 1988-1997 Sam Leffler, 1991-1997 Silicon Graphics, Inc.
Author:  Sam Leffler <sam@engr.sgi.com>
Version:  3.4.37-10

libz Data Compression Library

zlib 1.0.2 is a general purpose data compression library. All the code is reentrant (thread safe). The data format used by the zlib library is described by RFCs (Request for Comments) 1950 to 1952 in the files ftp://ds.internic.net/rfc/rfc1950.txt (zlib format), rfc1951.txt (deflate format) and rfc1952.txt (gzip format). These documents are also available in other formats from ftp://ftp.uu.net/graphics/png/documents/zlib/zdoc-index.html The sources are contained in the libgr source package

Contents: zlib-1.1.1

Copyright:  (c) Rob Hooft, Michael Mauldin, (see netpbm)
Author:  Rob Hooft <hooft@EMBL-Heidelberg.DE>, Michael Mauldin <mlm@cs.cmu.edu>
Version:  1.1.1-1

mlcreat !! HIER FEHLT DAS LABEL !!

Moonlight Creator is an OpenGL-based modeler and renderer (radiosity and raytracing). It is in early development stages, don’t expect too much!

Documentation is below /usr/doc/packages/mlcreat. In order to test the example scenes, please copy /usr/doc/packages/mlcreat/moonlight to your home directory.

Contents: ml-0.4.1 moonlight-db-0.3.1

Copyright:  (c) 1996-1998 By Stephane C.F. Rehel
Author:  Stephane C.F. Rehel <rehel@worldnet.fr>
Version:  0.4.1-1

mpegplay MPEG player

MPEG (Moving Pictures Experts Group) is a format for the compression of video and audio data. The compressed files can be played in real-time (hardware-decoder) or relatively slow (software solution) as a video sequence.

This MPEG player plays video sequences using a 8, 16, or 32 bit display under X11.

Contents: mpeg play-2.3-src

Copyright:  (c) 1992-93 University of California

Author:

Ketan Patel <tel@cs.berkeley.edu>, Brian Smith <bsmith@cs.berkeley.edu>, Henry Chi-To Ma <cma@cs.berkeley.edu>, Kim Man Liu <kliu@cs.berkeley.edu>

Version:  2.3-8

muscan Muscan, a scanner tool for Mustek scanners.

Muscan is a scanner program for Mustec scanners. Since there is a new project under development, muscan is not supported any more. Muscan contains Tkmuscan, a nice Tk tool which incorporates most of the features of a professional scanner tool. A list of supported scanners is to be found under: /usr/doc/packages/muscan Contents: muscan

Copyright:  (c) Copyright (C) 1995 Torsten Eichner
Author:  Thorsten Eichner <eichner@rhrk.uni-kl.de>
Version:  2.0.6-8

netpbm A powerful graphics conversion package.

The latest version of the Portable Bitmap Plus Utilities. This large package which is by now available for all hardware platforms provides tools for graphics conversion. Using these tools, images can be converted from virtually any format into any other format. A few of the supported formats include: GIF, PC-Paintbrush, IFF ILBM, Gould Scanner file, MTV ray tracer, Atari Degas .pi1 and .pi3, Macintosh PICT, HP Paintjet file, QRT raytracer, AUTOCAD slide, Atari Spectrum (compressed and uncompressed), Andrew Toolkit raster object, and many more. On top of that, man pages are included for all tools. The sources for this package are in the source-package of libgr. Contents: netpbm-1mar1994

Copyright:

(c) 1988-94 Alberto Accomazzi, Angus Duggan, Arthur David Olson, Bill Davidson, Bill Janssen, Bo Thide, Brian Moffet, Christos Zoulas, Dan Stromberg, David Beckemeyer, David Koblas, David K. Drum, David W. Sanderson, Doug Crabill, Douwe van der Schaaf, Frank Neumann, George Phillips, Ingo Wilken, James Darrell McCauley, Jamie Zawinski, Jef Poskanzer, John Tiller, Ken Yap, Mark Shand, Mark W. Snitily, Michael, Davidson, Michael Haberler, Mike Parker, Oliver Trepte, Paul Haeberli, Rainer Klute, ick Vinci, berg, Sam Leffler, Stephen Paul Lesniewski, Texas Agricultural Experiment Station, Tim Cook, Tony Hansen, Wilson Bent

Author:

Alberto  Accomazzi  <alberto@cfa.harvard.edu>,  Andre  Beck

<Andre Beck@IRS.Inf.TU-Dresden.de>, Angus Duggan <ajcd@dcs.ed.ac.uk>, Arnaud Le Hors <lehors@sophia.inria.fr>, Arthur David Olson <OLSONA@dc37a.nci.nih.gov>, Burkhard Neidecker-Lutz <neideck@nestvx.enet.dec.com>, Daniel Briggs <dbriggs@nrao.edu>, David Beckemeyer <bdt!david>, David Koblas <koblas@netcom.com>, Diomidis D. Spinellis <dds@cc.ic.ac.uk>, George Phillips <phillips@cs.ubc.ca>, Graeme W. Gill <graeme@labtam.oz.au>, Hadmut Danisch <danisch@ira.uka.de>, Ingo Wilken <Ingo.Wilken@informatik.uni-oldenburg.de>, James Darrell McCauley <mccauley@mcs.com>, John Tiller <tiller@galois.msfc.nasa.gov>, John Walker <kelvin@Autodesk.com>, Kanthan Pillay <svpillay@Princeton.EDU>, Ken Yap <ken@cs.rocester.edu>, Lyle Rains <lrains@netcom.com>, Marcel Wijkstra <wijkstra@fwi.uva.nl>, Marc Boucher <marc@PostImage.COM>, Oliver Trepte <oliver.trepte@dd.sigma.se>, Patrick J. Naughton <naughton@wind.sun.com>, Paul Haeberli <paul@manray.sgi.com>, Paul Breslaw <paul@mecazh.uu.ch>, S.Petra Zeidler <spz@specklec.mpifr-bonn.mpg.de>, Steve Belczyk <seb3@gte.com>, Warren Toomey <wkt@csadfa.cs.adfa.oz.au>, Wilson H. Bent. Jr. <whb@evtech.com>, Wolfgang Stuerzlinger <wrzl@gup.uni-linz.ac.at>

Version:  01mar94.p1-10

photocd PhotoCD image converter

This Programm includes an X11-Interface and programs for converting Photo CD images into TARGA or BITMAP (Windows 3.1) format. Documentation can be found in /usr/doc/misc/photocd.doc. Additionally included is the program ”hpcdtoppm” that converts Photo CD images int the PPM format. Contents: hpcdtoppm-0.6 photocd-1.12 xpcd-2.04

Copyright:  (c) 1994 Gerd Knorr
Author:  Gerd Knorr <kraxel@cs.tu-berlin.de>, Hadmut Danisch <danisch@ira.uka.de>
Version:  1.12-13

pictures Some pictures in GIF format.

So that you can try xv right away, we’ve put some pictures together. These pictures are installed under /usr/share/pictures. Please use xv’s special feature ”visual schnautzer” in the pictures’ directory to get an overview over the pictures. The example window manager menu contains a few items that load these pictures as desktop backgrounds.

Version:  98.3.13-0

pixmp Xpm pixel editor for X11

Pixmap is a programm which enables you to edit XPM-files (colour bitmaps). You can use them with every common used iconmanager and even incorporate them in your own desktop environment.

Contents: pixmap2.6

Copyright:  (c) Copyright 1991,1992,1993,1994 - Lionel Mallet
Author:  Lionel Mallet <pixmap@sophia.inria.fr>, Tim Wise <pixmap@sophia.inria.fr>
Version:  2.6-7

povray Ray tracer

The Persistence of Vision Ray tracer creates three-dimensional, photo-realistic images using a rendering technique called ray tracing. It reads in a text file containing information describing the objects and lighting in a scene and generates an image of that scene from the view point of a camera also described in the text file. Ray tracing is not a fast process by any means, (the generation of a complex image can take several hours) but it produces very high quality images with realistic reflections, shading, perspective, and other effects.

Copyright:  (c) 1996 POV-Team(tm)

Author:

Steve Anger <sanger@hookup.net>, Dieter Bayer <100255.3074@compuserve.com>, Chris Cason <Chris.Cason@oaks.com.au>, Andreas Dilger <adilger@enel.ucalgary.ca>, Alexander Enzmann <xander@mitre.com>, Dan Farmer <74431.1075@compuserve.com>, Lutz Kretzschmar <100023.2006@compuserve.com>, Mike Miller <70353.100@compuserve.com>, Joel NewKirk <102627.1152@compuserve.com>, Anton Raves <100022.2603@compuserve.com>, Tim Rowley <trowley@geom.umn.edu>, Eduard Schwan <71513.2161@compuserve.com>, Zsolt Szalavari <zsolt@cg.tuwien.ac.at>, Timothy Wegner <twegner@phoenix.net>, Chris Young <76702.1655@compuserve.com>

Version:  3.02-9

pstoedit PostScript and PDF converter to vector grafik formats

pstoedit converts Postscript(TM) and PDF files to other vector graphic formats so that they can be edited graphically. pstoedit supports: tgif: Tgif .obj format (for tgif version >= 3) rpl: Real3D Programming Language Format lwo: LightWave 3D Object Format rib: RenderMan Interface Bytestream mif: (Frame)Maker Intermediate Format fig: .fig format for xfig xfig: .fig format for xfig pdf: Adobe’s Portable Document Format gnuplot: gnuplot format ps: Flattened PostScript debug: for test purposes dump: for test purposes (same as debug) dxf: CAD exchange format java: java applet source code idraw: Interviews draw format and: cgm: CGM Format (binary) cgmt: CGM Format (textual) if you will use the ”CGM Draw” library by G. Edward Johnson <lorax@nist.gov>. Cf. the document CGM projects at NIST http://speckle.ncsl.nist.gov/˜ lsr/cgm.htm. pstoedit needs an installed Ghostscript interpreter.

Contents: cd1.3 pstoedit.2.60

Copyright:  (c) pstoedit: GPL, cd: Freely Redistributable

Author:

Wolfgang Glunz <Wolfgang.Glunz@mchp.siemens.de>, G. Edward Johnson <lorax@nist.gov>

Version:  2.60-3

sane SANE (Scanner Access Now Easy)

SANE (Scanner Access Now Easy) provides a graphical frontend to your Scanner. Various scanners are supported. Please notice that this project is still being developped.

Contents: sane-0.70

Copyright:  (c) 1996, 1997, Free Software Foundation, Inc.

Author:

Andreas Beck <becka@sunserver1.rz.uni-duesseldorf.de>, David Etherton <etherton@netcom.com>, David Mosberger <David.Mosberger@acm.org>, Gordon Matzigkeit <gord@gnu.ai.mit.edu>, Michael K. Johnson <johnsonm@redhat.com>, Oliver Rauch <Oliver.Rauch@Wolfsburg.DE>, Patrick Reynolds <patrickr@virginia.edu>, Tristan Tarrant <ttarrant@etnoteam.it>

Version:  0.70-5

sced Scene editor and animator for ray tracers

Sced is a tool for three dimensional modeling. These scenes might as well be animated. Moreover it is able to export animated sequences to a renderer (e.g. PovRay, Rayshade, RenderMan, Radiance, Genray and Genscan). Support for PovRay is set by default.

Documentation can be found under: /usr/doc/packages/sced.

Contents: sced-0.94-

Copyright:  (c) 1995, 1996 Steve Chenney
Author:  Steve Chenney <schenney@cs.berkeley.edu>
Version:  0.94-10

surf computer pictures of algebraic surfaces

This program was used to generate the pictures on the S.u.S.E. covers.

Contents: surf-0.81

Copyright:  (c) 1996 Friedrich-Alexander-Universitaet Erlangen-Nuernberg
Author:  Stephan Endrass <endrass@mi.uni-erlangen.de>
Version:  0.81-12

tgif Vector-oriented drawing program

This program has nothing to do with the GIF format. It is an extremely powerful and easy to use vector-oriented drawing program which has every feature you would expect in a drawing utility.

Contents: tgif-3.0-p12

Copyright:  (c) 1990-1996, William Chia-Wei Cheng
Author:  William Chia-Wei Cheng <william@cs.UCLA.edu>
Version:  3.0p12-8

tiff tools for converting from/to tiff

Library and support programs for the TIFF image format. The sources of this package are in the source-package of libgr.

Contents: tiff-v3.4beta036

Copyright:  (c) 1988-1996 Sam Leffler, 1991-1996 Silicon Graphics Inc.
Author:  Sam Leffler <sam@engr.sgi.com>
Version:  v3.4beta037-10

tkdraw tkdraw

drawing tool save in eps format

Contents: tkdraw

Copyright:  (c) (C) 1998 Sylvio Triebel (GPL)
Author:  Sylvio Triebel <silvio@inf-technik.tu-ilmenau.de>
Version:  1.1-0

xanim multiformat animation viewer for X

XAnim currently supports the following animation types: FLI, FLC, IFF, GIF87a, GIF89a, GIF89a animation extensions, DL, Amiga PFX, URT RLE, AVI (ULTI (16), CRAM (8 & 16), CVID (24), RGB (8), RLE8), Quicktime Animation (RAW (8), RLE (1,8,16 & 24), SMC (8), RPZA (16), CVID (16), YUV2 (24); and any combination of the above on the same command. After 11 pm it makes tea as well ;-) The program seems to be quite nice, anyway; there is even a small control panel included.

Contents: xanim27064

Copyright:  (c) 1992-95 Mark Podlipec

Author:

Mark  Podlipec  <podlipec@shell.portal.com>,  Mark  Podlipec

<podlipec@wellfleet.com>

Version:  2.70.64-2

xbmbrosr xbmbrowser

view and manage X bitmap and X pixmap files Contents: xbmbrowser5.1

Author:  Ashley Roll <ash@dstc.edu.au>, Anthony Thyssen <anthony@cit.gu.edu.au>
Version:  5.1-8

xgrabsc A screengrabber for X.

You need screenshots easily? xgrab and xgrabsc give you a powerful tools to save screenshots in files (ps, xpm, etc.) or to print them. xgrabsc is invoked from a shell an sends the rectangle grabbed on the screen to standard output. xgrab is a menu driven frontend to xgrabsc. Documentation may be obtained by man xgrab resp. man xgrabsc. The author’s original comments can be found in /usr/doc/packages/xgrabsc. :noitpircseD Notify: :Yfiton :Hsilgne Contents: xgrabsc-2.41

Copyright:

(c) 1990-95 Bruce Schuchardt, GemStone Systems Inc.; 1989 by the Massachusetts Institute of Technology

Author:

Bruce Schuchardt <bruce@slc.com>, Yves Arrouye, George Ferguson, David Lawrence, Alan Kent, Brian C. Smith, Johan Garpendal, Martin Birgmeier, Vasco Lopes Paulo, Hal R. Brand, Matt Freedman, Scott McCoy

Version:  2.41-8

xgrasp GRASP animation viewer

X11 GRASP animation viewer. Contents: xgrasp-1.7d

Copyright:  (c) 1991 Patrick J. Naughton
Author:  Patrick J. Naughton <naughton@sun.com>
Version:  1.7d-7

xli X11 Image Loading Utility

xli is a version of xloadimage. This utility will view several types of images under X11, or load images onto the X11 root window. xli can also be used on some (32 bit) MSDOS systems. The current version supports:

* CMU Window Manager raster files

* Faces Project images

* Fuzzy Bitmap (FBM) images

* GEM bit images

* GIF images (Including GIF89a compatibility)

* G3 FAX images

* JFIF style jpeg images

* McIDAS areafiles

* MacPaint images

* PC Paintbrush (PCX) images

* Portable Bitmap (PBM, PGM, PPM) images

* Sun monochrome rasterfiles

* Sun color RGB rasterfiles

* Targa (TGA) files

* Utah Raster Toolkit (RLE) files * X pixmap (XPM) files (Version 1, 2C and 3)

* X10 bitmap files

* X11 bitmap files

* X Window Dump (except TrueColor and DirectColor)

Contents: xli-1.16

Copyright:  (c) 1992, 1993 David Koblas
Author:  David Koblas <koblas@netcom.com>
Version:  1.16-8

xmgr Plotting Tool

A plotting tool under X. Is you want to use the help function, you need Mosaic or any other html browser to read the help files. The name of the help viewer is set to the environmental variable GR HELPVIEWER. Due to legal reasons, we are not allowed to put an X11 html browser on the CD. Your friendly neighbour (the one with access to the internet) will surely be able to help you.

Contents: xmgr-4.00beta.pl5

Copyright:  (c) 1995 Paul J Turner, Portland, OR
Author:  Paul J. Turner <pturner@teleport.com>
Version:  4.00pl5-9

xmorph A morphing program.

This morphing program reads two images in Targa format and computes the transitions from one image to the other (arbitrarily many steps). To avoid a simple fading effect, the program needs information about the shapes contained in the two images. This is done by using a grid consisting of connected checkpoints. The source grid is then transformed into the destination grid during the morphing process.

Contents: xmorph-22jul94

Copyright:  (c) 1994 Michael J. Gourlay
Author:  Michael J. Gourlay <mike@penguin.gatech.edu>
Version:  94.7.22-5

xmrm Multi Resolution Morphing with GUI

This program provides a comfortable way to morph pictures. You’ll find a HTML-Manual in /usr/doc/packages/xmrm/manual. Just load the file mrm manual.html into your browser.

Contents: xforms-0.81.ELF xmrm

Copyright:  (c) 1997 Gerhard Waldh”or, Andreas Artmann

Author:

Gerhard Waldh”or <xmrm@titus.cg.tuwien.ac.at>, Andreas Artmann <xmrm@titus.cg.tuwien.ac.at>

Version:  1.2-10

xpaint Xpaint

XPaint is a versatile bitmap/pixmap editing tool. The short features list: Usual paint operations: Brushes, Spray paint, Pencil, Lines, Arcs, Pattern Fill, Text, Boxes, Circles, Polygons. Works on multiple images simultaneously. Cut/Copy/Paste between all active images. Fatbits/Zoom on the image windows.

Contents: xpaint-2.4.5

Copyright:  (c) 1992, 1993, David Koblas
Author:  David Koblas <koblas@netcom.com>, Greg Roelofs <newt@uchicago.edu>
Version:  2.4.5-11

xpdf a pdf-file viewer under X11

This tool serves for viewing files in PDF (Portable Document Format). It is used, e.g., by Adobe PDF Software ”Acrobat”.

Contents: xpdf-0.7

Copyright:  (c) 1996 Derek B. Noonburg
Author:  Derek B. Noonburg <derekn@ece.cmu.edu>
Version:  0.7-9

xray XRay

KDE: K Desktop Environment

XRay is a program in which you can visualize 3D bodies with various optical properties. It can handle objects such as polygons, spheres, ellipsoids and torii. This tool uses the following light sources: point, ray and surface. You can represent your data using the following algorithms: z-buffer, raytracing, enhanced radiosity or a two phase method. The package contains two main programs. One generates the scene, and the other is the user interface for the raytracer. If you have installed the linuxthread package, you can use the mulithreaded version. Before starting please read the documentation in /usr/doc/packages/xray/doku/ . The documentation in PostScript format is in the package xrayps.tgz.

Contents: xray

Copyright:  (c) Helmut Fahrion
Author:  Helmut Fahrion <hf@suse.de>
Version:  97.12.10-3

xrayps XRay PostScript Documentation

This ist the documentation for XRay in PostScript format. This package is very large, 100Mb. Please install this package only if you want to print the documentation in book form.

Contents: xray

Copyright:  (c) Helmut Fahrion
Author:  Helmut Fahrion <hf@suse.de>
Version:  97.12.10-3

xwarp Warp an image (GIF format)

Warp an image (GIF format) defining 2 grids (source and goal) under X11. The grid can be modified with the mouse. Example files can be found in /usr/local/lib/XWarp-1.00-example/ .

Contents: XWarp-1.00

Author:  Luca Maranzano, Guido Cardino, Roberto Surlinelli, <calamaro@dist.dist.unige.it>
Version:  1.0-7

zgv Viewer (libvga)

Zgv is a GIF/JPEG/PNM/BMP graphic file viewer for VGA and SVGA displays. (It uses svgalib.) It provides a graphic-mode file selector to select file(s) to view. The program may be used as a command; in this case, the front end is not used.

Contents: zgv2.8-src

Copyright:  (c) 1993, 1994 Russell Marks
Author:  Russell Marks <mr216@gre.ac.uk>
Version:  2.8-9

0.1.9 KDE: K Desktop Environment

kadm KDE-Admin (Alpha!)

Alpha stage! KDE is still in an early development stage (beta). Thus, KDE being relatively stable, it may nevertheless happen that any of the KDE applications crashes or produces strange effects. This package is espescially concerned since it was built from not yet officially released programs. Newest KDE packages adepted to S.u.S.E. Linux may be found at ftp://ftp.suse.com/pub/suse update/KDE. General information for KDE can be found at the respective addresses in WWW (http://www.kde.org) or our support database (http://www.suse.de/Support/sdb). This package contains some additional admin tools. Documentation in: /usr/doc/packages/kde/kadm Contents: kadm klpc-0.1.3 Alpha stage!

KDE is still in an early development stage (beta). Thus, KDE being relatively stable, it may nevertheless happen that any of the KDE applications crashes or produces strange effects. This package is espescially concerned since it was built from not yet officially released programs. Newest KDE packages adepted to S.u.S.E. Linux may be found at ftp://ftp.suse.com/pub/suse update/KDE. General information for KDE can be found at the respective addresses in WWW (http://www.kde.org) or our support database (http://www.suse.de/Support/sdb).

Copyright:  (c) 1997,1998 The KDE-Team
Author:  The KDE Team <kde@kde.org>
Version:  alpha-5

kadmin KDE system administration

KDE is still in an early development stage (beta). Thus, KDE being relatively stable, it may nevertheless happen that any of the KDE applications crashes or produces strange effects. Newest KDE packages adepted to S.u.S.E. Linux may be found at ftp://ftp.suse.com/pub/suse update/KDE. General information for KDE can be found at the respective addresses in WWW (http://www.kde.org) or our support database (http://www.suse.de/Support/sdb). This package contains kdeadmin, some tools which are useful to administrate the system. Documentation in /usr/doc/packages/kde/kadmin Contents: kdeadmin-980211

Copyright:  (c) 1997,1998 The KDE-Team
Author:  The KDE Team <kde@kde.org>
Version:  beta3-7

kdesdk KDE SGML and HTML tools (Alpha!)

Alpha stage! KDE is still in an early development stage (beta). Thus, KDE being relatively stable, it may nevertheless happen that any of the KDE applications crashes or produces strange effects. This package is especially concerned since it was built from not yet officially released programs. Newest KDE packages adepted to S.u.S.E. Linux may be found at ftp://ftp.suse.com/pub/suse update/KDE. General information for KDE can be found at the respective addresses in WWW (http://www.kde.org) or our support database (http://www.suse.de/Support/sdb). Some tools which could be used to convert SGML documents into HTML documents. Documentation in: /usr/doc/packages/kde Contents: kdesdk-980204 Alpha stage! KDE is still in an early development stage (beta). Thus, KDE being relatively stable, it may nevertheless happen that any of the KDE applications crashes or produces strange effects. This package is espescially concerned since it was built from not yet officially released programs. Newest KDE packages adepted to S.u.S.E. Linux may be found at ftp://ftp.suse.com/pub/suse update/KDE. General information for KDE can be found at the respective addresses in WWW (http://www.kde.org) or our support database (http://www.suse.de/Support/sdb).

Copyright:  (c) 1997,1998 The KDE-Team

Author:

Bernd Wuebben <wuebben@kde.org>, Kalle Dalheimer <kalle@kde.org>, Martin Konold <konold@kde.org>, Martin R. Jones <mjones@kde.org>, Matthias Ettrich <ettrich@kde.org>, Stephan Kulow <coolo@kde.org>, Thorben Weis <weis@kde.org>

Version:  beta3-4

kdevel KDE development tools (Alpha!)

Alpha stage! KDE is still in an early development stage (beta). Thus, KDE being relatively stable, it may nevertheless happen that any of the KDE applications crashes or produces strange effects. This package is especially concerned since it was built from not yet officially released programs.

Newest KDE packages adepted to S.u.S.E. Linux may be found at ftp://ftp.suse.com/pub/suse update/KDE. General information for KDE can be found at the respective addresses in WWW (http://www.kde.org) or our support database (http://www.suse.de/Support/sdb). Contains some tools to administrate software projects. Documentation in: /usr/doc/packages/kde/kdevel Contents: kdevel Alpha stage! KDE is still in an early development stage (beta). Thus, KDE being relatively stable, it may nevertheless happen that any of the KDE applications crashes or produces strange effects. This package is especially concerned since it was built from not yet officially released programs. Newest KDE packages adepted to S.u.S.E. Linux may be found at ftp://ftp.suse.com/pub/suse update/KDE. General information for KDE can be found at the respective addresses in WWW (http://www.kde.org) or our support database (http://www.suse.de/Support/sdb).

Copyright:  (c) 1997,1998 The KDE-Team
Author:  The KDE Team <kde@kde.org>
Version:  alpha-5

kgames KDE games

KDE is still in an early development stage (beta). Thus, KDE being relatively stable, it may nevertheless happen that any of the KDE applications crashes or produces strange effects. Newest KDE packages adepted to S.u.S.E. Linux may be found at ftp://ftp.suse.com/pub/suse update/KDE. General information for KDE can be found at the respective addresses in WWW (http://www.kde.org) or our support database (http://www.suse.de/Support/sdb). Some nice games for KDE. Contents: kdegames-980211

Copyright:  (c) 1997,1998 The KDE-Team
Author:  The KDE Team <kde@kde.org>
Version:  beta3-8

kgraph KDE graphics programs

KDE is still in an early development stage (beta). Thus, KDE being relatively stable, it may nevertheless happen that any of the KDE applications crashes or produces strange effects. Newest KDE packages adepted to S.u.S.E. Linux may be found at ftp://ftp.suse.com/pub/suse update/KDE. General information for KDE can be found at the respective addresses in WWW (http://www.kde.org) or our support database (http://www.suse.de/Support/sdb). This package contains some graphics programs which sometimes can replace traditional programs: kdvi, kfax, kfract, kghostview, kpaint, kview Contents: kdegraphics-980209

Copyright:  (c) 1997,1998 The KDE-Team
Author:  The KDE Team <kde@kde.org>
Version:  beta3-7

klibs KDE base package (libs)

KDE is still in an early development stage (beta). Thus, KDE being relatively stable, it may nevertheless happen that any of the KDE applications crashes or produces strange effects. Newest KDE packages adepted to S.u.S.E. Linux may be found at ftp://ftp.suse.com/pub/suse update/KDE. General information for KDE can be found at the respective addresses in WWW (http://www.kde.org) or our support database (http://www.suse.de/Support/sdb). This package contains kdelibs, one of the basic packages of the K Desktop Environment. It contains necessary libraries the KDE desktop. This package is absolutely necessary for using KDE! Documentation may be found in:

/usr/doc/packages/kde/klibs Contents: kdelibs-980210

Copyright:  (c) 1997,1998 The KDE-Team
Author:  The KDE Team <kde@kde.org>
Version:  beta3-5

kmisc Some KDE tools (Alpha!)

Alpha stage! KDE is still in an early development stage (beta). Thus, KDE being relatively stable, it may nevertheless happen that any of the KDE applications crashes or produces strange effects. This package is especially concerned since it was built from not yet officially released programs. Newest KDE packages adepted to S.u.S.E. Linux may be found at ftp://ftp.suse.com/pub/suse update/KDE. General information for KDE can be found at the respective addresses in WWW (http://www.kde.org) or our support database (http://www.suse.de/Support/sdb). Contains some very practical tools. Documentation in: /usr/doc/packages/kde/kmisc Contents: kmodplayer-0.3.6 kmpg-0.3.2 konquest-0.0.3 kpov-0.1.5 krecord-0.3 ktamaga-0.2 Alpha stage! KDE is still in an early development stage (beta). Thus, KDE being relatively stable, it may nevertheless happen that any of the KDE applications crashes or produces strange effects. This package is especially concerned since it was built from not yet officially released programs. Newest KDE packages adepted to S.u.S.E. Linux may be found at ftp://ftp.suse.com/pub/suse update/KDE. General information for KDE can be found at the respective addresses in WWW (http://www.kde.org) or our support database (http://www.suse.de/Support/sdb).

Copyright:  (c) 1997,1998 The KDE-Team
Author:  The KDE Team <kde@kde.org>
Version:  alpha-7

kmulti KDE-multimedia package

KDE is still in an early development stage (beta). Thus, KDE being relatively stable, it may nevertheless happen that any of the KDE applications crashes or produces strange effects. Newest KDE packages adepted to S.u.S.E. Linux may be found at ftp://ftp.suse.com/pub/suse update/KDE. General information for KDE can be found at the respective addresses in WWW (http://www.kde.org) or our support database (http://www.suse.de/Support/sdb). This package contains some multimedia programs: kmix, kmedia, kscd Contents: kdemultimedia-980211

Copyright:  (c) 1997,1998 The KDE-Team
Author:  The KDE Team <kde@kde.org>
Version:  beta3-7

knet KDE networking package

In these days, KDE is in an early development stage (beta). Thus, KDE being relatively stable, it may nevertheless happen that any of the KDE applications crashes or produces strange effects. Moreover, we cannot give technical support for KDE matters directly but we’d like to appoint you to the respective addresses in WWW (http://www.kde.org) or in case to our support database (http://www.suse.de/Support/sdb). A mailer and a news reader for KDE. Contents: kdenetwork-980212

Copyright:  (c) 1997,1998 The KDE-Team
Author:  The KDE Team <kde@kde.org>
Version:  beta3-7

knetutil KDE network utilities (Alpha!)

Alpha stage! KDE is still in an early development stage (beta). Thus, KDE being relatively stable, it may nevertheless happen that any of the KDE applications crashes or produces strange effects. This package is especially concerned since it was built from not yet officially released programs. Newest KDE packages adepted to S.u.S.E. Linux may be found at ftp://ftp.suse.com/pub/suse update/KDE. General information for KDE can be found at the respective addresses in WWW (http://www.kde.org) or our support database (http://www.suse.de/Support/sdb). Contains some additional network tools. Documentation in: /usr/doc/packages/kde/knetutil Contents: kISDN-0.3.1 kfinger-0.6.1 khostchooser-0.2.2 kmodem0.0.3 knetutil ksamba-0.3 ktalkd-0.2.0 Alpha stage! KDE is still in an early development stage (beta). Thus, KDE being relatively stable, it may nevertheless happen that any of the KDE applications crashes or produces strange effects. This package is especially concerned since it was built from not yet officially released programs. Newest KDE packages adepted to S.u.S.E. Linux may be found at ftp://ftp.suse.com/pub/suse update/KDE. General information for KDE can be found at the respective addresses in WWW (http://www.kde.org) or our support database (http://www.suse.de/Support/sdb).

Copyright:  (c) 1997,1998 The KDE-Team
Author:  The KDE Team <kde@kde.org>
Version:  alpha-5

kscience KDE scientific tools (Alpha!)

Alpha stage! KDE is still in an early development stage (beta). Thus, KDE being relatively stable, it may nevertheless happen that any of the KDE applications crashes or produces strange effects. This package is especially concerned since it was built from not yet officially released programs. Newest KDE packages adepted to S.u.S.E. Linux may be found at ftp://ftp.suse.com/pub/suse update/KDE. General information for KDE can be found at the respective addresses in WWW (http://www.kde.org) or our support database (http://www.suse.de/Support/sdb). Some scientific tools. For example a simple fractal viewer. Documentation in: /usr/doc/packages/kde/kscience Contents: kplot-0.2 kscience Alpha stage! KDE is still in an early development stage (beta). Thus, KDE being relatively stable, it may nevertheless happen that any of the KDE applications crashes or produces strange effects. This package is especially concerned since it was built from not yet officially released programs. Newest KDE packages adepted to S.u.S.E. Linux may be found at ftp://ftp.suse.com/pub/suse update/KDE. General information for KDE can be found at the respective addresses in WWW (http://www.kde.org) or our support database (http://www.suse.de/Support/sdb).

Copyright:  (c) 1997,1998 The KDE-Team
Author:  The KDE Team <kde@kde.org>
Version:  alpha-5

ksupp KDE base package (support)

KDE is still in an early development stage (beta). Thus, KDE being relatively stable, it may nevertheless happen that any of the KDE applications crashes or produces strange effects. Newest KDE packages adepted to S.u.S.E. Linux may be found at ftp://ftp.suse.com/pub/suse update/KDE.

General information for KDE can be found at the respective addresses in WWW (http://www.kde.org) or our support database (http://www.suse.de/Support/sdb). This package contains kdesupport, one of the basic packages of the K Desktop Environment. It contains necessary libraries for managing graphics, mail and html processing. This package is absolutely necessary for using KDE! Documentation may be found in: /usr/doc/packages/kde/ksupp Contents: kdesupport-980204

Copyright:  (c) 1997,98 The KDE-Team
Author:  The KDE Team <kde@kde.org>
Version:  beta3-5

ktools KDE tools (Alpha!)

Alpha stage! KDE is still in an early development stage (beta). Thus, KDE being relatively stable, it may nevertheless happen that any of the KDE applications crashes or produces strange effects. This package is especially concerned since it was built from not yet officially released programs. Newest KDE packages adepted to S.u.S.E. Linux may be found at ftp://ftp.suse.com/pub/suse update/KDE. General information for KDE can be found at the respective addresses in WWW (http://www.kde.org) or our support database (http://www.suse.de/Support/sdb). Some practical tools which you may need every day. For example: kdayplan Documentation in: /usr/doc/packages/kde/ktools Contents: kdayplan kless ksnapshot-0.2.3 ktools ktop-0.9.3 kvmode-0.0.2 Alpha stage! KDE is still in an early development stage (beta). Thus, KDE being relatively stable, it may nevertheless happen that any of the KDE applications crashes or produces strange effects. This package is especially concerned since it was built from not yet officially released programs. Newest KDE packages adepted to S.u.S.E. Linux may be found at ftp://ftp.suse.com/pub/suse update/KDE. General information for KDE can be found at the respective addresses in WWW (http://www.kde.org) or our support database (http://www.suse.de/Support/sdb).

Copyright:  (c) 1997,1998 The KDE-Team
Author:  The KDE Team <kde@kde.org>
Version:  alpha-5

kutils KDE utility package

KDE is still in an early development stage (beta). Thus, KDE being relatively stable, it may nevertheless happen that any of the KDE applications crashes or produces strange effects. Newest KDE packages adepted to S.u.S.E. Linux may be found at ftp://ftp.suse.com/pub/suse update/KDE. General information for KDE can be found at the respective addresses in WWW (http://www.kde.org) or our support database (http://www.suse.de/Support/sdb). Some utilities for KDE: karm, kcalc, kedit, khexdit, kjots, knotes, kzip Contents: kdeutils-980212 kmke2fs

Copyright:  (c) 1997,1998 The KDE-Team
Author:  The KDE Team <kde@kde.org>
Version:  beta3-7

kbase KDE base package (base)

KDE is still in an early development stage (beta). Thus, KDE being relatively stable, it may nevertheless happen that any of the KDE applications crashes or produces strange effects. Newest KDE packages adepted to S.u.S.E. Linux may be found at ftp://ftp.suse.com/pub/suse update/KDE. General information for KDE can be found at the respective addresses in WWW (http://www.kde.org) or our support database (http://www.suse.de/Support/sdb).

N: Network-Support (TCP/IP, UUCP, Mail, News)

This package contains kdebase, one of the basic packages of the K Desktop Environment. It contains among others KWM, the KDE window manager; KFM, the KDE file manager; KDisplay, the desktop configuration program. This package is absolutely necessary for using KDE! Documentation may be found in: /usr/doc/packages/kde/kbase

Contents: kdebase-980212

Copyright:  (c) 1997,1998 The KDE-Team
Author:  The KDE Team <kde@kde.org>
Version:  beta3-8

0.1.10 N: Network-Support (TCP/IP, UUCP, Mail, News)

amanda Network Disk Archiver

Please note that this software is ALPHA.

Contents: amanda-2.4.0b6

Copyright:  (c) 1991, 1995 University of Maryland at College Park
Author:  James da Silva <jds@cs.umd.edu>
Version:  2.4.0b6-3

amd Automounter daemon

Automounter: mounts filesystems automatically. The daemon manages UFS and NFS mounts, and can mount and unmount filesystems based on demand. Contents: amd-upl102 This package is not maintained anymore, please use autofs instead.

Author:  Jan-Simon Pendry <jsp@doc.ic.ac.uk>
Version:  upl102-10

apache Apache httpd Web server

The Apache HTTP server. Very few programs have spread so quickly and attracted such a large following in such a short time as this HTTP server. According to Netcraft, Apache runs on about 42%% of all servers attached to the Internet. That represents about 510,000 systems. It has replaced within only a few months the NCSA server which, until then was the unchallenged favorite. This is undoubtedly because Apache was derived from the NCSA server sources and that its configuration is compatible to that of NCSA/1.3. The Apache project was established to provide a forum for the development of a freely available HTTP server. The goal is to develop a secure, efficient, and easily extendable server, which also closely parallels the development of the HTTP protocol. Thus Apache became a replacement to NCSA/1.3. Besides better performance, the removal of known errors, and better conformance to HTTP standards (HTTP/1.1), Apache offers the following useful features, which are often missing in other servers:

* password files, used to protect documents from unauthorized access, can be indexed in various ways. The search for an entry is as a result more efficient than in ordinary text files.

* client-request errors can be responded to in a flexible manner. CGI scripts can be used to analyze the error and generate more informative replies to both the client and the administrator than the HTTP status codes offer.

* there is no restriction on aliasing and redirection. The Alias and Redirect directives can be used in arbitrary number and depth.

* Apache can choose the best representation of a resource based on the browser-supplied preferences for media type, languages, character set and encoding (”Content Negotiation”).

* with Apache you can implement several virtual servers. Thus requests to different IP addresses (referring to the same computer) can be distinguished and, depending on the IP address, appropriately handled.

Further advantages are the module concept and the well-documented API (Application Programmers Interface). Extensions are easily developed and new features are easily integrated into the server. For example, an implementation of the Secure Socket Layer (SSL) from Netscape exists for Apache. This permits secure communication between server and client. Apache is installed in the following directories: /etc/httpd -> configuration files /usr/sbin/httpd -> server binary /var/log/httpd.* -> protocol files /var/run/httpd.pid -> server PID /usr/local/httpd -> server root /sbin/init.d/apache -> Start/Stop script Contents: apache 1.2.5+ssl 1.13 apache 1.2.5

Copyright:  (c) 1995 The Apache Group.

Author:

Rob McCool <robm@netscape.com>, Robert S. Thau <rst@ai.mit.edu>, Roy T. Fielding <fielding@avron.ics.uci.edu>, Brian Behlendorf <brian@wired.com>, Harald Hanche-Olsen <hanche@imf.unit.no>, David Robinson <kluft@amdahl.com>, Wietse Venema <drtr@ast.cam.ac.uk>, Keith Shafer <Keith.Shafer@SanDiegoCA.ATTGIS.COM>, Kevin Steves <stevesk@mayfield.hp.com>

Version:  1.2.5-3

archie Where do I find what?

Using this program, you can query archie servers without having to login with telnet. This version uses the server archie.funet.fi as default. You may specify another host using the command line switch ”-h”. Contents: archie-1.4.1

Copyright:  (c) 1989-91 University of Washington

Author:

Alan Emtage <bajan@cs.mcgill.ca>, Peter Deutsch <peterd@cs.mcgill.ca>, Bill Heelan <wheelan@cs.mcgill.ca>, Clifford Neuman<bcn@isi.edu>, Brendan Kehoe <brendan@cygnus.com>,

Version:  1.4.1-8

autofs autofs is a kernel-based automounter

autofs is a kernel-based automounter for Linux. It is still under development, although the features currently implemented seem to work well. There is not anywhere near enough documentation available yet; help in getting some written would be appreciated. This package replaces the amd package. Contents: autofs-0.3.14 This package replaces the amd package.

Copyright:  (c) GPL, 1997 Transmeta Corporation
Author:  H. Peter Anvin <hpa@transmeta.com>
Version:  0.3.14-8

bind BIND - Name Server Utilities

The named daemon and support utilities including: dig, dnsquery, host, nslookup, nsquery, and nstest. Documentation on setting up a name server can be found in /usr/doc/packages/bind . Contents: bind-4.9.6

Copyright:  (c) 1989 The Regents of the University of California.
Author:  Paul Vixie <vixie@vix.com>
Version:  4.9.6-9

bing point-to-point bandwidth measurement tool

Bing determines the real (raw, as opposed to available or average) throughput on a link by measuring ICMP echo requests roundtrip times for different packet sizes for each end of the link. Contents: bing-1.0.4

Copyright:  (c) GPL
Author:  Pierre Beyssac <pb@fasterix.freenix.fr>
Version:  1.0.4-6

bulkmail bulk mailer

This is a C program to do ”bulk” mailing. For input, it takes a file of recipient addresses (one address per line) and a message (with headers already attached) to be sent to the recipients. It sorts the recipient list by reversed domain (so similar ones sort together), splits up the recipients into several groups containing no more than N domains each, creates an SMTP envelope for each group of recipients, and feeds that envelope to ”/usr/lib/sendmail -bs”. Splitting the envelopes up allows sendmail to perform delivery in parallel, so instead of having one large queue entry (for which sendmail might take awhile to get around to attempting delivery for some recipients), it has several smaller queue entries. Depending on your point-of-view, this can still be considered ”cluttering up your mail queue”, but it does seem to deliver messages more quickly to most recipients. The core of this program was extracted from a somewhat strange mailing list manager called na-net; it was designed to efficiently send out mail to 5000 people at a time. I have used this program to attempt delivery of a message to over 12000 recipients around the world, within a few hours. I’m currently using bulk mailer as a back-end for several mailing lists of modest size. However, the program is not extensively tested, and may not work well in all environments. (In particular, if your system has per-user process quotas, or a small number of process table entries, you will want to modify this program to recover gracefully.) There’s no warranty on this, but you’re welcome to use it if you want. Contents: bulk mailer-1.5

Author:  Keith Moore <moore@cs.utk.edu>
Version:  1.5-5

cnews Cleanup Release of C News.

Controls the spooling and transmission of Usenet news. Cnews is typically used with UUCP -- for NNTP connections INN is usually a better choice. To activate the cnews-system, you have to edit the crontab belonging to root or news. An example crontab can be found under /var/lib/news/crontab. Please read the file /var/lib/news/README.SuSE as well. Contents: c-news nntp-1.5.12.1 Please note, that this package is not supported by our support team. Please use inn if you need support for news. Feedback and bug fixes for cnews are of course welcome at feedback@suse.de.

Copyright:  (c) 1985-92 University of Toronto
Author:  Geoffrey Collyer
Version:  1.G-5

deliver A program for the delivery of email

A small and simple program that delivers electronic mail to local users. If you are attached to a network, you should install the program ”sendmail” as well. Contents: deliver-2.1.13

Author:  Chip Salzenberg <chip@fin.uucp>
Version:  2.1.13-1

dhcp ISC DHCP server

Version 1of the ISC DHCP Distribution includes just a DHCP Server. Version 1 has been in feature freeze since late 1996, and is quite stable. This is the release that we would expect most sites to run in production. Please read the documentation in /usr/doc/packages/dhcp regarding configuration of the DHCP server. Contents: dhcp-1.0.0

Copyright:  (c) 1995, 1996, 1997 The Internet Software Consortium.
Author:  Ted Lemon <mellon@fugue.com>
Version:  1.0.0-3

diald Automatic dialer

Demand dialing daemon for SLIP and PPP links. Supports demand dialed links with link timeouts and redialing.

More info: /usr/doc/packages/diald Example configuration: /usr/doc/packages/diald/config

Contents: diald-0.14 diald-top-0.6

Copyright:  (c) 1994, 1995, 1996 Eric Schenk
Author:  Eric Schenk <schenk@cs.toronto.edu>
Version:  0.14-8

dochost Documentation server

With this package you are able to access a central documentation-server on which the S.u.S.E. onlinesupport system is stored. It should be installed on the server, too.

Contents: dochost-98.03.02

Copyright:  (c) Coypright (c) 1998 S.u.S.E. GmbH Fuerth, Germany
Version:  98.03.02-0

elm Menu-driven user mail program.

Elm is an interactive screen-oriented mailer program originally written by Dave Taylor. Elm is now maintained by the Elm Development Group.

Contents: elm2.4.ME+.37

Copyright:  (c) 1986,1987 Dave Taylor, 1988-1992 USENET Community Trust
Author:  Dave Taylor <taylor@intuitive.com>
Version:  2.4.37-7

emil Email filter

This program enables you to convert MIME, SUN mailtool and ’plain old style RFC822’. Could be interesting for ’elm addicted’ people..;-)

Contents: emil-2.1.0-beta9

Version:  2.1.0beta9-8

firewall Firewall and masquerading scripts

Two scripts to configure your machine as a firewall and/or masquerading router. See /usr/doc/packages/firewall for more details.

Copyright:  (c) 1997 S.u.S.E. GmbH
Author:  Bodo Bauer <bb@suse.de>, Volker Graf <volker.graf@student.uni-tuebingen.de> (fcii)
Version:  1.4.3-0

ftpdir A sample ”FTP configuration”

This package contains a ’change-root’-configuration for anonymous ftp.

Version:  97.6.29-2

gettyps Just another getty.

This getty is included because it supports ring-back. Documentation: man getty man gettydefs under /usr/doc/packages/gettyps you can find some examples.

Contents: getty ps-2.0.7i

Copyright:  (c) 1989,1990 by Paul Sutcliffe Jr.
Author:  Kris Gleason <gleasokr@boulder.colorado.edu>
Version:  2.0.7i-9

gup Group Update Program for INN and C-News

gup - A Group Update Program that accepts commands by mail to edit a newsgroup subscription file for subsequent use by news systems such as INN and C-News

Contents: gup

Copyright:  (c) 1993 Mark Delany & Andrew Herbert

Author:

Andrew Herbert <andrew@werple.apana.org.au>, Mark Delany <markd@bushwire.apana.org.au>

Version:  0.3-8

htdig WWW indexing and searching system

The ht://Dig system is a complete world wide web indexing and searching system for a small domain or intranet. This system is not meant to replace the need for powerful internet-wide search systems like Lycos, Infoseek, Webcrawler and AltaVista. Instead it is meant to cover the search needs for a single company, campus, or even a particular sub section of a web site. As opposed to some WAIS-based or web-server based search engines, ht://Dig can span several web servers at a site. The type of these different web servers doesn’t matter as long as they understand the HTTP 1.0 protocol.

Contents: htdig-3.0.8b2

Copyright:  (c) ht://Dig 1995, 1996, 1997 Andrew Scherpbier (GPL)
Author:  Andrew Scherpbier <andrew@contigo.com>
Version:  3.0.8b2-3

hylafax HylaFAX(tm) is a very complex and powerful fax server

HylaFAX(tm) is a very powerful fax server for UNIX systems. It includes the following features: Fax-Mail-Gateway (dispatching faxes to Email addresses via TSI) Mail-Fax-Gateway Adaptive Answer Support (dispatching of data or voice calls to external program’s like agetty, mgetty or vgetty) Outgoing documents can be any format Before executing the script ’faxsetup’ please DEFINITELY read /usr/doc/packages/hylafax/README.SuSE!!! Documentation can be found at /usr/doc/packages/hylafax/html. Contents: hylafax-v4.0pl2 latex-cover-1.04 Before you can use this package you have to run ‘faxsetup‘ as root.

Copyright:  (c) 1988-1996, 1991-1996 Sam Leffler, Silicon Graphics, Inc.
Author:  Sam Leffler <sam@engr.sgi.com>
Version:  v4.0pl2-3

i4l ISDN for Linux

You need these tools to use an ISDN card under Linux. An explicite explanation will be found under /usr/doc/packages/i4l/README.SuSE and in /usr/doc/packages/i4l/README.Quick.

Contents: isdn4k-utils-2.1b1 , isdnmon, isdnmonp, imon, xisdnload, isdnlog, imontty, isdnbar

Copyright:  (c) Fritz Elfert and more

Author:

Fritz Elfert <fritz@wuemaus.franken.de>, Karsten Keil <keil@temicech.spacenet.de>, Michael Hipp <Michael.Hipp@student.uni-tuebingen.de>, Andreas Kool <akool@Kool.f.EUnet.de>, Ullrich von Bassewitz <uz@musoftware.com>, Stefan Luethje <luethje@sl-gw.lake.de>, Carsten Paeth <calle@calle.in-berlin.de>

Version:  98.3.10-0

i4lfirm ISDN-firmware for active ISDN-cards

Firmware files for active ISDN-cards (ICN and AVM-B1)

Copyright:  (c) AVM, ICN
Version:  98.3.10-0

inetcfg Examples for SLIP and PPP-scripts

Some examples for DIP and PPP Scripts (Compuserve and German T-Online). The files live in /usr/doc/packages/inetcfg.

Contents: inetcfg

Copyright:  (c) S.u.S.E. GmbH, Fuerth, Germany
Author:  Klaus Franken <kfr@suse.de>
Version:  98.2.3-2

inn Inter Net News

Rich Salz’s InterNetNews news transport system.

Contents: inn-1.7

Copyright:  (c) 1983-1990 The Regents of the University of California.

Author:

Rich Salz <rsalz@osf.org>, Andreas Klemm <andreas@knobel.gun.de>, Vince Skahan <vince@atc.boeing.com>, Arjan de Vet <devet@adv.win.tue.nl>

Version:  1.7-6

ipfwadm Firewall and IP-accounting administration

Using this package you can activate, configure and maintain ’firewalling’ of the Linux kernel. Documentation: man ipfwadm

Contents: ipfwadm-2.3.0

Copyright:  (c) 1995,1996 by X/OS Experts in Open Systems BV
Author:  Jos Vos <jos@xos.nl>
Version:  2.3.0-8

ipxrip IPX routing daemon

This is a RIP/SAP daemon for Linux. You can turn your Linux machine into an IPX router.

Contents: ipxripd-0.7

Copyright:  (c) 1994, 1995 Ales Dryak; 1996, Volker Lendecke
Author:  Ales Dyrak <A.Dryak@sh.cvut.cz>, Volker Lendecke <lendecke@namu01.gwdg.de>
Version:  0.7-8

ircii Internet Relay Chat

This is a version of IRC (Internet Relay Chat) for systems with direct TCP/IP access (including SLIP/PPP). It also includes voice support and other extra features.

Contents: ircii-2.9

Copyright:  (c) 1990 Michael Sandrof

Author:

Michael  Sandrof  <ms5n+@andrew.cmu.edu>,  Matthew  Green

<mrg@mame.mu.oz.au>

Version:  2.9-9

irciihlp Help files for ircII

Version:  2.9-9

lprold BSD Print Spooling System

This is the Linux version of the BSD-compatible printer spooling system. The lpd binary included in this package expects to find TCP/IP running.

Contents: lpr-tlr-971016

Copyright:  (c) 1983, 1991 The Regents of the University of California.
Author:  Thomas Roessler <roessler@guug.de>
Version:  3.0.1-6

lynx A text-based html browser

Lynx is an easy to use browser for html documents. It is purely text based and therefore makes it possible to use WWW resources on text terminals. We must mention that this program does not fall under the GPL license and that the Copyright-owner maintains the right to require the program to become licensed if used for commercial use.

Contents: lynx2-7-2

Copyright:  (c) 1994 University of Kansas
Author:  Michael Grobe <grobe@kuhub.cc.ukans.edu>, Lou Montulli
Version:  2.7.2-0

mailx The BSD mailx mailer.

This program is a very good supplement to ”mail” and should be installed.

Contents: mailx-8.5.5

Copyright:  (c) 1980-90 Regents of the University of California.
Version:  8.5.5debian-8

marsnwe Novell server emulation

Complete emulation of a Novell server. Supports file-service, bindery-service, print-service and routing-service. Your kernel has to be configured according to /usr/doc/packages/marsnwe/INSTALL.

Contents: mars nwe-0.99.pl4

Copyright:  (c) Martin Stover
Author:  Martin Stover <mstover@freeway.de>
Version:  0.99.pl4-3

mdomo Majordomo

This is your majordomo. Provides an easy to configure means of maintaining mailing lists including user-administration. Contents: majordomo-1.94.4

Copyright:  (c) 1992-1995 D. Brent Chapman, John P. Rouillard

Author:

D. Brent Chapman <Brent@GreatCircle.COM>, Alan Millar <amillar@bolis.sfbay.org>, Paul Haas <paulh@hamjudo.mi.org>, Jon Luini <jon@luini.com>, John P. Rouillard <rouilj@dstar.iddis.com>

Version:  1.94.4-6

metamail MIME Mail-Handler

Metamail is required for reading multimedia mail messages (such as those using the Andrew toolkit) with elm. Contents: metamail-2.7-19

Copyright:  (c) 1991 Bell Communications Research, Inc.
Version:  2.7.19-8

mgetty mgetty

This package turns your computer into a fax machine. With a Zyxel modem, you can even use your computer as an answering machine. mgetty recognizes what kind of call it is receiving and does everything else automatically. You will find the documentation under /usr/doc/mgetty and TeX-Info files under /usr/info. The configuration files are in /usr/lib/mgetty+sendfax and /etc/mgetty+sendfax. Contents: mgetty-1.1.8

Copyright:  (c) 1993-1997 Gert D”oring
Author:  Gert D”oring <gert@greenie.muc.de>
Version:  1.1.8-11

mh Unix Mail Handler

Version:  6.8.4debian8-10

minicom minicom

Terminal program similar to Telix(tm) (program for calling other computers via modem) under MS-DOS. Contents: minicom-1.75

Copyright:  (c) 1991-1996 Miquel van Smoorenburg
Author:  Miquel van Smoorenburg <miquels@drinkel.ow.nl>
Version:  1.75-8

mininews Inews - post news from an NNTP client

Copyright:  (c) 1983-1990 The Regents of the University of California
Author:  Rich Salz <rsalz@osf.org>
Version:  1.25-1

mirror Perl scripts for mirroring ftp servers.

To use mirror, you must have done the following once: cd /usr/include; h2ph * sys/* linux/* Documentation: ”man mirror” and ”man mm”. Contents: mirror-2.8.f4

Copyright:  (c) 1993 Lee McLoughlin
Author:  Lee McLoughlin <lmjm@doc.ic.ac.uk>
Version:  2.8.f4-8

mrtg MRTG - The Mutli Router Traffic Grapher

MRTG - The Mutli Router Traffic Grapher A tool to visualise network traffic via a WebPage.

Contents: gd1.2 mrtg-2.5.2

Copyright:  (c) 1996 Tobias Oetiker and all the Contributers to MRTG
Author:  Tobias Oetiker <oetiker@ee.ethz.ch> , Dave Rand <dlr@bungi.com>
Version:  2.5.2-1

mutt Mail program

Copyright:  (c) 1996 Michael R. Elkins
Author:  Michael Elkins <me@cs.hmc.edu>
Version:  0.89.1-2

ncftp A comfortable ftp program

Author:  Mike Gleason <mgleason@cse.unl.edu>
Version:  2.4.2-8

ncpfs Tools for accessing Novell filesystems

You can mount Netware-server filesystems under Linux using this little tool. You can even print Linux documents on printers attached to the Novell server. Contents: ncpfs-2.0.10

Copyright:  (c) 1995 by Volker Lendecke
Author:  Volker Lendecke <lendecke@namu01.gwdg.de>
Version:  2.0.10-8

netatalk Netatalk - Appletalk for Linux

Copyright:  (c) 1990,1996 Regents of The University of Michigan
Author:  Wesley Craig <netatalk@umich.edu>, Ann Arbor <netatalk@umich.edu>
Version:  1.4b2-8

netcat a simple but powerful network tool

Netcat is a simple Unix utility which reads and writes data across network connections, using TCP or UDP protocol. It is designed to be a reliable ”back-end” tool that can be used directly or easily driven by other programs and scripts. At the same time, it is a feature-rich network debugging and exploration tool, since it can create almost any kind of connection you would need and has several interesting built-in capabilities. You can find the documentation in /usr/doc/packages/netcat/README. Contents: netcat-1.10

Copyright:  (c) hobbit@avian.org
Author:  hobbit@avian.org
Version:  1.10-0

nn The nn news reader

This is the nn news reader (NetNews). It is configured for use as NNTP news reader. Contents: nn-6.5.1

Copyright:  (c) 1990 Kim F. Storm
Author:  Kim F. Storm <storm@texas.dk>
Version:  6.5.1-9

nn spool The nn news reader

This is the nn news reader (NetNews). It is configured to read the local /var/spool/mail directory.

Copyright:  (c) 1990 Kim F. Storm
Author:  Kim F. Storm <storm@texas.dk>
Version:  6.5.1-5

nntpd NNTP server daemon.

This NNTP server is necessary if you plan reading news with the cnews package. It has been configured so that it accesses the /var/spool/news directory to supply news to the news reader. If you want to compile the sources yourself please keep in mind to install and compile cnews first. Please note, that this packages requires cnews. The packages cnews is not supported by our support team. Feedback and bug fixes for nntpd are of course welcome at feedback@suse.de.

Copyright:  (c) 1987 Regents of the University of California, 1989 Stan Barber, 1992 by Iain Lea

Author:

Brian Kantor <brian@ucsd.edu>, Stan Barber <nntp@academ.com>, Iain Lea <iain@ecrc.de>

Version:  1.5.12.1-5

pine The mail program Pine

A menu-driven user mail program. Originally based on Elm (Pine Is No-longer Elm), but with many additional features that make it easier to use. Contents: pine3.91 pine3.96

Copyright:  (c) 1986-94 University of Washington

Author:

Mike  Seibel  <mikes@cac.washington.edu>,  Steve  Hubert

<hubert@cac.washington.edu>, Laurence Lundblade <lgl@cac.washington.edu>

Version:  3.96-10

plp Print spooler

A newer printer spooling system for Linux. Configuration is located in /etc/printcap, /etc/plp.conf, and /etc/printer perms. If problems occur shutdown the spooling system by entering ”/sbin/init.d/lpd stop”. Then you can switch to the BSD spooling system by removing package plp in YaST and installing package lprold. Restart the system with ”/sbin/init.d/lpd start”. Contents: plp-4.0.3

Version:  4.0.3-8

pop POP Server/Client

This package contains (among others) fetchmail. Contents: fetchmail-4.3.8 imap-4.1.BETA pop3d-1.00.4 popclient-3.0b6 popsrc qpopper2.2 upop

Copyright:

(c) 1991 Regents of the University of California, 1993-95 by Carl Harris, Jr., 1991-2, RSA Data Security, Inc., 1989 Mark Davies, 1990 Andy Linton, 1990 Victoria University of Wellington

Author:

Katie Stevens <dkstevens@ucdavis.edu>, Carl Harris <ceharris@mal.com>, Mark Erikson <mark@qualcomm.com>, <msharov@bev.net>

Version:  97.6.29-11

ppp PPP (”Point to Point Protocol”) for Linux

PPP kernel driver and the pppd daemon. An alternative to SLIP. Contents: ppp-2.2.0f

Author:

Paul  Mackerras  <paul.mackerras@cs.anu.edu.au>,  Michael  Callahan

<callahan@maths.ox.ac.uk>, Al Longyear <longyear@netcom.com>, Brad Parker <brad@faster.fcr.com>, Greg Christy <gmc@premises1.Qutoron.COM

Version:  2.2.0f-9

procmail Local email delivery

Sendmail calls procmail to deliver email into a local folder. Procmail can be configured to store email into different folders. Contents: procmail-3.11pre7

Author:  Stephen R. van den Berg <srb@cuci.nl>
Version:  3.11pre7-8

radiusc Radius client-software

Radiusclient provides some useful programs for doing authentication and accounting via a (remote) radius-server. Contents: radiusclient-0.3-alpha-4

Copyright:  (c) see /usr/doc/packages/radiusc/COPYRIGHT
Author:  Lars Fenneberg <lf@elemental.net>
Version:  0.3alpha4-2

radiusd Radius-Server

radiusd is the Livingston radius-server. It provides remote authentication an accounting services; it e.g. use bei Livingston Portmaster products (you guessed) for login-authentication and logging. Contents: radius-1.16.1

Copyright:  (c) Copyright 1992 Livingston Enterprises, Inc.
Author:  Livingston Enterprises, Inc.
Version:  1.16.1-1

rdist ”Remote file distribution” utility

This utility makes it possible to distribute files to several computers within a network. If possible, file attributes such as: owner, group, mode, and mtime are retained. It’s even possible to overwrite running binaries. Contents: rdist-6.1.3

Copyright:  (c) 1993 Michael A. Cooper
Author:  Michael A. Cooper <Michael.Cooper@Sun.Com>
Version:  6.1.3-8

rsync replacement for rcp/mirror that has many more features

rsync uses the ”rsync algorithm” which provides a very fast method for bringing remote files into sync. It does this by sending just the differences in the files across the link, without requiring that both sets of files are present at one of the ends of the link beforehand. At first glance this may seem impossible because the calculation of diffs between two files normally requires local access to both files. A technical report describing the rsync algorithm is included with this package. Contents: rsync-1.6.3

Copyright:  (c) Andrew Tridgell and Paul Mackerras (GPL)
Author:  Andrew.Tridgell@anu.edu.au, paulus@cs.anu.edu.au
Version:  1.6.3-5

rzsz X-, Y- and Z-Modem data transfer protocols.

Allows you to Use ”sz filename” to send a file to your local system. Contents: lrzsz-0.12b

Version:  0.12b-8

samba Samba is a file server for Unix, similar to LanManager.

Samba is a suite of programs which work together to allow clients to access Unix filespace and printers via the SMB protocol (Seerver Message Block). CAUTION: The samba daemons are started by the init script /sbin/init.d/samba, not by inetd. The entries for /usr/sbin/smbd and /usr/sbin/nmbd must be commented out in /etc/inetd.conf. In practice, this means that you can redirect disks and printers to Unix disks and printers from LAN Manager clients, Windows for Workgroups 3.11 clients, Windows’95 clients, Windows NT clients and OS/2 clients. There is also a Unix client program supplied as part of the suite which allows Unix users to use an ftp-like interface to access filespace and printers on any other SMB server. Samba includes the following programs (in summary):

* smbd, the SMB server. This handles actual connections from clients.

* nmbd, the Netbios name server, which helps clients locate servers.

* smbclient, the Unix-hosted client program.

* smbrun, a little ’glue’ program to help the server run external programs.

* testprns, a program to test server access to printers.

* testparm, a program to test the Samba configuration file for correctness.

* smb.conf, the Samba configuration file.

* smbprint, a sample script to allow a Unix host to use smbclient to print to an SMB server. The suite is supplied with full source and is GPLed. This package expects its config file under /etc/smb.conf . Documentation: /usr/doc/packages/samba Contents: samba-1.9.18p3 smbfs-2.0.2

Copyright:  (c) 1992-95 Andrew Tridgell, Karl Auer, Jeremy Allison

Author:

Andrew  Tridgell  <Andrew.Tridgell@anu.edu.au>,  Karl  Auer

<Karl.Auer@anu.edu.au>, Jeremy Allison <jeremy@netcom.com>

Version:  1.9.18p3-1

sendfax sendfax part of mgetty

The sendfax part of mgetty. The sources are included in mgetty source package.

Copyright:  (c) 1993-1997 Gert D”oring
Author:  Gert D”oring <gert@greenie.muc.de>
Version:  1.1.8-11

sendmail BSD sendmail

The ”Unix System Administration Handbook” calls sendmail ”The most complex and complete mail delivery system in common use...” . Ready-made configuration files are included for systems connected by TCP/IP (with or without a name server) and for systems using UUCP. ’procmail’ is included as a local mail agent. Contents: check sendmail-8.8.8

Copyright:  (c) 1983 Eric P. Allman
Author:  Eric Allman <sendmail@sendmail.ORG>
Version:  8.8.8-10

seyon telecommunications under X

The best telecommunications package for X available at present; it has separate windows for control, terminal, telephone register, etc. Seyon supports several protocols (Zmodem, Kermit, etc.) and auto download (only Zmodem). Using the built-in script language, automatic login or up/downloading is no longer a problem. The actual transmission is provided by the several programs like sz/rz/kermit. Since we are not allowed to supply kermit with our distribution you have to get it yourself from the Internet (ftp: kermit.columbia.edu). Please look at the various sites as sunsite.unc.edu or any server which mirrors sunsite (ftp.gwdg.de under /pub/Linux/X11/xapps/comm or /pub/Linux/apps/comm). You have to configure ˜ /.seyon/protocol to use it. Contents: seyon-2.14c seyon-kermit-1.0

Copyright:  (c) 1992 Muhammad M. Saggaf
Author:  Muhammad M. Saggaf <alsaggaf@mit.edu>
Version:  2.14c-7

slip dial out/in programs

This package include the two dial out/in programs dip and sliplogin. Users who want to use dip, have to be in group ”dialout”. This is also necessary for ”diplogins”. Contents: dip-3.3.7p sliplogin-2.1.0

Copyright:

(c) 1988-1993 MicroWalt Corporation, 1991-1992 Ian Lance Taylor, 1990 The Regents of the University of California

Author:

Fred N. van Kempen <waltje@uwalt.nl.mugnet.org>, Sven Goldt <goldt@peagsus.inberlin.de>

Version:  97.10.27-4

slrn News reader.

Version:  0.9.4.3-5

smail The mail system ”Smail”.

Smail is a mail system which is easy to configure. A lot of examples are installed in addition to the program. Smail should NOT be used together with sendmail. Contents: smail-3.2 Please note, that this package is not supported by our support team. Please use sendmail if you need support for mail delivery. Feedback and bug fixes for smail are of course welcome at feedback@suse.de.

Author:  Ian Kluft (Portierung)
Version:  3.2-12

snmp CMU SNMP

SNMP is the Simple Network Management Protocol of the Internet defined as a standard in RFC 1155 and 1157 and 1901 up to 1908. The protocol allows mainly to retrieve and set variables, addressed as objects in the MIB (Management Information Base). The well-known MIB-2 is defined in RFC 1213 and contains variables related to the TCP/IP protocol suite. Contents: cmu-snmp-linux-3.2

Copyright:  (c) Copyright 1988, 1989 by Carnegie Mellon University
Author:  Erik Schoenfelder <schoenfr@gaertner.de>
Version:  3.2-8

squid squid: proxy server for WWW

Copyright:  (c) GPL
Version:  1.NOVM.20-1

strn News reader.

This version of strn is based on trn and allows scoring of a news article based on author and subject. Contents: strn

Version:  092q3-8

suck Reading News Offline

Author:

Tim Smith, Sven Goldt <goldt@math.tu-berlin.de>, Robert A. Yetman <boby@pixi.com>

Version:  3.9.1-1

susefax A fax client for the HylaFAX fax server

This is a fax-client for the HylaFax(tm) fax-server maintained and programmed by Sam Leffler (SGI). It is totally implemented in Java and requires the Java Developer Kit (JDK) at least version 1.1.1 (Package java, series d). It supports:

* integrated phone book

* sending faxes directly from the phone book

* setting up job-parameters before sending a job and during processing of the job in the job-queue

* removing jobs from the job-queue

* direct viewing of received faxes (needs an external tiffg3 viewer or converter, for Linux/UNIX the converting is done by fax2ps from libtiff and viewing by ghostview or gv automatically by a script that comes with this package) * creating of serial-fax lists out of the phone book entries

* automatic cover page generation (not very fast :-( )

* auto-spooling (that’s important for WINDOWS users) YES!!! This program is also running on WINDOWS and hopefully all the other platforms, that support Java by the JDK from Sun!!! It has been tested on WindowsNT 4.0 and it seems to run!!!!!

* password prompting if the administrator of the fax server wants the users to access the server via password Contents: SuSEFax

Copyright:  (c) 1997, 1998 S.u.S.E. GmbH, Fuerth
Author:  Carsten Hoeger <Carsten.Hoeger@suse.de>
Version:  0.83v2-0

suseppp SuSE PPP

Internet Provider Database including script to configure a PPP conection Contents: suseppp

Author:  Reinhold Sojer <reinhold@suse.de>
Version:  1-1

tcpdump packet sniffer

With this program you can ”read” all or only certain packets going over the ethernet. It can be used to debug specific network problems. Contents: tcpdump-3.4a6

Version:  3.4a6-1

tin The ”tin” News Reader

Tin is a full-screen, easy to use Netnews reader. It can read news locally (i.e. /var/spool/news) or remotely (rtin or tin -r option) via an NNTP (Network News Transport Protocol) server. Docu: man tin or man rtin. Contents: tin-980226

Copyright:  (c) 1991-93 Iain Lea
Author:  Iain Lea <iain@ecrc.de>
Version:  980226-1

trn News reader (conf. for NNTP)

This trn is compiled so it needs the Network News Transfer Protocol Daemon (nntpd) to get its news.

Contents: trn-3.6

Copyright:  (c) 1991-92 Stan Barber

Author:

Wayne Davison <davison@borland.com>, Larry Wall <lwall@jpldevvax.jpl.nasa.gov>, Stan Barber <sob@bcm.tmc.edu>

Version:  3.6-9

trn spl Threaded News Reader (conf. for spool directory)

Compiled for reading news from the local spool-directory which resides under /var/spool/news. If you would like another configuration you have to recompile trn. This shouldn’t be a problem as trn is effortlessly compilable under Linux.

Copyright:  (c) 1991-92 Stan Barber

Author:

Wayne Davison <davison@borland.com>, Larry Wall <lwall@jpldevvax.jpl.nasa.gov>, Stan Barber <sob@bcm.tmc.edu>

Version:  3.6-6

url get Fetch data from their URL-addresses.

With url get you may get data from their URL-address. HTML, Gopher and FTP are supported. For further documentation see /usr/doc/packages/url get.

Contents: url get

Author:  Jack Lund <j.lund@cc.utexas.edu>
Version:  98.2.13-2

uucp Taylor UUCP

Ian Taylor’s Unix to Unix copy: mail and news over modem lines. This is the standard UUCP package of the Free Software Foundation. Configured to use HoneyDanBer and/or Taylor configuration files. Remark: We did not include an uucp guest account. So, if you want to create a guest account please make sure the directory /var/spool/uucppublic exists.

Contents: uucp-1.06.1

Copyright:  (c) 1991-93 Ian Lance Taylor
Author:  Ian Lance Taylor <ian@cygnus.com>
Version:  1.06.1-8

uucpcfg Taylor UUCP - sample configuration

Sample configuration files for Taylor UUCP.

Contents: uucpcfg

Version:  98.2.3-2

vacation Automatic reply to incoming mails

This program answers when you’re lying on the beach ;-) Documentation: man vacation

Contents: vacation-1.1.1

Copyright:  (c) 1983, 1987 University of California, 1993 Harald Milz
Author:  Eric P. Allman, Harald Milz <hm@seneca.muc.de>
Version:  1.1.1-8

vbox vbox; an ISDN answering machine for Linux

Package for realizing an answering machine under Linux using its ISDN capabilities. Please keep in mind that you need a kernel with CONFIG ISDN AUDIO support. Of course the ISDN package (i4l) has to be installed as well. Please have a look at /usr/doc/packages/vbox Contents: vbox-1.1 Please be aware that for running this package you need a kernel with ISDN AUDIO SUPPORT compiled in.

Copyright:

(c) Copyright: Michael Herold, Chris Benndorf, Jan Schoenepauck, Joachim Gassen, Marc Eberhard

Author:

Michael  Herold  <michael@abadonna.franken.de>,  Chris  Benndorf

<chrisb@ganzfix.duew.eunet.de>, Jan Schoenepauck <schoenep@wrcs3.urz.uniwuppertal.de>, Joachim Gassen <gassen@uni-muenster.de>, Marc Eberhard <Marc.Eberhard@Uni-Duesseldorf.DE>

Version:  1.1-11

vbox2b4 Vbox 2b4 An ISDN answering machine

Vbox is an ISDN answering machine. Besides recording incoming calls it offers a wide variety of features. For configuring vbox please have a look at: /usr/doc/packages/vbox2b4 and check an adapt the shipped configuration files under: /etc/isdn to your needs. Contents: vbox2b4 This package is still considered beta, that’s why this package is not yet included in the installation support.

Copyright:  (c) Michael Herold <michael@abadonna.mayn.de>

Author:

Michael  Herold  <michael@abadonna.mayn.de>,  Fritz  Elfert

<fritz@wuemaus.franken.de>, Carsten Keil <keil@temic-ech.spacenet.de>, Andreas Kool <akool@Kool.f.EUnet.de>, Stefan L”uthje <luethje@sl-gw.lake.de>, Marc Eberhard <Marc.Eberhard@Uni-Duesseldorf.DE>, Gert Doering <gert@greenie.muc.de>

Version:  2b4-5

w3mir an administration tool

The idea is to get only files in the hierarchy that has changed since the last time the file was copied. This is a Good Idea. This saves considerable time and bandwidht. Contents: MD5-1.7 MIME-Base64-2.04 libnet-1.0605 libwww-perl-5.18 w3mir-0.99b2

Version:  0.99b2-2

webalize tool for webserver logfile analysis

The Webalizer is a web server log file analysis program which produces usage statistics in HTML format for viewing with a browser. The results are presented in both columnar and graphical format, which facilitates interpretation. Yearly, monthly, daily and hourly usage statistics are presented, along with the ability to display usage by site, URL, referrer, user agent (browser) and country (user agent and referrer are only available if your web server procduces Combined log format files).

Author:  Bradford L. Barrett <brad@usagl.net>
Version:  0.98.16-2

weblint a syntax and minimal style checker for HTML

Weblint is a syntax and minimal style checker for HTML: a perl script which picks fluff off html pages, much in the same way traditional lint picks fluff off C programms. Contents: weblint-1.020

Author:  Neil Bowers <neilb@cre.canon.co.uk>
Version:  1.020-2

wget C tool for mirroring FTP and HTTP servers

Wget enables you to retrieve WWW documents or FTP files from a server. This might be done in script files or via command line. Contents: wget-1.4.5

Copyright:  (c) GPL
Version:  1.4.5-9

whois whois client program (RIPE version)

This is intended for European users as it will query a RIPE whois server with extended whois capabilities rather than a standard whois protocol server. The whois program is fully compatible with the standard whois protocol and can thus replace your default whois program. Contents: ripe-whois-tools-2.2

Copyright:  (c) 1980 Regents of the University of California
Author:  <marten@ripe.net>
Version:  2.2-8

wuftpd wuftpd

This FTP daemon is a little different. It replaces the standard-in.ftpd included in series n. It is more flexible but more complicated to configure as well. The following manual pages are included too: ftpd (the programm) ftpshut (service-program) ftpcount (service-program) ftpwho (service-program) ftpaccess (explanation of /etc/ftpaccess) ftphosts (explanation of /etc/ftpaccess) xferlog (explanation of /var/log/xferlog) ftpconversions (explanation of /etc/ftpconversions) Furthermore there are some sample configurations which will be placed into /usr/doc/packages/wuftpd during installation. If you want to use it you have to replace the following line in /etc/services: ftp stream tcp nowait root /usr/sbin/tcpd in.ftpd with: ftp stream tcp nowait root /usr/sbin/tcpd wu.ftpd Then you have to send a HUP to inetd (kill(8)) or restart. Contents: wu-ftpd-2.4.2-beta-15

Copyright:  (c) 1994 Washington University in Saint Louis.
Author:  Anonymous (Internet, Washington University in Saint Louis)
Version:  2.4.2beta15-8

wwwoffle WWWOFFLE - World Wide Web Offline Explorer

A proxy HTTP server for computers with dial-up internet access.

- Cacheing of pages viewed while connected for review later.

- Browsing of cached pages while not connected, with the ability to follow links and mark other pages for download.

- Downloading of specified pages non-interactively.

- Multiple indices of pages stored in cache for easy selection.

- Interactive or command line option to select pages for fetching individually or recursively.

- All options controlled using a simple configuration file with a web page to edit it.

Copyright:  (c) Andrew M. Bishop 1996,97

Author:

Andrew M. Bishop <amb@gedanken.demon.co.uk>, Yannick Ver-

sley  <sa6z225@public.uni-hamburg.de>,  Axel  Rasmus  Wien-
berg  <2wienbe@informatik.uni-hamburg.de>,  Andreas  Dietrich

<quasi@baccus.franken.de>, Ullrich von Bassewitz <uz@wuschel.ibb.schwaben.com>

Version:  2.0-1

xinetd Another ’inetd’ with expanded functionality

xinetd takes the functionality of inetd and appends additional opportunities:

- Access Control

- Prevention of ’denial-of-access’ attacks

- Extensive logging abilities

- clear konfiguration file Contents: xinetd-2.2.1 Note that you have to set ’START INETD’ to ’no’ if you want to use xinetd.

Copyright:  (c) 1992 by Panagiotis Tsirigotis, 1995 by Charles Murcko

Author:

Panagiotis Tsirigotis <panos@cs.colorado.edu>, Chuck Murcko <chuck@topsail.org>, Jan Wedekind <jw@ping.de>, <jw@Germany.EU.net>

Version:  2.2.1-6

xntp Network Time Protocol daemon

xntpd is a daemon which sets and maintains a Unix system time-of-day in compliance with Internet standard time servers. xntpd is a complete implementation of the Net- work Time Protocol (NTP) version 3 standard, as defined by RFC 1305, but also retains compatibility with version 1 and 2 servers as defined by RFC 1059 and RFC 1119, respec- tively. Contents: xntp3-5.92

Author:  Dennis Ferguson, David Mills
Version:  3.5.92-0

xrn X11 news reader

Version:  8.01-8

ypclient YP - (NIS-) client

Use your Linux-PC as a NIS client (former YP-client) Included in this package are: ypbind, ypmatch, yppoll, yppasswd, ypset and ypwhich. You can configure a server or use a broadcast using ypbind and adapting /etc/yp.conf. Contents: ypbind-bsd-1.15-970316 ypclients-2.2-970318

Copyright:  (c) 1995 Theo de Raadt, Swen Th”ummler, Michael Rausch

Author:

Theo de Raadt <deraadt@fsa.ca>, Swen Th”ummler <swen@uni-paderborn.de>, Michael Rausch <rausch@julia.de>

Version:  2.2.970318-9

ypserv YP - (NIS)-server

Configure your Linux PC as YP-(NIS)-server using this package. Contents: ypserv-1.1.3

Copyright:

(c) 1989, 1991, Free Software Foundation, Inc., 1996, Florian La Roche, S.u.S.E. GmbH Fuerth, Peter Eriksson, Thorsten Kukuk, Swen Thuemmler, 1993,1994 Signum Support AB, Sweden, et.al.

Author:

Florian La Roche <florian@suse.de>, Peter Eriksson <pen@signum.se>, Thorsten Kukuk <kukuk@uni-paderborn.de>, Swen Thuemmler <swen@uni-paderborn.de>

Version:  1.2.4-9

0.1.11 PAY: Commercial Software

abuse Abuse - Arcade game

Abuse is an old style arcade game: find your way through the labyrinth, shoot every alien, and collect some items. To move the fighter, use the cursor keys. The mouse controls aim and shoot. This version contains only a demo level. However, a level editor is included. Abuse runs under X11 as well as on the console. Contents: abuse Please remember that this program is SHAREWARE.

Copyright:  (c) 1995 Crack dot Com
Author:  <abuse@crack.com>, <jc@crack.com>
Version:  1.10-3

ac3d dem Autocad look-alike graphic program

Ac3d is a graphic program which looks like autocad. Start the program with the command ac3d. Under /usr/doc/packages/ac3d/ you will find some demos in the directory acfiles/ and the documentation in html format in the directory manual/ . To read the documentation, start your WWW browser with the argument file://usr/doc/packages/ac3d/manual/ Please note the message in the dialog window at startup: this program is SHAREWARE and must be registered if you wish to use it. Contents: ac3d dem

Copyright:  (c) 1996 Andy Colebourne
Author:  Andy Colebourne <andy@comp.lancs.ac.uk>
Version:  1.56-3

PAY: Commercial Software

acroread Acrobat Reader for PDF files

Copyright:  (c) 1997 Adobe Systems Inc.
Version:  3.01-4

adabas Adabas D Personal Edition

This package contains Adabas D Personal Edition. Log on as user ”adabas”, start the X Window System and run the program ”panel”. This personal edition has the following restrictions:

- max. 24 MB data - max. 3 concurrent users Contents: adabas-10.0.1

Copyright:  (c) 1997 by Software AG
Version:  10.0.1-1

adadoc Adabas D English documentation

This package contains the English documentation for Adabas D personal edition.

Copyright:  (c) 1997 by Software AG
Version:  10.0.1-1

adadok Adabas D German documentation

This package contains the German documentation for Adabas D personal edition.

Copyright:  (c) 1997 by Software AG
Version:  10.0.1-1

adbsmydb MYDB demo database

This package contains the MYDB demo database for Adabas D personal edition. Start the program ”panel” as user ”adabas”. Now log on as ”control” with the password ”control” and enter ”MYDB” as the database.

Copyright:  (c) 1997 by Software AG
Version:  10.0.1-1

arcad ARCAD CAAD demo

Have a look at http://www.arcad.de Contents: arcad-0.068 This is a demo of the ARCAD CAAD System of ARCAD Systemhaus.

Copyright:  (c) ARCAD Systemhaus
Author:  Bernd Metzger <bernd@arcad.ic-marl.de>
Version:  0.068-2

arkeia ARKEIA network backup software

As root set the variable START ARKEIA in /etc/rc.config to yes and start the backup server: /sbin/init.d/arkeia start ARKEIA utilizes an exclusive multi-flow technology to deliver speeds that are 200 to 300% faster than rival software packages. Its unique transaction engine allows multiple backup and restore operations to be performed simutaneously with total reliabilty. This is especially valuable for multiple users backing up or restoring data on their own. (a multiple drive library is required). ARKEIA’s Java interface enables the system adminsitrator to manage multiple remote backup servers through internet as if they were local backups. ARKEIA is the most reliable software you can find. It has a powerful database capable of restoring sites to their original state, in the event of server disasters. ARKEIA integrates with all types of corporate environments and is available for UNIX or Windows NT servers. ARKEIA is the first software delivered with a JAVA interface, effectively neutralizing the problem of heterogeneous computer environments. Operators can use any Intranet station to access any remote network. Its universal interface makes it the first software able to backup via the Internet.

ARKEIA is compatible with a wide range of devices and provides the best ”plug and play” solution for effective data security. Contents: arkeia-4.0

Copyright:  (c) Knox Software Comm
Version:  4.0-1

bartels The Bartels AutoEngineer

Bartels AutoEngineer (BAE) is a fully integrated CAE/CAD/CAM system for the design of printed circuit boards. The system is based on the famous Bartels Autorouter which has been incorporated in most leading PCB layout systems throughout the world, setting new industrial standards of autorouting success and dramatically reducing the requirement for manual routing. Contents: bartels-4.0.beta

Copyright:  (c) Oliver Bartels F+E
Version:  4.0.beta-2

c16 demo CONZEPT 16 demo

CONZEPT 16 - Demo version CONZEPT 16 is a program development tool based on a relational database, which can be used under a variety of operating systems, on many different computers, and in every PC-networking environment. CONZEPT 16 is completely developed in Germany and is oriented to the latest generation of microcomputers. To exploit the capabilities of each PC generation, there are special versions adapted to the corresponding CPU (8086/88, 80286, 80386, 80486, and Pentium). CONZEPT 16 is not only an ideal development tool for programmers wanting to develop applications on a relational database platform, but also for those users who want to configure easily their personal environment. You can easily generate structures, masks,and lists using the menu choices, which means that you don’t have to learn a new programming language. CONZEPT 16 is presently available for MS-DOS, Novell Netware, UNIX System V/386, Solaris, OSF/1 and all PC networks which support DOS version 3.1 or higher (LAN-server, Netware-Lite, Network OS etc.). CONZEPT 16 has multiuser and multi-platform capability. Since data security is most important in this context, CONZEPT 16 uses automatic as well as user-driven transactions. For maximum process parallelization, CONZEPT 16 uses ”two-phase-locking”, not only on Novell but on all other networks, including OS/2 and UNIX. Give it a try and start the demo as root from text console: usr/lib/c16 demo/startdemo Contents: c16 demo

Copyright:  (c) vectorsoft GmbH
Version:  4.0.14-3

crisp A programmable editor

CRISP is an editor for programmers which has many improved features. Wherever possible, CRISP has no internal restrictions regarding the number of files that can be edited simultaneously, the maximum line length, etc. CRISP is only constrained by the main memory and the hard disk capacity of your system. As a result, CRISP combines some of the best features of other available editors such as vi, emacs, BRIEF. Please note that the program is SHAREWARE. Further information can be found in /usr/doc/packages/crisp . Contents: crisp

Copyright:  (c) 1996 Foxtrot Systems Ltd; Paul David Fox
Author:  Paul Fox fox@crisp.demon.co.uk
Version:  5.1.1b-4

cysched Synchronize server software and Cyberscheduler software

Edit /etc/rc.config and set START SYNCHRONIZE to ”yes”, then run ‘/sbin/init.d/synchro start‘ as root Go to the URL http://localhost/synchro/En US/login.html and log-on as ”demo”.

Product information: Industrial strength calendaring and scheduling for your changing enterprise. Synchronize is the scheduling solution that fits both your end-user and your IT requirements! Since 1990, enterprises of all sizes have relied on Synchronize for scalable, enterprise-wide, efficient real-time calendaring and scheduling. This proven client/server software supports enterprise-wide coordination and collaboration among team members, whether they’re located on different floors, multiple buildings, across the country, or around the world, bringing the enterprise closer together. The coordination and collaboration features designed into the Synchronize solution work together to enhance decision-making and communication. This helps your users increase their productivity, enhance teamwork, simplify collaborative efforts, and minimize project cycle times. Get the Synchronize advantage Designed specifically for enterprise scheduling, only Synchronize gives you all these important advantages: No hidden costs. Synchronize is framework independent, which means it does not require a proprietary messaging infrastructure and operates in mixed application or mixed framework environments. The broadest cross-platform support. No other calendaring and scheduling software supports as many different desktop environments and servers. Synchronize includes Web and native support for all these platformsWindows, Macintosh, UNIX/X and Character, in addition to the most complete UNIX and NT server line-up. The ability to support mixed Web and native desktops enables you to roll out your traditional native client and/or Intranet solution on the platforms you want and in the time frame that works best for your organization. Excellent technical support. Our Technical Support Department is staffed by a team of knowledgeable engineers trained in supporting cross-platform environments. CrossWind’s customers enjoy a rapid response with personal attention to each and every support question. Seamless UNIX & NT integration. Our servers are offered on both UNIX and NT platforms supporting seamless coexistence, or future integration, or migration plans. Scalability. A distributed database architecture lets you quickly and transparently add as many new users and user groups as you need. You can administer calendaring centrally or at multiple locations. No-hassle upgrades. Synchronize is the only enterprise calendaring and scheduling product that offers true backward compatibility. Our commitment to cross-release compatibility makes the upgrade process straightforward and discretionarynot an enterprise-wide crisis. Maximizing the productivity of every resource A powerful suite of real-time group calendaring and scheduling applications, Synchronize contains important coordination tools that save your organization time in scheduling meetings and company resources. Users can distribute agendas and memos, schedule meetings, allocate resources, assign and track tasks, and send remindersno matter what platform they’re running. Work teams in multi-vendor computing environments can achieve consensus on both short-term and ongoing issues through the use of shared notes in private or public forums … easily and transparently. Synchronize enables group and personal task management as well as flexible resource management capabilities. Users have access to individual and group calendars, shared notes, and updated to-do lists. Support for the Internet, printers and portable computers makes it easy to access or take your schedule and project notes with you wherever you go. Built for performance and simple administration Synchronize is based on a robust client/server architecture using multi-threaded servers. All schedules, meetings, deadlines, reserved resources, project notes, messages, and work assignments are maintained in a common database stored in the Synchronize server, enabling quick and easy sharing of vital data. Administrators can easily distribute Synchronize’s common database into many networked databases that take advantage of existing servers and networks. The ability to communicate directly across TCP/IP provides real-time access to critical information without the delays typically associated with file-based access methods, e-mail based scheduling, and slow transport protocols. Advanced architecture makes Synchronize highly scalable: It can be used for enterprise collaboration on LANs, WANs and corporate INTRANETs in organizations ranging from 5 to 50,000 users (or more). Synchronize’s distributed database architecture allows administrators to add new users easily and transparently, with no interruption in service. CrossWind is dedicated to building solutions based on industry standards. As a result Synchronize’s interoperability with Web servers and applications is simple and straightforward. As a member of the Internet Engineering Task Force working group, responsible for establishing Internet standards for calendaring and scheduling, CrossWind continues its commitment for interoperability through standards support. Technical Information Designed for cross-platform deployments, Synchronize runs on most all commercial UNIX and NT servers, with desktop access through Web browsers as well as native software clients for Microsoft Windows, Macintosh, X.11/Motif and ASCII desktops. Affordable with No Hidden Costs Although some competitive products appear to have a very low per-user pricea good look at the total cost of ownership will reveal that a substantial, continuing investment will be required to assure smooth operations for these products. Synchronize is easy to deploy and requires no additional expenses for integration into an enterprise’s existing environment. Ready ”off the shelf,” Synchronize does not require programming, extra database servers, or email/messaging platforms. All Synchronize clients and servers are available (at no charge) for unrestricted download from our ftp site 24 hours a day. You can reconfigure, migrate and ”webify” your system at anytime at no charge. Since upgrades are backward compatible, they can be done (or not) on a workgroup or departmental basis. Your organization will not be compelled to buy into new releases under the duress of lost support or lack of compatibility. Contents: cysched-2.1.0

Copyright:  (c) Crosswind Technologies
Version:  2.1.0-2

eloquen HP Eloquence (B1368B)

HP ELOQUENCE Product Description: HP Eloquence is a integrated environment for development and execution of commercial applications. It is easy to use and ideally suited for small to midsize companies. Designed to operate in a distributed network environment it supports fast, flexible application development, implementation and maintenance. HP Eloquence has a broad installed base on HP-UX systems worldwide with a large number of ready-to-use commercial applications. HP Eloquence is available for HP-UX, Windows NT and LINUX platforms. The product includes: - State-Of-The-Art Interactive Development Environment on Windows 95 and Windows NT platforms. - Standard, character-oriented Development Environment on HP-UX and LINUX platforms - Powerful Programming Language, a superset of HP Business Basic - Integrated Database Management System (IMAGE compatible) - Runtime Environment for Graphical User Interfaces HP Eloquence is a product of Hewlett-Packard GmbH, Germany. -------------------------------------------------------------------- Licence Information: The HP Eloquence ”Personal Edition” is available for free. It provides a free license for personal or evaluation usage. It is limited to two users and a maximum database size of 50 MB. Be sure to read and agree to the ”Personal Edition License Agreement” before you use the software. -------------------------------------------------------------------- Support Information: We provide 60 days free support via E-mail after registration on our Web-Page http://www.hpeloquence.com/support/register.html Information about commercial support for HP Eloquence are available on: http://www.hp-eloquence.com/info/ -------------------------------------------------------------------- Product and Price Information of the commercial version of HP Eloquence are available on : http://www.hp-eloquence.com/info/order.html If you need further information please contact: Marxmeier Software GmbH Phone: +49 202 243 1440 Fax: +49 202 243 1420 Email: info@hpeloquence.com Contents: eloquen-A.06.01 The license of this demo expires in 30 days!

Copyright:  (c) 1997 Marxmeier Software Entwicklung GmbH
Version:  A.06.01-1

empress Empress database

This package contains the version ”Personal Empress for Linux”, derived from the full product, Empress V6.8, and tailored for use as a standalone single-user Relational Database Management System. See the file /usr/doc/packages/empress/readme for further information. This package is not free software. Use is restricted by the license in /usr/doc/packages/empress/license .

Copyright:  (c) 1983, 1995 Empress Software Inc.
Version:  6.8-6

executor A demo of ARDIs Mac emulator.

A demo version of the Executor. It is limited to 15 minutes runtime per invocation and fully operational after installation (start with executor-demo). There is some Mac shareware included, e.g. a Lemmings demo. The emulator also supports sound. This is a demo version.

Copyright:  (c) 1986-1996 ARDI
Author:  ARDI <questions@ardi.com>
Version:  2.0-1

flagship Clipper compiler (DEMO) FlagShip for Linux

FlagShip is a database development and porting system for porting all applications based on the xBase language to UNIX. Completely source and file compatible to Clipper 5.2 (and previous versions) including macros, code blocks, objects, etc. The motto is ”the same sources and data for both DOS and UNIX”. (quote) You will find the documentation under /usr/doc/packages/flagship. To install, change to the directory /usr/lib/FlagShip and start the program FSinstCD. After installation the directory /usr/lib/FlagShip can be deleted. The full version is also included, you only need to order a key from S.u.S.E.

Copyright:  (c) 1996 Multisoft Datentechnik GmbH
Author:  multisoft Datentechnik GmbH <info@fship.com>
Version:  4.41-1

graffy graffyDOC DEMO - a package for technical design / CAD

Info: /usr/doc/packages/graffy Start: ’/usr/bin/graffy’

Copyright:  (c) DURST CAD/CONSULTING GmbH
Author:  Guenther <gl@durstcad.s.eunet.de>
Version:  1-5

heitml HeiTML

heitml (pronounced ”Hi”-TML) significantly extends and enhances the functionality of HTML by definable tags and full programming features. This makes dynamic content and database applications possible simply within the HTML world, without CGI and without external scripting or programming languages. This means you, as an HTML author, can embed applications in your web pages, simply by using some new tags without CGI and without programming. As an advanced user or programmer on the other hand you can create and program powerful tag libraries. This approach makes heitml suitable for HTML newcomers and professional programmers alike. Contents: heitml-1.2 HEITML (CGI) demo version, needs ADABAS D

Copyright:  (c) 1996-1997 H.E.I.
Version:  1.2-9

ivs demo I.V.S demo of a time registration and calculation program

This is a ready to be used demo version of I.V.S. You can calculate and organize projects during a certain period of time. We used to have the problem that by the end of a month we did not have an overview of the ”performance” of our employees. That made it difficult for us to calculate. I.V.S is thought for companies, independent professionals and all others. Input of hours-of-work is achieved as follows: Customer <-> Project <-> Hours-of-work The licensed version offers multiple customers, projects (customer-specific) and tasks. With the supplied demo version you do not have the builtin user-administration; therefore everybody is able to see the data of the others! This isn’t true with the full licensed version.

Copyright:  (c) K + R Informationssysteme Konstanz
Author:  K + R Informationssysteme Konstanz <info@iridium.de>
Version:  9604-1

ixware IXware - Enterprise Solution Server

IXware is a german enterprise solution software. English documentation is not available. Contents: ixware-2.2.2.1 This software is in German only!

Copyright:  (c) Steiner, Sch”afer, R”uhl GbR
Author:  Winfried Sch”afer <wschaefer@s2r-ixware.de>
Version:  2.2.2.1-2

mapledmo Demo Maple V R3

Maple V is a comprehensive problem-solving system that supports a wide variety of powerful mathematical operations such as symbolic analysis, numerical analysis, and graphics. Maple V also features a complete programming language that uses syntax similar to that of BASIC, FORTRAN, Pascal or C. This language is indispensable for those wishing to develop solutions to larger problems. This is a demo version of the well-known mathematical package. Maple V can be found in /usr/local/maple . The binary was compiled using X11R5 and therefore requires the old XFree86[tm] 2.x libraries. If you are interested in a licensed version, please contact S.u.S.E. GmbH. Contents: mapledmo

Copyright:  (c) 1993 Waterloo Maple Software
Author:  <info@maplesoft.on.ca>
Version:  3-3

mars e Nice Editor for console and for X.

This nice Editor works both on the console and under X. It is completely in German language. Please read the file /usr/lib/mars e/README for licensing terms! Start the Editor simply with ”e”!

Copyright:  (c) Martin Stover
Author:  Martin Stover <mstover@Kool.f.EUnet.de>
Version:  4.11b-1

mhitdemo WYSIWYG document editor

Magic HIT is a modern document editor. Each document is based on a document class which determines the layout and structure of its contents.

Additional pluses include: Easy to use, no interruptions in having to deal with formal or layout-related parameters. A modern, graphical user interface concept also makes it possible to integrate Magic HIT into a workflow installation. Magic HIT will be installed in the directory /usr/lib/magicHIT. Technical information is in the file /usr/lib/magicHIT/README. Magic HIT is integrated as an editor tool in the object oriented card and databank system, Xcard. Magic HIT runs in demo mode (no printing or saving). You can activate the full version with a key that you can get for free from Interface Connection GmbH. Documentation and a detailed description of the included document classes is available through Interface Connection GmbH. Also available: MagicHIT for Windows. InterFace Connection GmbH, Leipziger Strasse 16, 82008 Unterhaching, Telephone: +49 89 610 49-60, Telefax: +49 89 610 49-80

Copyright:  (c) 1996 InterFace Connection GmbH
Author:  InterFace Connection GmbH <service-center@ifconnection.de>
Version:  1.1-1

mind Motif Interface Designer

Mind is a GUI designer which generates C code for use with the Motif libraries. The graphical elements can be positioned directly on the screen and properties like look and behavior can be specified. The compiler can be started directly from Mind. For documentation/installation please see file /usr/doc/mind/INSTALL.SuSE. Contents: mind-3.0.r2 mind30 r2d

Copyright:  (c) flux GbR, Berlin
Author:  flux GbR, Berlin <flux.service@tonline.de>
Version:  3.0.r2-1

netscape Netscape (including script for starting netscape)

This package contains a start script for netscape. It makes it possible to set some variables an libraries for netscape. In this packages a netscape commuicator[tm] 4.04 is integrated. Note: the the program netscape is installed as netscape but the script will be called by its name Netscape. Attention: Please read the license conditions in /usr/doc/packages/netscape/ Contents: communicator-v404-export.x86-unknown-linux2.0 libXpm.4.6

Copyright:  (c) (C) 1994-1997 Netscape Communications
Version:  4.04-6

not demo NOTARIUS development tools with mask generator and DB connectivity

NOTARIUS is an development tool for UNIX systems with database access such as YARD, Informix and more. This is a demo version for Linux and offers the possibility to easily test the capabilities of NOTARIUS. More infos are to be found under: /usr/lib/not demo/README There is a PostScript document containing more than 100 pages.

Copyright:  (c) K + R Informationssysteme Konstanz
Author:  K + R Informationssysteme Konstanz <info@iridium.de>
Version:  1.2-1

simcity Demo version of the well known game.

The demo runs 15 minutes, or if you try and save the game. If you would like to purchase a key to unlock the game follow the directions when the demo ends. Start the game to get more information. Contents: simcity

Copyright:  (c) 1993 DUX Stoftware Corporation
Author:  Mark Mattes <RonnieW252@aol.com>
Version:  95.4.10-3

snavig Cygnus Source-Navigator

Source-Navigator, the source-code comprehension tool from Cygnus Solutions, helps you analyze, reuse, and develop software in C/C++, Java, Tcl, and FORTRAN. Source-Navigator dramatically reduces the time required to understand existing code and assists with the re-structuring and reengineering of existing software. Source-Navigator, with its intuitive graphical user interface, directs you through analysis, change, build, and debug cycles. Its built-in database enables true project scalability, even in projects exceeding one million lines of code. Source-Navigator Lite Try a version of Source-Navigator. Source-Navigator Lite is available for use free of charge and provides full browsing, navigation, and project editing capability. The Lite configuration does have some limitations to its scalability:

- Supports projects smaller than 50,000 lines of code - Does not expose the API for integration or customization - Allows limited project printing Source-Navigator Lite can be used free of charge, and provides an easy way to evaluate the functionality. If you decide to upgrade to a complete version of Source-Navigator, you simply need a license code from Cygnus, and your application and projects will be automatically updated. Contents: SN-3.3.1-linux The Lite version is resticted to 50,000 lines of source code

Copyright:  (c) 1997 Cygnus Solutions
Version:  3.3.1-3

sniff SNiFF+ Integrated Cross-Platform Development Environment

SNiFF+ is an open, extensible and scalable cross-platform programming environment for C, C++, FORTRAN, JavaTM, CORBA IDL and other languages. The main goal in developing SNiFF+ was to create an efficient and portable environment with a comfortable user interface. SNiFF+ is tailored for large multi-developers projects. Contents: sniff-2.4 sniff new To test this demo you need to obtain a license. Run /opt/sniff/bin/get licenseinfo and obtain a license key from sniff-license@takefive.co.at

Copyright:  (c) TakeFive Software GmbH
Version:  2.4-1

terroid Zap ’em arcade game.

TERROID is a classic-arcade game made to bring the quality of R-Type, SWIV, Wings of Death or Lethal Excess to your workstation. This is a DEMO of a shareware product.

Copyright:  (c) 1997 DIGITAL Nightmares
Author:  Michael Thayer <M.J.Thayer@ed.ac.uk>
Version:  1.3-1

uc X11 Unix commander.

”Unix Cockpit” (UC) is a new Unix file manager. UC integrates file browsers, a directory tree, and your standard shell into a useful, mouse driven tool. Using a source and target browser, file operations like copy, move, delete, or edit can be executed. Additional operations (compress, gzip, find, etc.) can be chosen via a menu which can also hold user defined commands. In a separated shell even more commands can be entered. The current directory tree can be displayed in an additional window, providing the possibilities to navigate through the file system, select the working directory, and move directories. Using a small control window, all used windows can be pushed down, popped up, or be hidden. UC also provides an online help system. Summarizing, UC is a good X11 shell. Please notice that this program is shareware. Contents: uc Please notice that this program is shareware.

Copyright:  (c) 1989 - 1995 Henrik A. Klagges
Author:  Henrik A. Klagges <henrik@UniX11.com>
Version:  2.03-5

vdx VDX - Demo version of a Motif Interface Designer

The View Designer/X is a motif interface builder, which enables application developers to interactively design user interfaces and generate C and C++ source code. The VDX provides following features: WYSIWYG and Tree View with drag & drop, Resource Editor, reusable components, templates, adaptable code generation. This package is a demo version. Contents: vdx-1.0.4

Copyright:  (c) Copyright 1996, Dirk Laessig
Author:  Dirk Laessig <lazy@de.ibm.com>
Version:  1.0.4-2

vse Visual SlickEdit - Demo

Demo version of the well-known ’Visual SlickEdit’ code editor. Invoke the demo by entering ’vsedemo’. Further information can be found under: /usr/doc/packages/vse Contents: vse-1.73

Copyright:  (c) 1993-1996 MicroEdge Inc.
Author:  Siegfried Hirsch <shirsch@infos.de>
Version:  1.73-2

wingz A easy to use spreadsheet program for X (SHAREWARE)

Wingz is a spreadsheet similar to Excel or Xess3, but it is shareware. Extensive graphics are possible. Wingz can totally be controlled by scripts. The registration fee is about $50,- per user. Wingz is available for several other UNIX systems and Mac/Windows machines and is extensible (real time versions, hyperscript). The manufacturer offers support for this product. Contents: wingz This program is SHAREWARE! Please read /usr/Wingz/SHAREWARE.

Copyright:  (c) Investment Intelligent Systems Corporation, USA
Author:  Investment Intelligent Systems Corporation <linux-bug@wingz.com>
Version:  1.4.1-4

xess3 Xess3 Spreadsheet - Demo

A spreadsheet program which can certainly be compared to the programs it was modelled after. Sample files can be found in /usr/local/lib/xess/examples. Xess also offers an API interface for C programs. Attention: This program will install into /usr/local.

Copyright:  (c) 1992-1996 Applied Information Systems, Inc. and others
Author:  <support@ais.com>
Version:  3.1.3b-4

xgal A Galaga-clone

This is a clone of the arcade hit Galaga. Contents: xgal Please remember that this program is shareware.

Copyright:  (c) 1995 Joe Rumsey
Author:  Joe Rumsey <ogre@atomic.com>
Version:  1.6c-8

xv GIF/TIFF/JPEG/PostScript image viewer

XV is an interactive image viewer for the X Window System. The XV program displays images in the GIF, JPEG, TIFF, PBM, PGM, PPM, X11 bitmap, Utah Raster Toolkit RLE, PDS/VICAR, Sun Rasterfile, BMP, PCX, IRIS RGB, XPM, Targa, XWD, possibly PostScript, and PM formats on workstations and terminals running the X Window System, Version 11. The XV program is also capable to manipulate pictures. It can zoom in and out of picture, and rotate and flip them. It also provides a wide variety of operations to modify colors. Surley, one of the most famous programs under Linux. You will find the documentation in /usr/doc/packages/xv/. Please remember that this program has shareware status for commercial use. Contents: xv-3.10a

Copyright:  (c) 1989, 1994 John Bradley (SHAREWARE)
Author:  John Bradley <bradley@dccs.upenn.edu>
Version:  3.10a-11

yardsql YARD-SQL Private Edition

This ”private edition” of YARD-SQL differs from the full version only in the amount of data which can be managed (in the private edition it is limited to 20 MB). YARD-SQL means: - standard SQL - dynamical SQL - CLient-Server architecture - Network capabilitiy - Referential integrity - Transaction security - embedded SQL for C - German dokumentation - Constraints - Interactive SQL YARD-SQL database server: Relational Client-Server architecture based database system for Linux providing ANSI/X/Open and SQL interface. Remote access via YARD-NET, extended SQL language and SELECT possibilities, 8 bit keymaps for national language support, Query Optimizer, support of different datatypes, data-integrity, power tools for database maintainement etc. YARD-ESQLC: SQL language embedding for C (without restrictions) YARD-NET: Distribution of Client-Server processes within one network via TCP/IP; including embedding of MS-Windows PCs via YARD-ODBC (without restrictions). YARD-X: Graphical user interface under OSF/Motif. YARD is also available for other operating systems; please contact us for questions or more informations. Please read the file /usr/lib/YARD/README. Contents: yardsql-4.04.0

Copyright:  (c) 1996-98 YARD Software GmbH
Author:  YARD Software GmbH <yard@yard.de>
Version:  4.04.0-1

0.1.12 SGM: Components for a SGML system

docbk30 DocBook: DTD and DSSSL-files

Copyright:

(c) Copyright 1992, 1993, 1994, 1995, 1996 HaL Computer Systems, Inc., O’Reilly & Associates, Inc., ArborText, Inc., and Fujitsu Software Corporation.

Author:

Eve L. Maler <elm@arbortext.com>, Terry Allen <tallen@sonic.net>, Norman Walsh <norm@berkshire.net>

Version:  3.0-12

gf gf - ”general formatter” for SGML documents (snafu.dtd, html.dtd)

‘gf’ from Gary Houston is short for ”general formatter”, i.e., it can work on documents which use the ISO ”general” document type definition (DTD). It can convert SGML documents conforming to a small number of DTDs into various output formats: LaTeX, ASCII, RTF and Texinfo. However not every output format can be generated for every DTD. Apart from the general DTD, gf supports the HTML DTD used in the WWW project and Gary’s Snafu DTD. ‘gf’ is not intended as a flexible system for hacking up a formatter for a random DTD, but as a usable document production system for a few DTDs.

SGM: Components for a SGML system

Contents: gf-0.46

Copyright:  (c) (c) Gary Houston, GPL
Author:  Gary Houston <ghouston@actrix.gen.nz>
Version:  0.46-11

html dtd HTML DTDs (Document Type Definitions) + Docs

Document Type Definitions (DTDs) for HTML 2.0 (standard), HTML 3.2 (recommendation of the W3 Consortium), HTML 4.0 (Specification of the W3 Consortium, Draft) and ISO-HTML (draft). The package also contains the documentations (v. /usr/doc/packages/html dtd).

Contents: html-3.2 html-spec-19950922 iso-html

Version:  0.3-2

hyperref Resolve LaTeX cross-references for the HyperTeX project

Copyright:  (c) Sebastian Rahtz
Author:  s.rahtz@elsevier.co.uk
Version:  6.14-2

iso ent Character Entity Sets f”ur ISO 1986:8879 (SGML)

Copyright:  (c) ISO
Version:  1.0-0

jade dsl DSSSL-Engine for SGML documents

Jade is an implementation of DSSSL (Document Style, Semantics and Specification Language); pronounce it as dissl -- it rimes with whistle.

Copyright:  (c) (c) James Clark, GPL
Author:  James Clark <jjc@jclark.com>, David Megginson <dmeggins@microstar.com>
Version:  1.1-0

jadetex JadeTeX - LaTeX macros to process TeX output from Jade (jade dsl)

With Sebastian Rahtz’ macro package ‘jadetex’ is is possible to process the output of the TeX backend of Jade (jade dsl). Resulting DVI files are viewable e.g., with ‘xdvi’ or printable like any other DVI file.

Contents: jadetex

Copyright:  (c) 1996,1997 Sebastian Rahtz <s.rahtz@elsevier.co.uk>
Author:  Sebastian Rahtz <s.rahtz@elsevier.co.uk>
Version:  0.57-2

linuxdoc linuxdoc-sgml

The file doc/guide.sgml is SGML source for the User’s Guide. As you probably don’t know how to format it yet, there is formatted plain ASCII in doc/guide.txt. There’s also guide.dvi and guide.ps, if you wish to print them instead. The first thing you should read is doc/guide.txt, which explains how to install the software, and how to start writing documents with it. The file doc/example.sgml is a simple example document which you can use as a tutorial. For real-life examples of uses of Linuxdoc-SGML, look at the Linux HOWTOs on sunsite.unc.edu:/pub/Linux/docs/HOWTO/other-formats/sgml. These HOWTOs are written in SGML and automatically translated into PostScript, dvi, and ASCII using this system. You can also look at the URL http://sunsite.unc.edu/mdw/HOWTO/ which has HTML formats of many Linux HOWTOs, produced using Linuxdoc-SGML.

Contents: linuxdoc-sgml-1.5

Author:  Matt Welsh <mdw@sunsite.unc.edu>, Greg Hankins <gregh@cc.gatech.edu>
Version:  1.5-11

sgmltool SGML-Tools - a text-formatting package

SGML-Tools is a text-formatting package based on SGML (Standard Generalized Markup Language), which allows you to produce LaTeX, HTML, GNU info, LyX, RTF, and plain ASCII (via groff) from a single source. This system is tailored for writing technical software documentation, an example of which are the Linux HOWTO documents. It should be useful for all kinds of printed and online documentation.

SGML-Tools is not able to process arbitrary SGML documents; in such a case, give jade dsl a try and write your own DSSSL scripts (take the docbk30 package as an example). Contents: sgml-tools-1.0.5 This package is considered as the successor of the Linuxdoc package. Production systems should probably prefere Linuxdoc for the moment.

Copyright:  (c) GPL

Author:

Cees A. de Groot <cg@pobox.com>, Arun Sharma <asharma@sco.com>, Andre Schoorl <aschoorl@engr.UVic.CA>, Daniele Giacomini <daniele@tv.shineline.it>, Dave Love <d.love@dl.ac.uk>, Dave Platt <dplatt@iq.TVSoft.com>, Eric S. Raymond <esr@thyrsus.com>, Francisco J. Montilla <pacopepe@nova.es>, Fabrizio Polacco <fpolacco@megabaud.fi>, Frank Pavageau <pavageau@imaginet.fr>, Jan Echternach <echter@informatik.uni-rostock.de>, Henry S. Thompson <ht@cogsci.ed.ac.uk>, Karsten Tinnefeld <tinnefeld@irb.informatik.uni-dortmund.de>, Kees de Bruin <kees de bruin@tasking.nl>, Ken MacLeod <kmacleod@midamerican.com>, Lee Marzke <marzke lee@si.com>, Marco Budde <Marco Budde@hqsys.fido.de>, Mitchell Perilstein <mnp@compass-da.com>, Sven Rudolph <sr1@os.inf.tudresden.de>, Tommy Marcus McGuire <mcguire@cs.utexas.edu>, Ulrich Windl <ulrich.windl@rz.uni-regensburg.de>

Version:  1.0.5-1

sgrep search a file for a structured pattern

Copyright:  (c) GPL
Author:  Jani Jaakkola <jjaakkol@cs.helsinki.fi>, Pekka Kilpelainen <kilpelai@cs.helsinki.fi>
Version:  0.99-9

sp SGML parser tools

The tools of this package provide the possibility to manage SGML and XML documents. It contains the parser ‘nsgmls’ and the supporting programs ‘sgmlnorm’, ‘spam’, ‘spent’, and ‘sx’. ‘sx’ is useful as a converting tool from SGML to XML, the comming WWW standard. You’ll find the documentation for all the programs under /usr/doc/packages/sp/.

Copyright:  (c) (c) James Clark, GPL
Author:  James Clark <jjc@jclark.com>
Version:  1.3-0

0.1.13 SND: Sound related stuff

aumix A curses and command line based mixer

With aumix you may set up the mixer parameters via script.

Contents: aumix-1.8

Copyright:  (c) 1993, 1996, 1997 the authors

Author:

Trevor Johnson <trevor@jpj.net>, Alessandro Rubini <alessandro.rubini@linux.it>, Kenn Humborg <kenn@wombat.ie>, Dan Fandrich <dan@fch.wimsey.bc.ca>, Jeff Spirko <spirko@topquark.physics.lehigh.edu>, Harley Silver <harley@widow.aracnet.net>, Rainer Canavan <canavan@cs.bonn.edu>

Version:  1.8-1

gmod Linux music module player

A module player which - similar to xmp - supports the Gravis Ultrasound and AWE soundcards. xgmod has a Qt-based frontend.

Contents: gmod-3.1

Copyright:  (c) 1996 by Andrew Robinson
Author:  Andrew J. Robinson <robinson@cnj.digex.net>
Version:  3.1-4

jazz Midi sequencer

Complete midi sequencer which supports recording and playing as a bundle of editing functions. There are two main windows. One displays the complete track the other further details. It requires either a midi card supported by /dev/sequencer or a complete MPU 401 compatible device.

SND: Sound related stuff

Contents: jazz-2.1d

Copyright:  (c) 1994 Andreas Voss

Author:

Andreas  Voss  <andreas@avix.rhein-neckar.de>,  Per  Sigmond

<Per.Sigmond@hiagder.no>

Version:  2.1d-10

mxv WAV-player, modify and analyze samples, sound-synthesis too

With mxv you may play and modify sound files of different formats (wav, au, etc.). Mxv also provides the function of recording via /dev/dsp.

Copyright:  (c) Douglas Scott
Author:  Douglas Scott <MixViews@ccmrc.ucsb.edu>
Version:  1.10-1

rplay Sound daemon

rplay is a sound package which handles local as well as network sound output. A lot of programs use this daemon for sound output. Contents: rplay-3.2.0b5

Copyright:  (c) 1993, 1994 Mark Boyns
Author:  Mark Boyns <boyns@sdsu.edu>
Version:  3.2.0b5-8

sapphire An Audio-Tracer

Sapphire is an audio-tracer. You could imagine this as a ”ray tracer for sounds”. Sapphire has the ability to compute whole music pieces given by a description language. There is a File andrew.sp under /usr/lib/sapphire. Let your PPro working its little heart out. :-) Contents: sapphire

Copyright:  (c) 1993-1995 J. Finnis
Author:  Jim Finnis <>
Version:  2.1-8

snd au Sound files in AU-format. (Sun Audio)

These are sound files in the AU format under /usr/share/sounds/au. So if you have a sound card and a kernel with sound support you can listen to those files with a simple ’cat <file> > /dev/audio’.

Version:  98.2.13-1

snd mod Some MOD-files (music on disc)

These are files for modplayers. They are containing sampled instruments and sequencer data which are interpreted by the players. One such player can be found in the package ”tracker”.

Copyright:  (c) Diverse
Version:  98.2.13-1

snd wav Soundfiles in WAV-format.

These are sound files in the WAV format under /usr/share/sounds/wav. So if you have a sound card and a kernel with sound support you can listen to those files with a simple ’play <file>’. Contents: sounds

Version:  98.2.13-1

speakf Voice Communication Over Data Networks

Voice Communication Over Data TCP/IP Networks. Supported compression algorithms are: GSM, ADPCM, and others. Sound can be encrypted either with a key supplied in a file, DES and/or IDEA with given key(s). Contents: speak freely-6.1c

Copyright:  (c) 1996 John Walker, Release 6.1b

Author:

John Walker <kelvin@fourmilab.ch>, ”Unmoderated list for discussion” <speak-freely@fourmilab.ch>, ”Periodic digest of messages” <speak-freely-digest@fourmilab.ch>, ”Announcements of general interest” <speak-freely-announce@fourmilab.ch>

Version:  6.1c-1

tcd Compact curses based CD player

Compact curses based CD player, with easier key control and more features than other CD players, such as track repeat, continuous play, track database, etc.

Contents: tcd-1.0

Copyright:  (c) GPL
Author:  Tim P. Gerla <timg@means.net>
Version:  1.0-5

tracker Tracker is a portable MOD-player for Linux.

Tracker allows to play standard-MOD files (MOD = Music On Disc) independent of the used sound card (/dev/audio support has to be in kernel). You’ll find examples of MOD-files in /usr/share/sounds/mod. Contents: tracker

Version:  4.3-9

wavplay Player / recorder for WAV files

Copyright:  (c) 1993 FF-Soft ’93
Author:  Andre Fuechsel <af1@irz.inf.tu-dresden.de>
Version:  0.21-8

workbone text-mode program for playing audio CD Roms

WorkBone is an interactive control panel in text-mode to play audio CDs under Linux. It is a text-modus interface to the ”hardware.c” module of the X11 CD-program WorkMan. WorkBone is designed to easily play audio CDs in text-mode under Linux. WorkBone by purpose doesn’t have all those features WorkMan has. It is thought for those people who want to listen to their music with as minimal expense as possible. WorkBone uses the number block as a control panel to enable jumping from track to track, play CDs, stop etc. Furthermore the actual playing time is permanently actualized and displayed. WorkBone works with most CD-ROM drives. But there still are some NEC CD-ROMs that don’t work because they us not standardized own audio commands which results in a malfunction of WorkBone. WorkBone expects to find the device: /dev/cdrom . If you don’t have this device, change to /dev and soft link your current cd rom device to ’cdrom’. Contents: WorkBone-2.3

Copyright:  (c) 1994 Thomas McWilliams
Author:  Thomas McWilliams <thomas.mcwilliams@f615.n109.z1.fidonet.org>
Version:  2.3-7

workman Workman

Audio CD player based on OpenLook. Contents: WorkMan-1.3

Author:  Steven Grimm <koreth@hyperion.com>
Version:  1.3-9

xmcd Motif(tm) CD Audio Player

xmcd is an X11/Motif-based CD player utility; and cda is a command-line driven, non-graphical CD audio player. After installation you still have to configure your CD-player. Go to the directory /usr/X11R6/lib/X11/xmcd/config and run configure.sh as root. Documentation: man xmcd Contents: xmcd-2.2

Copyright:  (c) 1995-1997 Ti Kan
Author:  Ti Kan <ti@amb.org>
Version:  2.2-8

xmitsumi CD-Player panel for Mitsumi CD-ROM drives

A simple CD-Player for Mitsumi CD-ROM drives. Contents: xmitsumi-0.75

Copyright:  (c) 1994 Liem Bahneman
Author:  Liem Bahneman <roland@cac.washington.edu>
Version:  0.75-10

xmix X11 interface for the Linux sound driver

If your sound card is supported by Linux, you can use this program to adjust the sound.

SPL: Spell checking utilities and databases

Copyright:  (c) 1992 Steve Haehnichen
Author:  Steve Haehnichen, Hal Brand
Version:  2.1-8

xmp eXtended Module Player

With xmp you may play several music formats: MOD, S3M, XM, 669, STM, MTM, PTM, OKT, FAR. Furthermore xmp uses the special hardware found in the GUS and AWE soundcards. Regular soundcards are only supported with the commercial OSS sound driver. Contents: xmp-0.99c

Copyright:  (c) 199[67] C Matsuoka, H Carraro Jr

Author:

Claudio Matsuoka <claudio@cpgei.cefetpr.br>, Takashi Iwai <iwai@dragon.mm.t.utokyo.ac.jp>

Version:  0.99c-2

xqmixer Sound Mixer

a mixer for X-Windows using qt-libraries Contents: XqMixer-1.4

Copyright:  (c) GPL
Author:  Rajko Albrecht <rajko.albrecht@informatik.tu-chemnitz.de>
Version:  1.4-3

0.1.14 SPL: Spell checking utilities and databases

dicts Ispell dictionary files

This package contains sources of eight ispell dictionaries: for German (ready compiled for ispell in package igerman) for Danish (ready compiled for ispell in package idanish) for Spanish (ready compiled for ispell in package ispanish) for French (ready compiled for ispell in package ifrench) for Italian (ready compiled for ispell in package iitalian) for Dutch (ready compiled for ispell in package idutch) for Swedish (ready compiled for ispell in package iswedish) for Norwegian (ready compiled for ispell in package inorsk) Please read the README’s under /usr/src/dicts/i*/ . Contents: dutch96 espa˜ nol-1.5 francais-IREQ-1.4 hk2-deutsch idanish-1.2 iswedish-1.2 italian norsk prepare-dicts

Author:

Heinz  Knutzen  <hk@informatik.uni-kiel.d400.de>,  Martin  Boyer

<mboyer@robot.ireq.ca>, Marco Roveri <marco@mrg.dist.unige.it>, G”oran Andersson <goran@debian.org>, Santiago Rodriguez <srodri@fi.upm.es>, Jesus Carretero <jesus@fi.upm.es>, Piet Tutelaers <rcpt@urc.tue.nl>, Anders Andersen <aandersen@acm.org>

Version:  1.2-2

idanish Danish ispell dictionary

This packages includes a ready danish dictionary for ispell. A short usage description for ispell is given in /usr/doc/packages/ispell/README of the packages ispell. The sources for this dictionary are included in the package dicts.

Copyright:  (c) 1997 G”oran Andersson
Author:  G”oran Andersson <goran@debian.org>
Version:  1.2-2

idutch Dutch ispell dictionary

This packages includes a ready dutch dictionary for ispell. A short usage description for ispell is given in /usr/doc/packages/ispell/README of the packages ispell. The sources for this dictionary are included in the package dicts.

Copyright:  (c) 1997 Piet Tutelaers
Author:  Piet Tutelaers <rcpt@urc.tue.nl>
Version:  1.2-2

ifrench French ispell dictionary

/usr/doc/packages/ispell/README of the packages ispell. The sources for this dictionary are included in the package dicts.

Copyright:  (c) 1993-96 Martin Boyer
Author:  Martin Boyer <mboyer@ireq-robot.hydro.qc.ca>
Version:  1.2-2

iitalian Italian ispell dictionary

This packages includes a ready italian dictionary for ispell. A short usage description for ispell is given in /usr/doc/packages/ispell/README of the packages ispell. The sources for this dictionary are included in the package dicts.

Copyright:  (c) 1997 Marco Roveri
Author:  Marco Roveri <marco@mrg.dist.unige.it>
Version:  1.2-2

inorsk Norwegian ispell dictionary

This packages includes a ready norwegian dictionary for ispell. A short usage description for ispell is given in /usr/doc/packages/ispell/README of the packages ispell. The sources for this dictionary are included in the package dicts.

Copyright:  (c) Anders Andersen, Ivar Aavatsmark(?)
Author:  Anders Andersen <aandersen@acm.org>, Ivar Aavatsmark
Version:  1.2-2

ispanish Spanish ispell dictionary

This packages includes a ready spanish dictionary for ispell. A short usage description for ispell is given in /usr/doc/packages/ispell/README of the packages ispell. The sources for this dictionary are included in the package dicts.

Copyright:  (c) 1994-96 Santiago Rodriguez, Jesus Carretero
Author:  Santiago Rodriguez <srodri@fi.upm.es>, Jesus Carretero <jesus@fi.upm.es>
Version:  1.2-2

ispell ispell

Ispell is a fast screen-oriented spelling checker that shows you your errors in the context of the original file, and suggests possible corrections when it can figure them out. Compared to UNIX spell, it is faster and much easier to use. Ispell can also handle languages other than English. Ispell has a long history, and many people have contributed to the current version - some of the major contributors include R. E. Gorin, Pace Willisson, Walt Buehring, and Geoff Kuenning. You find a short description in the directory /usr/doc/packages/ispell/: Contents: ispell-3.1.20

Copyright:  (c) 1983, by Pace Willisson, 1987 - 1993, 1995 by Geoff Kuenning, Granada Hills, CA.

Author:

R. E. Gorin, Pace Willisson <pace@ai.mit.edu>, Walt Buehring, Goeff Kuenning <geoff@ITcorp.com>

Version:  3.1.20-9

iswedish Swedish ispell dictionary

This packages includes a ready swedish dictionary for ispell. A short usage description for ispell is given in /usr/doc/packages/ispell/README of the packages ispell. The sources for this dictionary are included in the package dicts.

Copyright:  (c) 1997 G”oran Andersson
Author:  G”oran Andersson <goran@debian.org>
Version:  1.2-2

words English dictionary

This packages contains an English dictionary which will be installed as /usr/share/dict/words This dictionary is maybe used by look(1) and ispell(1). Contents: linux.words.2

Author:  Rik Faith <faith@cs.unc.edu>
Version:  words.2-9

TCL: Tcl/Tk/TclX, Tcl-Language and Tk-Toolkit for X

igerman German ispell dictionary

This packages includes a ready german dictionary for ispell. A short usage description is given in /usr/doc/packages/ispell/README of the packages ispell. The sources for this dictionary are included in the package dicts.

Copyright:  (c) 1993-95 Heinz Knutzen
Author:  Heinz Knutzen <hk@informatik.uni-kiel.d400.de>
Version:  1.2-2

0.1.15 TCL: Tcl/Tk/TclX, Tcl-Language and Tk-Toolkit for X

blt Tk extension

An add-on library for Tk. Does not require patching of the Tk and Tcl libraries. Provides some additional Widgets and Functions. These extensions only work with the tk old and tcl old libraries. Contents: blt-1.9

Author:  George Howlett <george.howlett@att.com>
Version:  1.9-11

convert convert

As you see when you ran Convert this is a unit conversion software. (over 700 conversions) Contents: convert-1.3

Copyright:  (c) (c) 1995 Costas Vlassi
Author:  Costas Vlassis <lonewolf@compulink.gr>
Version:  1.3-0

dejagnu dejagnu

DejaGnu contains an extension to tcl, called ’expect’. Dejagnu is a library of expect scripts that are used by GNU for running test suites on GNU tools. Expect is great for making e.g. login scripts for telnet, ftp or anything else. Or for any other automated programs that have to react to the output of an arbitrary program started by expect. Contents: dejagnu-970221

Copyright:  (c) Copyright (C) 1992-1995 Free Software Foundation, Inc.
Author:  Cygnus
Version:  970221-9

eddi eddi

X editor based on the TiX shell, with text highlighting and several other useful features. Contents: eddi-0.9b12

Copyright:  (c) Holger Schanz
Author:  Holger Schanz <holger.schanz@student.uni-ulm.de>
Version:  0.9b12-1

expect expect

Expect is a tool primarily for automating interactive applications such as telnet, ftp, passwd, fsck, rlogin, tip, etc. Expect really makes this stuff trivial. Expect is also useful for testing these same applications. Expect is described in many books, articles, papers, and FAQs. There is an entire book on it available from O’Reilly. Expect is necessary for dejagnu.

Copyright:  (c) public domain
Author:  libes@nist.gov
Version:  5.22-10

itcl object oriented extension for Tcl

incr Tcl adds object-oriented programming facilities to Tcl. It was NOT designed as yet another whiz-bang object-oriented programming language; indeed, it is patterned somewhat after C++. It was designed to support more structured programming in Tcl. Scripts that grow beyond a few thousand lines become extremely difficult to maintain. [incr Tcl] attacks this problem in the same way that any object- oriented programming language would, by providing mechanisms for data encapsulation behind well-defined interfaces. Please read the file /usr/doc/packages/itcl/README.SuSE.

Contents: itcl2.2

Copyright:  (c) 1993-1995 AT&T Bell Laboratories,

Author:

Michael J. McLennan <michael.mclennan@att.com>, Dr. John Ousterhout <john.ousterhout@eng.sun.com>

Version:  2.2-9

jstools jstools

This is version 4.2/4.5 of the jstools distribution. jstools is a suite of applications, and some libraries they share. The most interesting application is multi-mode editor with support for WYSIWYG generation of HTML and PostScript, among other goodies. There are also a number of small graphical utilities intended to be used from within shell scripts or window-manager configuration files. Contents: jstools-tk4.2v4.5 jstools src

Copyright:  (c) Jay Sekora
Author:  Jay Sekora <js@aq.org>
Version:  4.5-9

listexec List browser

With this little program one can list all files in a specified directory and e.g. hear them with a specified sound player or view them as pictures with a picture viewer. The syntax may be obtained by listexec -h Contents: listexec-0.4.3

Copyright:  (c) 1997, S.u.S.E. GmbH
Author:  Marcus Sch”afer <ms@suse.de>, Martin Scherbaum <maddin@suse.de>
Version:  0.4.3-5

omnimoni System Monitor

A Tcl/Tk based fully configurable system-monitor. See /usr/X11R6/lib/omnimoni for details. Contents: OmniMoni-1.0

Copyright:  (c) 1995 Rainer Mager
Author:  Rainer Mager <rvm@fore.com>
Version:  1.0-9

samt samt

SAmT is a tool for the System Administrator. It helps keeping a list of important system files without having to look for them each time you need them. It provides an easy means of rebooting the system or drive it into a certain runlevel respectively. Furthermore it can create a backup of your given files, thus giving you the opportunity of switching back to a ”running system”. Contents: samt Please be aware that this software is still considered beta!

Copyright:  (c) Michael Burghart S.u.S.E. GmbH 1997
Author:  Michael Burghart <mb@suse.de>
Version:  2.8-3

scotty Scotty - Tcl Extensions for Network Management Applications

Scotty is the name of a software package which allows to implement site specific network management software using high-level, string-based APIs. The software is based on the Tool Command Language which simplifies the development of portable network management scripts. The scotty source distribution includes two major components. The first one is the Tnm Tcl Extension which provides access to network management information sources. The second component is the Tkined network editor which provides a framework for an extensible network management system. Contents: scotty-2.1.7

Copyright:  (c) Juergen Schoenwaelder
Author:  J”urgen Sch”onw”alder <schoenw@cs.utwente.nl>
Version:  2.1.7-8

tcl The Tcl script language

Tcl is a simple to use text-based scripting language with many built-in features that make it especially nice for writing interactive scripts similar to perl.

The TCL shell is /usr/bin/tclsh7.6 This package is required if you want to install Tk. Contents: tcl7.6p2

Copyright:  (c) The Regents of the University of California, Sun Microsystems, Inc,
Author:  Dr. John Ousterhout <john.ousterhout@eng.sun.com>
Version:  7.6p2-10

tcl new The Tcl script language

Tcl is a simple to use text-based scripting language with many built-in features that make it especially nice for writing interactive scripts similar to perl. The TCL shell is /usr/bin/tclsh8.0 This package is required if you want to install Tk. The bytecode compiler is new. The core of the Tcl interpreter has been replaced with an on-the-fly compiler that translates Tcl scripts to byte codes. Contents: tcl8.0b2

Copyright:  (c) The Regents of the University of California, Sun Microsystems, Inc,
Author:  Dr. John Ousterhout <john.ousterhout@eng.sun.com>
Version:  8.0p1-6

tcl old The Tcl script language version 7.4

Tcl is a simple to use text-based scripting language with many built-in features that make it especially nice for writing interactive scripts similar to perl. The TCL shell is /usr/bin/tclsh7.4 This package is required if you want to install Tk old. Contents: tcl7.4p3

Copyright:  (c) The Regents of the University of California, Sun Microsystems, Inc,
Author:  Dr. John Ousterhout <john.ousterhout@eng.sun.com>
Version:  7.4p3-10

tclx TclX - Extended Tcl: Extended command set for Tcl

Extended Tcl is a superset of standard Tcl and is built alongside the standard Tcl sources. Extended Tcl has three basic functional areas: A set of new commands, a Tcl shell (i.e. a Unix shell-style command line and interactive environment), and a user-extensible library of useful Tcl procedures, any of which can be automatically loaded on the first attempt to execute it. In addition, a detailed help system is available for TCL/TK: tclhelp. Contents: tclX7.6.0

Copyright:  (c) Copyright 1992-1995 Karl Lehenbauer and Mark Diekhans.
Author:  Karl Lehenbauer and Mark Diekhan <info@NeoSoft.com>
Version:  7.6.0-10

tix Tools for tk

The Tix library has by far the greatest collection of widgets for programming with Tcl/Tk. Highlights include: Hierarchical List box, Directory List/Tree View, Spreadsheet, Tabular List box, Combo Box, Motif style FileSelectBox, MS Windows style FileSelectBox, PanedWindow, NoteBook, Spin Control widget .... and many more. With these new widgets, your applications will look great and interact with your users in intuitive ways. Contents: Tix4.1.0

Copyright:  (c) 1996, Expert Interface Technologies
Version:  4.1.0-10

tix new tix new

The Tix library has by far the greatest collection of widgets for programming with Tcl/Tk8.0. Highlights include: Hierarchical List box, Directory List/Tree View, Spreadsheet, Tabular List box, Combo Box, Motif style FileSelectBox, MS Windows style FileSelectBox, PanedWindow, NoteBook, Spin Control widget .... and many more. With these new widgets, your applications will look great and interact with your users in intuitive ways.

Copyright:  (c) Copyright (c) 1996, Expert Interface Technologies
Version:  4.1.0.006-2

tixlpq Tix-lpq

tix frontend for lpq Contents: tixlpq-0.1

Copyright:  (c) GPL
Version:  0.1-2

tk TK toolkit for TCL

Tk is an extension to Tcl that allows you to quickly and easily build X11 applications that have the look and feel of Motif apps. The TK-Shell is /usr/X11R6/bin/wish4.2 These applications might as well be run on Windows 95/NT as well as on apple. Demo: /usr/X11R6/lib/tk4.2/demo/widget Contents: tk4.2p2

Copyright:  (c) The Regents of the University of California. 1994-95 Sun Microsystems, Inc.
Author:  John Ousterhout <john.ousterhout@eng.sun.com>
Version:  4.2p2-9

tk new TK toolkit for TCL

Tk is an extension to Tcl that allows you to quickly and easily build X11 applications that have the look and feel of Motif apps. The TK-Shell is /usr/X11R6/bin/wish8.0 Native look and feel. The widgets have been rewritten to provide (nearly?) native look and feel on the Macintosh and PC. Demo: /usr/X11R6/lib/tk8.0/demo/widget Contents: tk8.0b2

Copyright:

(c) Copyright: The Regents of the University of California. 1994-95 Sun Microsystems Inc,

Author:  John Ousterhout <john.ousterhout@eng.sun.com>
Version:  8.0p1-6

tk old TK toolkit for TCL

Tk is an extension to Tcl that allows you to quickly and easily build X11 applications that have the look and feel of Motif apps. The TK-Shell is /usr/X11R6/bin/wish4.0 These applications might as well be run on Windows 95/NT as well as on apple. Demo: /usr/X11R6/lib/tk4.0/demo/widget Contents: tcl7.4p3 tk4.0p3

Copyright:  (c) The Regents of the University of California. 1994-95 Sun Microsystems, Inc.
Author:  John Ousterhout <john.ousterhout@eng.sun.com>
Version:  4.0p3-11

tkdesk TkDesk - an ultra comfortable graphical file/desktop manager

TkDesk is a graphical desktop and file manager for UNIX (especially Linux) and the X Window System. Compared with other file managers available, it offers the most complete set of file operations and services, plus gives the user the ability to configure most aspects of TkDesk in a powerful way. The reason for this is the use of Tcl/Tk as the configuration and (for the greatest part of TkDesk) implementation language. TkDesk has been influenced by various other systems and file managers: NeXT, for laying out the file browser windows, Apple Finder, for the idea of file annotations and, (shock horror), Windows 95, for some other (of course minor and unimportant ;-)) inspirations. This is a brief overview of the most prominent features of TkDesk: o Arbitrary number of automatically refreshed file browsers and file list windows, o Configurable file-specific pop-up-menus, o Drag and drop, o Files and directories may also be dropped onto the root window, o Configurable application bar, with several displays and cascadable pop-up menus for each button, files can also be dropped here, o History of visited directories, opened files, executed commands, and others, which is automatically saved to disk, o Find files through their annotation, name, contents, size or age, o Trash can for safe deletion of files and directories, o Calculation of disk usage for directory hierarchies, o All file operations (find, copy, disk usage, etc.) are carried out in the background, o Traversal of directory hierarchies through recursive cascaded menus, o Bookmarks, create menu entries for often used files/directories, o Comprehensive hypertextish online help, o Built-in multi-buffer and undo-capable editor, o Close coupling with Netscape and XEmacs, o Sound support, o Powerful configuration of nearly all aspect of TkDesk through Tcl/Tk, this also allows the Tcl-literate to extend TkDesk in arbitrary ways, o Free of charge! But see the file COPYING, or menu entry Help/License for information on usage and redistribution of TkDesk.

Contents: sounds tkdesk-1.0b4 tkdesk-1.0b4p1

Copyright:  (c) Christian Bolik
Author:  Christian Bolik <zzhibol@rrzn-user.uni-hannover.de>
Version:  1.0b4p1-10

tkfont tkfont

It is a program similar to xfontsel.

Contents: tkfont-1.1

Copyright:  (c) Neil Grant
Author:  Neil Grant <grantj@sfu.ca>
Version:  1.1-8

tkinfo tkinfo

TkInfo is a tk script to read GNU ”info” files and display them. TkInfo can be used stand alone (via WISH), or embedded within an application to provide integrated, on-line help.

Contents: tkinfo-1.4

Copyright:  (c) Axel Boldt
Author:  Axel Boldt <boldt@math.ucsb.edu>
Version:  0.8-8

tkinspec Browser for Tk programs

Tkinspect serves for supervising running Tk applications. All variables, resources and procedures can be viewed and changed.

Contents: tkinspect-5.1.6 tkinspect src

Author:  Sam Shen <SLShen@lbl.gov>
Version:  5.1.6-10

tkirc A pretty IRC client with GUI

Tkirc is a frontend for IRCII, written in Tcl/Tk. It can be customized with a file named .tkircrc in the users home-directory. An example .tkircrc can be found in /usr/doc/packages/tkirc/tkircrc-example

Contents: tkirc1.202

Copyright:  (c) 1997 Andreas Gelhausen
Author:  Andreas Gelhausen <atte@gecko.North.DE>
Version:  1.202-0

tkmail TkMail

TkMail is an X windows interface to mail built using Tcl/Tk. Reading, sending, and managing mail messages can almost all be done using only the mouse (except for the body of the message, of course). TkMail writes its folders in the format of the standard Unix ’Mail’ command. It also uses ’sendmail’ commands to do the mail delivery work.

Contents: tkmail-4.0beta9

Copyright:  (c) Christophe Le Bars
Author:  Christophe Le Bars <clebars@debian.org>
Version:  4.0beta8-16

tkman Manual-browser for X

A manual-browser for X with hyperlinks, history and more.

Contents: tkman-1.8

Copyright:  (c) 1993-1996 T.A. Phelps
Author:  T.A. Phelps <phelps@ACM.org>
Version:  1.8-10

tkman8 Manual-browser for X

A manual-browser for X with hyperlinks, history and more. Contents: tkman-2.0b4

Copyright:  (c) T.A. Phelps
Author:  T.A. Phelps <phelps@ACM.org>
Version:  2.0b4-11

tkppp tkppp

tkppp is a graphical front end to manually control a PPP connection. It presents in a small window a button for starting/quitting PPP, a small status message area, informational readouts and option buttons. Contents: tkppp-1.4-alpha

Copyright:  (c) Eric Jeschke
Author:  jeschke@cs.indiana.edu
Version:  1.4-9

tkrat TkRat

TkRat is a graphical Mail User Agent (MUA) which handles MIME. It is mainly written in C but the user interface is done in tcl/tk. Contents: tkrat-1.0.5

Copyright:  (c) Martin Forssen
Author:  maf@dtek.chalmers.se
Version:  1.0.5-3

tkstep TkSTEP

This is Tk 4.1p1 with some modifications to make it look like N*XTSTEP’s interface, plus some new widgets inspirated on N*XTSTEP. It also incorporates the DND drag and drop patches. Contents: tkstep4.2

Copyright:  (c) Alfredo K. Kojima
Author:  Alfredo K. Kojima <kojima@inf.ufrgs.br>
Version:  0.5b1-10

tkworld tkWorld

tkGrep: tkGrep is a Graphical User Interface to the Unix Grep command. It provides point and click access to Command Line Options, as well as a Toolbar for rapid execution of your grep command. tkSort: tkSort is just like tkGrep, except that it is built for the Unix sort command. tkREM: tkREM provides a Graphical User Interface for building Regular Expressions. The major modes (grep, Perl, Tcl, Python, Emacs, Sed & Awk) have buttons with meta-characters as their label. Contents: tkWorld-0.04 tkWorld src

Copyright:  (c) Wes Bailey
Author:  Wes Bailey <bailey@phys.ksu.edu>
Version:  0.04-8

tkxcd tkxcd

This is a diff front end with a look and feel based on Atria Clearcase xcleardiff. Both files are displayed in a window each and the differences are marked in different colors. Contents: tkxcd 1.1.0

Copyright:  (c) John C. Quillan
Author:  John C. Quillan <quillan@doitnow.com>
Version:  1.1.0-8

vtcl Visual Tcl

Visual Tcl is a freely-available, high-quality application development environment for UNIX, Windows and Macintosh platforms. Written entirely in Tcl and generating pure Tcl should make porting either unnecessary or trivial. Contents: vt-1.10 vtcl

Copyright:  (c) Stewart Allen
Author:  Stewart Allen <stewart@neuron.com>
Version:  1.10b-8

TEX: TeX/LaTeX and applications

xcdroast X-CD-Roast

CD-Writer-Program for X. Frontend for CD-related programs like ”cdwrite” and ”mkisofs”. Successor of the cdwtools-0.93. Tcl/Tk/Tix-based. Contents: xcdroast-0.96d-beta2 Please handle with extreme care! If a wrong SCSI device is entered this partition might be overwritten irreversibly (including partition tables and partitions belonging to other OS)

Copyright:  (c) Thomas Niederreiter
Author:  Thomas Niederreiter <p7003ad@sunmail.lrz-muenchen.de>
Version:  0.96b2-2

xf Interface builder for TCL/TK

XF allows you to interactively create and modify an user-interface based on the Tcl/Tk package. Tcl/tk (version 4.0 or later) is needed to run XF. XF creates a Tcl/Tk script as its output, which can then be interpreted using wish or any other Tcl/Tk intepreter. See: /usr/doc/packages/xf. A HTML-Documentation for XF and TCL/Tk is also included. The new version xf4.0 is configured for the use with Tcl/Tk 7.5/4.1. Contents: xf

Copyright:  (c) 1992-1996 Sven Delmas
Author:  Sven Delmas <sven@cimetrix.com>
Version:  4.0-11

xmclock a tcl based screen-locker

a menu based tcl program for creating various pictures which you will see after the screen locks. Contents: xmclock-1.1.2

Version:  1.1.2-3

zircon zircon

Zircon is an Tcl/Tk based IRC II client. This application is well designed and offers various features and possibilities. Contents: zircon-1.17p4b12

Copyright:  (c) Lindsay F. Marshall 1993
Author:  Lindsay F. Marshall <zircon@catless.newcastle.ac.uk>
Version:  1.17-10

0.1.16 TEX: TeX/LaTeX and applications

bibview X11 front-end for BibTeX databases

Using bibview, entries in several BiB databases can be moved, manipulated, created and searched. If you prefer German to English menus, all you have to do is copy the file var/X11R6/lib/appdefaults/BibView.ger to var/X11R6/lib/app-defaults/BibView . Contents: bibview-2.1

Copyright:  (c) 1992 TU MUENCHEN
Author:  Holger Martin, Peter Urban, Armin Liebl <liebla@informatik.tu-muenchen.de>
Version:  2.1-8

chess chess

This package contains (La)TeX macros (LaTeX2.09) and Fonts for typesetting chess figures within a (La)TeX document. Please have a look at the examples and documentation under /usr/lib/teTeX/texmf/doc/latex/chess/. Contents: chess12

Copyright:  (c) Febr. 1992, Piet Tutelaers
Author:  Piet Tutelaers <rcpt@urc.tue.nl>
Version:  1.2-13

colortbl colortbl - Add color to LaTeX tables

Contents: colortbl

Author:  David Carlisle
Version:  97.6.29-13

detex TeX -> ASCII Converter

This filter removes (La-)TeX macros.

Contents: deTeX-2.6

Copyright:  (c) 1986 Daniel Trinkle
Author:  Daniel Trinkle <trinkle@cs.purdue.edu>
Version:  2.6-8

dvi2tty A TeX to ASCII converter

dvi2tty converts a TeX-dvi file into a readable ASCII text file. You can then do full text searches without finding the TeX macros as well. For documentation see the man page (man dvi2tty) and /usr/doc/packages/dvi2tty.

Contents: dvi2tty-5.1

Copyright:  (c) 1989, 1995 Svante Lindahl / Marcel Mol
Author:  Svante Lindahl <zap@cs.kth.sunet>, Marcel Mol <marcel@duteca.et.tudelft.nl>
Version:  5.1-8

dviutils Tools for working on DVI files

This package contains some utilities for manipulating DVI documents. Page selection and rearrangement are supported, including arrangement into signatures for booklet printing, and page merging for n-up printing. dvibook rearranges pages into book signatures, adds blank pages as needed dviselect selects pages and page ranges dvitodvi performs general page rearrangement of selected pages dviconcat concatenation of DVI files Included you will find three simple scripts for scaling DIN A4 pages and rearranging two by two onto a DIN A4 page (a4toa5). But the number of pages must be even (which can be done by odd2even or dvibook).

Contents: Dviutils

Copyright:  (c) 1987, 1989 University of Maryland
Author:  Chris Torek (University of Maryland), Angus Duggan
Version:  94.3.19-5

l2h LaTeX2HTML converter

LaTeX2HTML lets you easily convert not too complex LaTeX-documents into the WWW format HTML. This enables one to have both a written and online version of even older LaTeX texts. Please look out for the corresponding variables in /etc/rc.config!!! Documentation may be found in: /usr/doc/packages/l2h

Contents: latex2html-96.1.h

Copyright:  (c) 1993 Nikos Drakos
Author:  Nikos Drakos <nikos@cbl.leeds.ac.uk>
Version:  96.1.h-9

l2hhtml Documentation for the l2h package

The LaTeX2html manual, in HTML, readable by any HTML browser. Call it e.g. by saying: arena file:/usr/doc/packages/l2h/manual/manual.html

Contents: l2hhtml-96.1

Copyright:  (c) 1993, Nikos Drakos
Author:  Nikos Drakos <nikos@cbl.leeds.ac.uk>
Version:  96.1-8

lyx LyX - LaTeX based WYSIWYG editor

LyX is a frontend for LaTeX under X11. It combines the comfortable use of a word processor with the high quality of LaTeX type setting. Documents are almost WYSIWYG-like displayed. Low level

TEX: TeX/LaTeX and applications

attributes (”large italic”) are not directly chosen, but high level layouts for each paragraph. Of course, low level formatting is still possible. Using the LaTeX compiler, the LyX document is translated into a .dvi file. The reverse process, to process an already existing LaTeX document with the LyX editor, is not possible (except as a plain ASCII file). Contents: lyx-0.12.0

Copyright:  (c) 1995-97 Matthias Ettrich
Author:  Matthias Ettrich <ettrich@informatik.uni-tuebingen.de>
Version:  0.12.0pre8-1

musictex MusiXTeX

This packages is only usable for TeX-experts with musical ambitions. It contains a music Tex format file musixtex for direct usage onto TeX music files. LaTeX users should use the corresponding style file. For a description you should look under /usr/lib/teTeX/texmf/doc/musixtex/. Contents: musixexa-T86 musixtex-T86

Copyright:  (c) 1995 Daniel Taupin

Author:

Daniel Taupin <taupin@lps.u-psud.fr>, Ross Mitchell <rmm@dar.csiro.au>, Andreas Egler <Andreas Egler@wob.maus.de>

Version:  5.15-8

psutils Tools for manipulating PostScript files

This archive contains some utilities for manipulating PostScript documents. Page selection and rearrangement are supported, including arrangement into signatures for booklet printing, and page merging for n-up printing. psbook rearranges pages into signatures psselect selects pages and page ranges pstops performs general page rearrangement and selection psnup put multiple pages per physical sheet of paper psresize alter document paper size epsffit fits an EPSF file to a given bounding box You will find a README under /usr/doc/packages/psutils/ which also describes several perl-scripts for importing PostScript files. A manual page for each ps utility is also included. Contents: psutils-p17

Copyright:  (c) 1991-1995 Angus J. C. Duggan
Author:  Angus J. C. Duggan <angus@harlequin.co.uk>
Version:  p17-4

tetex TeX, LaTeX, BibTeX, makeindex, MetaFont, MetaPost, dvips, xdvi

After installing teTeX you will find a large selection of documentation for TeX, LaTeX, and various layout styles under /usr/lib/teTeX/texmf/doc. TeX (pronounced tech) is an interpreter for text formating and was developed by Donald E. Knuth. It works with control and macro commands on a text file. Working with TeX is similar to good old typesetting methods. LaTeX or better LaTeX2e is a complex macro package which removes the cryptical TeX interface, and does most of the work for the user. TeX uses special fonts produced by the MetaFont program. Various printer drivers and an X11 viewer is also included in this package. The teTeX package is based on the standard TeX package of Karl Berry which makes configuration much easier. It’s also possible to use PostScript fonts; a real PostScript printer is required however. If the ghostscript (gs) package is installed all drivers for printing and viewing can use these fonts. Note, however that the fonts included in the ghostscript package are not identical to Adobe’s PostScript fonts. The copyright prohibits us from including them on the CD. Beside these features there are the programs MakeIndex (for producing indexes), BibTeX (for literature data-processing), and the script xtexsh (requires tcl/tk). Note the differences between LaTeX2.09 and LaTeX2e. LaTeX2e can process original LaTeX2.09 text but style files that require LaTeX2.09 internal macros may not work correctly. Under /usr/lib/teTeX/texmf/doc/latex/base/ you will find the documentation for LaTeX2e. New users should review the documentation stored there. The teTeX package from Thomas Esser includes a full texmf tree, many programs (tex, dvips, ...), shell scripts configuration, and a big collection of documentations. This package is simple configured by the script texconfig and has thereby multi-language options. Do NOT install this package together with ntex! teTeX is NOT an upgrade or update for the ntex package!! Contents: ec-ready-mf-tfm teTeX-lib-0.4pl8 teTeX-src-0.4pl7 texmf-src-0.4pl8

Author:

Karl Berry , Tim Morgan, Tomas Rokicki, Donald Ervin Knuth, Thomas Esser <te@informatik.uni-hannover.de>

Version:  0.4pl8-13

texcad Drawing program for LaTeX pictures

A really nice drawing program which is especially useful in cooperation with LaTeX. Please note the documentation in /usr/doc/packages/texcad and the file COPYRIGHT there in. Contents: xtexcad-2.4

Copyright:  (c) Klaus Zitzmann, 1993-1996 Johannes Sixt

Author:

Klaus  Zitzmann  <zitzmann@infko.uni-koblenz.de>,  J.K.Wight

<J.K.Wight@newcastle.ac.uk>, Byron Rakitzis <byron@netapp.com>, Johannes Sixt <Johannes.Sixt@eudaptics.co.at>, Notker Amann <amann@isr.uni-stuttgart.de>

Version:  2.4-8

tktex tktex

TkTeX is a Tk-based interface to the UN*X tools related to TeX, such as editors, previewers, printing programs, etc. Contents: tktex-2.2

Copyright:  (c) Norman Danner
Author:  Norman Danner <ndanner@indiana.edu>
Version:  2.2b1-10

transfig Graphic converter

TransFig is a set of tools for creating TeX documents with graphics which are portable, in the sense that they can be printed in a wide variety of environments. The transfig directory contains the source for the transfig command which generates a Makefile which translates Fig code to various graphics description languages using the fig2dev program. In previous releases, this command was implemented as a shell script. Documentation: man transfig Contents: transfig.3.2

Copyright:

(c) 1985-1994 Supoj Sutantavibul, Frank Schmuck, Konrad Kwock, Herbert Bauer, Micah Beck

Author:

Anthony Starks <ajs@merck.com>, George Ferguson <ferguson@cs.rochester.edu>, Herbert Bauer <heb@regent.e-technik.tu-muenchen.de>, Micah Beck <supoj@sally.utexas.edu>, Supoj Sutantavibul <beck@cs.utk.ecu>

Version:  3.2-9

ts Tcl/Tk based shell for LaTeX

X: Base X Window System - XFree86^{TM}3.3

My aim was writing a TeX environment similar to the DOS-TeXShell by Schlegelmilch. That is, you can use their short keys for editing (F3), composing (F6), or showing (F8). The complete list of keys is shown in the online help (press F1). Menus, Keys, and Warnings are in English, but you can use German or English online help. Keys, warnings end menus are in English (since they are shorter) but help is available in German and English. Will be installed in /usr/lib/ts and some READMEs under /usr/doc/packages/ts If you want to use the English help-system, change the symlink under /usr/lib/ts to the right path (there is no parameter): cd /usr/lib/ts ln -sf eng help Contents: ts-9710

Copyright:  (c) 1996 Jens P”onisch
Author:  Jens P”onisch <J.Poenisch@wirtschaft.tu-chemnitz.de>
Version:  9710-6

xfig Facility for Interactive Generation of figures under X11

Xfig is a menu-driven tool that allows the user to draw and manipulate objects interactively in an X window. The resulting pictures can be saved, printed on PostScript printers or converted to a variety of other formats (e.g. to allow inclusion in LaTeX documents). Documentation: /usr/doc/packages/xfig & man xfig Examples: /usr/doc/packages/xfig/Examples

Contents: xfig.3.2

Copyright:  (c) 1985 Supoj Sutanthavibul, 1991-94 Paul King, Brian V. Smith

Author:

Anthony Dekker <dekker@ACM.org>, Brian V. Smith <bvsmith@lbl.gov>, Jim Daley <jdaley@cix.compulink.co.uk>, Ross Martin <martin@trcsun3.eas.asu.edu>, Uwe Bonnes <bon@lte.e-technik.uni-erlangen.de>

Version:  3.2-9

xtem a TeX menu system for X-Window

xtem is a TeX environment under X. This package is rather powerful and offers a huge help and reference system which should help the TeX novices. TeX administrators estimate this environment. The authors of this package ask for feedbacl for continuing their work! For more information lookup /usr/doc/packages/xtem.

Contents: xtem src xtem texmenu.6.18

Copyright:  (c) 1996 G. Lamprecht, W. Lotz, R. Weibezahn; IWD, Bremen University
Author:  G. Lamprecht, W. Lotz, R. Weibezahn <weibezahn@iwd.uni-bremen.de>
Version:  6.18-2

0.1.17 X: Base X Window System - XFree86^{TM}3.3

xaw3d 3D Athena Widgets

This is a library that can be used instead of the standard Athena-Widget-Library. It was tried to keep the standard of the libXaw library. There are also programs which explicitely use this library (this is the reason why the library was included). NOTE: Do NOT replace /usr/X11R6/lib/libXaw.so.6.1!

Contents: Xaw3d-1.3

Copyright:  (c) 1989, 1994 X Consortium

Author:

Kaleb S. Keithley <kaleb@x.org>, Achille Hui <eillihca@drizzle.stanford.edu>, Robert Withrow <witr@rwwa.com>, Malcolm Strickland <chuckstrickland@orl.mmc.com>, David Flanagan <david@ora.com>, Jamie Zawinski <jwz@netscape.com>, Frank Terhaar-Yonkers <fty@bizarre.trpnc.epa.gov>, Mike Schulze <mike@cs.curtin.edu.au>, Gustaf Neumann <neumann@mohegan.wiinf.uni-essen.de>, Jerry Whelan <guru@stasi.bradley.edu>, Mark Rawling <mwr@mel.dit.csiro.au>, Tim Theisen <tim@cs.wisc.edu>, Mitch Trachtenberg <mitch@mta.com>, Keith Packard <keithp@ncd.com>, Dimitri P. Bouras <dimitri@ee.ubc.ca>, Uri Blumenthal <uri@watson.ibm.com>, Heiko Schroeder <heiko@pool.informatik.rwth-aachen.de>

Version:  1.3-8

xaw3dd 3D Athena Widgets as replacement for Athena Widgets

This version of libXaw3d will be installed in /usr/X11R6/lib/Xaw3d/ as libXaw.so.6.1. Note that this version is linked with the option ‘-soname libXaw.so.6’ and therefore it will be able to replace the regular libXaw.so.6.1 in /usr/X11R6/lib/. NOTE: Do NOT replace /usr/X11R6/lib/libXaw.so.6.1. All necessary changes in /etc/ld.so.conf will be done by S.u.S.E configuration done bei YaST. On any problem using or starting X11 programs please remove this package and delete any remaining symbolic link under /usr/X11R6/lib/Xaw3d/.

On any problem starting or using X11 programs, you should remove this package. Please read the package description.

Author:

Kaleb S. Keithley <kaleb@x.org>, Achille Hui <eillihca@drizzle.stanford.edu>, Robert Withrow <witr@rwwa.com>, Malcolm Strickland <chuckstrickland@orl.mmc.com>, David Flanagan <david@ora.com>, Jamie Zawinski <jwz@netscape.com>, Frank Terhaar-Yonkers <fty@bizarre.trpnc.epa.gov>, Mike Schulze <mike@cs.curtin.edu.au>, Gustaf Neumann <neumann@mohegan.wiinf.uni-essen.de>, Jerry Whelan <guru@stasi.bradley.edu>, Mark Rawling <mwr@mel.dit.csiro.au>, Tim Theisen <tim@cs.wisc.edu>, Mitch Trachtenberg <mitch@mta.com>, Keith Packard <keithp@ncd.com>, Dimitri P. Bouras <dimitri@ee.ubc.ca>, Uri Blumenthal <uri@watson.ibm.com>, Heiko Schroeder <heiko@pool.informatik.rwth-aachen.de>

Version:  1.3-8

xaw95 3D Athena Widgets with W95 look and feel

This is a library that can be used instead of the standard Athena-Widget-Library. It was tried to keep the standard of the libXaw library. There are two shared libraries: one with the name libXaw95.so.6.1 and one with the name libXaw.so.6.1 linked with ‘-soname libXaw.so.6’. This laste version is installed in /usr/X11R6/lib/Xaw95/ By setting the variable LD LIBRARY PATH it is possible to load this pseudo libXaw.so.6.1 first. NOTE: Do NOT replace /usr/X11R6/lib/libXaw.so.6.1. Contents: Xaw95-1.1

Copyright:  (c) 1989, 1994 X Consortium

Author:

Kaleb S. Keithley <kaleb@x.org>, Achille Hui <eillihca@drizzle.stanford.edu>, Robert Withrow <witr@rwwa.com>, Malcolm Strickland <chuckstrickland@orl.mmc.com>, David Flanagan <david@ora.com>, Jamie Zawinski <jwz@netscape.com>, Frank Terhaar-Yonkers <fty@bizarre.trpnc.epa.gov>, Mike Schulze <mike@cs.curtin.edu.au>, Gustaf Neumann <neumann@mohegan.wiinf.uni-essen.de>, Jerry Whelan <guru@stasi.bradley.edu>, Mark Rawling <mwr@mel.dit.csiro.au>, Tim Theisen <tim@cs.wisc.edu>, Mitch Trachtenberg <mitch@mta.com>, Keith Packard <keithp@ncd.com>, Dimitri P. Bouras <dimitri@ee.ubc.ca>, Uri Blumenthal <uri@watson.ibm.com>, Heiko Schroeder <heiko@pool.informatik.rwth-aachen.de>, iEddie Hiu-Fung, Lau <eddielau@cuhk.edu.hk>

Version:  1.1-7

xdevel Developing under XFree86^{TM}

This package contains the include files, libraries and the imake templates for the X-Window system. It is needed for the development and the compilation of X11 programs.

Version:  3.3.2-3

xdmbgrd S.u.S.E. Linux background

The S.u.S.E. Linux background for your xdm workstation. Contents: xdmbgrd

Copyright:  (c) Copyright (c) 1997 S.u.S.E. GmbH Fuerth
Author:  Werner Fink <werner@suse.de>
Version:  0.2-4

xdoc Japanese documentation of XFree86^{TM}

The English documentation is part of xf86.

Version:  3.3.2-3

xextra Two hardware independent X-servers, Xvfb and Xnest

Xvfb - a virtual framebuffer X server. Xvfb is an X server that can run on machines with no display hardware and no physical input devices. It emulates a dumb framebuffer using virtual memory. Xnest -- a nested X server. Xnest is a client of the real server which manages windows and graphics requests on its behalf. Xnest is a server to its own clients. Xnest manages windows and graphics requests on their behalf.

Version:  3.3.2-3

xf86 Basic X11 package

Contains XFree86^{TM}and some other programs from the contrib-directory. Contents: XFree86-3.3.2 latin2-fonts

Copyright:  (c) 1994, 1995, 1996 The XFree86 Project, Inc.

Author:

Alan  Hourihane  <alanh@fairlite.demon.co.uk>,  Bernhard  Bender

<br@elsa.mhs.compuserve.com>, Craig Groeschel <craig@metrolink.com> , David Dawes <dawes@XFree86.Org>, David Wexelblat <dwex@XFree86.Org>, Dirk Hohndel <hohndel@suse.de>, Erik Nygren <nygren@mit.edu>, Gertjan Akkerman <akkerman@dutiba.twi.tudelft.nl>, Glenn Lai <glennlai@netcom.com>, Hans Nasten <nasten@everyware.se>, Harald Koenig <koenig@tat.physik.unituebingen.de>, Harm Hanemaayer <hhanemaa@cs.ruu.nl>, Harry Langenbacher <harry@brain.jpl.nasa.gov>, Henrik Harmsen <harmsen@eritel.se>, Henry Worth <henry.worth@amail.amdahl.com>, Jim Tsillas <jtsilla@ccs.neu.edu>, Joe Moss <joe@morton.rain.com>, Jon Block <block@frc.com>, Jon Tombs <tombs@XFree86.org>, Kevin Martin <martin@cs.unc.edu>, Leonard N. Zubkoff <lnz@dandelion.com>, Marc Evans <marc@XFree86.org>, Marc La France <Marc.La-France@ualberta.ca> , Michael Rohleder <michael.rohleder@stadtfrankfurt.de>, Mike Hollick <hollick@graphics.cis.upenn.edu>, Norbert Distler <Norbert.Distler@physik.tu-muenchen.de> , Orest Zborowski <orestz@eskimo.com>, Paolo Severini <lendl@dist.dist.unige.it>, Pascal Haible <haible@izfm.uni-stuttgart.de>, Peter Trattler <peter@sbox.tu-graz.ac.at>, Rich Murphey <rich@XFree86.org>, Rik Faith <faith@cs.unc.edu>, Robin Cutsha, <robin@XFree86.org>, Simon Cooper <sc@sgi.com>, Stuart Anderson <anderson@metrolink.com>

Version:  3.3.2-3

xfnt100 100dpi screen fonts for the X-server

These fonts are compressed to save disk space.

Version:  3.3.2-3

xfntbig Big fonts for the X-server

These fonts are compressed to save disk space. These are the asiatic fonts (including Kanji, Jiskan) which the ’average’ european user won’t need..;-)

Version:  3.3.2-3

xfntcyr Cyrillic fonts

These fonts are compressed to save disk space.

Version:  3.3.2-3

xfntscl Scalable fonts for X11

Speedo and Type 1 fonts for XFree86^{TM}.

Version:  3.3.2-3

xfsetup New XFree86 configuration tool ”XF86Setup”.

Please also install the X server for your graphics board and invoke the configuration tool by entering XF86Setup...

Copyright:  (c) 1994, 1995, 1996 The XFree86 Project, Inc.
Version:  3.3.2-3

xlkit XFree86^{TM}Linkkit

If you want or need to rebuild an XFree86^{TM}server but don’t want to install the full source for XFree86-3.3.1, you need this package in addition to the base configuration and the programmers kit.

Version:  3.3.2-3

xman on-line manuals for XFree86^{TM}

Manpages for programs that come with XFree86^{TM}.

Version:  3.3.2-3

xpm The X-Pixmap libraries

The shared libraries are needed to run programs compiled with Xpm, such as xfilemanager and fvwm.

This package contains libXpm.so.4.

Contents: xpm-3.4j

Author:  Arnaud Le Hors <lehors@sophia.inria.fr>
Version:  3.4j-6

xshared X11 shared libs

Since some programs were linked with the X11 libraries but do work on a text console as well (e.g. emacs), we have put the X11 dynamic libraries into an extra package so you can use such programs without installing the entire X11 system.

Version:  3.3.2-3

xxprt X Print Server

Contains the X print server.

Copyright:  (c) 1994 - 1997 The XFree86 Project, Inc.
Version:  3.3.2-3

xdmsc Xterminal -- Usage S.u.S.E. Linux as X terminal

Some useful scripts for either using S.u.S.E. Linux as X terminal or to provide a comfortable xdm server under S.u.S.E. Linux. You will find the documentation in the following directory /usr/doc/packages/xdmsc/

Contents: Xterminal-0.1

Copyright:  (c) 1997 S.u.S.E GmbH
Author:  Werner Fink <werner@suse.de>
Version:  0.1-8

xvga16 server for vga cards (16 colors)

This server uses standard video modes but only supports 16 colors. It should work with almost any video-hardware.

Version:  3.3.2-3

x8514 server for 8514-based cards

Please only install if you have a fully 8514 compatible graphics card.

Version:  3.3.2-3

xagx server for agx-cards

X-Server for graphic cards based on the AGX chipset. If there is no card installed that uses this chipset you must not install this server.

Version:  3.3.2-3

xi128 Server for Number Nine Imagine 128 graphic cards

Version:  3.3.2-3

xmach32 Mach32 Server

X-Server for graphic cards based on the Mach32 chipset. If this card is not installed on your system you should not install this server.

Version:  3.3.2-3

xmach64 Mach64 Server

If you don’t have this special card installed in you computer you should not install this server.

Version:  3.3.2-3

xmach8 Mach8 server

If you don’t have this special card installed in you computer you should not install this server.

Version:  3.3.2-3

xmono X-monochrome server

If you don’t have a monochrome card installed in your computer you should not install this server.

Version:  3.3.2-3

xp9k Accelerated server for P9000-based cards

This server is designed for the following cards: Diamond Viper VLB, Diamond Viper PCI, Orchid P9000 and some clones (STAR 2000). Cards like Viper Pro and other P9100 and P9130 cards are not yet supported.

Version:  3.3.2-3

xs3 server for S3-based cards (excluding ViRGE and ViRGE/VX)

This server supports the following chipsets: 911, 924, 801/805, 928, 864, 964, Trio32, Trio64, Trio64V+, Trio64UV+, Aurora64V+, Trio64V2/DX and /GX, but not ViRGE and ViRGE/VX. It does not necessarily work with every card that is using one of the supported chipsets. For more information please refer to the chapter about X11 in our handbook.

Version:  3.3.2-3

xs3v server for S3 (ViRGE and ViRGE/VX)-based cards

This server is exclusively for cards with S3 ViRGE and ViRGE/VX. Only install it if you have such a card. Staring with XFree86-3.3 the SVGA server has S3V support as well and is to be prefered over the standalone S3V server.

Version:  3.3.2-3

xsvga server for svga cards

If you don’t have a SVGA card installed you shouldn’t install this server. Please check the handbook on details about the chipsets supported. This server contains experimental suport for the Matrox Millennium II.

Version:  3.3.2-3

xw32 server for W32 cards

This is the server for ET4000-W32 and ET6000 cards. If you don’t have such a card you shouldn’t install it. The SVGA server has better and faster support for most of these cards and offers high and true color on many W32 and all ET6000 cards.

Version:  3.3.2-3

xglint Accelerated server for GLINT/PERMEDIA/PERMEDIA-2 based cards

Currently the graphics boards Gloria-L/MX, GLoria-L, GLoria-S, GLoria Synergy and Winner 2000/Office from Elsa are supported! Only install it if you have one of these cards installed onto your system. Please notify, that this X server can ONLY be configured with ”xf86config”. Unfortunately the configuration with ”XF86Setup” is not possible.

Copyright:  (c) 1997-1998 S.u.S.E. GmbH / 1992-1998 The XFree86 Project, Inc.

Author:

Alan Hourihane <alanh@fairlite.demon.co.uk>, Dirk Hohndel <hohndel@suse.de>, Stefan Dirsch <sndirsch@suse.de>, Helmut Fahrion <hf@suse.de>,

Version:  3.02-2

oldlibs5 Libraries from XFree86^{TM}2.1.1 (X11R5)

This package installs libX11.so.3.1.0, libXaw.so.3.1.0 and libXt.so.3.1.0 libraries in the /usr/X11R6/lib directory. Almost all programs are linked with the new libraries, however there are a few binaries that require the old libraries. Contents: oldlibs5

Version:  97.6.29-3

oldlibs6 a.out (DLL) shared libraries for X11 binaries

These libraries are needed by X programs in the old a.out binary format. They will be installed in /usr/i486-linuxaout/lib. This package also contains the ”.sa” libraries which allow the generation of a.out binaries using ’gcc -b i486linuxaout’.

Version:  97.7.23-0

modegen Roll your own X video mode

Two ’sc’ spreadsheets to create X mode lines. Remember that this package is ’for experts only’ - normally XF86Setup (in package xfsetup) does everything you need. Contents: modegen

Copyright:  (c) 1994 Mike Jagdis
Author:  Mike Jagdis <jaggy@purplet.demon.co.uk>
Version:  98.2.3-2

0.1.18 XAP: X Applications

addrbo Address book

The program ”addressbook” lets you manage your addresses, phone numbers, e-mail addresses, and fax numbers. Contents: addrbo src addressbook-0.6.1

Copyright:  (c) Clemens Durka
Author:  Clemens Durka <durka@informatik.tu-muenchen.de>
Version:  0.6-9

arena WWW-Browser

This WWW browser is one of the few that do not fall under the restrictive copyright conditions of the HTML library. We greatly regret that we are not allowed to distribute popular browsers such as Netscape, Mosaic, and Chimera on our CDs, although these are freely available for private use in the Internet. The copyright conditions set by Netscape and Mosaic forbid us from doing so. Until Arena’s 1.0 version is released, this package will remain a binary only package. This is why the sources are not available. Contents: arena

Copyright:  (c) 1994 CERN

Author:

Hakon W Lie <arena@info.cern.ch>, Dave Raggett <arena@info.cern.ch> , Henrik <arena@info.cern.ch>, Phill <arena@info.cern.ch>

Version:  beta.1e-8

asedit An Editor

A simple ASCII text editor for X11/Motif. Mouse support, dialog boxes, hypertext online help, undo/redo. The help box for the online help has a little bug: it is too small, you must enlarge the box with the mouse. Contents: asedit-1.3.2

Copyright:  (c) 1994 Andrzej Stochniol
Author:  Andrzej Stochniol <A.Stochniol@ic.ac.uk>
Version:  1.3.2-12

auis ful Andrew User Interface System - Full System

AUIS is a collection of tools which enable you to design, view and mail multimedia documents. This is a complete AUIS Distribution. It is designed for users who want to create their own AUIS applications. If this isn’t what you want, you should consider installing the smaller packages. Please read also the README under /usr/andrew. Contents: auis1 auis2 auis3 auis4 contrib doc This package also includes the other AUIS packages. So please install either this package, or the other packages!

Copyright:  (c) 1992 Carnegie Mellon University & IBM
Author:  Andrew Consortium <info-andrew-request@andrew.cmu.edu>
Version:  635L1-0

auis ml Andrew User Interface System - Mail System

AUIS is a collection of tools which enable you to design, view and mail multimedia documents. This package is aimed at all those who want a ”better mail system” and are interested in a multi-media mail-interface (MIME).

XAP: X Applications

This package requires that the package ”auis-wp” be installed on your system. The X Window System, a running mail system and the ”csh” shell are required as well. Please read also the README files under /usr/andrew.

Copyright:  (c) 1992 Carnegie Mellon University & IBM
Author:  Andrew Consortium <info-andrew-request@andrew.cmu.edu>
Version:  635L1-0

auis src Andrew User Interface System - programming environment

AUIS is a collection of tools which enable you to design, view and mail multimedia documents. This package is designed for application programmers who want to design their own applications using AUIS. This is **NOT** the source code of AUIS! Please read also the README files under /usr/andrew.

Copyright:  (c) 1992 Carnegie Mellon University & IBM
Author:  Andrew Consortium <info-andrew-request@andrew.cmu.edu>
Version:  635L1-0

auis wp Andrew User Interface System - word processing

AUIS is a collection of tools which enable you to design, view and mail multimedia documents. This package contains the main word processor on which all the other packages of AUIS are based. Please read also the README files under /usr/andrew.

Copyright:  (c) 1992 Carnegie Mellon University & IBM
Author:  Andrew Consortium <info-andrew-request@andrew.cmu.edu>
Version:  635L1-0

axe A simple to use X11 editor

This simple to use editor features: multiple windows, multiple buffers, configurable menus and buttons, access to external filters, keyboard macros, comprehensive on-line help, ... This version of aXe-6.1.2 is usable only with the real shared Xaw library. For Xaw3d please use aXe from the axe3d package.

Contents: aXe3D-6.1.2

Copyright:  (c) 1992, 1994 The University of Newcastle upon Tyne
Author:  Jim Wight <j.k.wight@newcastle.ac.uk>
Version:  6.1.2-9

axe3d A simple to use X11 editor

This simple to use editor features: multiple windows, multiple buffers, configurable menus and buttons, access to external filters, keyboard macros, comprehensive on-line help, ... This version of aXe-6.1.2 is usable only with the real shared Xaw library. For Xaw3d please use aXe from the axe3d package.

Contents: aXe3D-6.1.2

Copyright:  (c) 1992, 1994 The University of Newcastle upon Tyne
Author:  Jim Wight <j.k.wight@newcastle.ac.uk>
Version:  6.1.2-8

babaya Tool to exit X11 gently

A tool to exit X11 gently. Have you always been annoyed by the error messages (... broken pipe) X11 produces when exiting ? Click the right mouse button in the babaya window for a ”clean” shutdown of X11. Documentation: /usr/doc/babaya

Contents: babaya-1.2b

Copyright:  (c) 1992-1993 - Muhammad M. Saggaf
Author:  Muhammad M. Saggaf <alsaggaf@mit.edu>
Version:  1.2b-8

calctool A calculator

A simple scientific calculator, which can also represent the values also in binary, hexadecimal or decimal format.

Contents: calctool-2.4.12

Copyright:  (c) Rich Burridge
Author:  Rich Burridge <richb@sunaus.oz.au>
Version:  2.4.12-8

cellular Cellular automate simulation program

automates named cellang and a compiler that makes programms for the X-Window-System from it. You can find the Documentation for cellang and a tutorial in /usr/doc/packages/cellular. Examples are in /usr/lib/cellular/examples. Documentation: man cellc, avcam, tpipe Contents: cellular-07apr94

Copyright:  (c) 1992, 1993, 1994 J Dana Eckart
Author:  J Dana Eckart <dana@rucs.faculty.cs.runet.edu>
Version:  94.4.7-5

conv Conversion utility (units of measurement)

Copyright:  (c) GPL
Author:  Stefan Heimers <sheimers@ezinfo.vmsmail.ethz.ch>
Version:  1997.08.11-7

coolmail Cool XBiff clone

Shows if new mail has arrived. The main advantage of this program is that you can click into the window to start the mailer. Documentation: man coolmail Contents: coolmail-1.3

Copyright:  (c) 1994 Byron C. Darrah
Author:  Byron C. Darrah <darrah@kaiwan.com>, Randall K. Sharpe <rsharpe@ncsa.uiuc.edu>
Version:  1.3-9

elem Table of chemical elements

Copyright:  (c) GPL
Author:  Stefan Heimers <sheimers@ezinfo.vmsmail.ethz.ch>
Version:  1998.01.20-2

fader Screen brightness adjuster

Adjusts the screen brightness. There is a version for X11 and one for the console. Contents: fader-1.01

Copyright:  (c) 1994 Klaus Thiele
Author:  Klaus Thiele <100520.634@compuserve.com>
Version:  1.01-7

felix Computer algebra system FELIX

More information about this package can be found in /usr/doc/felix-2.1. Unfortunately, the sources are not available. Contents: felix-2.1

Copyright:  (c) 1990-93 Joachim Apel & Uwe Klaus

Author:

Joachim Apel <japel@informatik.uni-leipzig.de>, Uwe Klaus <uklaus@informatik.unileipzig.de>

Version:  2.1-8

gnuplot gnuplot

Gnuplot is a command-line driven interactive function plotting utility. Gnuplot supports many different types of terminals, plotters, and printers (including many color devices, and pseudo-devices like LaTeX) and can easily be extended to include new devices. Contents: gnuplot-beta336

Copyright:  (c) 1986 - 1993 Thomas Williams, Colin Kelley

Author:

Mailingliste  <bug-gnuplot@dartmouth.edu>,  Mailingliste  <info-

gnuplot@dartmouth.edu>, Mailingliste <info-gnuplot-beta@dartmouth.edu>, Mailingliste <info-gnuplot-request@dartmouth.edu>

Version:  3.5beta6.336-9

gvim vi with GUI

README.SuSE for package ’gvim’.

This package includes only a new binary with GUI (3D-Athena-Widgets) and a resource file (/etc/gvimrc). The normal package ’vim’ is needed! Start: /usr/X11R6/bin/gvim Copy and modify /etc/gvimrc to ˜ /.gvimrc if needed.

Author:  Bram Moolenaar <mool@oce.nl>
Version:  4.5-8

knews Threaded news reader under X11

Knews is a threaded news reader with an X Window interface that uses NNTP to get news. It displays the threads in a graphical tree. For more information read the man page (”man knews”).

Contents: knews-0.9.8

Copyright:  (c) Copyright 1995 by Karl-Johan Johnsson
Author:  Karl-Johan Johnsson <mathematik.su.se>
Version:  0.9.8-9

mfm a file manager for X11

Copyright:  (c) 1995, Andrew T. Renalds
Author:  Andrew T. Renalds <renalds@abn.com>
Version:  1.3-12

mgdiff X-diff

Mgdiff is a graphical front end to the Unix diff command based upon X11 and the Motif widget set. It allows the user to select two files for comparison, runs the diff command, parses the output and presents the results graphically.

Contents: mgdiff-1.0

Copyright:  (c) Copyright (c) 1994 Daniel Williams
Author:  Daniel Williams <dan@sass.com>
Version:  1.0-13

mupad MuPAD - Multi Processing Algebra Data Tool

MuPad is a program system for doing computer algebra. It is being developed at the University of Paderborn, Germany. Computer algebra systems providei symbolic handling of mathematical formulae and, supplied with expert knowledge, make it easier to handle. They allow interactive manipulation of mathematical objects, effortless functional programming of particular mathematical problems. Their immense effect on mathematical lessons at universities and high schools has already been noted. Even schools are starting to use it as part of their lessons. MuPAD is one of the first universal computer algebra systems. It is already well known for its object oriented language as well as for its steady growing implementations of mathematical algorithms. An extraordinary comfortable GUI, an interactive on-line help system, the included debugger and a powerful graphics tool are the main advantages of MuPAD. In 1993 MuPAD has been awarded with the ”German-Austrian software price” followed by the First European Software Award in 1994. Before you start using MuPAD you should read the README file which is located under /usr/MuPAD. This file explains how to accustom your environment to fully use MuPAD.

Contents: mupad-1.3.1

Copyright:  (c) B. Fuchssteiner, Automath, University of Paderborn, Germany

Author:

Prof. Dr. Benno Fuchssteiner <benno@uni-paderborn.de> , Ralf Hillebrand <tonner@uni-paderborn.de>

Version:  1.3.1-2

nedit Full featured plain text ”Nirvana Editor”

NEdit is a GUI style plain text editor for workstations with X Windows and Motif. NEdit provides all of the standard menu, dialog, editing, mouse support, as well as macro extension language, Syntax highlighting, and a lot other nice features (and extensions for programmers).

Contents: nedit-5.0.1

Copyright:  (c) (c) 1991 Universities Research Association, Inc.
Author:  Mark Edel <edel@fnal.gov>
Version:  5.0.1-1

offix some programs with drag and drop features

This package contains a clipboard, editor, executor, file manager and a printer frontend with drag and drop features. Contents: offix-2.1.orig

Copyright:  (c) 1996 Csar Crusius, Randolf Werner, Andr Hentz
Author:  Csar Crusius <c.crusius@ieee.org>, Randolf Werner , Andr Hentz <a.hentz@ieee.org>
Version:  2.1.2debian-11

pan note manager

Post a Note. A quite useful note manager. Contents: pan-2.04

Copyright:  (c) 1992 Jeffrey W. Bailey
Author:  Jeffrey W. Bailey <jeff@rd1.interlan.com>
Version:  2.04-8

pcb Interactive printed circuit board design

Pcb is a handy tool for the X Window System built to design printed circuit boards. All coordinate units are 1/1000 inch. For details see the manual which can be found under /usr/doc/packages/pcb. There is also a man page available. Contents: pcb-1.5.3.ALPHA

Copyright:  (c) 1994-1996 Thomas Nau
Author:  Thomas Nau <Thomas.Nau@rz.uni-ulm.de>
Version:  1.5.3.ALPHA-7

pixedit Pixel Color Editor

A tool to modify colors using several scrollbars. The program to find the colors you ”really” like :-) Contents: pixedit-1994

Copyright:  (c) 1994 Mike Yang
Author:  Mike Yang <mikey@sgi.com>
Version:  1994-10

plan Clear and easy to use schedule planner

This is a really useful program. It is statically linked with the Motif library. If you have Motif installed on your system, you can recompile the program and save a lot of main memory. Contents: plan-1.6.1

Copyright:  (c) 1993-97 Thomas Driemeyer
Author:  Thomas Driemeyer <thomas@bitrot.in-berlin.de>
Version:  1.6.1-15

plotmtv A multipurpose X11 plotting program

PLOTMTV’s capabilities include 2D line and scatter plots (x-vs-y), contour plots, 3D surface, line and scatter plots as well as vector plots. The program has an rough but functional Graphical User Interface, through which it is possible to zoom in, zoom out, pan, toggle between 2D and 3D plots, and rotate 3D plots. Both color and gray-scale PostScript output are supported. Contents: Plotmtv1.4.1

Copyright:  (c) 1995 Kenny K.H. Toh
Author:  Kenny K.H. Toh <ktoh@td2cad.intel.com>
Version:  1.4.1-8

procmtr Display actual system parameters

With procmeter one can display various system parameters as e.g. processor load, network load, etc. Documentation may be found under /usr/doc/packages/procmtr Contents: procmeter-2.2

Copyright:  (c) 1994,95,96 Andrew M. Bishop
Author:  Andrew M. Bishop <amb@gedanken.demon.co.uk>
Version:  2.2-8

qddb relational database suit with tcl interface

Qddb is a data manipulation suite that allows you to add, modify, and delete records. Records may contain any ASCII characters and you can search on any fields or across fields with: * word ranges * numeric ranges * date ranges * words * numbers * dates * regular expressions You can also combine searches with intersection, union and binary exclusion. See the documentation in /usr/doc/packages/qddb for more details. Contents: qddb-1.43

Copyright:  (c) 1993, 1994 Herrin Software Development, Inc.
Author:  Eric H. Herrin II <eric@ms.uky.edu> , Raphael A. Finkel <raphael@ms.uky.edu>
Version:  1.43-12

qps visual process status monitor

qps is an X11 version of top or ps, with motif- or windows-like interface. Contents: qps-1.2

Author:  Mattias Engdegrd <f91-men@nada.kth.se>
Version:  1.2-6

rasmol Molecular Graphics Visualization tool

RasMol is an X Window System tool intended for the visualization of proteins and nucleic acids. It reads Brookhaven Protein Database (PDB) files and interactively renders them in a variety of formats on either an 8bit or 24/32bit color display. Documentation is in /usr/doc/rasmol & man rasmol Examples are in /usr/lib/rasmol Contents: RasMol-2.5.1

Copyright:  (c) 1992-1994 by Roger Sayle
Author:  Roger Sayle <ras32425@ggr.co.uk>
Version:  2.5.1-8

rxvt The ’other’ X-Terminal.

rxvt is smaller as xterm. That’s why many people like it. This package also includes rclock - it is now included in the rxvt source package. Contents: rxvt-2.4.5

Author:

Mark  Olesen  <olesen@me.QueensU.CA>,  Robert  Nation

<nation@rocket.sanders.lockheed.com>

Version:  2.4.5-2

scilab Scilab - large mathematical package

A large mathematical package containing a lot of demos. Please take a look at the license information in /usr/lib/scilab/notice.* . This program may not have a very appealing look, but it is as powerful as some commercial products. Scilab even includes a converter for Maple files. Contents: scilab-2.3.1-src

Copyright:  (c) 1994-97 INRIA, France
Author:  <Scilab@inria.fr>
Version:  2.3.1-8

snns The ”Stuttgart Neural Network Simulator”

Nice simulator of neural networks which has been developed at the university of Stuttgart. All parameters are presented in individual windows. To really use this program a fundamental background of neural networks is required. SNNS (Stuttgart Neural Network Simulator) is a software simulator for neural networks on Unix workstations developed on the university of Stuttgart. The aim of the project is the development of an flexible simulation interface for research and usage of neural networks. SNNS consists mainly of two parts: o simulator kernel written in C o GUI under X11 The simulator kernel works with the internal network-data structures of the neural network and performs all operations by learning or doing. It can also be used separately as a C program embedded in own applications. It supports multiple network topologies and almost like RCS supports the sites concept. SNNS can be extended by the individual user with user-specific activation functions, output functions, site-functions and learning procedures which can easily be connected as simple C programs with the simulation kernel. At the moment of this writing the following algorithms are supported: o back propagation o vanilla (online) BP o BP with Momentum Term and Flat Spot Elimination o Batch BP o Time Delay BP o Counter propagation o Quick prop o Back percolation 1 o Generalized radial-base-functions (RBF) o RProp o ART 1 o ART 2 o ARTMAP o Cascade Correlation o Return Cascade Correlation o Dynamic LVQ o Back propagation through time (for recursive networks) o Batch Back propagation through time (for recursive networks) o Quick propagation through time ” ” ” o Kohonen Feature Maps The GUI XGUI based on the kernel presents you 1/2/3-dimensional views of the neural network and controls the kernel during the runtime of the simulation. Additionally the 2D user-interface provides an integrated network-editor, which can be used to design, manipulate and present neural networks. A complete user manual can be found in the package snnsdoc. The current version of SNNS is to be obtained from ftp.informatik.uni-stuttgart.de /pub/SNNS Contents: SNNSinfo-1.03 SNNSv4.1.Manual.src SNNSv4.1

Copyright:

(c) 1990-96 SNNS-Gruppe, Institut f”ur Parallel und Verteilte H”ochstleistungsrechner (IPVR), Universit”at Stuttgart, Breitwiesenstr. 20-22, D-70565 Stuttgart

Author:

Andreas Zell <zell@informatik.uni-stuttgart.de>, G”unter Mamier, Michael Vogt, Niels Mache, Ralf H”ubner, Kai-Uwe Herrmann, Tobias Soyez, Michael Schmalzl, Tilman Sommer, Artemis Hatzigeorgiou, Sven D”oring, Dietmar Posselt, Tobias Schreiner

Version:  4.1-9

snnsdoc User Manual for SNNS

SNNS (Stuttgart Neural Network Simulator) User Manual, Version 3.3, Universitaet Stuttgart, Fakultaet Informatik, Report No. 3/94 (revised).

Copyright:

(c) 1990-95 SNNS-Gruppe, Institut f”ur Parallel und Verteilte H”ochstleistungsrechner (IPVR), Universit”at Stuttgart, Breitwiesenstr. 20-22, D-70565 Stuttgart

Author:

Andreas Zell <zell@informatik.uni-stuttgart.de>, G”unter Mamier , Michael Vogt , Niels Mache , Ralf H”ubner , Kai-Uwe Herrmann , Tobias Soyez , Michael Schmalzl , Tilman Sommer , Artemis Hatzigeorgiou , Sven D”oring , Dietmar Posselt , Tobias Schreiner

Version:  4.1-9

spice An analog circuit simulator

Spice is an analog circuit simulator Pieces of advice and examples at /usr/doc/packages/spice Contents: spice3f4

Copyright:  (c) 1990 Regents of the University of California
Version:  3f4-9

spicecad SPICECAD is a schematic entry for SPICE

SPICECAD is a graphical frontend to Spice. Spice does NOT have to be installed since it is statically linked to the SpiceCAD binary. Please read also /usr/doc/packages/spicecad/README and README.SuSE!!! Contents: spicecad-1.5

Copyright:  (c) Martin Maschmann

Author:

Martin  Maschmann  <martin.maschmann@t-online.de>,  Martin
Maschmann  <martin.maschmann@vlsi.com>,  Martin  Maschmann

<martin.maschmann@bigfoot.com>

Version:  1.5-9

susecurs S.u.S.E. logo as cursor in X11

With this package a script do chamaeleon is supplied which transform the boring X-cursor into the S.u.S.E. chamaeleon. Contents: do chameleon0.2

Version:  0.2-8

vidtext Viewer for Videotext

A program that allows you to view videotext pages under X11. The program is accompanied with a kernel loadable module that enables access to the hardware of the videotext decoder. Contents: VideoteXt-0.6 CAUTION: This precompiled module supports videotext cards (kits) as described in c’t 11/91. So, if you have another card installed you need to install the sources and to generate the module vtx.o. Information may be found under /usr/doc/packages/vidtext

Copyright:  (c) 1994-1996 Martin Buck
Author:  Martin Buck <martin-2.buck@student.uni-ulm.de>
Version:  0.6-11

x3270 IBM host access tool.

x3270 opens a telnet connection to an IBM host in an X window. Contents: x3270-3.1.1.6

Author:  Paul Mattes <pdm@cnt.com>
Version:  3.1.1.6-10

xarchie X11 tool to search archie servers

This program allows you to search the various archie servers automatically for a given file. You need an internet connection, though. Contents: xarchie-2.0.10

Copyright:  (c) 1989-91 University of Washington

Author:

Alan Emtage <bajan@cs.mcgill.com>, Bill Heelan <wheelan@cs.mcgill.ca>, Brendan Kehoe <brendan@cygnus.com>, George Ferguson <ferguson@cs.rochester.edu>, Peter Deutsch <peterd@bunyip.com>

Version:  2.0.10-8

xautolck lock the screen after a certain time automagically

This program uses another locker program to lock the screen automagically after an idle period. It may e.g. be called as follows: xautolock -locker ”xlock -mode blank” using xlock as locker program. Contents: xautolock

Copyright:  (c) Stefan De Troch, Michel Eyckmans

Author:

Stefan De Troch <stefan@specs.philips.com>, Michel Eyckmans <eyckmans@imec.be>

Version:  pl10-8

xbanner X-Background writings and images

How about your company’s name in the background of the login screen? Or your company’s logo? It will surely work with xbanner which is designed for the use with e.g. XDM login screens.

Many example configurations may be found in /usr/X11R6/lib/X11/xbanner and its subdirectories. Documentation lies in /usr/doc/packages/xbanner, in the subdir html you find a HTML version of the documents. Contents: xbanner-1.2

Copyright:  (c) 1996 Amit Margalit
Author:  Amit Margalit <amitm@netvision.net.il>
Version:  1.2-8

xbarcode program to create barcodes

Barcodes of the following types can be created, viewed, and saved in PPM, GIF, PCX, PBM, LaTeX or PostScript format: EAN13, EAN8, UPC, CodeABar, Code 128, Code 3/9, Code 2/5, Code 2/5 Matrix, Code 2/5 Interleaved. By using xbarcode, barcodes can be inserted into documents without any problems (and without doing real ”cut and paste”). A version which runs without X11 is also included (bar). Contents: xbarcode-2.11

Copyright:  (c) 1994 Martin Bauer, Jan Laitenberger
Author:  Martin Bauer <Martin Bauer@S2.MAUS.DE>, Jan Laitenberger
Version:  2.11-8

xbench benchmark for X11

A benchmark for X11. Contents: xbench-0.2

Copyright:  (c) 1989 Siemens
Author:  Claus Gittinger <unido!sinix!claus@uunet.uu.net>
Version:  0.2-6

xcb X11 cut&paste utility

Xcb provides access to the cut buffers built into every X server. It allows the buffers to be manipulated either via the command line, or with the mouse in a point and click manner. The buffers can be used as holding pens to store and retrieve arbitrary data fragments, so any number of different pieces of data can be saved and recalled later. The program is designed primarily for use with textual data. Contents: xcb-2.3

Copyright:  (c) 1992,1993,1994 by Farrell McKay
Author:  Farrell McKay <Farrell.McKay@mpx.com.au>
Version:  2.3-8

xcnf Make settings in X

With xcnf you may take some settings under X: - mouse - system beep - builtin X screen saver The settings may be stored in ˜ /.Xcnf and can be recalled every time starting X by calling xcnfset e.g. in ˜ /.xinitrc Documentation: xcnf --help xcnfset --help /usr/doc/packages/xcnf/README Contents: xcnf-0.1

Copyright:  (c) 1997, Martin Scherbaum, S.u.S.E. GmbH
Author:  Martin Scherbaum <maddin@suse.de>
Version:  0.1-8

xcolors displays colors defined in rgb.txt

Xcolorsel displays colors defined in rgb.txt. You can create an RGB file by redirecting the output of showrgb to a file. Contents: xcolors-04oct91

Copyright:  (c) 1989,1991 - Stephen Gildea
Author:  Stephen Gildea <gildea@intouchsys.com>
Version:  91.10.4-6

xcoral A multiwindow mouse-based text editor for the X Window System

Xcoral is a multiwindow mouse-based text editor for the X Window System. A built-in browser enables you to navigate through C functions, C++ classes, methods and files. A SMall Ansi C Interpreter (Smac) is also built-in to dynamically extend the editor’s possibilities (user functions, key bindings, modes etc). Xcoral provides variable width fonts, menus, scrollbars, buttons, search, regions, killbuffers, macros and undo. An on-line manual box, with a table of contents and an index, helps you to use and customize the editor. Commands are accessible from menus or key bindings. Xcoral is a direct Xlib client and runs on color/bw X Display. Contents: xcoral-3.1

Copyright:  (c) 1989-94 Lionel Fournigault, Bruno Pages and Dominique Leveque.
Author:  Lionel Fournigault <lf@lri.lri.fr>
Version:  3.1-8

xdf disk usage under X11

a graphical df (Disk free). Contents: xdf-1.3

Copyright:  (c) 1996 Georg Greve
Author:  Georg Greve <development@fusebox.hanse.de>
Version:  1.3-2

xdtm The X Desktop Manager

The X Desktop Manager is a graphical shell for the X Window System, it provides mechanisms for all the usual shell activities and is intended to reduce the training necessary for UNIX novice to become productive on your system. Contents: xdtm-2.5.8

Copyright:  (c) 1991 Edward Groenendaal

Author:

Edward Groenendaal, Ramon J. Santiago <rjs@pencom.com> , Lionel Mallet <Lionel.Mallet@sophia.inria.fr>

Version:  2.5.8-9

xephem Space tool

A program which displays all sorts of information about our planetary system. Lots of nice, colourful buttons. Contents: xephem-3.0

Copyright:  (c) 1990-93 Elwood Charles Downey
Author:  Elwood Charles Downey <ecdowney@noao.edu>
Version:  3.0-8

xfileman xfilemanager

Xfilemanager is an X11 based Unix file manager with full color displays. Contents: xfilemanager-0.4.1

Author:  Ove Kalkan <kalkan@ramz.ing.tu-bs.de>
Version:  0.4.1-8

xflame GUI for cdrecord und mkisofs

A little tool to create CD-Roms

Copyright:  (c) GPL
Author:  Dr. Frank Rieg <dr.frank.rieg@t-online.de>
Version:  1.0-2

xfm xfm, a file manager for X.

This is a powerful file and applications manager program written using the X toolkit. It contains virtually all of the features that you would expect in a file manager - move around your directory tree in multiple windows, and move, copy or delete files with simple mouse operations. The integrated application manager allows you to load files into your favorite applications from the file manager by dragging icons. You should invoke xfm.install before launching xfm for the first time. Contents: xfm-1.3.2

Copyright:  (c) 1990 - 1995 Simon Marlow, Albert Graef

Author:

Albert Graef <ag@MuwiInfb.Geschichte.Uni-Mainz.DE>, Simon Marlow <simonm@dcs.gla.ac.uk>

Version:  1.3.2-8

xfmail XForms based Email reader system

XFMail is an X-Windows application for sending and receiving electronic mail. Features include: MIME and PGP support, address books with mailing list functions, message filtering rules, and multilingual support. It retrieves mail both from POP servers and from a local mail spool. Mail can be sent using sendmail, an SMTP gateway, or a POP server. It also supports IMAP4 remote folders. Contents: xfmail-1.2 Please take a look at the Dokumentation for some limits.

Copyright:  (c) 1995, 1996, 1997 Gennady B. Sorokopud, 1995 Ugen J. S. Antsilevich
Author:  Gennady B. Sorokopud <gena@NetVision.net.il>
Version:  1.2-2

xformat xformat

Little X frontend to fdformat, mke2fs and mformat using the Xforms library.

Copyright:  (c) GPL
Author:  Christoph Neerfeld <Christoph.Neerfeld@mail.bonn.netsurf.de>
Version:  1.0.1-0

xftp X11 frontend to ’ftp’

Makes ftp much more easy!

Contents: llnlxftp-2.0.4

Copyright:  (c) 1993-1994 University of California
Author:  Neale G. Smith <llnlxftp@llnl.gov>
Version:  2.0.4-12

xgopher A gopher client for the X Window System

A gopher client for the X Window System. This program offers you a comfortable GUI for accessing gopher clients. You must be connected to the internet, though. Documentation can be found under /usr/doc/X11-misc/xgopher-1.3 & man xgopher

Contents: xgopher.1.3.3

Copyright:  (c) 1992, 1993 University of Illinois
Author:  Allan Tuchman <a-tuchman@uiuc.edu>
Version:  1.3.3-8

xgrok Graphical Resource Organizer Kit

Grok is a simple database manager and UI builder that can -

* keep phone and address lists (like a rolodex)

* store phone call logs

* keep todo lists

* manage any other database after simple GUI-driven customization More precisely, grok is a program for displaying and editing strings arranged in a grid of rows and columns. Each row is presented as a ”card” consisting of multiple columns, or ”fields”, that allow data entry. The presentation of the data is programmable; a user interface builder that allows the user to arrange fields on a card graphically is part of grok. Grok also supports a simple language that allows sophisticated queries and data retrieval. Grok comes with the above examples and a few others as pre-built applications.

Contents: xmbase-grok-1.4.1

Copyright:  (c) 1995 Thomas Driemeyer
Author:  Thomas Driemeyer <thomas@bitrot.in-berlin.de>
Version:  1.4-14

xgroups file manager - similar to the Windows program manager

This file or desktop manager is similar to the Windows program manager. Its configuration is somehow unconventional, but the program is useful nevertheless. The one or other will sureley like this program.

Contents: xgroups-1.3

Copyright:  (c) 1993 Giuseppe Zanetti
Author:  Giuseppe Zanetti <beppe@maya.dei.unipd.it>
Version:  1.3-8

xhtalk a talk program under X11

You should copy a sample configuration file as .xhtalk into your home directory before using xhtalk the first time. You find one in /usr/doc/packages/xhtalk/SAMPLE.xhtalk Contents: xhtalk-2.9p1-linux

Copyright:  (c) 1992 Klaus Hartenstein
Author:  Klaus Hartenstein <kdh@ira.uka.de>, Angelo Haritsis <ah@doc.ic.ac.uk>
Version:  2.9p1-8

xinfo X-Browser for GNU Info Pages

Makes reading of info files a lot easier. Contents: xinfo-1.01.01

Copyright:  (c) 1989, 1990 Jordan K. Hubbard
Author:  Jordan K. Hubbard <jkh@meepmeep.pcs.com>
Version:  1.01.01-8

xirc xirc -a chat programm under X11

An X-based irc-client

Copyright:  (c) 1996-1997 Joe Croft
Author:  Joe Croft <jcroft@unicomp.net>
Version:  2.0-10

xkeycaps A frontend to xmodmap

xkeycaps is a graphical frontend for xmodmap. A virtual keyboard is shown and you can redefine any key on runtime change your ˜ /.Xmodmap or use other applications (provided that you use Click-to-Focus). For layout you should choose ”PC 102 key keyboard #1 - Linux/XFree86 German”. Contents: xkeycaps-2.29

Copyright:  (c) 1991, 1992, 1993, 1994, 1995 Jamie Zawinski
Author:  Jamie Zawinski <jwz@netscape.com>
Version:  2.29-9

xled xled - a LED for vbox the ISDN answering machine for vbox.

xled is your interface to your answering machine. Clicking onto the icon launches your personal vbox. Contents: xled-1.3.1

Copyright:  (c) Jan Schoenepauck und Joachim Gassen
Author:  Jan Schoenepauck <schoenep@uni-wuppertal.de>
Version:  1.3.1-5

xless text browser

xless 1.7 - the well-known text browser. Necessary for displaying the various HOWTOs. Contents: xless-1.7

Copyright:  (c) 1994 Dave Glowacki
Author:  Dave Glowacki <dglo@cs.berkeley.edu>
Version:  1.7-9

xlock A screen saver

Patrick J. Naughton’s xlock screensaver/locker for X, with many more modes added by David A. Bagley. For security reasons we now have xlock sgid shadow instead of suid root. If you use a yp-server which only gives passwords to programs running as user ”root”, you have to change xlock back to suid-root. Contents: xlockmore-4.04

Copyright:  (c) 1988-1995 Patrick J. Naughton, David A. Bagley

Author:

David Albert Bagley <bagleyd@bigfoot.com>, Patrick J. Naughton <naughton@eng.sun.com>

Version:  4.04-7

xlogin xlogin, xtelnet

This package contains two scripts, which called with the name of a remote host, will open an xterm window on the local X display with a remote session or login shell. The sript xlogin starts by a rsh call (remote shell) an xterm on the remote host. If necessary the access will be allowd by sending the magic key (cookie) of the display to the remote host.

The script xtelnet starts a local xterm with a telnet session on the remote host. Contents: xlogin-0.1

Copyright:  (c) 1997 S.u.S.E. GmbH
Author:  Werner Fink <werner@suse.de>
Version:  0.1-8

xmail X Window System interface to the mail program

xmail is an X-based interface to the Berkeley mail program. This version of xmail includes support for decompressing and displaying the content of an X-Face mail header, if such a header exists in the message being read. This feature is a compile time option, and requires the existence of the compface library routines, which are NOT supplied with the xmail sources. X-Face headers are compressed bitmap images, typically of the face of the person owning such a header. The compressed header contains only printable characters, which allows it to be included in a mail message. The USENIX FaceSaver project is typically the source for these bitmaps, which are compressed into X-Face header format using the compface software tools written and distributed by James Ashton <jaa@cs.su.oz.au>. The compface library is also provided with the distri- bution of the faces software, written by Rich Burridge, and is available by email from his archive server <rbarchive-server@Aus.Sun.COM>. Contents: xmail-1.6

Copyright:  (c) 1990-94 by National Semiconductor Corporation
Author:  Michael C. Wagnitz <michael@berlioz.nsc.com>
Version:  1.6-8

xmftp xmftp

X/Motif FTP client, include static and dynamic version Contents: xmftp-1.0.4

Copyright:  (c) GPL
Author:  Viraj Alankar <valankar@bigfoot.com>
Version:  1.0.4-0

xnetload This program displays network traffic and uptime for the network

This program displays network traffic and uptime for the network connection named on the command line. Contents: xnetload-1.2

Author:  Roland Smith <rsmit06@ibm.net>
Version:  1.2-3

xosview info on the system load

A small program - similar to xsysinfo which is mostly configurable using resources via ˜ /.Xresources. Shows actual CPU, swap, memory, active interrupts and -if wanted- netpacket statistics. If you want to get informations about ’network traffic’ you still have to call ipfwadm (included in the same package) to activate this as root. This might look like: /sbin/ipfwadm -A -a -P all -S <own IP-address> -D 0/0 /sbin/ipfwadm -A -a -P all -S 0/0 -D <own IP-address> Please have a look at /usr/doc/packages/xosview/README.linux. For showing shared pages, the module memstat.o have to loaded with insmod. The sources of the module is located in /usr/doc/packages/xosview/memstat/. The installation is only after an upgrade of the kernel necessary. Contents: xosview-1.5.1 If you use the ipfwadm package from our distribution the ipfwadm will be invoked at boot time.

Copyright:  (c) 1994-97 Mike Romberg
Author:  Mike Romberg <romberg@md.fsl.noaa.gov>
Version:  1.5.1-1

xplot An interactive 3D plotting program

Xplot is an interactive 3D plotting program for X windows. It is capable of displaying simple objects that have easily described geometries. Currently, it only runs on 8bit pseudo color visuals. The way of using xplot is as follows. First one defines the object by specifying its geometry. Then plot it. Once an object is displayed on the screen, its appearance can be modified interactively by mouse input. Docu: /usr/doc/X11-misc/xplot-beta-09feb95 & man xplot Contents: xplot-beta-09feb95

Copyright:  (c) 1986-89 Zou Maorong, Thomas Williams, Colin Kelley
Author:  Zou Maorong, Thomas Williams, Colin Kelley
Version:  95.2.9-6

xplz German zip codes

This program displays the zip codes of German towns. For more information see /usr/doc/X11misc/xplz-3.4.README . Contents: xplz-3.4 xplz-data

Copyright:  (c) 1993 Rainer Klute
Author:  Rainer Klute <klute@irb.informatik.uni-dortmund.de>
Version:  3.4-9

xprompt small tool for prompting users

Small tool to ask the user for one or more responses (e.g. from batch files). Documentation: /usr/doc/X11-misc/xprompt-28sep91.README Contents: xprompt-28sep91

Copyright:  (c) 1989-91 - Barry Brachman & University of British Columbia
Author:  Barry Brachman <brachman@cs.ubc.ca>
Version:  91.9.28-5

xrolodex rolodex-like application

A ”virtual business card” manager. For more information see the xrolodex manpage and /usr/doc/packages/xrolodex. This is a demo version of a commercial product.

Copyright:  (c) 1991, 1992 Iris Computing Laboratories
Author:  Jerry Smith <jsmith@spectro.com>
Version:  1.5-8

xrpm X frontend for RPM

Author:  George A. Farris <george@gmsys.com>
Version:  2.1-4

xscrns a screensaver applikation

The xscreensaver program waits until the keyboard and mouse have been idle for a period, and then runs a graphics demo chosen at random. It turns off as soon as there is any mouse or keyboard activity. Contents: xscreensaver-2.14

Author:  Jamie Zawinski
Version:  2.14-2

xselect Manipulate the XSelection

With this little tool one can pipe the currently selected text under X into a file or vice versa. Help can be found by reading the manpage with man xselection Contents: xselection-1.6.1

Copyright:  (c) 1990 Richard Hesketh, Computing Lab. University of Kent at Canterbury, UK
Author:  Richard Hesketh <rlh2@ukc.ac.uk
Version:  1.6.1-8

xsession A session manager

The xsession program is a session manager. It is normally executed by your ˜ /.xinitrc (or ˜ /.xsession) script and controls your X Window session. As soon as it is started, xsession launches a window manager and some applications of your choice. At anytime during your session, you may switch to another window manager or execute some other applications from the xsession menus. Examples may be found under /usr/doc/packages/xsession/examples.

Contents: xsession-1.1

Copyright:  (c) 1993, 1994, Alain Nissen & Raphael Quinet

Author:

Alain  Nissen  <nissen@montefiore.ulg.ac.be>,  Raphael  Quinet

<quinet@stud.montefiore.ulg.ac.be>

Version:  1.1-7

xsky A Star Atlas

You find it under /usr/lib/xstar. Docu: man xstar

Contents: xsky-2.0.1

Copyright:  (c) 1992-93 Terry R. Friedrichsen
Author:  Terry R. Friedrichsen <terry@venus.sunquest.com>
Version:  2.0.1-7

xspline draws colorful spline functions in the root window

draws colorful spline functions in the root window

Contents: xspline-02may92

Copyright:  (c) 1992 Jef Poskanzer
Author:  Jef Poskanzer <jef@netcom.com>
Version:  92.5.2-5

xspread An X-Windows spreadsheet

A spreadsheet for the X Windows System.

Contents: xspread-2.3

Copyright:  (c) 1992 Board of Regents of the University of Wisconsin

Author:

Robert Parbs, James Cornelius, Michael Frey, Dan Gruber, Fang Wang, <softeng@cs.uwm.edu>

Version:  2.3-8

xsysinfo Displays the system load

This program displays information about the load of your system. CPU Load, CPU idle, Memory and Swap are displayed constantly.

Contents: xsysinfo-1.2 xsysinfo-1.5

Copyright:  (c) 1992, 1993 - Gabor Herr
Author:  Gabor Herr <herr@iti.informatik.th-darmstadt.de>
Version:  1.5-6

xtail X11 Version of tail

This tool (Motif, statically linked) may especially used as file monitor. Nice to view files.

Contents: xtail-1.03

Copyright:  (c) 1994 Klaus Thiele
Author:  Klaus Thiele <100520.634@compuserve.com>
Version:  1.03-12

xtalk X11 client for talk

XTalk is a talkclient and it was written for use under the graphical user interface X11. It is full useable with mouse or the keyboard via shortcuts. The important functions are also accessible over the buttonstrip with easy understand symbols. The functionality compared with the ”normal” talkcommand was expand considerable. It is possible to save the written and received text with an any name. Moreover you can scroll in the text. No informations are lost. The telephonbookfunction is very useful for frequent used talkpartner. The connection parameter will show permanent on the statusline under the buttonstrip.Besides you can open a second or third connection without great difficulty, by using the menupoint ”File->New XTalk”. Very useful is the build in FTP function. She is based on a own protocol and therefore only enabled, when the partner runs also a XTalk client. By construct a connection, XTalk looks, if XTalk is running on the other end. In this case, the transfer menu will enabled automatically.

Copyright:  (c) 1997 Stephan Pilz
Author:  Stephan Pilz <spilz@htwm.de>
Version:  1.2-7

xtar View contents of a tar file

XTar is a Motif program for viewing and manipulating tar archives (which may be compressed). The contents of the archive are displayed in the main window and various options are available from pulldown and popup menus. Docu: /usr/X11R6/lib/X11/xtar/xtar.doc & man xtar Contents: xtar-1.4

Copyright:  (c) 1993 University of Liverpool
Author:  Rik Turnbull <rik@csc.liv.ac.uk>
Version:  1.4-13

xtexit Prompt on exiting X

‘xtexit‘ sends an request to all clients to shut down. If the application still needs an user interaction (e.g. if a file should be saved) this is possible. If you answer by the affirmative all applications will be closed. This method is not fully water-proof but better than killing each and every client without being able to interfere. xterm applications anyway are killed immediately! If this package is installed, it will be automatically integrated into the sample user fvwm menu. Contents: xtexit

Copyright:  (c) 1990 Front End Oy

Author:

Klaus Franken <kfr@suse.de>, Teemu Torma, Front End Oy <tot@frend.fi>, Paul Raines <raines@slac.stanford.edu>

Version:  0.42-8

xvhelp A help system based on Openlook

This package includes a help browser (xvhelp - similar to the MS Windows Help system) and a compiler (hcc) to create your own help files. This version of xvhelp uses /usr/openwin/lib/help as the directory for help files. Documentation: /usr/doc/packages/xvhelp Contents: XVhelp-1.00

Copyright:  (c) Public Domain
Author:  Roberto Surlinelli, Luca Maranzano, Guido Cardino <calamaro@dist.dist.unige.it>
Version:  1.0-7

xvi Multi-window vi

Xvi is a portable multi-window version of vi. Different files can be displayed in separate windows (instead of displaying them in one window, separated by horizontal status lines). It is also possible to display different parts of one file in different windows. Contents: xvi-3.1

Copyright:  (c) 1990-92 Chris and John Downey
Author:  Chris Downey, John Downey
Version:  3.1-8

xvnews A News reader based on Openlook

If you find trn too obscure, this news reader might save you a lot of trouble when reading News. Requirements: access to an NNTP server. Contents: xvnews-2.3.2

Copyright:  (c) 1988-91 Ellen M. Sentovich, Rick L. Spickelmier
Author:  Ellen M. Sentovich , Rick L. Spickelmier, J.J. deGraaff <J.J.deGraaff@twi.tudelft.nl>
Version:  2.3-9

xvtdl A todo list manager

XVTDL has the following features:

* You can manage multiple todo lists; lists can have sublists.

* Each todo list item is prioritized; each list can be displayed in a combination of priority, chronological, and alphabetical order.

* List items, if you do not ”check them off”, will propagate from one day to the next.

* List items can be recurring items, i.e., items that will automatically pop up weekly, biweekly, monthly, etc. These items will propagate like ”normal” list items.

* List items can have deadlines associated with them, with several types of actions that execute on or after a deadline.

* You can set list items on any day, with easy calendar management.

* Lists can be printed on a ”plain” or PostScript printer.

* Printing/display supports ISO 8859.1 character set.

* List items can be logged when checked on/off with optional comments added to the log file.

* List items can be retained or deleted when checked off.

* Multiple lists can be combined for viewing or printing.

* Multiple lists from multiple files can be manipulated.

* Changes to the todo list can be made and are detected by the running programs.

Contents: xvtdl-5.2

Copyright:  (c) 1991-94 Mike Jipping, Hope College
Author:  Mike Jipping <jipping@cs.hope.edu>
Version:  5.2-6

xwpick X Windows Image Pick up Utility

Xwpick lets you pick an image from an arbitrary window or rectangular area of an X Window (DEC Window) server and write it to a file in a variety of formats.

Contents: xwpick-2.20

Copyright:  (c) 1993, 1994 by Evgeni Chernyaev
Author:  Evgeni Chernyaev <chernaev@mx.ihep.su>
Version:  2.20-8

xzap tool to view X11 programs

This very useful tool displays the running applications. Using this tool, applications which are not responding or can not be found on the desktop any more can be killed easily.

Contents: xzap-1.1

Copyright:  (c) 1993 The University of Newcastle upon Tyne
Author:  Gerry Tomlinson <gerry.tomlinson@newcastle.ac.uk>
Version:  1.1-9

xzet Graphical System Activity Monitor

XZET is an X windows application which is Zmans’ vesion of Xetat and Top, hence the name XZET. It provides 3 different types of real time system resource monitoring windows. The first is a global system activity window which displays stripcharts of global idle, user, system and nice percentages. It also displays stripcharts for free swap space and free physical memory. Finally, it has displays for global process counts and states, and free space monitoring of up to three file systems. A second window lists information on every process running in the system. The information varies based on whether a short or long format is selected. The long format is quite complete. The user can toggle the state of this list between ”frozen” and ”auto-updating”. If an entry in the list is selected with the mouse, a third and final window type will appear. This is an individual process window which features an updating tachometer widget for both the processes percent cpu load and the percent physical memory load. Information similar to that provided in the short format of the global process list is also enumerated.

Copyright:  (c) 1992-94 Loganville Technical Resources, Inc

Author:

Branko Lankester, Roger Binns, Michael K. Johnson <johnsonm@stolaf.edu>, Robert J. Nation <nation@rocket.sanders.lockheed.com>

Version:  1.0-7

0.1.19 XDEV: Development under X11

ddd Debugger with graphical user interface

The DDD debugger (Data Display Debugger) is a comfortable GUI to the well-known debuggers GDB and DBX. Data structures can be represented as graphs and interactively shown. You can debug programs in C, C++, Pascal, MODULA-2, FORTRAN and ADA, and even at the assembler code level!

XDEV: Development under X11

Contents: ddd-2.2.3

Copyright:  (c) 1995-98 Technische Universitaet Braunschweig, Germany.

Author:

Andreas Zeller <zeller@ips.cs.tu-bs.de>, Dorothea Luetkehaus <luetke@ips.cs.tubs.de>

Version:  2.2.3-0

dlgedit QT Dialog Editor

A nice small tool to design interactively input dialogs for the QT Toolkit. Documentation: /usr/doc/packages/dlgedit Contents: dlgedit-1.2

Copyright:  (c) 1997, Jeff Harris, Klaus Ebner
Author:  Jeff Harris <jharris@cis.ohio-state.edu>, Klaus Ebner <klaus@gaspode.ndh.com>
Version:  1.2-11

iv InterViews - a GUI.

Copyright:  (c) 1987-1991 Stanford University; 1991 Silicon Graphics, Inc.

Author:

Michael N. Lipp <mnl@dtro.e-technik.th-darmstadt.de>, Herbert Thielen <thielen@lpr.e-technik.tu-muenchen.de>

Version:  3.1g-1

mesa Mesa

Mesa is a 3-D graphics library with an API which is very similar to that of OpenGL*. To the extent that Mesa utilizes the OpenGL command syntax or state machine, it is being used with authorization from Silicon Graphics, Inc. Contents: MesaDemos-2.6 MesaLib-2.6

Copyright:  (c) Brian Paul
Author:  Brian Paul <brianp@elastic.avid.com>
Version:  2.6-1

qtdevel QT development environment

This package contains the QT runtime libraries. The former QT package was split into the following ones: qtlib runtime libraries qtdevel developpment environment qtext extensions If you just want to run KDE, it is sufficient to install the package qtlib. Please read in /usr/doc/packages/qt/LICENSE for licensing issues! Please read in /usr/doc/packages/qt/LICENSE for licensing issues!

Copyright:  (c) 1992-1997, Troll Tech AS
Author:  Troll Tech AS <info@troll.no>
Version:  1.32-3

qtext QT extensions

This package contains some QT extensions (Netscape plugin, ImageIO, OpenGL). The former QT package was split into the following ones: qtlib runtime libraries qtdevel developpment environment qtext extensions If you want to develop programs with QT or simply e.g. recompile KDE, then you have to install the qtdevel package. Please read in /usr/doc/packages/qt/LICENSE for licensing issues! Please read in /usr/doc/packages/qt/LICENSE for licensing issues!

Copyright:  (c) 1992-1997, Troll Tech AS
Author:  Troll Tech AS <info@troll.no>
Version:  1.32-3

qtlib QT runtime libraries

This package contains the QT runtime libraries. The former QT package was split into the following ones: qtlib runtime libraries qtdevel developpment environment qtext extensions If you want to develop programs with QT or simply e.g. recompile KDE, then you have to install the qtdevel package. Please read in /usr/doc/packages/qt/LICENSE for licensing issues! Contents: qt-1.32 Please read in /usr/doc/packages/qt/LICENSE for licensing issues!

Copyright:  (c) 1992-1997, Troll Tech AS
Author:  Troll Tech AS <info@troll.no>
Version:  1.32-3

vtk Visualization ToolKit

The Visualization ToolKit (vtk) is a software system for 3D Computer Graphics and Visualization. Vtk includes a textbook published by Prentice-Hall (”The Visualization Toolkit, An Object-Oriented Approach To 3D Graphics, 2nd edition” ISBN 0-13-954694-4), a C++ class library, and several interpreted interface layers including Tcl/Tk, Java, and Python. Vtk has been implemented on nearly every Unix-based platform and PC’s (Windows NT and Windows95). The design and implementation of the library has been strongly influenced by object-oriented principles. Contents: vtk20p1src

Copyright:  (c) Ken Martin, Will Schroeder, Bill Lorensen
Version:  2.0-3

vtkdata Visualization ToolKit - Datafiles

Datafiles for the Visualization ToolKit. Contents: vtk20data.part1 vtk20data.part2 vtk20data.part3 vtk20data.part4 vtk20data.readme vtk20manhtml

Copyright:  (c) Ken Martin, Will Schroeder, Bill Lorensen
Version:  2.0-1

wxwdemo Demo programs for wxwin

Some examples to show the possibilities of wxwindows. These programs have been compiled using OpenLook.

Copyright:  (c) 1993-97, AIAI, University of Edinburgh
Author:  Julian Smart <J.Smart@ed.ac.uk>
Version:  1.67-6

wxwdocht Documentation for wxwin in HTML

Documentation for wxwin in HTML.

Copyright:  (c) 1993-97, AIAI, University of Edinburgh
Author:  Julian Smart <J.Smart@ed.ac.uk>
Version:  1.67-6

XDEV: Development under X11

wxwdocps Documentation of wxwin in PostScript

Documentation of wxwin in PostScript

Copyright:  (c) 1993-97, AIAI, University of Edinburgh
Author:  Julian Smart <J.Smart@ed.ac.uk>
Version:  1.67-6

wxwin GUI Library for Motif, Openlook and MS-Windows

wxWindows is a Class Library for C++, which provides a GUI for more than one platform. The current version supports OpenLook (XView), Motif and Windows 3.1. WxWindows has originally been developed at the Artificial Intelligence Applications Institute at the University of Edinburgh. This version has been compiled for the use with OpenLook but there is a MS-Windows version as well. The whole package can be found under /usr/lib/wxwin.

Copyright:  (c) 1993-97, AIAI, University of Edinburgh
Author:  Julian Smart <J.Smart@ed.ac.uk>
Version:  1.67-6

wxwstat Static libraries for package wxwin

This package contains the static libraries for package wxwin.

Copyright:  (c) 1993-97, AIAI, University of Edinburgh
Author:  Julian Smart <J.Smart@ed.ac.uk>
Version:  1.67-6

wxxt Port of wxwin with motif look-and-feel

The package wxxt is a port of wxwin that imitates the look and functionality of a motif interface but does not need the motif libraries to be compiled or run.

Copyright:  (c) 1995, AIAI, University of Edinburgh (Julian), 1995, GNU (Markus)
Author:  Markus Holzem <mho@comnets.rwth-aachen.de>
Version:  1.66d-6

xforms Forms Library

XForms is a GUI toolkit based on Xlib for X Window Systems. It features a rich set of objects, such as buttons, sliders, and menus etc. integrated into an easy and efficient object/event callback execution model that allows fast and easy construction of X-applications. In addition, the library is extensible and new objects can easily be created and added to the library. Contents: xforms-0.88

Copyright:  (c) 1995 T.C. Zhao and Mark Overmars
Author:  T.C. Zhao <zhao@bragg.phys.uwm.edu>, Mark Overmars
Version:  0.88-1

xtpanel Interface Builder using scripts

This program allows you to build (complex) programs using a simple scripting language. Included are a tutor program and some sample scripts (xt-colors, xt-lister, xt-load, xt-lpq, xt-page). Documentation: /usr/doc/packages/xtpanel Contents: xtpanel-3.01

Copyright:  (c) 1992, 1993 - Board of Trustees - Leland Stanford Junior University
Author:  Steve Cole <steve@sep.stanford.edu>, Dave Nichols <dave@sep.stanford.edu>
Version:  3.01-5

xwb software development environment for openwin

Graphical user interface using the Openwin look. Supports editing, compiling and launching of C programs with a mouse-click. Contents: xwb v3.0

Copyright:  (c) 1996 by Michael Sohmen
Author:  Michael Sohmen <SOMI0011@RZ03.FH-Karlsruhe.de>
Version:  3.0-9

xwpe A X-window programming environment

xwpe is a X-window programming environment designed to use on UNIX-systems. It is similar to ’Borland C++ or Turbo Pascal’ environment. The difference between the programming environments from Borland an xwpe is that many compilers and linkers may be started. From a menu three different debuggers can be chosen. And of course not only key strokes can select these options but you can also use a mouse. Contents: xwpe-1.4.2

Copyright:  (c) 1994 Fred Kruse
Author:  Fred Kruse <kruse@rrzn-nis.uni-hannover.de>
Version:  1.4.2-9

xxgdb xxgdb, user interface to the GNU debugger under X

Xxgdb is a graphical user interface to the gdb debugger under the X Window System. Contents: xxgdb-1.12

Copyright:

(c) 1989 The University of Texas at Austin, 1990 Microelectronics and Computer Technology Corporation

Author:  Po Cheung <cheung@sw.mcc.com>
Version:  1.12-8

ygl a GL graphics library with sample file

Besides the Ygl library the glwidget library is included. You can check what to do with this library by trying one of the examples that come along: coltest, lines, lissa, rgbtest, simple and smile Contents: Ygl-3.1

Copyright:  (c) 1994 Fred Hucht, Michael Staats

Author:

Fred  Hucht  <fred@thp.Uni-Duisburg.DE>  ,  Michael  Staats

<michael@hal6000.thp.Uni-Duisburg.DE>

Version:  3.1-9

0.1.20 XV: XView (OpenLook, Applications)

ssbase Slingshot extensions

These are extensions for XView programming; it contains static and dynamic libraries in elf-format. Also included are the include files and the documentation. Contents: SlingShot2.1

Copyright:  (c) 1991 Sun Microsystems, Inc.
Version:  2.1-10

ssexmp Slingshot examples

A little collection of Slingshot examples. You can see the use and the advantage of the Slingshot extension package.

Copyright:  (c) 1991 Sun Microsystems, Inc.
Version:  2.1-10

xtoolpl xtoolplaces and xtoolwait

This program offers you information about every window existing on your current desktop; providing a means of starting it at exactly the same position. Furthermore there is the program xtoolwait that just waits after an X application has been launched and presents its window. Contents: xtoolplaces-1.3.1 xtoolwait-0.3

Copyright:  (c) 1993 Xerox Corporation.

Author:

Adam  Stein  <adam@iset.scan.mc.xerox.com>,  Richard  Huveneers

<richard@hekkihek.hacom.nl>

Version:  1.3.1-8

xv32aout Shared XView Libraries (a.out)

These are the a.out libraries of the XView system.

Version:  3.2-1

xv32base XView (OpenLook) Base system

Programs, dynamic libraries, configuration files and documentation for XView 3.2 for Linux. It contains: xgettext, msgfmt, olwm, svenv, toolwait, winsysck, owplaces, openwin, xcenter, clock, cmdtool, olwmslave, props, textedit, meminfo Contents: xview-3.2p1.4.src

Copyright:  (c) 1990-1992 Sun Microsystems, Inc.
Version:  3.2p1.4-6

XWM: Window Manager and Desktop

xv32dev XView (OpenLook) development kit

Imake-configuration and include files for developing under XView 3.2.

Copyright:  (c) 1990-1992 Sun Microsystems, Inc.
Version:  3.2p1.4-6

xv32exmp Sample programs for Xview

Sample programs for Xview which demonstrate the Slingshot and UIT extensions, libraries which greatly simplify programming a user interface under X.

Copyright:  (c) 1990-1992 Sun Microsystems, Inc.
Version:  3.2p1.4-6

xv32slib XView version 3.2 static libraries

Static libraries for XView 3.2

Copyright:  (c) 1990-1992 Sun Microsystems, Inc.
Version:  3.2p1.4-6

0.1.21 XWM: Window Manager and Desktop

3dpixm 3d pixmap icons for fvwm

Very nice 3d pixmaps for the fvwm window manager, in two sizes, for menus and button bars. You should install this package.

Contents: 3dpixm

Author:  Joerg Mertin <joerg@pc50.zrz.tu-berlin.de>
Version:  98.2.4-2

3dpixms Small 3d pixmap icons for fvwm

Very nice 3d pixmaps for the fvwm window manager. These pixmaps are used for the pop-up menus.

Author:  Joerg Mertin <joerg@pc50.zrz.tu-berlin.de>
Version:  98.2.4-2

afterstp NeXTStep-like window manager

This window manager follows the tradition of NeXTStep. The functionality of the bowman window manager was enhanced with some configuration options. As well, there are shadowed menus, the button-bar Wharf, animated windows, a proper sound support, and many more features. It should be started with a color depth of at least 16 bits! Documentation: /usr/doc/packages/afterstp

Contents: AfterStep-1.0

Copyright:

(c) 1988 Evans & Sutherland Computer Corporation, 1989 Massachusetts Institute of Technology,1994 Robert Nation, 1996 by Beat Christen

Author:

Frank Fejes <frank@ssax.com> , Dan Weeks <dan@mango.sfasu.edu>, Jonathan B. Leffert <jonl@geom.umn.edu>, Bo Yang <eric@coeus.ucsd.edu>, Alfredo Kengi Kojima, Rob Nation, Dave Adkins, Beat Christen, Robert Depenbrock, Amadeus Foudray, Martin Scherbaum <maddin@suse.de>

Version:  1.0-7

amiwm AMIGA-like window manager

For those who still want their AMIGA desktop :-) Documentation: /usr/doc/packages/amiwm

Contents: amiwm-0.20pl28

Copyright:  (c) 1996, Marcus Comstedt
Author:  Marcus Comstedt <marcus@lysator.liu.se>
Version:  0.20pl28-8

bowman Window manager with the NeXT-Step ”look and feel”

Based on the fvwm window manager, imitating the NeXTSTEP GUI. Can be installed without problem with fvwm and fvwm95. Documentation: /usr/doc/packages/bowman Contents: bowman-0.93e

Copyright:  (c) Bo Yang
Author:  Bo Yang <eric@coeus.ucsd.edu>
Version:  0.93e-10

cdesim CDE simulator

CDEsim simulates a window manager. With the help of fvwm (version 2), it simulates the ”look and feel” of the popular CDE user interface. More the look than the feel. It is fully integrated into the susewm window manager configuration tool, and will be automatically preconfigured. Documentation and original configuration files under: /usr/doc/packages/cdesim Please read /usr/doc/packages/cdesim/README.cdesim (also available from the cdesim menu!).

Copyright:  (c) October 9, 1996, Michael A. Haas
Author:  Michael A. Haas <gnosis@infinet.com>, Martin Scherbaum <maddin@suse.de>
Version:  0.42-8

ctwm Window Manager

A very popular window manager among users with fast video cards. ctwm is an improvement of twm (Tab Window Manager). If you launch ctwm from the fvwm2 menu, it will return to fvwm2 when you click on the restart menu button in the ctwm root menu. Documentation: man ctwm or in /usr/doc/packages/ctwm Contents: ctwm-3.4pl2

Copyright:  (c) 1992 Claude Lecommandeur

Author:

Claude Lecommandeur <lecom@sic.epfl.ch>, Dave Payne, Dave Sternlicht, Steve Pitschke, Tom LaStrange

Version:  3.4pl2-8

enlight The Enlightenment windowmanager (beta)

The Enlightenment windowmanager by Carsten ”Rasterman” Haitzler. Included are the themes (configurations) DEFAULT and DEFAULT small. Attention! This version (DR-0.12) is still BETA! Contents: EFAQ-1.0 enl DR-0.12

Copyright:  (c) 1997 Carsten Haitzler (Rasterman)
Version:  DR 0.12-6

fvwm THE window manager for Linux

The fvwm window manager (version 2) is the standard window manager for the X window system under Linux. fvwm is highly configurable and has a similar look-and-feel to the Motif standard. The personal configuration is stored in the file ˜ /.fvwm2rc in the user’s home directory. The system wide configuration is generated at installation time by SuSEconfig and may be found in /usr/X11R6/lib/X11/fvwm2 . Documentation: man fvwm2 /usr/doc/packages/fvwm Contents: cdesim-0.42 fvwm-2.0.46

Copyright:  (c) 1994 Robert Nation, Mark Boyns, Mark Scott

Author:

Charles  Hines  <chuck hines@vnet.ibm.com>,  Robert  Nation

<fvwm@wonderland.org>, Andrew Davison <davison@cs.monash.edu.au>, Drew Moseley <ddm@cis.ufl.edu>, Jason L Tibbitts <tibbs@tcamc.uh.edu>, John Heidemann <johnh@ficus.CS.UCLA.EDU>, Mark Boyns <boyns@sdsu.edu>, Mark Scott <mscott@mcd.mot.com>, Mike Finger <mfinger@mermaid.micro.umn.edu>, Mr. Per Persson <pp@solace.mh.se>, Nobutaka Suzuki <nobuta-s@is.aist-nara.ac.jp>

Version:  2.0.46-10

fvwm1 THE window manager for Linux (old version)

Since versions 1.x and 2.x are not compatible, we have kept the old version in the distribution. This version is configured with the system-wide configuration file /usr/X11R6/lib/X11/fvwm/system.fvwmrc, as well as the user’s personal configuration file ˜ /.fvwmrc in the home directory. The program susewm is able to configure this window manager,too.

XWM: Window Manager and Desktop

Important! Please read the file /usr/doc/packages/fvwm1/README.SuSE! Documentation: man fvwm fvwmman <manpage> /usr/doc/packages/fvwm1 Contents: fvwm-1.24r

Copyright:  (c) 1994 Robert Nation, Mark Boyns, Mark Scott

Author:

Andrew  Davison  <davison@cs.monash.edu.au>,  Drew  Moseley

<ddm@cis.ufl.edu>, Jason L Tibbitts <tibbs@tcamc.uh.edu>, John Heidemann <johnh@ficus.CS.UCLA.EDU>, Mark Boyns <boyns@sdsu.edu>, Mark Scott <mscott@mcd.mot.com>, Mike Finger <mfinger@mermaid.micro.umn.edu>, Mr. Per Persson <pp@solace.mh.se>, Nobutaka Suzuki <nobuta-s@is.aist-nara.ac.jp>, Robert Nation <fvwm@wonderland.org>

Version:  1.24r-11

fvwm95 Virtual window manager, look’n’feel of Win95

The look’n’feel of Win95 without Win95! With icons in the menus, frames and handling like win95, a task-bar with start button a.m.o.m. By installing the package susewm one can configure fvwm95 similar to fvwm, fvwm2 or bowman. Important1! The config file’s name was changed to .fvwm95rc ! Important2! Please read the file /usr/doc/packages/fvwm95/README.SuSE! Documentation: /usr/doc/packages/fvwm95 man fvwm95 fvwm95man <man page> Contents: fvwm95-2.0.43a-Imake

Copyright:  (c) 1993, Robert Nation
Author:  Rob Nation <fvwm@wonderland.org>, Hector Peraza <peraza@mitac11.uia.ac.be>
Version:  2.0.43a-12

icons Lots of icons

An extensive collection of more than 1800 icons for any purpose. Among them black/white and color ones, small and huge ones, simple and artistic ones. Documentation in: /usr/doc/packages/icons Contents: icons

Version:  98.2.4-2

lxuser Configuration for the sample user

This package contains the main dependencies for the sample user.

Copyright:  (c) 1995,1996,1997 S.u.S.E. GmbH
Version:  97.6.29-2

mlvwm MacIntosh like window manager

Mac users will feel right at home! This is a very early version. Documentation can be found: /usr/doc/packages/mlvwm man mlvwm Contents: mlvwm-current

Copyright:  (c) 1996, Takashi HASEGAWA / Uchikawa-Lab.
Author:  Takashi HASEGAWA <tak@bioele.nuee.nagoya-u.ac.jp>
Version:  0.7.0-7

olvwm olvwm - OpenLook Virtual Window Manager

This is a virtual window manager based on the older OpenLook window manager olwm with exactly the same ”look & feel”.

Copyright:  (c) (c) Copyright 1989 Sun Microsystems, Inc.
Author:  Scott Oaks <scott.oaks@East.Sun.COM>
Version:  3.2p1.4-6

pixmaps Icons and pixmaps for window managers

In this package you find all icons of the window managers of the S.u.S.E. Linux CD. The window managers concerned are:

- fvwm (package fvwm1) - fvwm2 (package fvwm) - fvwm95 (package fvwm95) - bowman (package bowman) - AfterStep (package afterstp) - CDEsim (package cdesim) - ctwm (package ctwm) - qvwm (package qvwm) - mlvwm (package mlvwm) - lxuser (former package lxuser) Also you find in this package the program xpmroot, formerly contained in the fvwm* window manager family. The reason why we put the icons together in this package is to avoid conflicts and redundancy whilst installing or removing packages. If there were icons in the former package with identical name but different contents these icons were renamed (icon1,icon2,iconN). Nevertheless, you may find the original pixmaps in the source packages (package s) of the above mentioned window managers. Documentation: /usr/doc/packages/pixmaps/README.SuSE man xpmroot Contents: pixmaps-0.1

Copyright:  (c) 1997, S.u.S.E. GmbH, Fuerth
Author:  Martin Scherbaum <maddin@suse.de>
Version:  0.1-11

qvwm Win95 like window manager

Q V stands for 9 5 in Japanese. It is a window manager that resembles the Win95 GUI Documentation in: /usr/doc/packages/qvwm Contents: qvwm-1.0b6g

Copyright:  (c) 1996, KOURAI Kenichi
Author:  KOURAI Kenichi <kourai@is.s.u-tokyo.ac.jp>
Version:  1.0b6g-8

susewm Window manager configuration tool

This package allows you to create in one step the configuration files for the fvwm, fvwm2, fvwm95, bowman, afterstep, cdesim, mwm, and ctwm window managers. In the menus you’ll only find programs which are really installed in the system. The desktop of the seven window managers will look the same (considering the particularities of the window managers). In version 1.5 for mwm and ctwm there is only support for keeping the menus actual since these two are very different from the five other window managers. The icons used in the fvwm2 / fvwm95 menus are taken from the package 3dpixms which themselves are generated from the 3dpixm package. You can find some information in /usr/doc/packages/susewm. The documentation is also accessible by the desktop menu. Contents: susewm-1.6

Copyright:  (c) 1996-1997 S.u.S.E. GmbH, Fuerth, Germany
Author:  Martin Scherbaum <maddin@suse.de>
Version:  1.6-11

wm2 A minimalist window manager

For all who hate popup menus and icons. Just set WINDOWMANAGER in ˜ /.bashrc to wm2. Dokumentation in: /usr/doc/packages/wm2 Contents: wm2-4

Copyright:  (c) 1996, Chris Cannam
Author:  Chris Cannam <cannam@zands.demon.co.uk>
Version:  4-9

0.2. Disk usage of the packages

0.2 Disk usage of the packages

Package  /  /usr/X11  /usr  /var  Σ

Series a1

aaa base  431  16  98  21  566
aaa dir  28  17  60  23  128
base  96  0  288  0  385
bash  403  0  748  0  1151
bash1  329  0  40  0  369
bdflush  6  0  1  0  8
compress  0  0  17  0  17
cpio  0  0  75  0  75
cron  0  0  45  2  48
db  0  0  539  0  539
devs  45  0  0  0  45
diff  0  0  118  0  118
eazy  24  0  1  0  25
elvis  0  0  745  0  745
ext2fs  352  0  109  0  461
file  0  0  173  0  173
fileutil  167  0  187  0  355
find  0  0  110  0  110
gawk  0  0  420  0  420
gdbm  0  0  470  0  470
gppshare  0  0  1760  0  1760
groff  0  0  2670  0  2670
gzip  44  0  22  0  66
kbd  41  0  1462  1  1504
kernmod  4463  0  0  0  4463
ldso  334  0  110  0  445
less  0  0  101  0  102
libaout  620  0  1507  0  2128
lilo  63  0  486  0  550
loadlin  152  0  0  0  152
localedb  0  0  3680  0  3680
man  2  0  209  65  277
modules  183  0  10  0  193
ncurses  1312  0  4356  0  5668
net tool  99  0  35  0  135
netcfg  14  0  0  1  15
nkita  12  0  595  0  607
nkitb  12  0  761  0  774
nvi  0  0  373  0  373
pcmcia  127  15  192  0  335
perl  0  0  7095  0  7095
ps  43  17  229  0  290
rpm  523  0  611  1  1136
sh utils  111  0  335  0  447

Package  /  /usr/X11  /usr  /var  Σ
shadow  306  0  3346  0  3652
shlibs  1849  0  463  0  2312
syslogd  1  0  53  0  54
sysvinit  191  0  17  0  209
termcap  0  0  186  0  186
terminfo  0  0  884  0  884
texinfo  0  0  489  0  489
textutil  8  0  320  0  329
timezone  0  0  364  0  364
umsprogs  35  0  3  0  38
util  342  0  406  0  748
vim  1  0  1162  0  1164
yast  2637  0  368  0  3006
Σ  15406  65 38904  114  54508

Series a2

oldshlbs  1820  0  9  0  1829
Σ  1820  0  9  0  1829

Series ap1

a2ps  0  0  30  0  30
alien  0  0  100  1  101
apmd  0  60  107  0  167
aps  0  0  31  128  159
arc  0  0  122  0  122
ash  60  0  33  0  93
at  0  0  55  1  57
autolog  0  0  19  0  20
bc  0  0  214  0  214
buffer  0  0  13  0  13
bzip  0  0  121  0  121
cdcons  0  0  21  0  21
cdrecord  0  0  208  0  208
cdwrite  0  0  104  0  104
cdwtools  0  0  315  0  315
cfengine  0  0  764  0  764
chipmunk  0  0  265  0  265
compat  0  0  522  0  522
ctags  0  0  29  0  29
deco  0  0  216  0  216
dialog  0  0  34  0  34
dosfsck  90  0  1  0  92
duconv  0  0  4  0  4
dump  119  0  11  0  131
ed  0  0  57  0  57
edy  0  0  70  0  70
enscript  4  0  1745  0  1749
faxprint  0  0  1  1  3
gcal  0  0  1535  0  1535

0.2. Disk usage of the packages

Package  /  /usr/X11  /usr  /var  Σ
gdtmon  294  0  0  0  294
git  0  0  406  0  406
glimpse  0  0  1086  0  1086
gpm  2  0  177  1  181
gs both  0  0  1309  0  1309
gs fonto  0  0  1488  0  1488
gs fonts  0  0  2154  0  2154
gs serv  0  0  1268  0  1268
gs vga  0  0  1273  0  1273
gs x11  0  0  1304  0  1304
hello  0  0  34  0  34
id utils  0  0  291  2  294
idled  0  0  43  0  43
ingres  0  0  2376  0  2376
ingrtool  0  0  312  0  312
inlinem  0  0  44  0  44
isapnp  36  0  60  0  96
jed  0  270  1088  0  1358
joe  0  0  262  0  262
jove  0  0  411  3  415
lha  0  0  46  0  46
lsof  0  0  363  0  363
makewhat  0  0  2  0  2
mbedit  0  0  353  0  353
mc  0  0  716  0  716
mdutils  21  0  27  0  48
mkdosfs  15  0  5  0  21
mkisofs  0  0  154  0  154
mktemp  3  0  2  0  5
mmv  0  0  111  0  111
mpegutil  0  0  93  0  93
mt st  0  0  14  0  14
mtools  1  0  724  0  725
nc  0  0  853  0  853
obst  0  0  9935  0  9935
octave  0  0 47594  0  47594
onyx  0  0  557  0  557
pdksh  0  0  307  0  307
pico  0  0  123  0  123
postgres  0  1533  2871  836  5242
quota  0  0  58  0  58
recode  0  0  199  0  199
rman  0  0  132  0  132
rpm2html  0  0  174  6  181
sc  0  0  213  0  213
screen  0  0  255  0  255
scsiinfo  0  0  163  0  163
seejpeg  0  0  439  0  439

Package  /  /usr/X11  /usr  /var  Σ
sharutil  0  0  571  0  571
splitvt  0  0  83  0  83
sudo  0  0  337  1  338
svgatext  44  0  584  1  630
taper  294  0  23  0  317
tcsh  245  0  65  0  311
teapot  0  0  657  0  657
unarj  0  0  125  0  125
units  0  0  208  0  208
unzip  0  0  170  0  170
xdelta  0  0  191  0  191
yamm  0  0  47  0  47
ytree  0  0  101  0  101
z88  0  4708  1131  0  5840
zip  0  0  124  0  124
ziptool  0  0  8  0  8
zoo  0  0  104  0  104
zsh  0  0  646  0  646
Σ  1228  6571 93791  981  102586

Series ap2

gs lib  0  0  2261  0  2261
Σ  0  0  2261  0  2261

Series d1

autoconf  0  0  557  0  557
automake  0  0  526  0  526
beg  0  0  1709  0  1709
bin86  0  0  71  0  71
binuti o  0  0  6984  0  6984
binutils  0  0  2040  0  2040
bison  0  0  227  0  227
bissawt  0  0  2548  0  2548
bwbasic  0  0  149  0  149
clisp  0  0 12954  0  12954
cvs  0  0  739  0  739
efence  0  0  41  0  41
f2c  0  0  812  0  812
f2caout  0  0  693  0  693
flex  0  0  183  0  183
fpk  0  0  2857  0  2857
g77  0  0  2387  0  2387
gcc  0  0  2986  0  2986
gcc o  0  0 10482  0  10482
gcl  0  0  3476  0  3476
gdb  0  0  1154  0  1154
gdlib  0  0  182  0  182
gettext  0  0  945  0  945
gforth  0  0  934  0  934

Package  /  /usr/X11  /usr  /var  Σ
gmp  0  0  300  0  300
gnat  0  0 15576  0  15576
gnats  0  0  1129  8  1137
gobjc  0  0  2574  0  2574
gpc  0  0  2604  0  2604
gpp  0  0  2897  0  2897
guavac  0  0  2719  0  2719
guile  0  0  1608  0  1608
icmake  0  0  221  0  221
indent  0  0  148  0  148
java  0  0 26096  0  26096
javadoc  0  0 11030  0  11030
javarunt  0  0  9477  0  9477
jdk115  0  0 32701  0  32701
jresep  0  0  6066  0  6066
kaffe  0  0 10671  0  10671
lclint  0  0  2394  0  2394
leda  0  0 14939  0  14939
libc  0  0  2324  0  2324
libd  0  0  5743  0  5743
libgpp  0  0  3744  0  3744
libobjc  0  0  2852  0  2852
linclude  0  0  2350  0  2350
linux  0  0 22122  0  22122
lthreads  55  0  113  0  169
lx hack  0  0 59228  0  59228
lx suse  0  0 25310  0  25310
make  0  0  223  0  223
minigcc  0  0  78  0  78
mocka  0  0  1620  0  1620
mtc  0  0  586  0  586
p2c  0  0  593  0  593
patch  0  0  134  0  134
pmake  0  0  124  0  124
prcs  0  0  490  0  490
python  0  0 14144  0  14144
rcs  0  0  452  0  452
scheme  0  0  162  0  162
slang  0  0  292  0  292
smaltalk  0  0  1393  0  1393
smeiffel  0  0  1016  0  1016
strace  0  0  89  0  89
svgaaout  0  0  166  0  166
svgalib  9  0  1073  0  1083
swi pl  0  0  1855  0  1855
toolpack  0  0  5585  0  5585
tvision  0  0  4708  0  4708
tya  0  0  114  0  114

Package  /  /usr/X11  /usr  /var  Σ
xclass  0  5461  40  0  5501
xscheme  0  0  218  0  218
yacc  0  0  72  0  72
Σ  64  5461 357799  8  363334

Series d2

joystick  0  0  44  0  44
Σ  0  0  44  0  44

Series doc1

allman  0  0 11497  0  11497
books  0  0 23873  0  23873
booksfnt  0  0  0  3705  3705
cdb  0  0  1447  0  1447
changes  0  0  1332  0  1332
dns faq  0  0  605  0  605
elispint  0  0  1265  0  1265
elispman  0  0  5247  0  5247
fhs  0  0  363  0  363
gawkdoc  0  0  1550  0  1550
gccinfo  0  0  399  0  399
ge info  0  0  893  0  893
ger man  0  0  253  0  253
gnuhtml  0  0  9958  0  9958
gppinfo  0  0  101  0  101
howto  0  0  5428  0  5428
howtodvi  0  0  8577  0  8577
howtohtm  0  0  9711  0  9711
howtoinf  0  0  445  0  445
howtops  0  0 18540  0  18540
howtosgm  0  0  7250  0  7250
i4ldoc  0  0  1185  0  1185
inf2htm  0  0  26  0  26
kcs  0  0  559  0  559
ldp  0  0 13700  0  13700
ldpman  0  0  749  0  749
libcinfo  0  0  430  0  430
man9  0  0  52  0  52
manyfaqs  0  0  4887  0  4887
odshtml  0  0  307  0  307
perlref  0  0  875  0  875
ppp nt  0  0  63  0  63
rfc  0  0 25400  0  25400
sdb  0  0 11017  0  11017
sdb cgi  0  0  12  0  12
selfhtml  0  0  3859  0  3859
suse lx  0  0  7875  0  7875
suse lxe  0  0 15303  0  15303
susebuch  0  0  784  0  784

Package  /  /usr/X11  /usr  /var  Σ
susehilf  0  0  4114  0  4114
susepak  0  0 18735  0  18735
susepake  0  0 18688  0  18688
toppp  0  0  8  0  8
uucpdoc  0  0  1536  0  1536
wie dvi  0  0  1591  0  1591
wie geht  0  0  387  0  387
wie html  0  0  1410  0  1410
wie info  0  0  54  0  54
wie pdf  0  0  8088  0  8088
wie ps  0  0  2558  0  2558
wie sgml  0  0  1136  0  1136
xe info  0  2442  0  0  2442
xf86html  0  0  6933  0  6933
Σ  0  2442 261055  3705  267202

Series e1

elib  0  0  188  0  188
ge auc  0  0  1406  0  1406
ge exec  0  0 12513  3  12516
ge lisp  0  0 11556  0  11556
ge site  14  0  974  0  988
lacheck  0  0  51  0  51
psgml  0  0  720  0  720
xe exec  0  33926  0  2  33929
xe lisp  0  26059  0  0  26059
Σ  14  59985 27408  5  87413

Series emu1

adamem  0  143  22  0  166
dosemu  8  7  2120  626  2762
ei  0  0  33  0  33
ibcs2  200  0  581  0  781
ines  0  96  84  0  181
stella  0  113  177  0  290
stonx  0  911  931  2  1845
stonxsft  0  0  2 10001  10004
uae  0  0  687  0  687
vgb  0  151  245  0  396
vice  0  0  1598  0  1598
wine  1  0 10703  0  10704
xzx  0  345  121  0  466
z81  0  56  180  0  236
Σ  209  1822 17484 10629  30149

Series fun1

3d chess  0  41  5  0  47
antipoli  0  0  103  0  103
asclock  0  24  5  0  29
astrolog  0  257  234  0  491

Package  /  /usr/X11  /usr  /var  Σ
batalion  0  521  1607  0  2129
battball  0  286  11  0  297
bsdgames  0  0  1487  14  1502
craft  0  20489  489  0  20978
crosfire  0  24432  3178  0  27610
daliclck  0  63  0  0  63
emiclock  0  245  0  0  245
empire  0  0  105  0  105
fishtank  0  387  9  0  396
flying  0  154  0  0  154
fortune  0  0  3574  0  3574
freeciv  0  241  1377  0  1618
frisk  0  87  67  0  155
glchess  0  22  116  0  139
gnuchess  0  0  1812  0  1812
gnugo  0  0  38  0  38
gshogi  0  0  558  0  558
hextris  0  783  0  1  785
imaze  0  1135  517  0  1652
lincity  0  1567  1613  0  3181
maelstr  0  253  2246  0  2499
manix  0  17  12  1  31
meltflip  0  18  0  0  18
mirrma  0  74  800  21  896
moontool  0  786  1  0  788
net3d  0  201  0  0  201
nethack  0  0  1633  0  1633
oneko  0  35  7  0  43
oonsoo  0  474  4  0  479
pacman  0  183  0  0  183
pipeman  0  38  0  0  38
playone  0  1011  157  0  1168
puzzle  0  21  0  0  21
rockdiam  0  123  1241  248  1613
sastroid  0  0  311  2  313
space  0  0  1818  0  1818
thrust  0  0  256  0  256
vgacard  0  0  120  0  120
x3d  0  1061  6  0  1067
xabacus  0  13  0  0  13
xancur  0  270  14  0  284
xaos  0  0  279  0  279
xball  0  1133  16  0  1149
xbill  0  33  0  152  186
xbl  0  502  13  0  515
xblast  0  340  0  0  340
xboard  0  293  30  0  324
xboing  0  599  378  0  978

Package  /  /usr/X11  /usr  /var  Σ
xbomb  0  22  0  0  22
xbombs  0  18  0  1  19
xdemine  0  27  0  0  27
xearth  0  173  0  0  173
xengine  0  11  0  0  11
xfract  0  1216  4  0  1221
xgammon  0  3274  0  0  3274
xgas  0  34  0  0  34
xjewel  0  45  0  0  45
xlife  0  467  65  0  533
xlyap  0  29  0  0  29
xmahjong  0  47  7  0  54
xmemory  0  27  2  0  29
xmine  0  24  0  0  24
xmines  0  20  0  0  20
xmountns  0  35  5  0  41
xmris  0  185  0  0  185
xpat2  0  475  0  0  475
xphoon  0  88  0  0  88
xpilot  0  580  822  0  1403
xpinguin  0  181  0  0  181
xpuzzles  0  614  0  0  614
xroach  0  18  0  0  18
xrobots  0  20  0  0  20
xshogi  0  110  0  0  110
xskat  0  180  2  0  182
xsnow  0  24  5  0  30
xsok  0  56  59  5  121
xsol  0  29  0  0  29
xspringi  0  59  412  0  471
xtacy  0  67  6  0  74
xtartan  0  59  0  0  59
xteddy  0  68  8  0  77
xtetris  0  40  0  0  40
xthing  0  3891  3  0  3894
xtron  0  117  0  0  117
xvier  0  37  0  0  37
yahtzee  0  0  19  2  21
Σ  0  70579 27666  447  98716

Series gra1

bttv  0  1115  89  0  1204
cgmdraw  0  0  132  0  132
faxedit  0  490  66  0  557
films  0  0 18452  0  18452
freetype  0  0  1249  0  1249
gd  0  0  225  0  225
gimp  0  0 29568  0  29568

Package  /  /usr/X11  /usr  /var  Σ
gimpman  0  0 18843  0  18843
gsview  0  123  0  0  123
gtk  0  0  9706  0  9706
gv  0  278  123  0  402
html2ps  0  36  300  0  337
imagemag  0  2658  717  0  3376
imfilm  0  0  4434  0  4434
imlib  2  98  0  0  100
impict  0  0  529  0  529
jpeg  0  88  164  0  252
libfbm  0  167  14  0  182
libgr  0  0  9  0  9
libgrout  0  0  1291  0  1291
libjpeg  0  352  0  0  352
libmpeg  0  0  320  0  320
libnetpb  0  182  0  0  182
libpng  0  369  0  0  369
librle  0  153  0  0  153
libtiff  0  532  0  0  532
libz  0  121  205  0  326
mlcreat  0  1360  2654  0  4015
mpegplay  0  126  4  0  131
muscan  0  23  53  0  77
netpbm  0  1582  103  0  1685
photocd  0  53  214  0  268
pictures  0  0  4014  0  4014
pixmp  0  196  0  0  196
povray  0  573  5878  0  6451
pstoedit  0  0  226  0  226
sane  1  138  1432  0  1571
sced  0  497  415  0  912
surf  0  0  1815  0  1815
tgif  0  1943  14  0  1958
tiff  0  410  329  0  740
tkdraw  0  294  3  0  297
xanim  0  517  828  0  1345
xbmbrosr  0  65  0  0  65
xgrabsc  0  80  13  0  94
xgrasp  0  56  140  0  197
xli  0  235  0  0  235
xmgr  0  1889  1893  0  3783
xmorph  0  1121  8  0  1129
xmrm  0  299  1155  0  1454
xpaint  0  469  18  0  487
xpdf  0  256  330  0  587
xray  0  8632 14434  0  23066
xrayps  0  0 115351  0  115351
xwarp  0  259  5  0  265

Package  /  /usr/X11  /usr  /var  Σ
zgv  0  0  225  0  225
Σ  3  27835 237990  0  265844

Series kde1

kadm  692  0  3  0  695
kadmin  461  0  3  0  464
kdesdk  56  0  0  0  56
kdevel  116  0  1  0  117
kgames  4490  0  25  0  4515
kgraph  1562  0  0  0  1562
klibs  2327  0  31  0  2359
kmisc  3139  0  0  0  3139
kmulti  1205  0  28  0  1233
knet  2322  0  25  0  2347
knetutil  1613  0  6  0  1619
kscience  230  0  0  0  230
ksupp  1119  0  26  0  1145
ktools  982  0  5  0  987
kutils  2291  0  27  0  2318
Σ  22605  0  180  0  22786

Series kde2

kbase  5175  2  37  1  5216
Σ  5175  2  37  1  5216

Series n1

amanda  1  0  1685  3  1689
amd  1  0  776  0  778
apache  19  0  1052  0  1072
archie  0  0  39  0  39
autofs  4  0  92  0  96
bind  2  0  1674  3  1680
bing  0  0  52  0  52
bulkmail  0  0  27  0  27
cnews  0  0  973  30  1004
deliver  0  0  82  1  83
dhcp  5  0  689  0  694
diald  0  0  220  0  220
dochost  8  0  9  1  19
elm  0  0  584  0  584
emil  0  0  342  0  342
firewall  20  0  228  0  249
ftpdir  0  0  765  0  765
gettyps  0  0  84  0  84
gup  0  0  38  5  44
htdig  2635  0  1381  0  4017
hylafax  3  0  4382  410  4796
i4l  436  19  882  4  1343
i4lfirm  0  0  644  0  644
inetcfg  0  0  38  0  38

Package  /  /usr/X11  /usr  /var  Σ
inn  0  0  1382  95  1477
ipfwadm  25  0  10  0  35
ipxrip  0  0  32  0  32
ircii  0  0  508  0  508
irciihlp  0  0  357  0  357
lprold  2  0  238  1  242
lynx  0  0  1889  0  1889
mailx  0  0  74  0  74
marsnwe  32  0  597  0  630
mdomo  10  0  691  5  706
metamail  0  0  371  0  371
mgetty  20  0  1035  3  1058
mh  0  0  4507  0  4507
minicom  1  0  154  1  156
mininews  0  0  36  0  36
mirror  0  0  154  0  154
mrtg  0  0  262  0  262
mutt  0  0  582  0  582
ncftp  0  0  237  0  237
ncpfs  356  0  7268  0  7624
netatalk  5  0  503  0  509
netcat  0  0  121  0  121
nn  0  0  666  0  666
nn spool  0  0  656  0  656
nntpd  0  0  150  0  151
pine  0  0  1641  0  1641
plp  3  0  1216  0  1220
pop  0  0  2174  0  2174
ppp  1  0  130  0  131
procmail  0  0  179  0  179
radiusc  24  0  555  0  579
radiusd  9  0  55  2  67
rdist  0  0  108  0  108
rsync  0  0  469  0  469
rzsz  0  0  84  0  84
samba  1  0  2648  0  2650
sendfax  3  0  131  3  137
sendmail  119  0  1527  4  1651
seyon  0  428  2  0  430
slip  5  0  196  0  202
slrn  0  0  301  18  319
smail  11  0  580  5  596
snmp  2  0  425  0  428
squid  40  0  577  2  620
strn  0  0  797  0  797
suck  0  0  146  0  146
susefax  0  2  443  0  445
suseppp  23  0  22  1  47

Package  /  /usr/X11  /usr  /var  Σ
tcpdump  0  0  626  0  626
tin  0  0  438  0  438
trn  0  0  315  0  315
trn spl  0  0  293  0  293
url get  0  0  53  0  53
uucp  0  0  1084  4  1089
uucpcfg  0  0  0  17  17
vacation  0  0  35  0  35
vbox  2  16  389  1  409
vbox2b4  1  0  416  0  417
w3mir  0  0  1190  0  1190
webalize  9  0  175  0  184
weblint  0  0  126  0  126
wget  3  0  159  0  163
whois  0  0  21  0  21
wuftpd  1  0  154  0  156
wwwoffle  18  0  528  2  548
xinetd  5  0  164  0  170
xntp  0  0  1136  0  1137
xrn  0  507  0  0  507
ypclient  1  0  113  1  116
ypserv  1  0  249  15  266
Σ  3867  972 61288  637  66795

Series pay1

abuse  0  0  6080  0  6080
ac3d dem  0  1721  2495  0  4217
acroread  0  8366  25  0  8391
adabas  1  0 83520  0  83522
adadoc  0  0 14210  0  14210
adadok  0  0 22535  0  22535
adbsmydb  0  0 41458  0  41458
arcad  33945  0  0  0  33945
arkeia  0  0 21573  0  21574
bartels  53240  0  1  0  53241
c16 demo  0  0  5717  0  5717
crisp  0  1690  8559  0  10249
cysched  1633  0  880  0  2514
eloquen  9078  0  1730  5121  15929
empress  0  0 56548  0  56548
executor  0  0  7242  0  7242
flagship  0  0 18637  0  18637
graffy  3899  0  0  0  3899
heitml  0  0  3961  0  3961
ivs demo  0  0  548  0  548
ixware  0  0 17737  0  17737
mapledmo  0  134  4512  0  4646
mars e  0  0  252  0  252

Package  /  /usr/X11  /usr  /var  Σ
mhitdemo  0  80 11323  2  11406
mind  0  0 15569  0  15569
netscape  17279  3  1  0  17283
not demo  0  0  3669  0  3669
simcity  0  0  4664  0  4664
snavig  8989  0  12  0  9001
sniff  42024  0  110  0  42135
terroid  0  0  1971  0  1971
uc  0  0  1394  0  1394
vdx  0  2580  0  0  2580
vse  0  0  6018  0  6018
wingz  0  1  3647  0  3649
xess3  0  0  4003  0  4003
xgal  0  198  165  0  364
xv  0  855  1069  0  1924
yardsql  0  0 20648  0  20648
Σ  170088  15628 392483  5123  583330

Series sgm1

docbk30  0  0  4701  0  4701
gf  0  0  463  0  463
html dtd  0  0  3178  0  3178
hyperref  0  0  615  0  615
iso ent  0  0  97  0  97
jade dsl  0  0  1852  0  1852
jadetex  0  0  1337  0  1337
linuxdoc  0  0  753  0  753
sgmltool  0  0  783  0  783
sgrep  6  67  98  0  171
sp  0  0  4072  0  4072
Σ  6  67 17949  0  18022

Series snd1

aumix  0  0  53  0  53
gmod  0  205  188  0  394
jazz  0  345  98  0  444
mxv  0  2180  57  0  2237
rplay  1  0  1009  0  1010
sapphire  0  0  1407  0  1407
snd au  0  0  1513  0  1513
snd mod  0  0  827  0  827
snd wav  0  0  2692  0  2692
speakf  0  0  307  0  307
tcd  0  0  38  0  38
tracker  0  0  402  0  402
wavplay  0  33  0  0  33
workbone  0  0  36  0  36
workman  0  0  307  0  307
xmcd  0  2151  76  0  2228

Package  /  /usr/X11  /usr  /var  Σ
xmitsumi  0  909  2  0  912
xmix  0  14  0  0  14
xmp  0  0  73  0  73
xqmixer  0  44  2  0  46
Σ  1  5881  9087  0  14973

Series spl1

dicts  0  0 10555  0  10555
idanish  0  0  1650  0  1650
idutch  0  0  3121  0  3121
ifrench  0  0  1503  0  1503
iitalian  0  0  1654  0  1654
inorsk  0  0  1091  0  1091
ispanish  0  0  2813  0  2813
ispell  0  0  6001  0  6001
iswedish  0  0  788  0  788
words  0  0  812  0  812
Σ  0  0 29988  0  29988

Series spl2

igerman  0  0  3593  0  3593
Σ  0  0  3593  0  3593

Series tcl1

blt  0  0  823  0  823
convert  0  158  5  0  164
dejagnu  0  0  424  0  424
eddi  0  270  29  0  300
expect  0  0  698  0  698
itcl  0  0 17816  0  17816
jstools  0  32  2548  0  2580
listexec  0  14  0  0  14
omnimoni  0  0  238  0  238
samt  1  17  212  0  232
scotty  0  1  7614  0  7615
tcl  0  0  1338  0  1338
tcl new  0  0  1991  0  1991
tcl old  0  0  951  0  951
tclx  0  1102  1194  0  2297
tix  0  5380  1510  0  6890
tix new  0  2789  1896  0  4686
tixlpq  0  12  3  0  16
tk  0  2843  286  0  3129
tk new  0  2340  745  0  3085
tk old  0  1197  1303  0  2500
tkdesk  0  1541  460  0  2002
tkfont  0  12  28  0  41
tkinfo  0  131  10  0  142
tkinspec  0  145  0  0  145
tkirc  0  302  86  0  389

Package  /  /usr/X11  /usr  /var  Σ
tkmail  0  863  19  0  883
tkman  0  219  41  0  261
tkman8  0  290  22  0  313
tkppp  0  23  9  0  32
tkrat  0  1976  95  0  2071
tkstep  0  1588  160  0  1748
tkworld  0  622  32  0  654
tkxcd  0  21  25  0  46
vtcl  0  0  419  0  419
xcdroast  0  874  86  0  960
xf  0  243  6233  0  6476
xmclock  0  69  0  0  69
zircon  0  1  489  0  490
Σ  1  25075 49838  0  74928

Series tex1

bibview  0  219  0  0  219
chess  0  0  745  0  745
colortbl  0  0  90  0  90
detex  0  0  20  0  20
dvi2tty  0  0  46  0  46
dviutils  0  0  255  0  255
l2h  2  0  1506  223  1732
l2hhtml  0  0  635  0  635
lyx  0  1247  2928  0  4175
musictex  0  0  4231  0  4231
psutils  0  0  191  0  191
tetex  1  0 40175  3  40179
texcad  0  82  29  0  112
tktex  0  36  117  0  154
transfig  0  222  94  0  316
ts  0  511  0  0  511
xfig  0  648  1222  0  1871
xtem  0  2754  154  0  2908
Σ  3  5719 52438  226  58390

Series x1

xaw3d  0  957  10  0  967
xaw3dd  0  295  0  0  296
xaw95  0  1286  10  0  1297
xdevel  0  5407  0  0  5407
xdmbgrd  0  12  0  0  12
xdoc  0  376  0  0  376
xextra  0  5279  0  0  5279
xf86  0  8904  5  8  8917
xfnt100  0  2088  0  0  2088
xfntbig  0  2440  0  0  2440
xfntcyr  0  370  0  0  370
xfntscl  0  1599  0  0  1599

Package  /  /usr/X11  /usr  /var  Σ
xfsetup  0  1227  0  0  1227
xlkit  0  12994  0  0  12994
xman  0  1232  0  0  1232
xpm  0  174  0  0  174
xshared  0  1916  0  0  1916
xxprt  0  2928  0  0  2928
Σ  0  49484  25  8  49519

Series x2

xdmsc  3  3  7  1  15
Σ  3  3  7  1  15

Series x3

xvga16  0  2033  0  0  2033
Σ  0  2033  0  0  2033

Series x4

x8514  0  1790  0  0  1790
xagx  0  1980  0  0  1980
xi128  0  2244  0  0  2244
xmach32  0  1941  0  0  1941
xmach64  0  2073  0  0  2073
xmach8  0  1803  0  0  1803
xmono  0  2106  0  0  2106
xp9k  0  2004  0  0  2004
xs3  0  2503  0  0  2503
xs3v  0  2222  0  0  2222
xsvga  0  3184  0  0  3184
xw32  0  1843  0  0  1843
Σ  0  25693  0  0  25693

Series x5

xglint  0  2305  1  5  2311
Σ  0  2305  1  5  2311

Series x6

oldlibs5  0  0  1150  0  1150
oldlibs6  0  0  2817  0  2817
Σ  0  0  3967  0  3967

Series x7

modegen  0  21  1  0  22
Σ  0  21  1  0  22

Series xap1

addrbo  0  116  330  0  446
arena  0  1204  0  0  1204
asedit  0  1578  0  0  1578
auis ful  0  0 38371  0  38371
auis ml  0  0  5018  0  5018
auis src  0  0  1786  0  1786
auis wp  0  0  5674  0  5674

Package  /  /usr/X11  /usr  /var  Σ
axe  0  203  28  0  232
axe3d  0  204  28  0  232
babaya  0  251  0  0  251
calctool  0  57  0  0  57
cellular  0  0  667  0  667
conv  0  6  18  0  25
coolmail  0  33  1  0  35
elem  0  17  19  0  36
fader  0  6  5  0  11
felix  0  0  1060  0  1060
gnuplot  0  0  3267  0  3267
gvim  5  390  1  0  398
knews  0  687  0  0  687
mfm  0  1250  143  0  1394
mgdiff  0  1158  0  0  1158
mupad  0  0 31775  0  31775
nedit  0  2193  210  0  2403
offix  0  402  78  0  481
pan  0  1028  3  0  1031
pcb  0  695  0  0  695
pixedit  0  2055  2  0  2057
plan  0  2145  235  0  2380
plotmtv  0  1954  1008  0  2962
procmtr  0  25  33  0  58
qddb  0  0  3370  0  3370
qps  0  1434  30  0  1464
rasmol  0  206  1292  0  1498
rxvt  0  167  175  0  342
scilab  0  51 24972  0  25023
snns  0  0  5539  0  5539
snnsdoc  0  0  4134  0  4134
spice  0  0  1851  0  1851
spicecad  0  0  6868  0  6868
susecurs  0  5  0  0  5
vidtext  0  136  189  56  381
x3270  0  455  282  0  738
xarchie  0  281  0  0  281
xautolck  0  18  0  0  18
xbanner  0  76  163  0  240
xbarcode  0  78  2  0  81
xbench  0  98  4  0  103
xcb  0  16  0  0  16
xcnf  0  34  2  0  37
xcolors  0  9  0  0  9
xcoral  0  519  6225  0  6744
xdf  0  16  4  0  20
xdtm  0  143  110  0  254
xephem  0  4597  52  0  4649

Package  /  /usr/X11  /usr  /var  Σ
xfileman  0  99  60  0  160
xflame  0  904  1  0  905
xfm  0  583  0  0  583
xfmail  0  727  56  0  784
xformat  0  26  24  0  50
xftp  0  1448  10  0  1459
xgopher  0  260  0  0  260
xgrok  0  2  1739  0  1742
xgroups  0  60  18  0  79
xhtalk  0  58  6  0  64
xinfo  0  38  5  0  44
xirc  0  639  60  0  699
xkeycaps  0  626  0  0  626
xled  0  25  1  0  27
xless  0  37  0  0  37
xlock  0  1781  0  0  1781
xlogin  0  4  0  0  4
xmail  0  135  0  0  135
xmftp  0  3356  12  0  3368
xnetload  0  7  21  0  29
xosview  0  65  21  0  87
xplot  0  0  374  0  374
xplz  0  11076  45  0  11122
xprompt  0  16  0  0  16
xrolodex  0  1170  0  0  1170
xrpm  0  0  254  1  255
xscrns  0  10607  32  0  10640
xselect  0  9  0  0  9
xsession  0  48  38  0  87
xsky  0  25379  0  0  25379
xspline  0  11  0  0  11
xspread  0  176  433  0  610
xsysinfo  0  15  9  0  24
xtail  0  1034  1  0  1036
xtalk  0  1176  660  0  1836
xtar  0  1140  7  0  1148
xtexit  0  10  3  0  13
xvhelp  0  48  46  0  94
xvi  0  0  477  0  477
xvnews  0  150  0  0  150
xvtdl  0  222  0  0  222
xwpick  0  41  0  0  41
xzap  0  39  0  0  39
xzet  0  50  37  0  87
Σ  5  89293 149474  57  238857

Series xdev1

ddd  0  224  8736  0  8960

Package  /  /usr/X11  /usr  /var  Σ
dlgedit  0  836  23  0  859
iv  0  0  6970  0  6970
mesa  0  0 11010  0  11010
qtdevel  0  2366  3610  0  5976
qtext  0  597  0  0  597
qtlib  0  1197  42  0  1240
vtk  0  76107  912  0  77020
vtkdata  0  0 128476  0  128476
wxwdemo  0  0  1472  0  1472
wxwdocht  0  0  1994  0  1994
wxwdocps  0  0  4533  0  4533
wxwin  0  0  4614  0  4614
wxwstat  0  0  3336  0  3336
wxxt  0  3490  3037  0  6527
xforms  0  1934  2011  0  3945
xtpanel  0  342  15  0  357
xwb  0  79  25  0  104
xwpe  0  0  570  0  570
xxgdb  0  85  0  0  85
ygl  0  891  51  0  943
Σ  0  88148 181437  0  269588

Series xv1

ssbase  0  0  1404  0  1404
ssexmp  0  0  430  0  430
xtoolpl  0  30  0  0  30
xv32aout  0  0  3205  0  3205
xv32base  0  7  2228  1  2237
xv32dev  0  50  1160  0  1211
xv32exmp  0  0  1115  0  1115
xv32slib  0  0  1941  0  1941
Σ  0  87 11483  1  11573

Series xwm1

3dpixm  0  2425  0  0  2425
3dpixms  0  1195  0  0  1195
afterstp  0  595  157  0  753
amiwm  0  115  7  0  122
bowman  0  162  75  0  238
cdesim  0  87  119  0  207
ctwm  0  258  60  0  319
enlight  0  2227  39  0  2266
fvwm  0  1310  295  0  1605
fvwm1  0  390  234  0  624
fvwm95  0  767  171  0  939
icons  0  5233  2  0  5236
lxuser  0  0  0  1  1
mlvwm  0  118  44  0  163
olvwm  0  0  288  0  288

0.3. Disk usage of the series

Package  /  /usr/X11  /usr  /var  Σ
pixmaps  0  2561  1  0  2563
qvwm  0  124  58  0  183
susewm  7  491  290  2  792
wm2  0  65  7  0  72
Σ  7  18123  1847  3  19991

0.3 Disk usage of the series

Serie  /  /usr/X11  /usr  /var  Σ
a1  15406  65  38904  114  54508
a2  1820  0  9  0  1829
ap1  1228  6571  93791  981  102586
ap2  0  0  2261  0  2261
d1  64  5461  357799  8  363334
d2  0  0  44  0  44
doc1  0  2442  261055  3705  267202
e1  14  59985  27408  5  87413
emu1  209  1822  17484 10629  30149
fun1  0  70579  27666  447  98716
gra1  3  27835  237990  0  265844
kde1  22605  0  180  0  22786
kde2  5175  2  37  1  5216
n1  3867  972  61288  637  66795
pay1  170088  15628  392483  5123  583330
sgm1  6  67  17949  0  18022
snd1  1  5881  9087  0  14973
spl1  0  0  29988  0  29988
spl2  0  0  3593  0  3593
tcl1  1  25075  49838  0  74928
tex1  3  5719  52438  226  58390
x1  0  49484  25  8  49519
x2  3  3  7  1  15
x3  0  2033  0  0  2033
x4  0  25693  0  0  25693
x5  0  2305  1  5  2311
x6  0  0  3967  0  3967
x7  0  21  1  0  22
xap1  5  89293  149474  57  238857
xdev1  0  88148  181437  0  269588
xv1  0  87  11483  1  11573
xwm1  7  18123  1847  3  19991
Σ  220505k  503294k 2029534k 21951k 2775476k

0.4 Which series contains which package?

Package  Series Page Package Series Page Package Series Page
3d chess  fun1 56 babaya  xap1 131 compress a1  3
3dpixm  xwm1 151 bartels  pay1 100 conv  xap1 132
3dpixms xwm1 151 base  a1  1  convert  tcl1 115
a2ps  ap1  12 bash  a1  2  coolmail xap1 132
aaa base  a1  1  bash1  a1  2  cpio  a1  3
aaa dir  a1  1  batalion  fun1 56 craft  fun1 57
abuse  pay1 98 battball  fun1 56 crisp  pay1 100
ac3d dem pay1 98 bc  ap1  14 cron  a1  3
acroread  pay1 99 bdflush  a1  2  crosfire  fun1 57
adabas  pay1 99 beg  d1  28 ctags  ap1  15
adadoc  pay1 99 bibview  tex1 121 ctwm  xwm1 152
adadok  pay1 99 bin86  d1  28 cvs  d1  30
adamem  emu1 51 bind  n1  84 cysched  pay1 100
adbsmydb pay1 99 bing  n1  84 daliclck  fun1 57
addrbo  xap1 130 binuti o  d1  28 db  a1  3
afterstp  xwm1 151 binutils  d1  28 ddd  xdev1 146
alien  ap1  12 bison  d1  29 deco  ap1  15
allman  doc1 43 bissawt  d1  29 dejagnu  tcl1 115
amanda  n1  83 blt  tcl1 115 deliver  n1  85
amd  n1  83 books  doc1 43 detex  tex1 122
amiwm  xwm1 151 booksfnt  doc1 44 devs  a1  3
antipoli  fun1 56 bowman xwm1 151 dhcp  n1  85
apache  n1  83 bsdgames fun1 56 diald  n1  85
apmd  ap1  13 bttv  gra1  67 dialog  ap1  15
aps  ap1  13 buffer  ap1  14 dicts  spl1 113
arc  ap1  13 bulkmail  n1  85 diff  a1  3
arcad  pay1 99 bwbasic  d1  29 dlgedit  xdev1 147
archie  n1  84 bzip  ap1  14 dns faq  doc1 44
arena  xap1 130 c16 demo pay1 100 docbk30 sgm1 108
arkeia  pay1 99 calctool  xap1 131 dochost  n1  86
asclock  fun1 56 cdb  doc1 44 dosemu  emu1 51
asedit  xap1 130 cdcons  ap1  14 dosfsck  ap1  16
ash  ap1  13 cdesim  xwm1 152 duconv  ap1  16
astrolog  fun1 56 cdrecord  ap1  14 dump  ap1  16
at  ap1  13 cdwrite  ap1  14 dvi2tty  tex1 122
auis ful  xap1 130 cdwtools  ap1  15 dviutils  tex1 122
auis ml  xap1 130 cellular  xap1 132 eazy  a1  3
auis src  xap1 131 cfengine  ap1  15 ed  ap1  16
auis wp  xap1 131 cgmdraw gra1  67 eddi  tcl1 115
aumix  snd1 110 changes  doc1 44 edy  ap1  16
autoconf  d1  27 chess  tex1 121 efence  d1  30
autofs  n1  84 chipmunk ap1  15 ei  emu1 52
autolog  ap1  14 clisp  d1  29 elem  xap1 132
automake  d1  28 cnews  n1  85 elib  e1  49
axe  xap1 131 colortbl  tex1 121 elispint  doc1 44
axe3d  xap1 131 compat  ap1  15 elispman doc1 44

0.4. Which series contains which package?

Package Series Page Package Series Page Package  Series Page
elm  n1  86 gdb  d1  31 guile  d1  33
eloquen pay1 102 gdbm  a1  4  gup  n1  86
elvis  a1  3  gdlib  d1  32 gv  gra1  69
emiclock fun1 57 gdtmon  ap1  17 gvim  xap1 132
emil  n1  86 ge auc  e1  50 gzip  a1  5
empire  fun1 57 ge exec  e1  50 heitml  pay1 103
empress pay1 103 ge info  doc1 45 hello  ap1  18
enlight  xwm1 152 ge lisp  e1  50 hextris  fun1 59
enscript  ap1  16 ge site  e1  50 howto  doc1 45
executor pay1 103 ger man doc1 45 howtodvi doc1 46
expect  tcl1 115 gettext  d1  32 howtohtm doc1 46
ext2fs  a1  4  gettyps  n1  86 howtoinf  doc1 46
f2c  d1  30 gf  sgm1 108 howtops  doc1 46
f2caout  d1  30 gforth  d1  32 howtosgm doc1 46
fader  xap1 132 gimp  gra1 68 htdig  n1  87
faxedit  gra1  68 gimpman gra1 68 html2ps  gra1  69
faxprint  ap1  16 git  ap1  17 html dtd  sgm1 109
felix  xap1 132 glchess  fun1 58 hylafax  n1  87
fhs  doc1 44 glimpse  ap1  17 hyperref  sgm1 109
file  a1  4  gmod  snd1 110 i4l  n1  87
fileutil  a1  4  gmp  d1  32 i4ldoc  doc1 46
films  gra1  68 gnat  d1  32 i4lfirm  n1  87
find  a1  4  gnats  d1  32 ibcs2  emu1 53
firewall  n1  86 gnuchess fun1 58 icmake  d1  33
fishtank fun1 57 gnugo  fun1 58 icons  xwm1 153
flagship pay1 103 gnuhtml doc1 45 id utils  ap1  18
flex  d1  30 gnuplot  xap1 132 idanish  spl1 113
flying  fun1 57 gobjc  d1  33 idled  ap1  18
fortune  fun1 58 gpc  d1  33 idutch  spl1 113
fpk  d1  30 gpm  ap1  17 ifrench  spl1 114
freeciv  fun1 58 gpp  d1  33 igerman  spl2 115
freetype gra1  68 gppinfo  doc1 45 iitalian  spl1 114
frisk  fun1 58 gppshare  a1  4  imagemag gra1  69
ftpdir  n1  86 graffy  pay1 103 imaze  fun1 59
fvwm  xwm1 152 groff  a1  4  imfilm  gra1  69
fvwm1  xwm1 152 gs both  ap1  17 imlib  gra1  69
fvwm95 xwm1 153 gs fonto  ap1  17 impict  gra1  69
g77  d1  31 gs fonts  ap1  17 indent  d1  33
gawk  a1  4  gs lib  ap2  27 ines  emu1 53
gawkdoc doc1 44 gs serv  ap1  17 inetcfg  n1  87
gcal  ap1  16 gs vga  ap1  17 inf2htm  doc1 46
gcc  d1  31 gs x11  ap1  18 ingres  ap1  18
gcc o  d1  31 gshogi  fun1 58 ingrtool  ap1  18
gccinfo  doc1 45 gsview  gra1 68 inlinem  ap1  19
gcl  d1  31 gtk  gra1 68 inn  n1  87
gd  gra1  68 guavac  d1  33 inorsk  spl1 114

Package Series Page Package Series Page Package  Series Page
ipfwadm  n1  88 ktools  kde1 82 make  d1  39
ipxrip  n1  88 kutils  kde1 82 makewhat ap1  20
ircii  n1  88 l2h  tex1 122 man  a1  6
irciihlp  n1  88 l2hhtml tex1 122 man9  doc1 47
isapnp  ap1  19 lacheck  e1  51 manix  fun1 59
iso ent  sgm1 109 lclint  d1  34 manyfaqs doc1 47
ispanish  spl1 114 ldp  doc1 46 mapledmo pay1 104
ispell  spl1 114 ldpman doc1 46 mars e  pay1 104
iswedish spl1 114 ldso  a1  5  marsnwe  n1  88
itcl  tcl1 115 leda  d1  34 mbedit  ap1  20
iv  xdev1 147 less  a1  5  mc  ap1  21
ivs demo pay1 104 lha  ap1  20 mdomo  n1  89
ixware  pay1 104 libaout  a1  5  mdutils  ap1  21
jade dsl sgm1 109 libc  d1  35 meltflip  fun1 59
jadetex  sgm1 109 libcinfo doc1 47 mesa  xdev1 147
java  d1  33 libd  d1  36 metamail  n1  89
javadoc  d1  34 libfbm  gra1  70 mfm  xap1 133
javarunt  d1  34 libgpp  d1  36 mgdiff  xap1 133
jazz  snd1 110 libgr  gra1  70 mgetty  n1  89
jdk115  d1  34 libgrout gra1  70 mh  n1  89
jed  ap1  19 libjpeg  gra1  70 mhitdemo pay1 104
joe  ap1  19 libmpeg gra1  70 mind  pay1 105
jove  ap1  20 libnetpb gra1  71 minicom  n1  89
joystick  d2  43 libobjc  d1  36 minigcc  d1  39
jpeg  gra1 69 libpng  gra1  71 mininews  n1  89
jresep  d1  34 librle  gra1  71 mirrma  fun1 59
jstools  tcl1 116 libtiff  gra1  71 mirror  n1  89
kadm  kde1 77 libz  gra1  71 mkdosfs  ap1  21
kadmin  kde1 78 lilo  a1  5  mkisofs  ap1  21
kaffe  d1  34 lincity  fun1 59 mktemp  ap1  21
kbase  kde2 82 linclude  d1  36 mlcreat  gra1  71
kbd  a1  5  linux  d1  37 mlvwm  xwm1 153
kcs  doc1 46 linuxdoc sgm1 109 mmv  ap1  21
kdesdk  kde1 78 listexec  tcl1 116 mocka  d1  39
kdevel  kde1 78 loadlin  a1  6  modegen  x7  130
kernmod  a1  5  localedb  a1  6  modules  a1  6
kgames  kde1 79 lprold  n1  88 moontool fun1 60
kgraph  kde1 79 lsof  ap1  20 mpegplay gra1  71
klibs  kde1 79 lthreads  d1  39 mpegutil  ap1  22
kmisc  kde1 80 lx hack  d1  39 mrtg  n1  89
kmulti  kde1 80 lx suse  d1  39 mt st  ap1  22
knet  kde1 80 lxuser  xwm1 153 mtc  d1  39
knetutil  kde1 81 lynx  n1  88 mtools  ap1  22
knews  xap1 133 lyx  tex1 122 mupad  xap1 133
kscience kde1 81 maelstr  fun1 59 muscan  gra1  72
ksupp  kde1 81 mailx  n1  88 musictex  tex1 123

Package  Series Page Package Series Page Package Series Page
mutt  n1  90 pine  n1  91 samba  n1  92
mxv  snd1 111 pipeman  fun1 60 samt  tcl1 116
nc  ap1  22 pixedit  xap1 134 sane  gra1 74
ncftp  n1  90 pixmaps xwm1 154 sapphire snd1 111
ncpfs  n1  90 pixmp  gra1  73 sastroid  fun1 61
ncurses  a1  6  plan  xap1 134 sc  ap1  24
nedit  xap1 133 playone  fun1 60 sced  gra1 74
net3d  fun1 60 plotmtv  xap1 134 scheme  d1  40
net tool  a1  7  plp  n1  91 scilab  xap1 135
netatalk  n1  90 pmake  d1  40 scotty  tcl1 116
netcat  n1  90 pop  n1  91 screen  ap1  24
netcfg  a1  7  postgres  ap1  23 scsiinfo  ap1  24
nethack  fun1 60 povray  gra1  73 sdb  doc1 48
netpbm  gra1  72 ppp  n1  91 sdb cgi  doc1 48
netscape  pay1 105 ppp nt  doc1 47 seejpeg  ap1  24
nkita  a1  7  prcs  d1  40 selfhtml doc1 48
nkitb  a1  7  procmail  n1  91 sendfax  n1  92
nn  n1  90 procmtr  xap1 134 sendmail n1  93
nn spool  n1  90 ps  a1  8  seyon  n1  93
nntpd  n1  90 psgml  e1  51 sgmltool sgm1 109
not demo pay1 105 pstoedit  gra1  73 sgrep  sgm1 110
nvi  a1  7  psutils  tex1 123 sh utils  a1  9
obst  ap1  22 puzzle  fun1 60 shadow  a1  9
octave  ap1  22 python  d1  40 sharutil  ap1  24
odshtml  doc1 47 qddb  xap1 135 shlibs  a1  9
offix  xap1 134 qps  xap1 135 simcity  pay1 106
oldlibs5  x6  129 qtdevel  xdev1 147 slang  d1  40
oldlibs6  x6  129 qtext  xdev1 147 slip  n1  93
oldshlbs  a2  11 qtlib  xdev1 148 slrn  n1  93
olvwm  xwm1 153 quota  ap1  23 smail  n1  93
omnimoni tcl1 116 qvwm  xwm1 154 smaltalk  d1  40
oneko  fun1 60 radiusc  n1  91 smeiffel  d1  41
onyx  ap1  22 radiusd  n1  91 snavig  pay1 106
oonsoo  fun1 60 rasmol  xap1 135 snd au  snd1 111
p2c  d1  39 rcs  d1  40 snd mod snd1 111
pacman  fun1 60 rdist  n1  91 snd wav snd1 111
pan  xap1 134 recode  ap1  23 sniff  pay1 106
patch  d1  39 rfc  doc1 47 snmp  n1  93
pcb  xap1 134 rman  ap1  23 snns  xap1 135
pcmcia  a1  7  rockdiam fun1 61 snnsdoc xap1 136
pdksh  ap1  23 rplay  snd1 111 sp  sgm1 110
perl  a1  8  rpm  a1  8  space  fun1 61
perlref  doc1 47 rpm2html ap1  23 speakf  snd1 111
photocd  gra1  72 rsync  n1  92 spice  xap1 136
pico  ap1  23 rxvt  xap1 135 spicecad xap1 137
pictures  gra1  73 rzsz  n1  92 splitvt  ap1  24

Package Series Page Package Series Page Package  Series Page
squid  n1  94 timezone  a1  10 vdx  pay1 107
ssbase  xv1 150 tin  n1  94 vgacard  fun1 61
ssexmp  xv1 150 tix  tcl1 117 vgb  emu1 54
stella  emu1 53 tix new  tcl1 117 vice  emu1 54
stonx  emu1 53 tixlpq  tcl1 118 vidtext  xap1 137
stonxsft emu1 53 tk  tcl1 118 vim  a1  10
strace  d1  41 tk new  tcl1 118 vse  pay1 107
strn  n1  94 tk old  tcl1 118 vtcl  tcl1 120
suck  n1  94 tkdesk  tcl1 118 vtk  xdev1 148
sudo  ap1  25 tkdraw  gra1 74 vtkdata  xdev1 148
surf  gra1  74 tkfont  tcl1 119 w3mir  n1  96
suse lx  doc1 48 tkinfo  tcl1 119 wavplay  snd1 112
suse lxe doc1 48 tkinspec  tcl1 119 webalize  n1  96
susebuch doc1 48 tkirc  tcl1 119 weblint  n1  96
susecurs xap1 137 tkmail  tcl1 119 wget  n1  96
susefax  n1  94 tkman  tcl1 119 whois  n1  96
susehilf  doc1 48 tkman8  tcl1 120 wie dvi  doc1 49
susepak doc1 48 tkppp  tcl1 120 wie geht  doc1 49
susepake doc1 48 tkrat  tcl1 120 wie html  doc1 49
suseppp  n1  94 tkstep  tcl1 120 wie info  doc1 49
susewm xwm1 154 tktex  tex1 124 wie pdf  doc1 49
svgaaout  d1  41 tkworld  tcl1 120 wie ps  doc1 49
svgalib  d1  42 tkxcd  tcl1 120 wie sgml  doc1 49
svgatext  ap1  25 toolpack  d1  42 wine  emu1 54
swi pl  d1  42 toppp  doc1 48 wingz  pay1 107
syslogd  a1  9  tracker  snd1 112 wm2  xwm1 154
sysvinit  a1  9  transfig  tex1 124 words  spl1 114
taper  ap1  25 trn  n1  95 workbone snd1 112
tcd  snd1 111 trn spl  n1  95 workman  snd1 112
tcl  tcl1 116 ts  tex1 124 wuftpd  n1  97
tcl new  tcl1 117 tvision  d1  42 wwwoffle  n1  97
tcl old  tcl1 117 tya  d1  42 wxwdemo xdev1 148
tclx  tcl1 117 uae  emu1 53 wxwdocht xdev1 148
tcpdump  n1  94 uc  pay1 106 wxwdocps xdev1 149
tcsh  ap1  25 umsprogs a1  10 wxwin  xdev1 149
teapot  ap1  25 unarj  ap1  26 wxwstat  xdev1 149
termcap  a1  9  units  ap1  26 wxxt  xdev1 149
terminfo  a1  9  unzip  ap1  26 x3270  xap1 137
terroid  pay1 106 url get  n1  95 x3d  fun1 61
tetex  tex1 123 util  a1  10 x8514  x4  128
texcad  tex1 124 uucp  n1  95 xabacus  fun1 61
texinfo  a1  10 uucpcfg  n1  95 xagx  x4  128
textutil  a1  10 uucpdoc doc1 49 xancur  fun1 62
tgif  gra1  74 vacation  n1  95 xanim  gra1  74
thrust  fun1 61 vbox  n1  95 xaos  fun1 62
tiff  gra1  74 vbox2b4  n1  96 xarchie  xap1 137

Package Series Page Package Series Page Package Series Page
xautolck xap1 137 xfnt100  x1  127 xmitsumi snd1 112
xaw3d  x1  125 xfntbig  x1  127 xmix  snd1 112
xaw3dd  x1  125 xfntcyr  x1  127 xmono  x4  128
xaw95  x1  126 xfntscl  x1  127 xmorph  gra1 76
xball  fun1 62 xformat  xap1 140 xmountns fun1 64
xbanner  xap1 137 xforms  xdev1 149 xmp  snd1 113
xbarcode xap1 138 xfract  fun1 63 xmris  fun1 65
xbench  xap1 138 xfsetup  x1  127 xmrm  gra1 76
xbill  fun1 62 xftp  xap1 140 xnetload xap1 142
xbl  fun1 62 xgal  pay1 107 xntp  n1  98
xblast  fun1 62 xgammon fun1 63 xosview  xap1 142
xbmbrosr gra1 75 xgas  fun1 64 xp9k  x4  129
xboard  fun1 62 xglint  x5  129 xpaint  gra1 76
xboing  fun1 63 xgopher  xap1 140 xpat2  fun1 65
xbomb  fun1 63 xgrabsc  gra1 75 xpdf  gra1 76
xbombs  fun1 63 xgrasp  gra1 75 xphoon  fun1 65
xcb  xap1 138 xgrok  xap1 140 xpilot  fun1 65
xcdroast  tcl1 121 xgroups  xap1 140 xpinguin fun1 65
xclass  d1  42 xhtalk  xap1 141 xplot  xap1 142
xcnf  xap1 138 xi128  x4  128 xplz  xap1 143
xcolors  xap1 138 xinetd  n1  97 xpm  x1  127
xcoral  xap1 138 xinfo  xap1 141 xprompt  xap1 143
xdelta  ap1  26 xirc  xap1 141 xpuzzles fun1 65
xdemine fun1 63 xjewel  fun1 64 xqmixer  snd1 113
xdevel  x1  126 xkeycaps xap1 141 xray  gra1 76
xdf  xap1 139 xled  xap1 141 xrayps  gra1 77
xdmbgrd  x1  126 xless  xap1 141 xrn  n1  98
xdmsc  x2  128 xli  gra1 75 xroach  fun1 65
xdoc  x1  126 xlife  fun1 64 xrobots  fun1 65
xdtm  xap1 139 xlkit  x1  127 xrolodex xap1 143
xe exec  e1  51 xlock  xap1 141 xrpm  xap1 143
xe info  doc1 49 xlogin  xap1 141 xs3  x4  129
xe lisp  e1  51 xlyap  fun1 64 xs3v  x4  129
xearth  fun1 63 xmach32  x4  128 xscheme  d1  42
xengine  fun1 63 xmach64  x4  128 xscrns  xap1 143
xephem  xap1 139 xmach8  x4  128 xselect  xap1 143
xess3  pay1 107 xmahjong fun1 64 xsession  xap1 143
xextra  x1  126 xmail  xap1 142 xshared  x1  128
xf  tcl1 121 xman  x1  127 xshogi  fun1 66
xf86  x1  127 xmcd  snd1 112 xskat  fun1 66
xf86html doc1 49 xmclock  tcl1 121 xsky  xap1 144
xfig  tex1 125 xmemory fun1 64 xsnow  fun1 66
xfileman xap1 139 xmftp  xap1 142 xsok  fun1 66
xflame  xap1 139 xmgr  gra1 76 xsol  fun1 66
xfm  xap1 139 xmine  fun1 64 xspline  xap1 144
xfmail  xap1 140 xmines  fun1 64 xspread  xap1 144

Package  Series Page Package Series Page Package Series Page
xspringi  fun1 66 ytree  ap1  26
xsvga  x4  129 z81  emu1 55
xsysinfo  xap1 144 z88  ap1  26
xtacy  fun1 66 zgv  gra1 77
xtail  xap1 144 zip  ap1  27
xtalk  xap1 144 ziptool  ap1  27
xtar  xap1 145 zircon  tcl1 121
xtartan  fun1 66 zoo  ap1  27
xteddy  fun1 67 zsh  ap1  27
xtem  tex1 125
xtetris  fun1 67
xtexit  xap1 145
xthing  fun1 67
xtoolpl  xv1 150
xtpanel  xdev1 149
xtron  fun1 67
xv  pay1 108
xv32aout  xv1 150
xv32base  xv1 150
xv32dev  xv1 151
xv32exmp xv1 151
xv32slib  xv1 151
xvga16 x3 128
xvhelp  xap1 145
xvi  xap1 145
xvier  fun1 67
xvnews  xap1 145
xvtdl  xap1 145
xw32 x4 129
xwarp  gra1 77
xwb  xdev1 149
xwpe  xdev1 149
xwpick  xap1 146
xxgdb  xdev1 150
xxprt x1 128
xzap  xap1 146
xzet  xap1 146
xzx  emu1 55
yacc d1 43
yahtzee  fun1 67
yamm ap1 26
yardsql  pay1 108
yast a1 11
ygl  xdev1 150
ypclient  n1  98
ypserv  n1  98